A Selected Bibliography of Publications by, and about, Erwin Schrödinger

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

12 June 2024 Version 1.139

Title word cross-reference

1 [Sch55g, Sch55h]. $1.75 [Bun53, LH53a]. $2.00 [Bun53]. $2.75 [Bel51a]. $24.99 [Ish20]. $29.95 [Whe17]. $39.50 [Kil90]. $6 [Hie68]. $6.50 [Kle68a]. · [dB65]. g [KS14]. ≫ [Ano11c, Lar93, Pol96]. H [KS26]. ≪ [Ano11c, Lar93, Pol96]. n [Sch44h]. Ψ [Bal68, Ken31].

-curve [KS26]. -fold [Sch44h]. -Kurve [KS26]. -rays [KS14]. -Strahlen [KS14].

// [Ano04c]. /6 [Kai85a, Kai85b]. /Herder [RFS^{+}91]. /Weisbuch [RFS^{+}91].

1 [MR87b, SvM11]. 1. [Gal03, Shi99, Ste02a, YKZ03]. 10 [Ano03a, Gaw97, Jon97, Lie97, Zei97]. 10. [Kar03, Zur99]. 100 [JS87]. 100th [BRZ88, GS87, SKD^{+}87, Sch87c]. 10th [Ano03a, TU77]. 11.

[Asp02, CD03, Thi99]. 110th [SS97]. 11th [Sch26e]. 12. [Ber02, Bla99, Kur03a]. 12th [Sch28b, TU77]. 13. [Har99, Rai03, Rau02]. 13.März [Jon03]. 14. [Fry02, Kim99, Sim03]. 14.März [Osk03]. 14th [Sch28b]. 15. [Bla03, Zei02, Zol99]. 16. [Bar03, Rau99]. 17. [Hor99, Koc02].

18. [Lei02, Orl99]. 1887 [Ano11a]. 1887-1961 [Sch61b]. 19. [Raj02, Yng99]. 1908 [Sch13j]. 1910 [Sch12b]. 1911 [Meh75, Sch12d, Sch13c, Sch13e, Sch14c]. 1912 [Sch13f, Sch13i, Sch13b, Sch14b]. 1913 [Sch13h, Sch14a, Sch20b]. 1914 [Sch14e]. 1915 [Sch17c]. 1928 [Sch28b, Sch28e]. 1929 [Pla29, Sch29a]. 1933 [Krö98]. 1933/Chadwick [Ano51]. 1935/1936 [Ano51]. 1944 [Sch44j, Sch52e]. 1952 [Sch52g, Sch98]. 1954/Mladjenovic [ARK^{+}99]. 1956 [Sch56d, Sch58b]. 1960s [ARK^{+}99]. 1960s/Frenkel [ARK^{+}99]. 1961 [Hei61b, Mat75]. 1976 [TU77]. 1985 [CW85]. 1992 [GM94]. 1993 [Ano93]. 1993/04/20 [Ano93]. 1994 [Mik94]. 1995 [Krö98]. 1997 [Gaw97, GHKO99, Jon97, Lie97, Zei97]. 1997/04/10 [Gaw97, Jon97, Lie97, Zei97]. 1997/04/12 [Pie97]. 1998 [Ano99, Ben99, Bla99, Bou99, Che99, Ghi99, Got99, GRZ99, Har99, Hil99, Hor99, How99, Kim99, LL99, Mic98a, Oes99, Omn99, Orl99, Rau99, Shi99, Thi99, Yng99, Zei99, Zol99, Zur99].

2 [MR01]. 2. [Nar03, Shi02, Zei99, Zur03]. 20 [Ano93]. 20. [Bou99, Sut02]. 2000 [Asp02, Ben02, Ber02, Cla02, Eke02, Eng02, Fry02, Ghi02, Gis02, tH02, Hor02, Jac02, Kle02, Koc02, Lei02, Pen02, Per02, Raj02, Rau02, Sel02, Shi02, Ste02b, Sut02, Wei02, Whi02, Zei02, Zic02, d’E02, dR02]. 2003 [AP03, Ano03c, Ano03a, Arn03, Bar03, Bla03, CD03, FF03, Frö03, Gal03, Jon03, Kar03, Kie03, Kur03b, Kur03a, Nar03, Osk03, Rai03, SK03, Sim03, Ste03, YKZ03, Zur03]. 2006 [GRY06]. 2006-06-06 [Ano06]. 20th [Sch96a].

21. [Ben02, Oes99]. 22. [Ben99]. 23. [Ano99]. 24. [Sel02]. 2400 [Sch48c]. 25 [Sch14e]. 25. [Wei02]. 25th [KS86]. 26. [Gis02]. 27. [Ghi02]. 28. [Per02]. 28th [Sch29a]. 29. [Eke02]. 2nd [Cha13].

3-540-96284-0 [Hen88]. 3-Color [Sch94c]. 3. [Cla02, Jon03, Kur03b, LL99]. 3/2 [Sch26e]. 30. [Pen02]. 3rd [Kle84].

4 [Ano93, Pla29, Sch30a]. 4-Color [Sch94c]. 4. [AP03, Frö03, How99]. 4.1.1961 [Hei61a]. 4/20/1993 [Ano93].

5 [MR87b, MR01, Sch34a]. 5. [Che99, FF03, Osk03, Whi02]. 50 [Ben34]. 50th [DS83]. 52 [Bal86]. 5th [Sch28b].

6. [Got99, Kie03, d’E02]. 60 [Sch13d]. 60th [Wes57]. 68 [Bun53].

7 [Whe17]. 7. [Ghi99, SK03, Zic02]. 7.80 [Deh36]. 70th [Gaw97, Jon97, Lie97, Pie97, Zei97]. 75 [Ano04a, GM94]. 7th [Sch28b].

8. [Arn03, Hil99, Hor02]. 80th [Ano07].

9. [tH02, Omn99, Ste03]. 96377-4 [Hen88]. 978 [Whe17]. 978-0-19-067963-7 [Whe17]. 978-1-108-81978-7 [Ish20]. 978-1-78634-490-8 [Bus20]. 9th [Sch20a].

= [BD92, Sch84b].

abandonment [Sch86i]. Abb [Deh36]. abermals [Haa34]. Abhandlungen [Pla28, Sch13e, Bor63a, Sch27a, Sch78, Sch84c, Sch84d, Sch84e, Sch84f, Sch84b]. Abraham [Sch14b]. Absolute [BS35]. abundance [Sch87i]. academic [Sch13f]. Academy [KK75, LMOS15]. accelerated [Lei02]. Accidental [Sch44h]. according [Sch27c]. accumulators [Sch13a]. achievements [SS97]. acoustics [Sch17e, Sch17d]. active [Hil99]. Actors [FW20]. actuelle [Sch53d]. Adam [DT18]. Adami [Sch20a]. adaptation [SM35]. Additions [Gis06]. Address [Sch56a, Sch29b, Sch29c, Sch32a, Sch29a]. Adolf [Jah56, CL83]. Adresse [Sch29a]. Advanced [Sch52e, Ish20]. adventure [MB11b]. aesthetics [Mil03]. Affine [MS45, Sch44f, Sch45b, Sch48a, Sch47b, Sch48e, Sch48f, Sch46a]. Affinity [Sch47e, Sch52a]. afraid [MZP97]. African [Arm46]. After [Ano11c, Kae31, RY13, Ano97, Ano11a, AS96, Frö11, GGW94, Kae48, Krö98, Per02]. Afterlife [Bou12]. Age [Pur18, Stu18, Arm46, Han77a, Sch54e, Gil09]. air [Sch10]. Akademie [KK75]. Akademische [Deh36]. akademischen [Sch13f]. Akkumulatoren [Sch13a]. Aktualität [GB08]. Akustik [Sch17e, Sch17d]. Albert [Adl03, Ano05, Mic68, Kle05, Meh87c, Prz86, Prz11]. Alfons [Krö98]. Alfred [Adl03]. Algebra [BS41]. algebras [Jon97]. algorithm [Vis91]. alive [Wra09, Bus13]. alle [Sta04]. Allgemein [Sch84f, Sch18e]. Allgemeine [Aud87]. Allgemeinen [KPMN20]. Am [Bal86, Ano04a, Ano06, Asp02, Ben02, Ber02, Cla02, Eke02, Eng02, Frö03, Fry02, Gal03, Ghi02, Gis02, tH02, Hor02, Jac02, Jon03, KPMN20, Kle02, Koc02, Lei02, Nar03, Osk03, Pen02, Per02, Raj02, Rau02, Sch14e, Sch29a, Sel02, Shi02, Ste02b, Sut02, Wei02, Whi02, Zei02, Zic02, d’E02, dR02]. ambiguity [Sch38a]. Ambition [Jon08]. Amherst [Kle68a]. among [Leg90]. amplitude [Sch22c, Sch22c]. Analyses [BD95, Bit93]. Analysis [Sch20c]. analytical [BWE^{+}29]. analytischen [BWE^{+}29]. Anderson [Ano51]. Angewandte [Sch12b]. Angle [Sch54c]. anläßlich [KS86]. Anmerkungen [Sch33a]. anniversary [Ano03a, BRZ88, DS83, KS86, Sch29a, Wes57]. Annus [Frö11]. anomalen [Sch13g]. anomalous [Sch13g]. anomaly [Jac02, Sut02]. Anpassung [SM35]. Anschaulichkeit [dR97]. anthropological [Bar00]. anthropologische [Bar00]. Antrittsrede [Sch32a, Sch56a, Pla29, Sch29b, Sch29c, SP29, Sch61b]. Anwendung [Sch26f]. Apology [Pra93]. App [Sal14]. appearance [Oes99]. application [Cer93, Sch26f]. Applications [Hen88, PW35, RFS^{+}91, Meh88, MR01]. Applied [Sch12b, BJK03]. approach [BJK03, Sch89i]. approximate [SM35]. Arbeit [Sch31a]. Arbeiten [ARK^{+}99, Sch24a]. Archimedes [Pic08]. areas [GSL26]. argon [Sch21c]. arithmetic [Sch86a]. arithmetische [Sch86a]. Arnold [Han62]. Art [Sch35o, Sch96d, Whe17, Sch30h]. Arthur [Adl03, Deh36]. Article [Sch26d, Sch17d]. Articles [Sch21c]. Ash [Krö98]. Ash/Alfons [Krö98]. Aspect [Ano45a, Ano45b, Del45, Sch44l, Sch45e, Sch62b, Sch67, Sch86e, Sch86j, Sch92c, Sch12]. aspects [Meh75, Par88, Sch35j]. Assessment [Whe57]. assisted [Zur03]. Assuming [Gri22]. astounding [Haw11]. astronomy [Sea87]. Atmosphäre [Sch13h, Sch17e, Sch17d]. atmosphärischen [Sch12a, Sch12e]. atmosphere [Sch13h, Sch17e, Sch17d]. atmospheric [Sch12a, Sch12e]. Atom [Ano35b, Dro20, MMKW96, Rai03, Rod19, Sch17a, Sch17b, Smi23, Tau96, LKS^{+}05, Sch24e]. Atom- [Sch17a, Sch17b]. Atom/mesoscopic [Rai03]. Atombau [SW32]. Atombegriffs [Sch56a]. Atome [Sch24e, Sch33f]. Atomenergie [Sch55a]. Atomic [Ano96, ARK^{+}99, BBSR69, FR13c, HSD34a, SW32, Sch56a, Kle02, Sch17a, Sch17b, Sch25c, Thi87, vM92, Bro36, Sch55a, Sch57a]. Atomism [Rüg88]. Atomisten [Sch55b]. atomists [Sch55b]. Atomrumpfes [Sch25c]. Atoms [Sch26l, AP03, Bla99, Har99, Kim99, Sch33f]. Atomtheorie [Ben34, HSD34a, Wil36, vM92, Sch57a]. attempt [Dor87, Sch32c, Sch21b]. attempts [Sch13c]. attitude [Kra92]. attuale [Sch59b]. auch [Wes57]. audit [Sch13e]. Audretsch [RFS^{+}91]. Aufgeben [Sch86i]. Aufl [Deh36, Sch20a, Haa34, Sch20c]. Auflage [Sch12b, Sch13f, Sch14b, Sch26e, SW32]. Aufnahme [KK75]. Aufsatz [Sch17d]. Aufsätze [Sch84f]. auftretendes [Sch18d]. Augen [SM89, Sch99b]. Auges [Sch24g]. August [Bla03, GM94, Mik94]. Auguste [Dic85]. Ausgewählte [Bor63a]. außerhalb [Kai85a, Kai85b]. außerordentlicher [Hen12, SvM11]. Ausstellung [Ano06, Ano11b, KS86, KDK^{+}98, KDK^{+}99]. Ausstellungseröffnung [Ano11b]. Austria [BRZ88]. Austrian [Ano06, Ano11b, McL99, Sch45a, Sch45c, Sch45d, Sch94a]. auteur [Sch33c]. Authorized [Sch12d, Sch13i]. Autobiographical [Pra93, Sch92c, Sch12, Sch35i]. Autobiographie [Sch06]. autobiography [Sch06]. Autorisierte [Sch12d, Sch13i]. autres [Kle05, Ano05]. Autumn [Sch89d]. avant [Sch33c]. avant-propos [Sch33c]. Award [Ano33].

B [Bus20, Kai85a, Kai85b, Sch12b, Sch12d, Sch14b, Sch17c, Sch33e, KS14]. baby [McG99]. ball [Fel06]. Ballroom [Ano11b]. Balmer [Sch26f]. Balmerlinien [Sch26f]. Band [ARK^{+}99, Kai85a, Kai85b, Sch13i, SvM11, Sch84c, Sch84d, Sch84e, Sch84f]. Bär [Sch41e]. Bargmann [MS45]. barrières [Dat38]. barriers [Dat38]. based [Haa34, Sch21b, Sch87b]. bases [Sch35n]. basic [Sch35e, Sch86b]. basis [Fri89]. Battled [Sau16, Hal15]. Bd [Sch12b, Sch12d, Sch13e, Sch20a, Sti60]. Bde [Hen12]. Be [Far03, BS55, BS56, EPR35, Far02, Sch58a]. Beacon [Bun53]. beams [Sch24f]. Bearbeitnng [Sch13i]. Bearbeitung [Sch12d]. Beautiful [Far03, Far02]. Becker [DT18]. Become [Kae31]. Bedeutung [Sch55c]. Bedingung [Mül40]. Been [Kae31]. before [Han79a, Meh87a, vM92]. Began [Kae48]. Begegnung [Sch54a]. beginning [Gef14b]. behalf [How99]. Behandlung [Sch13a]. Behavior [Ano03b, RGS32]. behind [Pic08]. beim [Sch13f, Sch25a]. Beiträge [Sch90, Sch79, Sch84c, Sch84d, Sch84e, Sch97, Sch08b, vM92, Sch12a]. belief [Bre84b]. Bell [Asp02, Ber02, BZ02, Cla02, Eke02, Fry02, Ghi02, Got99, Hor02, Per02, Sch14a, Shi02, Ste02a, Wei02, Zei02, Zic02, d’E02, Ste02a]. Belong [Yox79]. bemerkenswerte [Sch23b]. Bemerkung [Sch24a, Sch31a]. Bemerkungen [Sch25a]. Bericht [BWE^{+}29]. Berlin [Hen88, Sch17c, Ano11b, KK75, LMOS15]. Berliner [KK75]. Berlins [LMOS15]. Bermbach [Sch13a]. Bertotti [Pol96]. Beruht [Sch87b]. Besonderheit [Sch48d]. besonders [Sch19b]. Besprechungen [GNK^{+}29, HF46, HSE31, KSP32, KGS25, SWK^{+}32, SvMFT35]. Bestimmt [Ein27]. betrachtet [SM89, Sch99b]. Betrachtungsweise [Sch89i]. betreffend [Sch19b]. Between [Stu18, Bal92, Bit96a, Che92, Che99, Kox13, Nef92, Sal01, Sch24f, Sch35g, Sch36c, Sch47e, Sch55f]. Bewegung [Ein27, Sch30e, Sch15b]. Bewegungen [Gar53]. Bewußtseins [Sch86a]. Beyond [Rod19, Bla03, Dro20]. Big [Hür22]. Bilder [SKD^{+}87, KDK87, KBG11]. Biographie [Moo12, Moo15, Paw12]. Biographien [Kai85a, Kai85b]. Biographies [FW20, McL99]. biography [Bra92]. biologie [Sch86d, Sch86f]. Biology [Yox79, HD13, Pau87, Per87, Sch86d, SF11, Wha11, MO95]. biomolecules [Kar87]. Birth [Bal68, BD95, BD92, DL82b, BRZ88]. birthday [Ano07, Gaw97, Jon97, Lie97, Pie97, SKD^{+}87, Sch87c, Wes57, Zei97, GS87, SS97]. birthday/The [Wes57]. bis [KK75]. Bitbol [BD95, Cus96, Lar90, Lar93]. blade [Bel07]. Blueprints [Bou12]. bodies [Sch13i]. body [Wig67b]. Bohm [Hil99]. Bohr [Adl03, ARK^{+}99, Bro36, Che99, How99, Lak96, Deg89, Che92, CG14, Kra11, Kra13, MR82a, MR82b, Meh87c, Plo10, Sch22b, Sch24b, Shi83, Shi86, Wei14, vCK17]. Bohrs [Sch24b]. Bohrsche [Sch22b]. bold [Kur03a]. Boltzmann [Ano06, GRY06, Fla87, GRY06]. Boltzmanngasse [Ano06]. Book [Ano90, Bal68, Bar59, Bas56, Bel51a, Bus20, Cor52, Cus96, DT18, Dro20, Fre12, Hei68, Hen88, Hen12, Hie68, Ish20, Kil90, Kle65, Kle68a, Klo63, LH53a, Lin57, M.47, Mic68, Moo90, MF91, Pol96, Pol59, Sau16, Sch25d, Sch26e, Sch29h, Sch30a, Sch31c, Sch32c, SW32, Sch33e, Sch34a, Sch35j, Sch40b, Sco67a, Tem54, Ter90, Tho55, Wal54, Wes91, Whe17, Ano87c, SS53, HF46]. Books [Bur65, L.88, AFG90, Sch20a]. Boolean [BS41]. Borda [Sch13c]. Born [ABdB^{+}53, Sch17c, Sch26k, Sch66b, Sch35c, Sch43a, Sig96, dG14]. bosons [Lew87]. Boston [Bun53, CW85]. Boundary [Fro38]. brain [Loc89]. brief [How99]. Briefe [Klo63, OGT92, Prz63, PS63, Kil65, Kle65]. Briefen [Dic85]. Briefwechsel [ARK^{+}99, Hen12, SvM11]. Brillouin [Fow29]. Bringing [Yam97, qYL17]. Broadening [Sch55f]. Broglie [Ano35a, Deh36, Fow29, Gar53, Haa34, RFS^{+}91, Ano35a, BJK03, Cro01b, Dar11, Dar13, Gar53, Haa34, Han77b, Hou16, RF69, Rei44a, Rei98a]. Broschur [Kai85a, Kai85b]. Brownian [Sch15b]. Brown’scher [Sch15b]. Brückner [RFS^{+}91]. Bruno [Bus13, Pol96]. Bruxelles [Far01]. BSB [Kai85a, Kai85b]. bubbles [Sch15a]. Buchbesprechung [Deh36, RFS^{+}91, Wil36]. Bücher [Ben34, Sch20a, ARK^{+}99]. bucket [BP95]. Build [Sal14]. Bündeln [Sch20i]. Bunge [BM01, Mah01]. Bunin [Ano33]. Buonaiuti [Jah56]. butterfly [Sch87c].

C [Adl03, Bur65, Bus13, Jah56, Sch21c]. Cage [Ano12]. calculating [Bre84b]. Calculus [Eck26]. called [Che99]. Cambridge [Bel51a, Bun53, Ish20, Kil90, Krö98, LH53a, Sch58b, SS53, Ter90]. campo [Sch35p]. Can [EPR35, He20, Mer08, Sch87c, Sch24e]. capillary [Sch15a]. carbon [Sch21c, Sch54e]. care [Sut08]. Carnap [Nef92]. Carvisy [Sch14c]. Case [Han79a, Joh13, Ben92, kA13]. cases [Cec01]. Cassini [Sch13c]. Cat [Ahl96, Ano03b, Ano15, Bru80, Bun01, Bus13, CS96, Gis06, Hoa96, Kle96, Koc78, MMKW96, Mor96, Sau16, Sei96, Tau96, Tes86, Tes00, Tri80, WZ51, Ano96, Ano10, Axo89, Bal90, Bal99, Bal00, Bal03, Bal11, Bal86, Bla00, Bli01, Car13, D’A09, Ebe09, Gri84b, Gri84a, Hal15, He20, KKS^{+}11, Leg84, Leg90, Leg00, Leg09, Leg13, LKS^{+}05, Loc92, Loe84, MZP97, Mer08, Per91, Pie11, Vil86, Wil99, Win13a, Win13b, Wra09, qYL17, Ald99, Ano12, SvM11, Yam97]. catalog [KS86]. caterpillar [Sch87c]. catflap [Ste91]. Cats [Mac24, Gri13a, Kni92, OJTBG07, kA13]. Causality [Yng99, Sch33a]. Causation [Sch29d]. cavity [Har99, Kim99]. ce [Miq07, Sch86d, Sch86e, Sch86f]. Cecily [Sco67a]. celebration [Kil87b]. celebre [D’A09]. cell [Sch44l, Sch45e, Sch62b, Sch67, Sch86e, Sch86j, SM89, Sch92c, Sch99b, Sch12, Ano45a, Ano45b, Del45]. cellule [Sch86e]. centenary [Kil87b]. Central [Ano06, Ano11b, KDK^{+}98, KDK^{+}99, Gri22]. century [Bra09, D’A01b, D’A00, Sch96a]. ceremony [Ano11c]. cette [Sch33c]. Chadwick [Ano51]. Chair [ABdB^{+}53]. challenges [Zic02]. Chance [Sch29d]. Change [Krö98, AS96]. changes [RGS32, Sch13b]. Channels [Col98]. chaos [Zur99]. characterisation [Eng02]. Charge [Gao18, PS51, Sch44g, Sch54b, Raj02]. Che [Sch88a]. chemical [Fuk87]. Chemie [BWE^{+}29]. Chemistry [BSM06, Joh13, NM12, PW35, Sch14c, Sch48b, Buc87, Len14, Pau87, BWE^{+}29]. Chemists [Joh13, Sut08]. Chernogolovka [Mik94]. childhood [Sch89c]. Chimerical [Bru80]. chimie [Sch14c]. chiral [Jac02, Sut02]. Christopher [Bus13]. Chromosome [Wil16]. chromosomes [Bre84b]. cipher [WH17]. circus [Kim99]. civilization [Arm46]. Clairaut [Sch20b]. Clara [Wes57]. classic [HT10]. Classical [Sch95b, Ben02, Oes99, Orl99, Sch25b]. Classics [Pra93, Sch13c, Sch20b]. classification [Car93]. clock [Ano06, Ano11b, Sch54e]. cloning [Zur00]. closer [Bal03]. Closing [Ano99, Pen02, YKZ03]. clöture [Lar93]. Cloud [Gao18]. clusters [Sch20i]. Co [Sch60a]. Code [Pag82, WH17]. code-script [WH17]. Coherence [Kle02, Kur03a, AP03, Kar03, Wha11]. coherent [FF03]. Coin [FR13a].

Coincidences [Sch44h]. collaboration [Ber02]. collapse [Gis02, Orl99]. Collected [Sch28a, Sch82a, Sch84c, Sch84e, Sch84f, Sch84b, Sch03, Bel87b, Sch13e, Sch78, Sch84d, Fow29, Kle84]. Collection [Sch20c, Mic68]. collective [FF03]. College [Sch58b]. Colloque [BD95]. Colloquium [CW85]. Color [Sch94c, Sch20e, Sch20f, Sch20g, Sch25g, Sch26d, Sch20d, Sch23a]. colorimetry [Sch25i]. colors [Sch25h]. Combinatorial [Sch51b]. Combined [Sch54b]. comes [Bal03]. coming [Han77a]. Commemoration [BZ02, GRY06]. Comment [Bal86, Sch31a]. commentaries [Yan13a]. Comments [Kle65, Sch33a]. Commilitonen [Sch59a]. communication [Zol99]. communications [BD95]. commutative [Nar03]. Complementarity [Hor99, Mac24, Meh87c]. complete [EPR35, Rau02]. completely [Ein27]. Completeness [D’A92a, D’A92b]. Complex [Ano03c, RFS^{+}91, Arn03, Kur03b, Kur03a, RA11, Yan87, Yan13b]. component [Sch18a]. compound [Loc89]. comprehensive [Sch04]. Compton [Sch27e]. Comptoneffekt [Sch27e]. Computation [Ben95]. computational [Len14]. Computer [Sal14]. computers [Len14]. con [DL82b]. concepción [Sch88c]. Concept [Sch56a, Sch58a, Bit92, Sch30c]. Conception [Han62, Sch53e, Sch82d]. concepts [Kie03, Yng99]. Conceptual [Sch29f]. concerning [Sch19b]. condense [Lew87]. Condensed [ARK^{+}99, CD03, Kar03]. condition [Mül40, Sch22b]. Conditions [Fro38, Ken31]. conduction [Sch10, Sch12f]. conductivity [Sch35m, Sch36b]. Conference [BZ02, Mik94, Sch56d, GM94, Far01, BV09]. conferences [Meh75]. conflict [Che92]. conflit [Che92]. connaissance [Sch35n]. Connection [Sch45b]. Connections [MS45]. Connexion [Sch44f]. consciousness [Sch86a]. consensus [Lev17b]. Consequence [Gri22]. Consequences [Smi23, Thi87]. Conservation [Sch44k, Sch27b]. considerations [CD03]. considered [EPR35]. consistency [Sch20i]. Constant [BS35]. contemporanea [Sch87d]. contemporary [Sch87d]. content [Sch13h, Sch19a]. contingency [Sig96]. Continuity [D’A92a, D’A92b]. continuous [Sch26a, Sch29g]. Contributions [Pie97, Sch35c, Pau87, Sch84d, vM92, Sch12a, Sch79, Sch84c, Sch84e, Sch97, Sch08b]. Control [AP03]. Controlled [Kur03b]. Controlling [Ano03b, GGW94]. controverse [Nef92]. Controversy [Bit96a, DL82b, Nef92]. Cook [Bus20]. Copenhagen [Che99, Bel92, Bit96a, DL82b, He20]. core [Sch25c]. corpusculaire [Sch41d]. corpuscular [Sch41d]. correlations [Kle02, dB65]. correspondance [Bal92]. correspondence [Bal92, Han79b, Prz11, Sch27d]. cosè [Sch88a]. Cosmic [Pag82, RFS^{+}91, Sch51b]. Cosmogony [Bur65]. Cosmological [ARK^{+}99]. Cosmology [Rüg88, Kib87]. Cosmos [Sch56b]. Coulomb [Sut08]. Counting [Sch51b]. coupled [Sch14f]. course [Sch44j, Sch46b, Sch52e]. cousins [Leg84, Leg09, Leg13, Ebe09]. covariant [Sch18e]. Crazy [Arr06]. Create [Sau16, Hal15]. Creating [SF11]. Creation [Bur65, FF03, Hen88, LKS^{+}05, Pur18, Meh87a, Meh87b, MR01, vM92]. Creativity [ADS83]. Creator [Ful15]. creators [Str11]. Crisis [Sch56a].

Critique [Ano05]. crossroads [BV09, ET07]. Crucial [Kra82]. cryptography [Eke02]. crystal [Sch17c]. crystals [SK03]. Császár [Sch24a]. cuántica [HBS96]. cuánticas [W^{+}87, WdC02]. Cuestiones [W^{+}87, WdC02]. cultural [HY87]. Culture [Sch31d]. cura [Pol96]. Curi [Pol96]. Curies [Kae48]. Curiosity [Ald99]. Current [Sch54b, Sch53d, Sch59b]. curve [KS26]. curves [Sch24g]. Cyclotrons [Row09]. czastka [Sch60a]. Czech [SS53].

D [Bus20, Sch21c]. Dämmerungsempfindung [Sch25h]. dancing [Zuk79]. Daniel [Ish20]. dans [Sch33b]. Darrigol [BD95]. daughter [Gef14b]. dawn [Wei14]. daylight [Sch20e, Sch20f, Sch20g]. days [Blo76a, Blo76b]. DDR [Kai85a, Kai85b]. d’E. [SB90b]. Dead [Bun01, Bus13]. Death [Bar00, GRY06, KS86, Wes57]. debate [BPW^{+}71, Got99, Kox13]. Debye [Sch14g]. Debyeeffekts [Sch14g]. decay [Sch18c]. deceased [Sch25f]. December [GHKO99, Sch52g]. decimal [Sch13c]. Decline [Sch35d]. Decoherence [Ano03c, Kie03, Omn99, Zur99, Arn03, Bal00, CD03, FF03, Kar03]. d’Eddington [Sch38d]. dedicated [Ano11a]. Deeper [Sei05]. defects [Kib87]. defense [Sch89i]. Defining [ARK^{+}99]. definition [Sch25a]. definitive [Fry02]. deformable [Sch13i]. deformierbaren [Sch13i]. degeneracy [Sch24d]. degrees [Sch31a]. d’Einstein [dB65]. del [D’A09, Sch26m, Sch35p, Sch86g, Sch87d, Sch88c, W^{+}87, WdC02]. Delambre [Sch13c]. Delft [Bla03]. delivered [Sch28b, Sch52e, Sch58b, Sch44j]. della [DL82b, Pol96, Sch59b, Sch87d]. dell’universo [Sch37b]. demonstration [Ano10]. Denkanstöße [Sta04]. Denken [Sch86i]. Density [Gao18]. dependent [Vis91]. Derivation [Rei44a, Rei98a]. Descendants [Jac02]. Description [Bro36, EPR35]. descriptive [Wes83]. Dessauer [Jah56]. detail [Sch87i]. detection [Sch29g]. détermination [Che92]. determine [Ein27]. Determining [Sch41b]. determinism [tH02]. deutsche [Sch12d, Sch13i]. deutscher [Sta36]. Deutung [Sch21b, Sch32c]. Developed [RF69]. Development [Hen88, MR87b, Sch87i, Wes91, Meh75, MR01, WH17]. device [Sch19b]. Devices [Tes86]. Dezember [Sch52g]. dezimalen [Sch13c]. diachronic [Mel88]. diagrams [Car93]. dialogue [Bel92, Bel99, Kra13, Sch75b]. Dice [Ano15, Sau16, Hal15, MB11b]. Dick [Dic85]. Dictionary [DG99]. did [Dar11, Dar13]. Dielectric [Sch18b]. dielectrics [Sch12c]. Dielektrika [Sch12c]. Dielektrizität [Sch18b]. Dienstag [Ano06]. Dieter [Kai85b, Kai85a]. difference [Che99]. different [GSL26]. Differential [Sch35j, Sch52a]. Differential- [Sch35j]. difficulties [MR82a, MR82b]. Diffusion [ARK^{+}99]. Digital [MZ13]. Dimensionen [Bar00]. dimensions [Bar00]. Dirac [Ano51, Ben34, Deh36, Haa34, Lak96, Wil36, Ano33, Ano35a, DL82a, DS83, Haa34, RA11, Sch32b, Sch33b, Sch41a, Sch52d, SM35]. Dirac-Gleichung [SM35]. Dirac’sches [Sch32b]. directions [Sch28d, Sch29j, Sch32d]. Discipline [BSM06]. Disciplines [Len14]. discontinuity [Kra13]. discourse [Sal01]. discover [Dar11, Dar13]. discovered [Dar11, Dar13]. discoveries [Bra09]. discovery [Kae48, SvM11]. Discussion [Sch35g, BS56, Kur03a]. Discussione [HSBA60]. Discussions [Meh87c, HSBA60]. Diskurs [Sal01]. dispersion [Sch13g, Sch13g]. dispute [Sal01]. Dissertation [KPMN20]. Dissipative [Frö03]. Distant [Sch45b]. distribution [Sch12e]. Divide [Bus13]. DM [Hen88]. Do [FS13, Lew87, Sch49a, Sei05]. d’object [Bit92]. Dobzhansky [Adl03, Cec01]. doctorate [Sch29a]. Documents [OGT92, SKD^{+}87]. Does [Yox79, MB11b, Ein27]. dog [Mon09]. Doktorjubiläum [Sch29a]. Dokumentation [Sch90]. Dokumente [KDK87, KBG11, OGT92, SKD^{+}87]. dominance [dR02]. Doppler [Sch22b]. Dopplerprinzip [Sch22b]. double [Ano04b]. Doubles [Tau96]. d’ouvrages [BD95, Bit93]. Dowling [Sal14]. Dr [Sch13a, Sch13j, Sch14b]. Dr. [Ano34]. dreams [Haw11, Kur03a]. Dreifarbentheorie [Sch25g]. Dresden [Sch13j]. Dritte [Sch26f]. Drop [Kae31]. Dualism [Sch86i]. Dualismus [Sch86i]. Duality [NM12]. Dublin [Cus96, Sch52e, Kir95, McC87b, Sch95a]. Duhem [Sch13b]. Duino [GRY06]. d’un [Sch82d]. d’une [Nef92]. durch [Sch29g]. durchdringenden [Sch12e]. during [GRY06, Hal87]. duties [Sch30f]. dynamic [Ste03]. dynamical [Ghi99, Ghi02]. Dynamics [Eck26, RFS^{+}91, Sch43a, Göt92a, Kar87, Sch13i, Sch14f, Sch14e, Sch18d, Sch31g, Sch17c]. Dynamics/Audretsch [RFS^{+}91]. Dynamik [Sch13i, Sch14f, Sch14e, Sch17c, Sch18d].

E. [Bas56, Ber88, DL82b, SB90b, Klo63]. Eamon [McC87b]. Early [Cla02, Hen88, Jam74a, Blo76a, Blo76b, SF11, Whi02, Meh88, MR01]. Earth [Sch20b, Sch54e, Sch44a]. ebb [Sch13e]. Ebbe [Sch13e]. écrites [Sch33c]. ecuaciones [Sch35p]. Ed [Hie68, Sch20a, Kle84]. Ed. [Sch20c]. Eddington [Adl03, Sch37b, Sch29h, Sch32c, Sch38d, Sch40b, Sch44k]. Eddington’sche [Sch37b]. Edinburgh [ABdB^{+}53]. Edited [Sch13e, Sch20b, Kil65, Kle65, Mic68]. Edition [Sch12b, Sch26e, Sch13f, Sch14b, SW32]. edn [Cha13]. Education [Whi02]. Effect [Bar03, MS44, RGS32, Sch14g, Sch26f, Sch27e, Eps26]. effects [Kle02]. Ehrenfest [KS26]. Ehrenfestsche [KS26]. Eigenfunctions [Sch41b]. Eigenschaft [Sch23b]. Eigenschwingungen [Sch38b]. eigentlich [Sch30g]. Eigenvalue [Sch41c, Sch77, Sch26f, Sch26g, Sch26h, Sch26i, Sch29k]. Eigenvalues [Kir95, Sch41b]. Eigenwertproblem [Sch26f, Sch26g, Sch26h, Sch26i]. Eigenwertspektra [Sch29k]. eighty [Frö11]. einatomigen [Sch26b]. Einfluß [Sch12f]. Einführ [Ano35a]. Einführ-ung [Ano35a]. Einführung [Deh36, Haa34, RFS^{+}91, Sch14b]. Einführung/Gaisser [RFS^{+}91]. Einheit [Sch86a]. Einiges [Sch24c]. Einladung [Ano06]. einleitende [Ano11b]. Einstein [Adl03, Ano05, ARK^{+}99, Hie68, Kil65, Kle65, Klo63, Mic68, Prz63, Prz86, Sch26m, Acz01, Ano15, Deg89, Bal92, Bel64, CG14, Fin86, Gef14b, Hal15, Han79b, Hou16, Kle05, MS45, MR82a, MR82b, Meh87c, Prz67, Prz11, RA11, Sau16, Sch26m, Sch29i, Sch54b, Sch86g, Sci61, WB00, dB65, vCK17].

Einstein-Fields [Sch54b]. Einsteinschen [Sch26m]. einzelnen [Sch23b]. either [Bus13]. elastic [Sch14f, Sch14e]. elastisch [Sch14f]. elastischen [Sch14e]. Electric [Sch54b, Sch13g]. Electricity [Sch14b, Sch10, Sch12a, Sch30g, Sch20a]. Electrodynamics [Sch52d]. Electromagnetic [RFS^{+}91, Sch35c, Sch21c, Sch35p]. electromagnético [Sch35p]. Electromagnetism [Sch44e]. Electron [Gao18, Sch43a, Smi23, Sch23b, Sch32b, Sch32e, Sch33b, Sch32e, Sch31a, Sch31g]. Electrons [Sch37a, Sch49a, Lei02, Sch12f, Sch23b]. elektrischen [Sch13g]. Elektrizität [Sch10, Sch14b, Sch20a, Sch30g, Sch12a]. Elektron [Sch32b]. Elektrons [Sch31a, Sch31g]. element [Rig02]. Elementarmathematik [Sch12b]. elementarna [Sch60a]. Elementarteilchen [Sch51d]. Elementary [Sch12b, Sch50c, Sch51d, Sch51e, Sch53b, Sch60a]. elements [Bra92]. Elemér [Sch24a]. Elision [SB90b]. Emigration [Krö98]. Emigré [Krö98, AS96]. emission [Sch21c]. Empfindlichkeitskurven [Sch24g]. Encyclopedia [Sch12b]. endlicher [Sch22c]. Energieaustausch [Sch27c]. Energieimpulssatz [Sch27b]. Energieinhalt [Sch19a]. Energiekomponenten [Sch18a]. Energien [LMOS15]. Energies [LMOS15]. Energiesatz [Sch24b]. Energiestufen [Sch26b]. Energy [Sch58a, Seg85, Sch18a, Sch19a, Sch24b, Sch26b, Sch27b, Sch27c, Sch55a]. Engel [Sch13e]. Engelmann [Sch13c]. Engendered [Sch54b]. engineering [Har99]. enlarged [Sch13f]. ensayo [Cam88]. ensnared [Bal90]. ensuing [Hal87]. entangled [Ano10]. Entanglement [Acz01, Ano03c, Bou99, CD03, TWD03, Eng02, Kur03b, Kur03a, Rai03, Wha11, Win13a, Win13b, Gil09]. Entäuschte [Sch89f]. Entdeckung [Hen12, SvM11]. Entgegennahme [Wil36, HSD34a]. entre [Bal92, Che92, Nef92]. Entropie [Sch00]. Entropiedefinition [Sch25a]. entropy [Moo87, Sch25a, Sch00]. Entstehung [vM92]. Environment [Zur03, PS33]. enyémhez [Sch66b]. Enzyklopädie [Sch12b]. épaisseur [dB65]. Epicuro [Cam88]. Epicurus [Cam88]. episodes [Bra09]. Epistemological [Oes99, GRZ99]. Epistemology [Plo10, Che99]. Equality [Sch35h, Sch35k]. Equation [Aja15, ARK^{+}99, Cha13, FR13b, Fle18, Fro38, Gri22, Kra82, Sch42a, Sch55h, Ano04a, Dar11, Dar13, Dat38, Dor87, Fri89, Haw87, Mik94, Mil03, Osk03, Rei44a, Rei98a, Sch55g, Ses13, SM35, TU77, Thi87, Vis91, Dat38, Cer93, Ish20]. Equation/Gaponenko [ARK^{+}99]. Equations [Eck26, Far03, Sch41a, Sch42c, BJK03, Far02, Gar53, Hou16, RA11, Sch35j, Sch35p, Sch18e]. equilibrium [Sch24f]. Equivalence [Dir64, Per64, dG14, Gal03]. equivalent [Sel02]. Eranos [Jah56]. Erdalter [Sch54e]. Erdgestalt [Sch20b]. Erfassung [Sch29g]. Erfüllbarkeit [Sch25b]. Ergebnisse [Sch17a, Sch17b]. ergodic [Nar03]. Erinnerungen [Sch87a]. erkenntnistheoretische [Sch29f]. Ernesto [Jah56]. Ernst [BIT01, FS06]. Eröffnung [Ano06, Ano11b, Sch56d, Ano93]. Erotica [Mil03]. Erste [Sch26g]. Erwiderung [Pla29, SP29, Sch61b]. Erwin [Ack55, Adl03, Ano35b, Ano45b, Ano87c, Ano93, Ano03a, Ano04c, Ano06, Ano11a, Aud87, Bec54, Bel51a, Bel51b, BD95, Ben48, Bro36, Bun53, Bur65, Bur55, C.35, Cam88, Car45, Dav57, Dic85, Dir61, Ede55, Fin53, Fre61, Fre12, Ful15, GRY06, GM94, GHKO99, GB08, Han62, HF46, Hei63, Hen88, Jah56, Kai85a, Kir55, KK75, Lar90, Lar93, LH53a, LH53b, Lon52, M.47, McC51, McC56, Mic68, Moo90, N.46, Nag52, Opp59, Pla28, Pol96, Pos62, Pra93, Prz86, Rob56, Sch90, Sei61, Spr55, Sti60, Tem54, Thi11, Wes91, Wes57, Wig55, Yan56, Ano61, Ano87a, Ano87b, Ano87c, Ano88, Ano11b, Ano11c, Atk87, Aud87, Bal68, Bar59, BC94, BD92, BG61, Bor63b, Bra92, Cam88, Cor52, Cro01b, Cus96, Dic85, Ful15, GGW94]. Erwin [Göt92a, Göt92b, Gri12, Gri13b, GHHL10, G^{+}11, Hal87, Han77a, Han77b, Han79a, Hei61a, Hei61b, Hof84, Kai85b, KS86, KDK87, KDK^{+}98, KDK^{+}99, KBG11, Kra82, Lin57, Mal12, Mat75, McC87b, McL99, Meh87a, Meh87b, MR87a, MR87b, Meh87c, Meh88, MR01, Mel88, MB11a, Moo94, Moo96, Moo12, Moo15, OGT92, Pér20, Per87, Pla28, Pol59, Prz11, RY13, Rüg88, Sch27d, Sch35i, SKD^{+}87, Sch87c, Sco67a, Sco67b, Sof14, Thi13, Tho55, Wal54, Wes83, Wes57, Wil16, Yan87, Yan13b, dR97, dR01, BD95, Hei68, Kle68a, Kle68b, Paw12]. Erwin-Schrödinger-Institute [GHKO99]. Erwina [SS53]. erzählt [FS06]. erzeugten [Sch14d]. escritos [Sch01, W^{+}87, WdC02]. ESI [Ano11a, Ano04a, Ano04c]. Esquisses [Bit92]. Essai [SB90b]. Essay [SB90b, D’A09]. Essays [ADS83, CL83, D’A00, Mah01, Sch96a, BM01, D’A01b, Pla28, Sch56e, Sch56f, Sch95a, Wig67d, Sch27a, Cus96]. Essen [Bar00]. essential [Rig02]. estate [Sch13e]. estupor [Cam88]. était [Ano05, Kle05]. etc. [Sch22a]. ethics [Sch30d]. Ethik [Sch30d]. Europe [Dro20, Rod19]. Ever [Fra05]. Everett [Whe57]. everyday [Mil97]. everything [Gef14b]. éves [Sch61a]. Exact [Sch44c, Ses13, Sch13c]. exactly [Sch30g]. exakten [Sch13c]. example [AP03, Che92, Ghi99]. examples [Kie03]. Exchange [SH42, Sch27c]. excitations [FF03, Lei02]. exemple [Che92]. exhibit [KDK^{+}98, KDK^{+}99]. exhibition [Ano06, Ano11b, KS86]. Exile [Ano33, Sig96]. Exiles [Fra05]. Existence [ST65]. Exner [Han79a]. Expanding [Sch41a, Sch56c, Sch10, Sch39b, Bas56, Lin57, Dav57, McC56, Rob56, Yan56]. Experiment [Sch55e, Bli01, Bru09, Sch25e]. Experimental [Ano10, Gav28, Sch13d, GRZ99, RGS32, Sch18d, Shi99]. experimentalist [Asp02]. Experimentalphysik [Sch13d]. experimentellen [Sch18d]. Experimenteller [RGS32]. Experimenting [Zei97]. experiments [Kar03, Rau02]. Explain [Kae31]. explained [Gri84b, Sch29i]. explanation [Sch13b]. explicit [Vis91]. Exploded [Kae31]. explosive [Mül40]. Extract [Sch37b]. extraordinary [SvM11]. eye [Sch24g]. eyes [SM89, Sch99b].

F [Bur65, Sch13d, Sch20a, Sch26e]. Facial [Sch26c]. Factorization [Sch42a, Sch41c]. Facuity [Ano07]. failed [HY87]. Fall [Sch15b]. Fall- [Sch15b]. falls [Sch13j]. Family [Gis06]. famosos [W^{+}87, WdC02]. famous [D’A09, W^{+}87, WdC02]. fantastic [Kim99]. Faraday [Ano12]. Farbe [Sch23a]. Farbenmessung [Sch25i]. Farbenmetrik [Sch20e, Sch20f, Sch20g, Sch20d]. Farbenmischungsaufgaben [Sch26d]. fast [Vis91]. fat [Bla00, Car13]. father [Gef14b]. Fattening [Bal11]. Feldänderungen [RGS32]. Feldtheorie [Sch46a, Sch84d]. féle [Sch66b]. Feller [RFS^{+}91]. Fellow [Sch59a]. femtosecond [CD03]. Fest [Ano07]. fester [Sch22c, Sch22a]. Festival [Ano87c]. Festkörper [Sch19a]. Festrede [Sch56d]. Festsaal [Ano11b]. Festschrift [Ano87c]. Festvortrag [Ano11b]. feuchter [Sch10]. few [Dar11, Dar13]. Feynman [Car93]. fictional [Sal01]. Field [BS35, PS51, Sch35c, Sch44a, Sch44f, Sch44g, Sch44k, Sch48a, Sch47b, Sch48e, Sch48f, Sci61, Hit87, Rai03, RGS32, Sch18a, Sch32b, Sch35p, Sch46a, Sch84d]. Fields [MS44, Sch44a, Sch44b, Sch44e, Sch54b, Kib87]. Fifth [SW32, Sch56d]. fiftieth [Sch29a]. Fifty [Kae48, MO95]. Fig [Wil36]. figure [Sch87d]. fiktiver [Sal01]. filosofico [Sch87d]. filosofo [Pol96, BC94]. Final [Sch47b, Sch48e, Sch48f, Sch17b]. finally [FS13]. finesse [Kim99]. finite [Sch22c]. finite-amplitude [Sch22c]. First [Ano33, Sal14, Sch12b, Sch26g, DL82b, Sch13e, Stu18]. Fischer [FS06]. Fisher [Fri89]. fisica [HSBA60, Sch87d, HBS96]. físicos [W^{+}87, WdC02]. Five [Pie97, CW85, Frö11, Gaw97, Jon97, Lie97, Mic98a, Zei97]. Fleisch [Ish20]. flow [Sch13e]. fluctuations [Sch19b]. fluid [Gal03]. Flut [Sch13e]. flying [Kim99]. Fock [Sch31a]. Fois [Ano05, Kle05]. fold [Sch44h]. followed [Sch88c]. Food [Bar00, Sta04]. Foppl [Sch14b, Sch14b]. For.Mem.R.S. [Dir61]. force [Sch30e]. force-free [Sch30e]. Forced [AS96, Kae31, Krö98]. Forces [Cou62]. Foreign [Cla23]. Foreword [Sch29l]. Form [Aja15, Sch55h, Bit92, Sch55g]. forme [Bit92]. Formulation [Eve57, Whe57]. Forschung [Sch17a, Sch17b, Sch19a]. Fortschritte [BWE^{+}29]. Found [Sch21c]. Foundation [Sch47c, Sch47d, Ano03a, MR82a, MR82b]. Foundations [GL23, Rei44b, Ano35a, Rei98b, Sch35n]. Founders [ADS83]. Four [Sch28b, Sch25g, Sch82a, Sch03, Fow29, Sch28e]. four-color [Sch25g]. Fourier [Sch53b]. Fourth [Sch14b, Sch26h]. Fractional [Raj02]. Fragment [Sch75b]. frame [Gis02]. France [Ano33]. Frank [Bun53, Sch13b, Sch35j]. Franz [Han79a]. Free [Sch36a, Sch38c, FF03, Sch24d, Sch30e]. Freedom [EST36, Sch35h, Sch31a, Sch35k]. freie [Sch24d]. Freiheit [Sch35k]. Freiheitsgrade [Sch31a]. French [Bal92, Bit92, Che92, Dat38, Kle05, Miq07, Nef92, Sch28c, Sch32e, Sch33b, Sch35n, Sch38d, Sch41d, Sch55d, Sch82d, Sch86d, Sch86e, Sch88b, SB90a, SB90b, Sch92a, Sch92b, SB11, dB65]. Frenkel [ARK^{+}99]. Frequency [Sch55f, Sch22b]. Frequency-Shift [Sch55f]. Frequenzbedingung [Sch22b]. Friday [GRY06]. Friedrich [ARK^{+}99, Jah56, Sch13e]. fringe [Hor99]. Frisch [Pra93]. Fritz [Jah56]. Fueki [RFS^{+}91]. Fulfillment [Sch95b]. Function [Sch42c, Sch38a]. functions [GHKO99, Sch26d, Sch29g]. Fundamental [Aja15, Hob17, Sch44e, Whe17, Sch99a]. Fundamentals [Sch13c, Sch13d]. Fünf [Mic98a]. fünften [Sch56d]. fünfzigjährigen [Sch29a]. Funktionen [Sch29g]. Further [Sch41c]. Future [Sch50b, MO95].

G [Adl03, Jah56, Kai85a, Kai85b, Krö98, Pra93, Sch12b, Sch12d, Sch14b, Sch17c, Sch21c]. Gaisser [RFS^{+}91]. Galileo [Cro01a, Sch75b]. game [Bel05, Bel06, Bel07, Fin86, Sch30h]. Games [Col00a]. Gamow [Pra93]. ganz [Hen12, SvM11]. Gaponenko [ARK^{+}99]. gas [Sch15a, Sch25a, Sch26m, Sch26b, Sch86g, Sch24d, Sch25a]. Gasblasen [Sch15a]. Gasentartung [Sch24d]. gases [Han77a, Sch26m]. Gasmodells [Sch26b]. Gastheorie [Sch26m]. gatto [D’A09]. Gauge [Kib87]. geb [Deh36]. Gebieten [GSL26]. Geborgen [Sch89f]. Gebrauche [Sch13f]. Geburtstag [Sch87c, GS87, SS97]. Geburtstages [SKD^{+}87]. Gedankenexperiment [He20]. Gedenkworte [Hei63]. Gedichte [Sch49b]. gegenwärtige [Sch35f, Sch35a, Sch35b, Sch84a, Sch96b]. geh [Ben34]. Gehalt [Sch13h]. gehalten [Sch28e, Sch56d]. gehaltenen [HSD34a, Wil36]. gehetzte [Sch30b]. Geist [Sch47a, Sch66a, Sch86i, GB08, GHHL10, G^{+}11, Sch89b, Sti60, Sei61]. Geistesgeschichte [Sch57d]. gekoppelter [Sch14f]. General [NM12, Sch14c, Sch44b, Sch48a, Bus13, Sal14, Aud87, Sch33b, Sch40c, Sch53c, Sch84f]. générale [Sch14c, Sch33b]. Generalization [SH51]. Generalizations [DD13]. generalized [GHKO99, HS51]. generally [Sch18e]. generally-covariant [Sch18e]. generated [Sch14d]. Generating [OTBLG06]. Generation [Hür22, OJTBG07]. genesis [Ren13]. genetic [WH17]. gentler [Bli01]. geometrical [Sch26d]. Geometrie [Sch34c]. geometrische [Sch26d]. Geometry [Koc02, Sch34c]. Georg [Ano11b]. George [Ano11b, Pra93]. German [Hie68, Krö98, Sch12d, Sch13a, Sch13i, Sco67a, Ano35a, Ano87c, Ano93, Ano04a, Ano06, Ano11b, Arr06, AS96, Aud87, Bar00, Bel07, BWE^{+}29, BEvL^{+}28, Bor63a, Dic85, Ein27, FS06, Ful15, Gar53, GNK^{+}29, Gri13a, GSL26, GS87, Haa34, HSD34a, HF46, Hei61a, Hei63, HSE31, Joh13, KS86, KDK^{+}98, KDK^{+}99, KSP32, KK75, KS14, KS26, KGS25, LMOS15, Mic98a, Mül40, Neu96, OGT92, Pla28, Pla29, PS63, Prz63, RGS32, RW11, Sal01, Sch96a, SS97, Sch10, Sch12a, Sch12b, Sch12e, Sch12f, Sch12c, Sch12d, Sch13c, Sch13d, Sch13e, Sch13f, Sch13g, Sch13h, Sch13i, Sch13b, Sch13j, Sch14f, Sch14e, Sch14d, Sch14c, Sch14g, Sch14b, Sch15a, Sch15b, Sch17e, Sch17c, Sch17a, Sch17b, Sch17d, Sch18b]. German [Sch18a, Sch18d, Sch18e, Sch18c, Sch19a, Sch19b, Sch20a, Sch20d, Sch20e, Sch20f, Sch20g, Sch20i, Sch20b, Sch20h, Sch20c, Sch21a, Sch21b, Sch22b, Sch22c, Sch22a, Sch23b, Sch23c, Sch23a, Sch24b, Sch24a, Sch24c, Sch24d, Sch24e, Sch24h, Sch24f, Sch24g, Sch25a, Sch25b, Sch25i, Sch25d, Sch25e, Sch25f, Sch25h, Sch25g, Sch25c, Sch26m, Sch26b, Sch26c, Sch26d, Sch26e, Sch26f, Sch26g, Sch26h, Sch26i, Sch26a, Sch26j, Sch26k, Sch27a, Sch27e, Sch27b, Sch27c, Sch28d, Sch28e, Sch29a, Sch29b, Sch29c, SP29, Sch29g, Sch29f, Sch29d, Sch29j, Sch29e, Sch29k, Sch29l, Sch29m, Sch30a, Sch30b, Sch30i, Sch30e, Sch30d, Sch30f, Sch30g, Sch30h, Sch30c, Sch31a]. German [Sch31g, Sch31e, Sch31f, Sch32a, Sch33a, SWK^{+}32, Sch32b, Sch32c, Sch32f, Sch32d, SW32, Sch33e, Sch33f, Sch33d, Sch34a, Sch34b, Sch34c, SvMFT35, Sch35j, Sch35k, Sch35f, Sch35a, Sch35b, Sch35e, Sch35q, Sch37b, Sch38b, Sch38a, Sch45c, Sch45d, Sch46a, Sch47a, Sch48d, Sch48c, Sch48g, Sch49b, Sch51d, Sch52f, Sch52g, Sch53d, Sch53e, Sch54a, Sch54e, Sch54i, Sch55b, Sch55a, Sch55c, Sch56d, Sch56a, Sch56b, Sch57a, Sch57d, Sch59a, Sch61b, Sch61e, Sch63, Sch66a, Sch79, Sch84a, Sch85a, Sch86a, Sch86c, Sch86b, Sch86h, Sch86i, Sch87b, SKD^{+}87, Sch87a, Sch87f, Sch87g, Sch87c, Sch89c, Sch89b, Sch89d, Sch89e, Sch89a, Sch89i, SM89, Sch89h, Sch93]. German [Sch96b, Sch97, Sch98, Sch99b, Sch00, Sch02, Sch04, Sch06, Sch08b, SvM11, SM35, Sta04, Sta36, Wes57, vM92]. German-Speaking [Krö98, AS96]. Germany [Gol90, Joh13]. Gerry [Bus13]. Gershenfeld [ARK^{+}99]. gesammelte [Sch13e, Sch84c, Sch84d, Sch84e, Sch84f, Sch84b]. Geschichte [LMOS15]. geschliffenen [Bel07]. Gesetz [Sch29d]. Gesetzen [Sch87b]. Gesichtsempfindungen [Sch26c]. Get [Ano33]. Gets [Ano33]. GHZ [Wei02]. Gibbs [Sch21a]. Gibbssches [Sch21a]. Gleichgewicht [Sch24f]. Gleichheit [Sch35k]. Gleichung [Ano04a, SM35]. Gleichungen [Gar53]. Glückliche [Sch89c]. Go [Fra05]. goals [Kur03a]. God [DC09, MB11b]. Gödelian [Orl99]. Goes [Ano33]. Goeschen [Sch20c]. Goettingen [DL82b]. Going [Bro36]. golden [Arm46]. Gordon [BJK03]. Göschen [Sch20c]. Göttingen [Bit96a, DL82b]. Göttinger [Bel92]. Goudsmit [Lak96]. Grant [Ano33]. Grassmann [Sch13e, Sch13e]. Graßmanns [Sch13e]. Gravitation [Sch44e, SH51, HS51, Pen02]. gravitational [Hal87, Sch18a, Sch18e, Sch32b]. Gravitationsfeldes [Sch18a]. Gravitationsgleichungen [Sch18e]. gravity [BP95]. Graz [Ano38]. Great [Far03, HT10, Sau16, Far02, Hal15, Kae48, Pic08, Sch20h, Sch31b, AFG90, Cro01a]. Greatest [Acz01, Hür22]. Grecs [Lar93, Sch92a]. Greeks [Sch54d, Sch56b, Sch89a, Sch92a, Sch96c, Ack55, Bec54, Bur55, Ede55, Kir55, Spr55, Tem54, Tho55, Wig55, Egl56]. grid [Cer93]. Griechen [Sch56b, Sch89a]. Grimsehl [Sch13f]. Grosse [ARK^{+}99]. größter [Sch20h]. Group [Far01, Sch33e]. group-theoretic [Sch33e]. Growth [Bal68]. Grub [Sch13d]. Grummt [RFS^{+}91]. Grünbaum [CL83]. Grund [Ano35a, Deh36, Haa34]. Grundansicht [Sch86b]. Gründe [Sch86i]. Grundgedanke [Sch34b, Sch35e]. Grundlagen [Sch13c]. Grundlinien [Sch20e, Sch20f, Sch20g]. Grundzüge [Sch13d]. gruppentheoretische [Sch33e]. Guenther [Sch20a]. Guide [Dro20, Fle18, Rod19, Ish20]. Guild [Sch26d]. Günther [Sch20a].

H [Adl03, C.35, Klo63, Mic68, Pra93, RGS32, Sch12d, Sch13i, Sch21c, Sch30a, Sch31c, Sch34a, Kar87]. H. [Prz11]. H2 [Kar87]. Haas [Deh36]. Hadô [Sch82b]. Hall [Ano07]. Halpern [Sau16]. Hamiltonian [Sch55g, Sch55h, Sut08]. Hand [Ahl96, CS96, Hoa96, Mor96, Sei96]. Handbook [Sch25d]. Handbuch [Sch25d]. handling [Sch13a]. happiness [Tuc11]. Happy [Sch89c]. Harald [Cha13]. hardback [Bus20]. hardcover [Hen88, Whe17]. Hardy [Adl03, Pra93]. harried [Sch30b]. harvest [Bra09]. Hastings [Sco67a]. Hauptsatz [Sch33d]. Hawking [Cro01a, Pic08]. Heat [Sch26e, Sch22c, Sch22a, Sch24h, Sch26j]. heats [Sch17b, Sch17a, Sch24a]. Heidelberg [Hen88]. Heisenberg [Adl03, Ano35a, Ano51, ARK^{+}99, Ben34, Deh36, Ful15, Haa34, Han62, Lak96, RFS^{+}91, Sch26k, Sch30i, Wil36, Ano33, Ano35a, Blo76a, Blo76b, Cla23, CG14, Dir64, DD13, FR13a, Ful15, Haa34, Han79a, Meh87c, Per64, Plo10, Sal01, Sch26k, Sch66b, Shi83, Shi86, ST65, dG14]. Heisenbergschen [Sch30i]. Heitjans [ARK^{+}99]. held [BRZ88, GM94, Sch28e, Sch56d]. helium [Sch21c]. helix [Ano04b]. Helmholtz [KK75]. Helmut [Hen88, Sch87i, Wes91]. Hendrik [Prz86, Kox13]. Henri [Adl03]. her [Ebe09, Leg84, Leg09, Leg13]. Herausgegeben [Klo63, Sch13e, Sch20b]. Herbst [Sch89d]. Herder [RFS^{+}91]. heretics [DT18]. Hermann [KK75, Sch13d, Sch14c, Sch13e]. Heroic [Pur18]. Herrn [Pla29, Sch24a, Sch31a]. Herstellung [Sch13a]. hervorragender [Kai85a, Kai85b]. Herzfeld [Sch26e]. Hess [Ano51]. high [Sch22c]. Hilbert [Osk03, Zei97]. Hirzel [Ben34, Wil36]. Historical [Bru80, GL23, Hen88, Jam74b, MR87b, Seg85, Wes91, MR01, Sch87i]. History [D’A00, FW20, HR74, Hun74, LMOS15, WP85, Yox79, Cla02, D’A01b, ET07, KHFA67, Sch57d]. Hobson [Whe17]. Hoffmann [Kai85a, Kai85b]. Höhenverteilung [Sch12e]. hoher [Sch22c]. Honour [CL83]. Houtermans [ARK^{+}99]. Hrn [Sch29a, Sch29b, Sch29c, Sch32a]. Hrsgg [Krö98]. Hüfner [RFS^{+}91]. Hugo [Jah56]. Human [Ano35b, Bro36, C.35, SMJ35a, SMJ35b]. Humanism [Bun53, Fin53, LH53a, LH53b, Lon52, Nag52, Sch51c, Sch61f, Sch88a, Sch96c, SS53]. Humanisme [Ano59]. Hungarian [Sch61a, Sch66b, Sch70]. Hydrodynamics [Sch14a]. Hydrogen [Lak96, Rig02, Smi23, Kle02, Sch21c, Sch24h, Sch25c]. hydrogen-like [Sch25c]. hydromechanics [Sch14a]. hyper [Ano10]. hyper-entangled [Ano10]. Hypergeometric [Sch42a]. Hypothesis [Gao18, Sch24b]. hypothetical [HY87].

Idea [Kae31, Sch35e, Sch99a]. ideal [Han77a, Sch25a, Sch26b]. idealen [Sch25a, Sch26b]. Idealistic [Kae31]. Ideas [D’A00, RF69, D’A01b, HT10, MZP97]. Identities [Sch44k, Sch52a]. ignorance [Zur03]. ihre [Sch13a]. II [ARK^{+}99, Sch13f, Sch13i, Sch14a, HS51, Meh87b, Sch47d, Sch48e, Sch52c, Sch55h, Yan13a]. III [Sch12b, Sch13e, Meh88, Sch48f]. illus [Hie68, Kil90]. im [Ano06, Ein27, Sch13j, Sch19a, Sch20e, Sch20f, Sch20g, Sch28e, Sch32b, Sch34c, Sch54e]. Image [Sch61c, Sch61d, Sch62a, Sch64b]. images [Sch14d]. Implications [NM12]. importance [SvM11]. important [Sta04]. improved [Sch13f]. inapplicability [Sch34c]. Inaugural [Sch56a, Pla29, Sch29b, Sch29c, SP29, Sch32a, Sch61b]. Including [Sch42c, Mic98b]. Indeterminacy [Han79a]. Indeterminism [Sch36a, Orl99, PS33, Sch32f]. Indeterminismus [Sch32f]. indeterminizm [PS33]. inédites [Sch33c]. inediti [Sch87d]. inequalities [Fry02]. iner [Sch20c]. Inference [Sch31c]. Infinitesimal [MS45]. influence [Fla87, Sch12f]. Influential [Rog10]. Information [Ben95, BZ02, OTBLG06, Sch90, Ste02b, Ben99, Ben02, Fri89, Hil99, Zei99, Zei02, Zur03]. informational [Orl99]. inneren [Sch31a]. Innocence [Stu18]. innovation [HD13]. Innsbruck [Col98, OGT92]. insight [Rai03]. Institut [Ano06]. Institute [Ano93, Ano03a, Ano04c, Ano06, Ano11b, GRY06, GM94, GHKO99, Sch52e, McC87b]. Institution [Sch28b, Sch28e]. Institutions [FW20]. insulators [Sch10]. Integral [Sch35j, Cha13]. Integralgleichungen [Sch35j]. Intellectual [Fre12, Mal12, G^{+}11, Sch57d]. interaction [d’E02]. Interactions [Hor02, Thi99]. Interatomic [Cou62]. Interest [L.88, Bus13, Sal14]. interference [Arn03, Sch14d]. Interferenzbilder [Sch14d]. interferometry [BRZ88]. internal [Sch31a]. International [Ano93, Ano04c, BRZ88, GM94, GRZ99, TU77, GRY06]. Interpretation [Cus96, D’A92a, D’A01a, He20, Hug89, RFS^{+}91, Sci61, Ano11a, Che99, Coo18, Cra86, DL82b, Dor87, Hil99, Sch21b, Sch32e, Sch95a, Sch32e, Sch32c, Bus20]. Interpretations [Deg89, Jam74a, Jam74b]. Interpreting [Joh07]. Interviews [Sch31b, GS87]. Introducing [Axo89]. Introduction [Bal68, Hei68, Kil87a, Kle68a, Kle68b, PW35, Sch14b, Ano35a, Haa34, Hie68, Sco67b, Cha13]. introductory [Ano11b]. invariance [Zur03]. inventory [KHFA67]. Invitation [Ano06]. iodine [Sch21c]. ion [SK03]. Ionics [RFS^{+}91]. Ionics/Kovetz [RFS^{+}91]. Ionization [Sch21c]. ions [MZ13]. Ireland [MW03]. Irene [Wil16]. Irreversibility [Sch51a, Sch82c, Omn99]. ISBN [Bus20, Hen88, Ish20, Whe17]. Isolatoren [Sch10]. Isotopie [Sch21a]. Isotopy [Sch21a]. isotropen [Sch13j]. isotropic [Sch13j]. Issues [PS33, W^{+}87, WdC02]. Italian [BES86, D’A09, DL82b, HSBA60, Sch26m, Sch37b, Sch59b, Sch86g, Sch87d, Sch88a]. IV [ARK^{+}99, Far01]. IV/Teil [ARK^{+}99].

J [Bal86, Cha13, Ebe09, Han62, Hie68, Mic68, Paw12, Sch26d]. Jagdish [Sch87i, Wes91, Hen88]. Jahre [Ano04a, Mic98a, Sch48c, Sch98, Wes57]. Jahre/Zur [Wes57]. James [Ano35b, Bro36, C.35]. Janossy [Moo90]. January [Ano11a, Sch44j, Sch52e]. Japanese [Sch82b, Sch87h]. Jeffreys [Sch31c]. jest [Sch60a]. jiden [Sch87h]. John [Ber02, BZ02, Ghi02, Got99, Hor02, Per02, Ste02a, Zic02, d’E02, Réd05]. Johnston [C.35]. Jonathan [Sal14]. Jones [Ses13]. Jordan [Deg89, Sch26k, Sch66b, dG14]. Jordan-féle [Sch66b]. Jordanschen [Sch26k]. Josephson [Bar03]. Jourdain [Sch20b]. Journey [FR13c, Lev17b]. Jr. [Sch20a]. juggling [Kim99]. Juli [Pla29]. July [Mik94, Pla29]. Jumps [Sch52b, Sch52c, Bel87a]. jun [Sch20a]. June [GRY06, Sch14e, Sch29a, TU77]. Jung [Jah56]. Juni [Sch14e, Sch29a, Sch89f]. Jürgen [Ano11b]. Justus [Sch13e].

Kaegi [Jah56]. Kampf [Sch29d]. kann [Sch87c, Sch24e]. Kapillardruck [Sch15a]. Kärger [ARK^{+}99]. Karl [Hen12, Kle65]. kart [Wil36]. Katalog [KS86]. Katze [Gri13a, RFS^{+}91]. Katzenparadoxon [SvM11, Hen12]. Kausalproblem [Sch33a]. keep [MZP97]. Keil [BG61]. keiseishi [Sch82b]. Kenntnis [Sch12a]. Kerenyi [Jah56]. KF [Sch26e]. killed [Bal00]. Killer [Sal14]. Kimberley [Bus13]. kinder [Bli01]. Kindheit [Sch89c]. kinds [Wig67c]. kinematics [Sch12d]. Kinematik [Sch12d]. Kinetic [Sch26e, Sch12f]. kinetics [Sch12c]. Kinetik [Sch12c]. Kinetische [Sch26e]. kinetischen [Sch12f]. Kirdly [Moo90]. Kirsten [Cha13]. Kitten [Tau96]. Kittens [OTBLG06, Gri95b, Gri95a]. Klassiker [Sch13c, Sch20b]. klassischen [Sch25b]. Klein [Hie68, Mic68, Sch33f, BJK03]. Kleinen [Sch34c]. Kleiner [Ano11b]. Klingen [Bel07]. knives [Bel05, Bel06]. knize [SS53]. knots [Car93]. know [Sta04, Sut08]. knowledge [Göt92a, Sch12a, Sch35n, Sch04]. Kohärenz [Sch20i]. Kohlenstoff [Sch54e]. Kohlenstoff-Uhr [Sch54e]. Köhler [Köh76]. Kohn [Woo06]. kontinuierliche [Sch29g]. Körper [Sch13i, Sch22c, Sch22a]. Kosmos [Sch56b]. kovarianten [Sch18e]. Kovetz [RFS^{+}91]. kräftefreie [Sch30e]. Krise [Sch56a]. Kristallgitter [Sch17c]. Kudo [RFS^{+}91]. Kunst [Sch30h]. Kurve [KS26]. kvantumelmélet [Sch61a]. kvantummechanika [Sch66b].

L [Fow29, Sch21c, Sch33e]. laboratory [Axo89, Ebe09, Leg84, Leg09, Leg13]. Lajos [Moo90, JS87]. Länge [Sch13c]. Language [Pag82]. large [dR02]. large-Nc [dR02]. last [Hal87]. later [Hit87]. lattices [Sch17c]. laureate [SS97]. lavori [DL82b]. Law [Sch29d, Sch33d, Sch44i, Sch29m, Sch79, Sch97, Sch08b, Zur99]. lawn [Gef14b]. Laws [Gri22, Sch48a, Sch47b, Sch48e, Sch48f, Pic08, Sch29g, Sch31f, Sch87b]. leading [Cro01a, Sch87d]. Learning [EST36]. learnt [Loc92]. Leben [ARK^{+}99, GB08, RFS^{+}91, Sch85a, Sch87b, Sch89e, SM89, Sch99b, Sch06, Bar00, Sch00]. lebende [SM89, Sch99b]. Lecture [Ano07, Sch52g, Ano11b, HSD34a, Sch98, Win13a]. Lectures [Fow29, Sch30a, Sch34a, Sch40b, Sch58b, Ano03a, Ano04a, Gaw97, Jon97, Lie97, Pie97, Sch13f, Sch28b, Sch28e, Sch44j, Sch46b, Sch52e, Sch82a, Sch03, Zei97]. legacy [Ano11a, Atk87, GRY06]. Leggett [Ebe09]. Lehrbuch [Sch26e, Sch13f]. Leibniz [LMOS15]. Leibnizschen [LMOS15]. Leiden [Sch30a, Sch34a]. Leipzig [Ben34, Deh36, Kai85a, Kai85b, Sch12b, Sch12d, Sch13a, Sch13c, Sch13f, Sch13b, Sch13j, Sch17c, Sch20a, Sch20b, Wil36]. Leitung [Sch10]. Leitungselektronen [Sch12f]. L’électron [Sch33b]. L’Élision [SB90b]. Lenard [Sta36, Sta36]. Length [Sch54c, Sch13c]. Lennard [Ses13]. Lennard-Jones [Ses13]. Leo [Fra05]. L’Esprit [SB90b, SB90a, SB11, Lar90]. Letters [OGT92, Prz67, Prz86, Prz11, Dic85, Réd05, SvM11, PS63, Prz63, Hie68, Mic68]. lettura [D’A09]. Leuchtkraft [Sch20h]. Level [Bus20, Ish20, Bus13, Cha13, Sal14]. levels [Sch26b]. levelse [Moo90]. L’évolution [Sch13b]. Lexikon [RFS^{+}91]. Li [Zuk79]. Library [Ano06, Ano11b, Hie68, KDK^{+}98, KDK^{+}99]. Licht [Sch13j, Sch24f, Sch30b]. Licht- [Sch24f]. Lichte [Sch19a]. Lichts [Sch29e]. Liebeslied [Sch89f]. liest [Dic85]. Life [Ano45a, Ano45b, Ano97, Ben48, Car45, Del45, Fre12, GHHL10, Kil90, M.47, Mal12, Miq07, Moo96, N.46, Per87, Pra93, Sch86d, Sch86e, Sch89e, Sym87, Ter90, Yam97, Yox79, Ano90, Ano12, Bit93, Cro01a, GGW94, G^{+}11, Leg00, Mil97, Moo89, Moo94, MO95, Olb71, Sch44l, Sch45e, Sch56e, Sch56f, Sch62b, Sch67, Sch85a, Sch86j, Sch87b, Sch87d, Sch88a, SM89, Sch92c, Sch99b, Sch06, Sch12, Sig96, qYL17, Bar00, MF91, Sch88c, Sch00]. Light [BRZ88, HS51, MS88, Sch13j, Sch19a, Sch24f, dB24, Sch29e, Sch30b]. light- [Sch24f]. like [Sch25c, Wei02]. L’image [Sch53d]. Limit [Gri22]. L’immagine [Sch59b]. lineales [Sch35p]. Linear [Sch42b, Sch44c, Sch35p]. lines [Sch26f, SW32]. Links [Kir95]. Lise [McL99]. Literatur [BWE^{+}29]. literature [HT10, BWE^{+}29]. Little [Pra93]. Liverpool [Row09]. Living [Ano45a, Ano45b, Del45, Moo87, Sch44l, Sch45e, Sch62b, Sch67, Sch86e, Sch86j, SM89, Sch92c, Sch99b, Sch12]. livre [Ano05]. local [Yng99]. Lochak [RFS^{+}91]. logic [SK03]. London [Bun53, Sch14a, Sch28b, Sch28e]. Lone [Ter90]. Lorentz [Kil65, Klo63, Prz63, Kle65, Kox13, Prz63, Prz67, Prz86, Prz11, Sch30a, Sch34a, Hie68, Mic68]. Lösung [Sch26d, SM35]. Lösungssystem [Sch18e]. Louis [Cro01b, Dar11, Dar13, Han77b]. love [CG14]. Ludwig [Ano06, GRY06]. Luft [Sch10]. luminosity [Sch20h]. Lummer [Sch25f].

M [Ben34, Bus13, Kai85a, Kai85b, Klo63, Sch14b, Wil36, Wil36]. Ma [Sch82d]. Mach [BP95, BIT01]. machines [Mil97]. macro [Sch26a]. macro-mechanics [Sch26a]. Macroscopic [Thi99, Oes99]. made [Haw11]. Magic [Ber02]. Magnetic [MS44, Sch44a]. Magnetism [Kir95, Sch12f]. Magnetismus [Sch12f]. Mainzer [RFS^{+}91]. making [Bel99]. Makromechanik [Sch26a]. man [SR92, Sch24e, Sch57b, Sch57c, Pos62, SF11]. Manton [Wil16]. manufacture [Sch13a]. Many [Bru04, SR92]. Manybody [Gaw97]. March [Ano03a, Sch28b, Sch28e, Sch44j, Sch52e]. Marcolongo [Sch12d, Sch13i]. Maria [Moo90]. Mario [BM01, Mah01]. Mark [Sch13d]. Marks [Sch13d]. married [Tuc11]. Martin [ARK^{+}99, Hie68, Mic68]. Marx [Dat87, Sch25d]. Mary [Ste02a]. März [Frö03, Gal03, Nar03, Sch28e]. más [W^{+}87, WdC02]. Mass [BS55, BS56, Sch42c]. Massachusetts [Kle68a]. masses [Sch53b]. Masson [Jah56]. Masson-Oursel [Jah56]. masters [Zuk79]. materia [BES86, GHHL10, Sch59b]. Materialien [KDK87, KBG11, SKD^{+}87]. Materials [Pie11, SKD^{+}87]. Materie [GB08, Sch52g, Sch86h, Sch98, Sch02, GHHL10, G^{+}11, Sch53d, Sch53e, Sch54i, Sch86i, Sch89b, Sei61, Sti60]. Materiewellen [Deh36, Ano35a, Haa34, Sch27b]. Mathematical [Ano93, Ano04c, ARK^{+}99, GRY06, GM94, Sch12b, Sch13e]. Mathematician [Pra93]. Mathematicians [Acz01, Sut08]. Mathematics [Sch12b, Car93]. Mathematische [Sch12b, Sch13e]. mathematischen [Sch13e]. matiére [SB11, Sch53d, SB90a, SB90b, Lar90]. Matrices [Sch43b, Sch44d]. Matter [ARK^{+}99, BRZ88, GHHL10, Pra93, Sch53f, Sch60b, Sch61c, Sch61d, Sch62a, Sch64b, Sch89b, SB90a, SB90b, Sch02, BES86, CD03, Kar03, Lie97, Sch27b, Sch52g, Sch53d, Sch58b, Sch59b, Sch67, Sch86j, Sch86h, Sch98, SB11, Sch12, Ano35a, BRZ88, Haa34, Han62, Sch53e, Sch54i, Sch86i, Sch92c, Bar59, Fre61, Opp59, Pol59]. Matter/Friedrich [ARK^{+}99]. Max [Adl03, Ano11b, ABdB^{+}53, Bro36, Jah56, Mic68, Prz86, Sch17c, Sch29a, Sch61b, Wes57, Prz11, Wes57]. Maxwell [Sch14b, Cou62, Hou16, RA11, Sch41a, Sch54b]. Maxwellsche [Sch14b]. May [Ano07]. Mean [Sch38c, Sch24d]. Meaning [DT18, Sch55c, Sch55d, Gef14b, Ste02b]. Measurement [Sch54c, FF03, Oes99, Orl99, Wig67a]. measuring [Sch19b]. mecánica [Sch35l, Sch01]. mécanique [BD95, BD92, Sch13b, Sch28c, Sch32e, Sch33c, Sch55d, Sch88b]. meccanica [DL82b]. Méchain [Sch13c]. Mechanical [Kae31, Sch41b, Ber88, EPR35, Sch13b]. Mechanics [BRZ88, Bit96b, Bru80, Cha13, Cro01b, Cus96, D’A92a, D’A01a, DL82a, Eve57, GGK02, GL23, HSD34b, Hen88, Hie68, Hug89, Jam74b, Mic68, MB11a, PW35, PS63, Prz63, Prz67, Pur18, RFS^{+}91, Rei44b, Rüg88, Sch26l, Sch40c, Sch54c, Sch55c, Sch55d, Sch63, Sch75a, Sch94b, Sch95b, Ter90, Tri80, Ano35a, Aud87, BPW^{+}71, Bel87b, BD95, BD92, Blo76a, Blo76b, Bus20, Coo18, Cra86, DL82b, DeW70, Ein27, ET07, Frö11, FS13, Gal03, Haa34, Han79b, tH02, Hor99, Joh07, Koc02, Kox13, Kra92, Meh87a, Meh87b, MR87a, MR87b, Meh88, MR01, Oes99, Pla28, Prz86, Prz11, Rei98b, RW11, Ren13, Sch12d, Sch13i, Sch13b, Sch25b, Sch26k, Sch27a, Sch27c, Sch28a, Sch28b, Sch28c, Sch28e, Sch30e, Sch35j, Sch35l, Sch53c]. mechanics [Sch66b, Sch78, Sch82a, Sch84c, Sch84a, Sch88b, Sch95a, Sch99a, Sch01, Sch03, SvM11, Sch87i, Wes79, qYL17, vM92, Fow29, Ano11a, Sch26a, Sch31e, Sch32e, Sch33e, Sch35f, Sch35a, Sch35b, Sch35e, Sch96b, Fow29, Kle84, Wes91]. Mechanics/Spatschek [RFS^{+}91]. Mechanik [Sch12d, Sch13i, Sch13b, Sch35j, Sch25b, Sch84c]. mechanische [Sch13b]. Medaille [BEvL^{+}28]. Medal [BEvL^{+}28]. Medium [Sch13j]. Mediziner [Kai85a, Kai85b]. meeting [Pla29, Sch14e, Sch54a]. Meetings [GNK^{+}29, HSE31, KSP32, KGS25, SWK^{+}32, SvMFT35]. mehr [Sta04]. Mehra [Hen88, Sch87i, Wes91]. Mehrdeutigkeit [Sch38a]. Meier [Jah56]. Mein [Sch85a, Sch89e, Sch06]. meine [Sch85a, Sch89e, Sch06, Sco67a]. meinem [Sch26k]. meinen [Sch17d]. Meitner [McL99]. melting [Sch12c]. Membership [Cla23]. Mémoires [Sch88b, Sch33c]. Memoirs [Kle96, Sch88b]. Memorial [Hei63]. memoriam [JS87]. memories [FF03, Sch87a]. Menschen [Sch57d]. Mente [GHHL10]. mercury [Sch21c]. merely [Sch58a]. Mérite [Wes57, Wes57]. Meson [Sch43b, Sch44d, Sch44e, Sch46a, dR02, Sch46a]. Meson-Matrices [Sch43b, Sch44d]. mesoscopic [Rai03]. Meßanordnung [Sch19b]. Metaphysics [Sch94b]. metastable [Kle02]. Method [Sch41b, Sch33e, SM35]. Methode [Sch33e]. methods [Cer93]. Metric [Sch47e, Sch13c]. metrics [Sch20d, Sch20e, Sch20f, Sch20g]. metrischen [Sch13c]. Meÿenn [ARK^{+}99, Hen12]. Mi [Sch88c]. Michel [BD95, Cus96, Lar90, Lar93]. Michelson [Sch25e]. Michelsonscher [Sch25e]. micro [Sch26a]. micro- [Sch26a]. microbe [qYL17]. Microphysics [D’A01a]. Might [Sch58a]. migration [AS96]. Mikro [Sch26a]. Mikro- [Sch26a]. Mind [Hür22, Loc89, Sch58b, Sch67, Sch12, Sch86j, Sch86i, Wig67b, Bar59, GHHL10, Han62, Pol59, Pra93, Sch89b, SB90a, SB90b, Sch92c, SB11, Fre61, Opp59]. Minds [Sau16, Hal15, Pic08]. mine [Sch26k, Sch66b]. minimalist [Frö11]. minus [Yan87, Yan13b]. Mirabilis [Frö11]. Mirovozzrenie [Sch94a, Sch94b]. mirror [Sim03]. Mises [Sch35j]. Missing [Joh13]. místicos [W^{+}87, WdC02]. Mitarbeit [KPMN20]. Mitchell [Krö98]. Mitteilung [Sch26f, Sch26g, Sch26h, Sch26i]. Mitteilungen [GSL26]. mixture [Sch26d]. Mladjenovic [ARK^{+}99]. Mobilization [Joh13]. model [Gaw97, KS26, Sch21b, Sch92b, Sch26b]. model-based [Sch21b]. Modelers [Lak96]. Modeling [ARK^{+}99]. Modeling/Origin [ARK^{+}99]. Modell [KS26]. modellmäßigen [Sch21b]. Modellvorstellungen [Sch29f]. Models [Joh13, Lak96, Sch29f, Gal03, Len14]. Modern [ADS83, Far03, Han62, HSD34b, Sch29d, TWD03, Bra09, Far02, HSBA60, HBSA61, PS33, SS97, WB00, HSD34a]. moderna [HSBA60]. moderne [Ben34, HSD34a, Wil36]. modernen [SS97, Sch29d]. Modernizing [Joh13]. Modified [BJK03]. Moe [Sch94a, Sch94b]. moist [Sch10]. Molecular [Joh13, Kar87, Yox79, Per87, Sch17a, Sch17b, SF11, Thi87]. Molecules [Sch26l, Arn03]. Molekularwärmen [Sch17b, Sch17a]. moment [CG14]. moments [Ber02]. momentum [Sch21c, Sch27b]. monatomic [Sch26b]. monde [Sch38d, Sch82d, Sch92b]. mondo [Sch87d]. monograph [Bus20]. monoxide [Sch21c]. moons [RA11]. Moore [Kil90, Ter90, Paw12, Bit93]. Most [Hob17, Rog10, Whe17, Haw11, Sta04, W^{+}87, WdC02]. motion [Ein27, Sch15b, Sch30e]. mousetrap [Coh05, Gar05, Gee05, GGGG05, Mat05, Spa05, Ste05, Tut05, Ven05, Zeg05]. movements [Gar53]. moving [Gis02]. MPI [Ano11b]. Mr [Pla29, Pra93, Sch24a, Sch29a, Sch29b, Sch29c, Sch31a, Sch32a]. Mr. [Pla29]. Mueller [Sch26e]. Mueller-Pouillets [Sch26e]. Müller [Sch26e, Cha13]. Müller-Kirsten [Cha13]. Müller-Pouillets [Sch26e]. multiparticle [Ben02]. Multiple [Fra05]. mundo [Sch88c, W^{+}87, WdC02]. Murphy [Bro36, C.35, Ano35b, Bro36]. music [Pie97]. Must [BS55, BS56, Far03, Far02]. My [Got99, Sch64a, Sch83, Sch87d, Sch89e, Sch08a, d’E02, Sch17d, Sch85a, Sch06, Sch82d, Sch85a, Sch87d, Sch88c, Sch89e, Sch06, Sco67a, Bur65, Tho66]. mysteries [Gri95b]. Mystery [Acz01, FR13c, Sch29e]. mystical [Deg89, W^{+}87, WdC02]. mysticism [Coh92].

nach [Gri13a, Sch27c, Sch04]. Nachlasse [Sch13e]. Nachtrag [Sch17d]. näherungsweisen [SM35]. naissance [BD95, BD92]. naive [Asp02]. Named [Ano33]. Nanocrystals [ARK^{+}99]. Nanocrystals/Gershenfeld [ARK^{+}99]. Narayne [Moo90]. nascita [DL82b]. Nation [Ano33]. national [HY87, RW94]. Natur [Sch56b, Sch89a]. Natural [ABdB^{+}53, PS33, Sta36, Sch38b]. Nature [ARK^{+}99, Jah56, Pag82, Plo10, Sch29h, Sch39a, Sch44i, Sch54d, Sch96c, Lar93, MS88, Sch13b, Sch29m, Sch79, Sch92a, Sch97, Sch08b, Bro36, Sch56b, Sch89a, Sch92a, Tho55, B^{+}54, Sch31f, Han62, Ack55, Bec54, Bur55, Ede55, Egl56, Kir55, Spr55, Tem54, Wig55]. Naturerklärung [Sch13b]. Naturforscher [Sta36]. Naturgesetz [Sch29m, Sch79, Sch97, Sch08b]. Naturgesetze [Ber88, Sch31f]. Naturwissenschaft [Sch20a, Sch30d, Sch47a, Sch48d, Sch61e, Sch66a, Sch86c]. Naturwissenschaften [Sch29d]. Naturwissenschaftler [Kai85a, Kai85b]. naturwissenschaftlichen [Sch79, Sch97, Sch08b]. nauki [PS33]. Nazi [Joh13]. Nc [dR02]. near [GRY06]. Nebenserien [Sch21b]. Nebular [Sch39a]. need [MZP97]. negative [Sch00]. negativer [Sch00]. Neglected [Sch30f]. Nernst [Sch14c]. neue [Sch24b, Ben34, Sch28d, Sch29j, Sch32d]. neueren [Sch17a, Sch17b, Sch19a]. Neumann [Bre84a, Réd05]. neurons [Leg90]. neutron [CD03, Kar03, Rau02]. Newton [BP95]. next [MO95]. nicht [Sch87c]. Niels [Adl03, Bro36, How99, Meh87c, Wei14]. nineteenth [D’A01b, D’A00]. nitrogen [Sch21c]. No [Sch13c, Sch82b]. Nobel [HSD34a, SS97, Ano33, DS83, HSD34b, McL99, Sch87a, Win13a]. Nobelpreis [Ano51, Sch87a]. Nobelpreises [HSD34a, Wil36]. Nobelpreisträgers [SS97]. Nominations [KK75]. Non [Nar03, PS51, Sch42b, Sch42c, Sch44c, SH51, Bus13, WB00]. Non-commutative [Nar03]. Non-Linear [Sch42b, Sch44c]. non-scientists [WB00]. non-specialists [Bus13]. Non-Symmetric [SH51, PS51]. Non-Vanishing [Sch42c]. Nonlinear [GHKO99, McC87a, Mik94, GHKO99]. Nonlocality [Sei05, Hor02]. North [Adl03]. Note [Sch13g, Sch15a, Sch18c]. notes [Coh92, Sch33c, SS53]. nothing [Gef14b]. Notiz [Sch13g, Sch15a, Sch18c]. notre [Ano59]. novel [Bel07, Fel06, Fry02]. November [Asp02, Ben02, Ber02, BZ02, Cla02, Eke02, Eng02, Fry02, Ghi02, Gis02, tH02, Hor02, Jac02, Kle02, Koc02, Lei02, Pen02, Per02, Raj02, Rau02, Sel02, Shi02, Ste02b, Sut02, Wei02, Whi02, Zei02, Zic02, d’E02, dR02]. Nr [Sch13c]. Nuclear [ARK^{+}99, Sch48b, Seg85, Stu18, WZ51, Kae48]. Nucleus [FR13c]. nueva [Sch35l, Sch01]. number [OJTBG07]. Numerical [Cer93]. Nur [RFS^{+}91, Ein27].

Oberfläche [Sch10]. Obituary [Ano61, Dir61, Sch41e]. object [Bit92]. Objects [FW20, Oes99]. observers [Ghi99]. occasion [Ano03a, Ano07, BRZ88, Gaw97, Jon97, KS86, Lie97, Pie97, Zei97]. occurring [Sch18d]. October [Far01, GHKO99, Sch58b]. Oettingen [Sch20b]. Offentliche [Pla29]. Olivier [BD95]. onde [BES86]. ondes [Sch28c]. ondulatoire [Sch33c, Sch55d, Sch88b]. ondulatoria [DL82b, Sch35l, Sch01]. One [Kae31, Sch24e, Yan87, Yan13b]. ongoing [Got99]. only [Ein27]. open [MZ13, Sch20i]. Opened [Col98]. Opening [Ano11c, YKZ03, Ano06, Ano11b, Sch56d, Ste02a, Ano93]. Operator [Eck26]. Optical [ARK^{+}99, OTBLG06, RFS^{+}91, OJTBG07]. Optics [Sch42b, Sch44c, Bla99, MS88, McC87a]. Optomechanical [kA13]. Orbital [Fuk87]. Orbiting [RA11]. Order [Mer38, Bar03, Sch18c]. Ordnung [Sch18c]. organism [Bre84b]. organisms [Moo87]. Orientation [Sch54e].

Orientierung [Sch54e]. Origin [ARK^{+}99, Sch24g]. Original [Sch02, D’A09, Dor87, Sch02]. Original-Tonaufnahmen [Sch02]. originale [D’A09]. Origins [Meh87c]. oscillation [Rai03]. oscillations [Sch38b, Sch22c]. oscillatory [Osk03]. Österreichische [Sch45c, Sch45d]. Österreichischen [Ano06, Ano11b]. Ostwald [Sch13c, Sch20b]. Ostwalds [Sch13c, Sch20b]. Other [Cus96, L.88, Sch56e, Sch56f, Sch95a, Sch01]. otros [Sch01]. Öttingen [Sch20b]. Otto [KPMN20, Pra93, Sch25f]. Our [Bun53, Fin53, LH53a, LH53b, Lon52, Nag52, Sch61c, Sch61d, Sch62a, Sch64b, Sch51c, Sch61f, Sch53e, Sch54i]. Oursel [Jah56]. Outline [Sch20e, Sch20f, Sch20g]. Outsider [HD13]. Overview [Sal87, Zuk79]. Oxford [Whe17, HY87].

P [Ben34, Sal14, Sch13b, Wig67d, Wil36]. P. [Sch14a]. Paarläufe [FS06]. Padova [Pol96]. pair [FS06]. Paper [Tri80, Ber88, SF11]. Paperback [Ish20, Pra93, Hen88]. Papers [Fow29, Jah56, Kle84, Köh76, ABdB^{+}53, Bel87b, Haw11, Sch13e, Sch24a, Sch28a, Sch78, Sch82a, Sch84c, Sch84d, Sch84e, Sch84f, Sch84b, Sch03, Yan13a]. Parabel [Sch54a, Sch87e, Sch87f, Sch87g, Sch89f]. Parable [Sch54a, Sch87f, Sch87g]. Paradigm [Woo06]. Paradox [Tri80, Bel64, Sch86a, Vil86, dB65, Sch21a, SvM11]. paradoxe [dB65]. Paradoxes [Joh13]. Paradoxon [Sch86a, Sch21a]. parameter [Bar03]. paranormal [Deg89]. Paris [Sch13c, Sch14c]. Part [Hen88, MR87a, Sch12b, RGS32, Sch26j, MR87b, MR01, Sch52b, Sch52c]. Particle [ARK^{+}99, RFS^{+}91, Hor99, NM12, Sal87, Sch50c, Sch51d, Sch51e, Sch60a]. particles [SS97, Sch15b]. particularly [Sch19b]. Passion [Jon08, MW03]. past [Orl99]. Path [Cha13, Sch38c, Len14, Sch24d]. Paul [Jah56, KPMN20, Sau16]. Pauli [ARK^{+}99, Deg89]. Peacock [ARK^{+}99]. pendulum [Sch13c]. penetrating [Sch12e]. pensiero [Sch87d]. Pentads [Sch43b]. people [Sch57d]. Percorsi [Pol96]. perhaps [Sch58a]. period [Mic98b]. Permanent [Sch44a]. Personal [Ste02a, Thi11, Thi13]. Perspective [Bru80, Jam74b, Seg85, TWD03]. perspectives [Bal86, ET07, GRZ99, Loe84, Shi99]. Pertinence [Fre12, Mal12, G^{+}11]. Perturbation [Fro38, Sch26f]. Peter [Moo90]. Pflichten [Sch30f]. Ph [Sch35j]. Phase [Kra82]. phases [Yan87, Yan13b]. phenomena [Sch21c]. Phenomenalism [BIT01]. Phenomenological [Sch35m, Sch36b]. Philip [Sch13b]. Philipp [Bun53, Sch13b, Sch20a, Sta36]. Philosopher [Sch94a]. Philosophic [Rei44b, Rei98b]. Philosophical [Hie68, Sch29f, Shi99, Göt92b, Sch32c, Sch87d, Sch06]. Philosophie [Sch90, BD95, BD92, SB90b]. philosophische [Sch06]. philosophischen [Sch32c]. Philosophy [ABdB^{+}53, BSM06, BD95, Bit96b, BIT01, Bru80, CL83, CW85, Han62, Jam74b, Neu96, Sch40b, Sch40a, Sch55e, Sch94a, Sch94b, Shi83, Shi86, Bel87b, BD92, ET07, Mel88, SB90b]. Philosphers [Acz01]. Photo [Ano87a, Ano87b, Ano88]. photograph [Sch24e]. photographieren [Sch24e]. Photon [BS55, BS56, Kle02, OJTBG07]. photons [FF03, Har99, Kim99, Zei97]. Phys [Bal86]. Physic [Hob17, Whe17]. Physical [Ano45a, Ano45b, D’A92a, Del45, GL23, Gri22, Sch29h, Sch40b, Sch40a, Sch44b, Sch44f, D’A92b, EPR35, Sch13e, Sch30c, Sch35j, Sch44l, Sch45e, Sch62b, Sch67, Sch86e, Sch86j, Sch87b, Sch92c, Sch12, SF11, Ste02b]. physicien [Sch82d]. Physicist [BZ02, Han62, Kle68b, Sal01, Sch82d, SM89, Sch99b, SS53]. Physicists [Ano33, Bit96a, Bru09, Cho00, Hür22, KK75, MW03, McL99, Bel92, Cro01a, W^{+}87, WdC02, KK75]. Physics [Acz01, Ano93, Ano07, Ano11b, ARK^{+}99, BBSR69, Bun53, CL83, D’A00, Fin53, FW20, GRY06, Gil09, GM94, HSD34a, KDK^{+}98, KDK^{+}99, LH53a, LH53b, Lon52, Nag52, Pag82, RW64, RFS^{+}91, Sch12b, Sch13d, Sch26e, Sch29f, Sch52e, Sch56b, Sig96, Stu18, vCK17, Deg89, Bra09, Car93, D’A01b, Dar92, ET07, Ful15, Göt92b, GRZ99, Gri84b, Gri84a, Gri13a, Hal87, Hal15, Haw11, HSBA60, HBSA61, HBS96, Kae48, KHFA67, Meh75, Rau99, Sal87, SS97, Sch13e, Sch13f, Sch30a, Sch32c, Sch32f, Sch34a, Sch35j, Sch51c, Sch61f, Sch86d, Sch87d, Sch92b, Shi99, Str11, Thi87, WP85, WB00, Wra09, Yng99, Zic02, Zuk79, Zur03, Ano04c, Han62, Sch44j, Ano06, Sch28d, Sch29j, Sch32d, DT18, Sau16]. Physics/Grosse [ARK^{+}99]. Physics/Hüfner [RFS^{+}91]. Physics/Peacock [ARK^{+}99]. Physik [Ano51, Ano11b, KDK^{+}98, KDK^{+}99, RFS^{+}91, SS97, Sch12b, Sch13e, Sch13f, Sch26e, Sch30a, Sch32c, Sch34a, Sch35j, Ano06, Sch28d, Sch29j, Sch32f, Sch32d, Sch56b]. Physik/Kudo [RFS^{+}91]. physikalische [Sch13e]. physikalischen [Sch30c, Sch87b]. Physikalischer [Sch35j, Sch29f]. Physiker [KK75, KK75]. Physikern [KK75]. Physikers [SM89, Sch99b]. Physikerstreit [Sal01]. Physique [Ano59, Sch92b, Sch86d, Sch86e, Sch86f]. Pictures [Dir64, Per64, ST65, dG14, SKD^{+}87]. Pierre [Sch13b]. piezoelectricity [Sch12c]. Piezoelektrizität [Sch12c]. Pigmente [Sch20h]. pigments [Sch20h]. pioneering [SS97]. Pionierleistungen [SS97]. Planar [Jon97]. Planck [Adl03, Ano11b, Bro36, Hie68, tH02, Kil65, Kle65, Klo63, Mic68, Prz63, Prz86, Sch29a, Wes57, BEvL^{+}28, CG14, MR82a, MR82b, Pla29, Prz67, Prz11, SP29, Sch61b, Wes57]. Planck-Medaille [BEvL^{+}28]. Planck/Der [Wes57]. Planck/The [Wes57]. Planetary [MS44]. Plasmaphysik [RFS^{+}91]. Play [Sch35o, Sch96d, MB11b]. plötzlichen [RGS32]. Pluto [RA11]. Podolsky [Bel64]. Poems [Sch87d]. Poesie [Sch87d]. Poet [Ano33]. Poetry [Sof14, Sch49b]. Poincaré [Adl03]. Point [PS51, Sch44g, Sch12c, Sch14f, Sch14e]. Point-Charge [PS51]. Polarisierbarkeit [Sch25c]. polarizability [Sch25c]. polemiche [DL82b]. poligrafo [Pol96]. Polish [PS33, Sch60a]. polymath [Kil87b]. Pomerans [Han62]. Popper [Pop89]. popular [Sch84f]. populäre [Sch84f]. Portmann [Jah56]. Portrait [CW85, SS97]. Portraits [LMOS15, Far01]. Porträts [LMOS15]. position [Sch57d]. Possibility [Sch95b]. postgraduate [Cha13]. potential [Dat38, Ses13]. potentials [Sch21c]. potentiel [Dat38]. Pouillets [Sch26e]. Pour [Wes57, Sch33c, Wes57]. Power [Sch56d, Sch43a].

Poznámky [SS53]. Pp [Bun53, Hen88, Bel51a, Bus20, Hie68, Kil90, Kle68a, LH53a, Pol96, Whe17]. Practical [Kni92]. practitioners [Sch48g]. Praktiker [Sch48g]. pre [Gol90]. pre-war [Gol90]. Précédé [SB90b]. Preceded [SB90b]. precision [MW03]. Precursors [Sut02]. prehistory [Mic98b]. Preis [Ben34, Deh36, Wil36]. Present [Tri80, Sch35f, Sch35a, Sch35b, Sch84a, Sch96b]. presentation [Fuk87]. presented [ABdB^{+}53, Sch14e]. présentées [BD95]. President [Ano11b]. Press [Bel51a, Bun53, Ish20, Kil90, Kle68a, Krö98, LH53a, SS53, Ter90]. pressure [Sch15a]. Price [Whe17]. primi [DL82b]. Princeton [Ano34]. principle [BP95, Rei44a, Rei98a, Sch22b, Sch30i]. Principles [RFS^{+}91, Sei05, Meh87c]. Prior [Gri22]. Prize [HSD34a, Ano33, DS83, Sch87a]. Probabilistic [RFS^{+}91, Sch19b]. Probability [BS41, Plo10, Sch36c, Sch48b, Sch47c, Sch47d, Par88, Sch35g]. probe [Bar03]. Problem [Sch18d, Sch29d, Sch51b, Sch77, DKF92, Dat38, Hor02, Olb71, Sch26f, Sch26g, Sch26h, Sch26i, Sch33a, Wig67a, Wig67b]. problème [Dat38]. problemi [Pol96]. Problems [Sch41c, Sch48b, Gaw97, Oes99]. proceedings [BRZ88, GM94, GHKO99, Mik94, TU77]. Processing [OTBLG06]. Prof [Sch13a, Sch13d]. Prof. [Ano38, Ano49, Dir61, Sch13d, Sch31b, Sch41e]. profesora [SS53]. Professor [ARK^{+}99, SS53]. Profs [Ano33]. program [Ghi99, Ghi02]. progress [BWE^{+}29]. Project [Bit96c]. proper [Sch39b]. Properties [ARK^{+}99]. property [Sch23b]. propos [Nef92, Sch33c]. propose [Bru09]. protagonista [Sch87d]. protection [AP03]. Protons [Sch37a, Sch38c, CD03, Kar03]. Proved [Acz01]. Prüfungsarbeit [Sch13e]. Przibram [Hie68, Klo63, Kil65, Kle65, Mic68]. przyrodnicze [PS33]. Psychoanalysis [CL83]. Public [Pla29]. Pulver [Jah56]. Punktreihe [Sch14e]. Punktsysteme [Sch14f]. pure [Car93]. purification [Thi99]. Put [Ano12, Kae31]. Puzzle [Wil16]. Pyro [Sch12c]. pyro- [Sch12c].

QCD [dR02]. QED [Kim99]. QUACS [Bla03]. Qualität [Sch25h]. quality [Sch25h]. Quandary [Kae31]. quanglement [Pen02]. Quanta [vCK17, Kra92, dB24]. Quantelung [Sch22c]. Quantenbahnen [Sch23b]. Quantendynamik [Sch31g]. Quantengesetze [Sch29g]. Quantenmechanik [Ano35a, Deh36, Haa34, Sch26k, Sch33e, Sch84a, Sch30e, Sch31e, Sch35f, Sch35a, Sch35b, Sch96b]. Quantenphilosophie [Neu96]. Quantenphysik [Ful15, Gri13a]. Quantenspringerei [Kra11]. Quantentheorie [Klo63, Sch48c, Sch84e]. Quantenwelt [Arr06]. Quantification [Eng02]. quantique [BD95, BD92, Sch32e, Sch92b, dB65]. Quantisation [D’A01a]. Quantisierung [Sch26f, Sch26g, Sch26h, Sch26i]. quantistica [BES86, DL82b]. Quantization [Sch77, dG14, Sch22c, Sch26f, Sch26g, Sch26h, Sch26i]. Quantum [Ano03c, Arn03, Arr06, BPW^{+}71, Bel99, Ben95, Ben99, BZ02, Bit96b, Bla99, Bou99, Bru04, Bru80, Buc87, Bus13, Cha13, Col98, Cus96, D’A92a, DL82a, DeW70, Eck26, Eps26, ET07, Eve57, GGK02, Gil09, GL23, Han77b, Har99, HSD34b, Hen88, HW86, Hob17, tH02, Hug89, HR74, Hun74, Hür22, Jam74b, Joh13, Jon08, Kar03, KKS^{+}11, Kle65, Lev17a, LL99, Mac24, MR87b, Orl99, OTBLG06, Pag82, PW35, Pen02, Plo10, PS63, Pur18, Rau99, Rei44b, Sal14, Sau16, SK03, Sch21c, Sch41b, Sch52b, Sch52c, Sch53a, Sch54c, Sch54f, Sch94a, Sch94b, Sea87, Sei05, Ter90, TWD03, Tri80, Wes91, Wha11, Whe57, Whe17, Wra09, Zei99, Zol99, Zur00, vCK17, AP03, Ano35a, Ano11a, Deg89, BP95, Bel87b, Bel87a, BD95, Ben02]. quantum [Ber88, BD92, Blo76a, Blo76b, Che99, Cra86, CG14, DT18, DL82b, EPR35, Eke02, FF03, Fin86, Frö11, FS13, Ful15, Ghi99, Gis02, GRZ99, Gri84b, Gri84a, Gri95b, Gri12, Gri13b, Gri13a, Haa34, Hal15, Han77a, Haw11, Hor99, Joh07, Koc02, Kra92, Kra13, KHFA67, Len14, Loc89, MR82a, MR82b, MR01, Mil97, MB11b, MZ13, Oes99, Per91, Rau02, Rei98b, Sch23b, Sch26k, Sch29g, Sch30e, Sch31g, Sch31e, Sch32e, Sch33e, Sch35f, Sch35a, Sch35b, Sch48c, Sch61a, Sch66b, Sch84e, Sch84a, Sch88b, Sch95a, Sch96b, SvM11, Sch87i, Shi99, Sim03, Ste03, Str11, Wei14, WB00, qYL17, Yng99, Zei02, Zur03, dB65, dR97, BV09, BZ02, BES86, HBS96, Neu96, Sch92b, W^{+}87, WdC02]. Quantum-Mechanical [Sch41b]. Quantum-Theoretical [Plo10]. Quasi [FF03]. Quasi-decoherence [FF03]. qubit [Ano10]. que [Miq07, Sch86d, Sch86e, Sch86f, GHHL10]. Quelques [Sch35n]. Qu’est [Miq07, Sch86d, Sch86e, Sch86f, DT18, Sch04]. Qu’est-ce [Miq07, Sch86d, Sch86e, Sch86f]. quite [SvM11].

R [RGS32, Sch35j]. Rabbits [Bru04]. Rabi [Rai03]. Race [Sal14]. radiation [Kle02, KS14, Sch12e, Sch21c, Sch24b, Sch24a]. Radiologie [Sch25d]. Radiology [Sch25d, Sch25d]. Radium [Kae48, Sch13h, Sch23c, Sch18d]. Radium-A [Sch13h]. Radium-A-Gehalt [Sch13h]. Radiumforschung [Sch18d]. Radon [GM94]. Rahner [Jah56]. raising [Win13a, Win13b]. Ramsey [Sch14a]. random [Sch18c]. Randomness [Sau16, Hal15]. Ranger [Ter90]. Rate [Sch44h]. Rats [Gef14a]. Rätsel [Sch29e]. Raum [Aud87, Sch54e, Sch93]. Raum-Zeit [Sch93]. Raum-Zeit-Struktur [Aud87]. Raumes [Sch38b]. Raupe [Sch87c]. Rausch [RGS32]. Rays [RFS^{+}91, Sch51b, Sch14d, KS14]. Re [Pra93]. Re-Reading [Pra93]. Reaction [Han77b]. reactions [Fuk87]. reader [HT10]. readership [Sal14]. Reading [Pra93, D’A09]. reads [Dic85]. Ready [Fra05]. reaffirmation [MS88]. Real [DT18, vCK17, Sch86i]. Realism [Mah01, Ben92, BM01, Fin86]. realistic [Joh07, Kur03a]. reality [DeW70, EPR35, Göt92a, Gri84a, Gri95b, Gri95a, Gri13a, Wig67c]. Realization [Tes86, Ano96]. realm [Kae48]. reasons [Dar11, Dar13, Sch86i]. Reborn [Gil09]. receiving [KK75]. Rechenberg [Hen88, Sch87i, Wes91]. reciprocity [Sch53b]. Reclam [Sch20a]. Recollections [Shi02, Ste02a]. reconsidering [BV09]. Reconstruction [NM12]. recordings [Sch02]. Recueil [BD95]. Red [Sch39a]. Red-Shift [Sch39a]. reduction [Ghi99, Ghi02, Pen02]. Reexamination [Gao18]. reference [Gis02]. Reflections [Shi83, Shi86, Shi02, Wig67d]. Refugee [Seg85].

Reincarnation [Mer08]. Rektor [Ano11b]. related [Dar92]. Relation [Sch44k, Sch47e, Sch55f, Sch25g, Sch26k, Sch41d, Sch66b]. relations [DD13, Par88, Sch35g, Sch36c]. Relationship [Sch94c]. Relative [Eve57, Whe57]. relativiste [Sch32e]. Relativistic [Sch53b, Sch54f, Smi23, Gar53, Gis02, Kra92, Sch30e, Sch32e, Sch53a]. relativistischen [Gar53, Sch30e]. Relativität [Sch35k]. Relativitätsforderung [Sch25b]. Relativitätstheorie [KPMN20, Sch31e, Aud87, Sch25e]. relativité [Sch33b]. Relativity [Sch25e, Sch31e, Sch35h, Sch95b, Sel02, Kra92, Sch25b, Sch35k, WB00, Aud87, Sch33b, Sch37a, Sch40c, Sch53c, Bun53]. Relevance [Bur65]. Religion [Sch61e, Sch86c]. Remark [Sch24a]. remarkable [Sch23b]. Remarks [Wig67b, Sch35n, YKZ03, Sch25a]. remarques [Sch35n]. Remember [Pra93]. Reminiscences [Blo76b, Ste02a, Thi11, Thi13]. Renn [Ano11b]. repeaters [Zol99]. Reply [Sch61b, Pla29, SP29]. Report [BWE^{+}29, Ano04c, KHFA67, Mic98b, Mic03, Sch26f, Sch26g, Sch26h, Sch26i]. Reports [HSD34b, GSL26]. représentation [Sch92b, Lar93]. Reprint [SP29]. Republic [Pér20]. Requirement [Sch95b, Sch25b]. rescue [Dor87]. research [Hal87, Sch17a, Sch17b, Sch18d, Sch19a]. reshaping [Mil97]. résolution [Dat38]. Resolved [RFS^{+}91]. Response [Hen88, Meh88, MR01]. Rest [Sch42c]. Rest-Mass [Sch42c]. result [CD03]. results [Frö03, Sch17a, Sch17b]. retirement [ABdB^{+}53]. retrieval [Ben02]. Return [D’A92a, SKD^{+}87]. reversal [Ste03]. reverse [Sch31f]. Review [Ack55, Ano35b, Ano45a, Ano45b, Ano59, Bal68, Bar59, Bas56, Bec54, Bel51a, Bel51b, Ben48, Bro36, Bun53, Bur55, Bus20, C.35, Car45, Cor52, Cus96, DT18, Dav57, Del45, Dro20, Ede55, Egl56, Fin53, Fow29, Fre61, Fre12, Han62, Hei68, Hen88, Hen12, Hie68, Ish20, Jah56, Kil65, Kil90, Kir55, Kle65, Kle68a, Kle68b, Kle84, Klo63, Lar90, Lar93, LH53a, LH53b, Lin57, Lon52, M.47, McC51, McC56, Mic68, Moo90, MF91, N.46, Nag52, Opp59, Pol96, Pol59, Pos62, Pra93, Rob56, Sau16, Sch21c, Sco67a, Sei61, Spr55, Sti60, Tem54, Ter90, Tho66, Tho55, Wal54, Wes91, Whe17, Wig55, Yan56, Cha13, Sch25d, Sch26d, Sch26e, Sch29h, Sch30a, Sch31c, Sch32c, SW32, Sch33e, Sch34a, Sch35j, Sch40b]. reviews [Ano90, HF46]. revised [Sch14b]. Revolution [Han62, Kae48, Bel99, Gri12, Gri13b, Kle05, WB00, Ano05, Kle05]. Rezension [Kai85b]. Rezensionen [Krö98]. rhetoric [Cec01]. rich [Atk87]. Richard [Sch41e]. Richer [Bun53]. right [Bre84a]. rikigaku [Sch82b]. Rise [Hen88, Wes91, MR82a, MR82b, Meh87a, Meh87b, MR87a, MR87b, Meh88, MR01, Sch15b]. road [RW11]. Robert [Sch12d, Sch13i]. role [Göt92b]. Roman [Bel07]. Rome [Col98]. Röntgenstrahlen [Sch14d]. root [Yan87, Yan13b]. Rosen [Bel64]. rotation [Sch24h]. Rotationswärme [Sch24h]. Roth [Sch13d]. Round [Kur03a]. route [Wes79]. routes [HD13]. Royal [Sch28e, Cla23, Sch28b]. runs [FS06]. Russian [Ano33]. Rutherford [Lak96]. Rydberg [AP03].

S [Ben34, BZ02, Deh36, Sch14a, Sch20c, Sch29h, Sch32c, Sch40b, Wil36]. S.

[Klo63]. saggio [D’A09]. Sallhofer [Lak96]. Same [FR13a]. Sammlung [Sch20c]. satisfiability [Sch25b]. save [Mer08]. scale [tH02]. Scattering [Sch43a, CD03, Kar03]. Scattering-power [Sch43a]. Schallstrahlen [Sch24f]. Schärfe [Sch14d]. scharfen [Sch21b]. Schicksal [ARK^{+}99]. Schicksal/Kärger [ARK^{+}99]. Schizophrenic [Tau96]. Schluß [Sch17b]. Schmelzpunkt [Sch12c]. Schmetterling [Sch87c]. Scholars [Krö98, AS96]. school [DL82b, Sch44j, Sch52e]. Schöpfer [Ful15]. schreitet [Sch13j]. Schrödinger [Adl03, Ano35a, Ano35b, Ano51, Ano93, Ano03a, Ano04a, Ano04c, Ano06, Ano11b, Ano11a, Ano11c, ARK^{+}99, Aud87, BRZ88, Bal68, Bal86, Bar59, Bas56, Ben34, Ber88, BC94, BG61, Bro36, Bun53, Bur65, C.35, Cam88, Cor52, Cus96, Deh36, Dir61, Fow29, Frö11, GRY06, Gar53, GM94, GHKO99, Haa34, Hei68, HF46, Hei61a, Hei63, Hei61b, Hen88, Hie68, Hou16, Ish20, Jah56, Kai85a, KS86, KDK87, KDK^{+}98, KDK^{+}99, KBG11, Kil65, Kle68a, Kle68b, Klo63, Lak96, Lar90, Lar93, Lin57, Mat75, Mik94, Moo12, Moo15, OGT92, Osk03, Paw12, Pla29, Pol59, Pra93, Prz86, Sal01, Sau16, Sco67a, SM35, TU77, Thi11, Tho55, Wal54, Wil36, Ack55, Ahl96, Aja15, Ald99, Ano35a, Ano38, Ano45a, Ano45b, Ano49, Ano59, Ano61]. Schrödinger [Ano87a, Ano87b, Ano87c, Ano88, Ano96, Ano97, Ano03b, Ano04b, Ano10, Ano12, Ano15, Deg89, Atk87, Aud87, Axo89, AH01, BJK03, Bal92, Bal90, Bal99, Bal00, Bal03, Bal11, Bec54, Bel51a, Bel51b, Bel92, BD95, Ben92, Ben48, Bit92, BD92, Bit96a, Bit96b, Bit96c, Bla00, Bli01, Bor63b, Bra92, Bre84b, Bre84a, Bru04, Bru09, Bun01, Bur55, Bus13, Cam88, Car45, Car13, Cec01, Cer93, Cha13, Che92, Cho00, Cla23, Coh92, Coh05, Col00a, Col00b, Cou62, Cro01b, CS96, D’A92a, D’A92b, D’A01a, D’A09, DKF92, DC09, Dar92, Dar11, Dar13, Dat87, Dat38, Dav57, DL82b, DL82a, Dic85, Dir64, DS83, Dor87, DD13, Ede55, Egl56, Ein27, Eps26, Fel06, FR13a, FR13b, Fin53, Fla87, Fle18, Fow29, Fre61, Fri89, Fro38, Ful15]. Schrödinger [Gao18, Gar53, Gar05, Gav28, Gee05, Gef14a, Gis06, GGW94, GGGG05, Gol90, Göt92a, Göt92b, Gri84b, Gri84a, Gri95b, Gri95a, Gri12, Gri13b, Gri13a, Gri22, GB08, GHHL10, G^{+}11, Haa34, Hal87, Hal15, Han77a, Han79a, Han79b, Han62, Haw87, He20, Hei63, Hit87, Hoa96, HY87, Hof84, Joh07, Kai85b, Kat90, Ken31, KS86, Kil87b, KKS^{+}11, Kir55, KK75, Kle65, Kle84, Kle96, Kni92, Koc78, Kox13, Kra82, Kra92, Kra11, Kra13, LH53a, LH53b, Leg84, Leg90, Leg00, Leg09, Leg13, LKS^{+}05, Loc92, Loe84, Lon52, Mac24, Mal12, MZP97, Mat05, McC87a, McC51, McC56, McC87b, McG99, McL99, Meh87a, Meh87b, MR87a, MR87b, Meh87c, Meh88, MR01, Mel88, Mer08, Mil97, Mil03, MMKW96, Mon09, MB11a, Moo87, Moo89, Moo90, MF91, Moo94, Moo96, Mor96]. Schrödinger [N.46, Nag52, Nef92, OGT92, Olb71, Opp59, OTBLG06, OJTBG07, Par88, Pau87, Per91, Pér20, Per64, Per87, Pie11, Pla28, Plo10, Pol96, Pop89, Pos62, Prz63, Prz67, Prz11, RF69, RA11, Rei44a, Rei98a, RY13, RW11, Ren13, Rob56, Rüg88, Sal01, SR92, Sch27d, Sch29b, Sch29c, SP29, Sch31b, Sch32a, Sch35i, SKD^{+}87, Sch87c, SB90b, SvM11, Sci61, Sco67b, Sei61, Sei96, Ses13, Shi83, Shi86, Sig96, ST65, Smi23, Sof14, SM35, Spa05, Spr55, Ste91, Ste05, Sti60, Sut08, SS53, Sym87, Tau96, Ter90, Tes86, Tes00, Thi87, Thi13, Tho66, Tri80, Tut05, Ven05, Vil86, Vis91, WH17, Wes79, Wes83, Wes91, Wes57, Wig55, Wil16, Wil99, Win13b, WZ51, Woo06, Wra09, Yam97, Yan56, Yan87, Yan13b, qYL17]. Schrödinger [Yox79, Zeg05, Zur00, dB65, dG14, dR97, dR01, kA13, vM92, Coo18, Tem54, Ano90, Bit93, Bus20, Ebe09, Sal14, Fre12]. Schrödinger- [SM35]. Schrödinger-cat [KKS^{+}11]. Schrödinger-Gleichung [Ano04a]. Schrödinger-Symposiums [Ano11b]. Schrödinger. [Kil90]. Schrödinger/Wilhelm [BG61]. Schrödingers [Ein27, Gri13a, Hen12, KPMN20, RFS^{+}91, Sch90, SvM11, Han77b, Win13a, RW11]. Schroedinger [Ano34, Mic68, Ano33, D’A09, M.47]. Schroedingera [SS53]. Schüler [KPMN20]. Schwankungen [Sch19b]. Schweidler [Sch19b]. Schweidler’sche [Sch19b]. Schwerefeld [Sch32b]. Schwingungen [Sch22c]. Science [ADS83, BIT01, Bro36, Bur65, CW85, Egl56, Far03, GGK02, Jon08, Kae31, LMOS15, Mer38, RW94, Sch96a, Sch31d, Sch35o, SMJ35a, SMJ35b, Sch40b, Sch40a, Sch45a, Sch51c, Sch54h, Sch57b, Sch57c, Sch61f, Sch96c, Sch96d, Cec01, Far02, FS06, HT10, Lev17b, MZP97, Pic08, PS33, Sch20a, Bro36, Bun53, Han62, Kae31, Sch29d, Sch30d, Sch30h, Sch86c, Sch88a, Sch45c, Sch45d, Sch47a, Sch61e, Sch66a, Ano35b, Ano59, C.35, Fin53, LH53a, LH53b, Lon52, Nag52, Pos62]. sciences [Sch13c]. Scientific [Ano04c, ADS83, BM01, Bus20, Dro20, Kir95, Krö98, Mah01, Mic98b, Mic03, Rod19, Sch31c, AS96, Haw11, Meh87a, Sch35n, Sch48d, Sch56e, Sch56f, Sch79, Sch84f, Sch97, Sch08b, Wig67d, ABdB^{+}53]. scientifique [Sch35n]. scientist [Sta36]. Scientists [Acz01, DG99, Kae31, Krö98, Rog10, Seg85, AS96, HD13, Pie97, Sch31b, Str11, WB00]. scienza [Pol96, Sch88a]. scienziato [BC94, Pol96]. Scope [Bus20, Ish20, Bus13, Cha13, Sal14]. Scott [Bal68, Kle68a, Hei68, Kle68b]. script [WH17]. scritti [Sch87d]. scuola [DL82b]. search [Gri84b, Gri84a, Gri95b, Gri95a, Gri13a]. Second [D’A01a, Sch12b, Sch13f, Sch13i, Sch20a, Sch20c, Sch26i, Sch33d, Zur99, Ano03c, Pér20, Stu18]. Second-Quantisation [D’A01a]. secondary [KS14]. seconds [Sch13c]. Secretary [Pla29]. see [Pie11]. Seeham [Sch13h, Sch13h]. seen [Hit87, Kar03, SM89, Sch99b]. seguido [Sch88c]. Sein [Sch86i]. seine [KPMN20]. seiner [Sch32c, vM92]. seines [KS86]. sekaikan [Sch87h]. Sekretars [Pla29]. Sekundärstrahlung [KS14]. Sekundenpendels [Sch13c]. Selbststudium [Sch13f]. Selected [Mah01, Yan13a, BM01, Réd05, Bor63a, Fow29, Sch70]. self [Sch13f]. self-study [Sch13f]. Semiclassical [Jam74a]. Semiconductor [ARK^{+}99]. seminar [Sch44j, Sch46b, Sch52e]. Seminars [Cus96, Sch95a]. sensation [Sch25h]. sensations [Sch26c]. sense [Ein27]. sensitivity [Sch24g]. separated [Sch35g, Sch36c]. separation [Wei02]. Sept [Ano99, Ben99, Bla99, Bou99, Che99, Ghi99, Got99, Har99, Hil99, Hor99, How99, Kim99, LL99, Oes99, Omn99, Orl99, Rau99, Shi99, Thi99, Yng99, Zei99, Zol99, Zur99, Ano05, Kle05]. September [AP03, Ano03c, Arn03, BRZ88, Bar03, Bla03, CD03, FF03, GRY06, GM94, GRZ99, Kar03, Kie03, Kur03b, Kur03a, Rai03, SK03, Sim03, Ste03, YKZ03, Zur03]. series [Ano03a, Sch14e, Sch18c, Sch21b]. session [Ano99]. seven [Kle05]. seven-times [Kle05]. Seventh [Far01]. shôden [Sch82b]. shaky [Fin86]. Shaping [Cec01]. Share [Ano33]. sharp [Sch21b]. sharpened [Bel05, Bel06, Bel07]. sharpness [Sch14d]. sheep [Zur00]. Shielding [MS44]. Shift [Sch39a, Sch55f]. shokan [Sch82b]. shook [Haw11]. shore [Sch89f]. Short [Ano33]. Short-Story [Ano33]. should [Loc92, Sut08]. Shurêdingâ [Sch82b]. Sides [FR13a]. signification [Sch55d]. Simplifying [vCK17]. Simson [Wes57, Wes57]. simulation [MZ13]. since [Meh75]. Singapore [Bus20]. Single [Hür22, Bla99, Sch23b]. Sinne [Ein27]. Sir [Adl03]. Situation [Sch35f, Sch35a, Sch35b, Sch84a, Sch96b, Tri80]. Sitzung [Pla29, Sch14e]. six [LKS^{+}05]. six-atom [LKS^{+}05]. Sixty [FR13c]. sketch [Sch35i]. Sketches [Bit92, Sch92c, Sch12, Pra93]. skin [Bal99]. Skurrile [Arr06]. Slurred [Sch29k]. Small [Ano11b, Sch33f, Sch34c]. so-called [Che99]. sobre [Cam88, GHHL10]. Social [Mer38]. Socialism [RW94]. Society [GGK02, Cla23]. soft [KS14]. Software [ARK^{+}99]. Solid [RFS^{+}91]. solids [Sch19a, Sch22c, Sch22a]. Söllner [Krö98]. Solution [Eck26, Fro38, Sch44c, Dat38, Hou16, Sch18e, Sch26d, Ses13, SM35]. Solutions [Sch42c, RA11, Sut08]. Solvay [BV09, Far01, Meh75]. Solving [Sch41c, Gri95b]. Some [Coh92, L.88, Bra92, Dar92, Hal87, Sch24c, Sch35n]. Sommaire [Sch53d]. Sommerfeld [Lak96, MR82a, MR82b, SW32]. Son [Sch14a, Sch35p]. Sonderdruck [SP29]. sound [Sch24f, Sch23a]. Source [Sch42c]. Source-Function [Sch42c]. Sources [Egl56, KHFA67, Han79b]. sous [Che92]. sous-détermination [Che92]. Space [Rüg88, Sch50a, Sch54g, Sch85b, Aud87, Sch38b, Sch54e, Sch93, Wei02, Zei97, kA13, Cor52, Sch54e, Bel51a, Bel51b, McC51]. space-like [Wei02]. Spain [SR92]. Spanish [Pér20, Sch88c, Cam88, GHHL10, HBS96, Moo96, Sch35l, Sch35p, Sch01, W^{+}87, WdC02]. Spatschek [RFS^{+}91]. Speakable [Bel87b, Per02]. speakables [BZ02]. Speaking [Krö98, AS96]. Special [Sel02, Sch31e]. spécialement [Sch33c]. specialist [Bus20]. specialists [Bus13]. specific [Sch22c, Sch22a, Sch24a, Sch26j]. specificity [Sch48d]. spectra [Sch25c, Sch29k]. spectral [SW32]. spectroscopy [KKS^{+}11, Pie11, RFS^{+}91]. speculations [MO95]. speech [Pla29, SP29, Sch61b, Sch56d]. Spektrallinien [SW32]. Spektren [Sch25c]. Spezielle [Sch31e]. spezifische [Sch22c, Sch22a, Sch24a, Sch26j]. sphärischen [Sch38b]. spherical [Sch38b]. Spiel [Bel07, Sch30h]. Spin [SH42, Sch55h, Sch55g]. Spirit [Sch54h, Sch47a, Sch66a, Sch86i, B^{+}54, Jah56]. Spookiest [Acz01]. Sprengstoff [Mül40]. Springer [Hen88, Klo63]. Springer-Verlag [Hen88, Klo63]. Springs [ADS83]. Square [Yan87, Yan13b]. Squid [Col00b, Cho00]. srodowiska [PS33]. SS [Kai85a, Kai85b]. Stabilität [Gar53]. Stability [Lie97, Gar53]. standpoint [Sch25c]. Standpunkte [Sch25c]. star [Sch25h]. Stark [RGS32, Sch26f, Eps26]. Starkeffekt [Sch26f]. Starkeffekts [RGS32]. stars [Sch24c]. startling [Gri84b]. State [Eve57, MMKW96, RFS^{+}91, Whe57, Ano10, Gis02, Kle02, LKS^{+}05, Pen02]. states [FF03, KKS^{+}11, OJTBG07, Thi99]. statics [Sch12d]. Statik [Sch12d]. statistica [BES86]. Statistical [Gal03, Sch44i, Sch44j, Sch46b, Sch52e, Sch58a, Sch89g, Dar92, Sch18d, Sch25a, Sch84c, HF46, Sch52f, Wal54]. statistics [BES86, Ein27, Han77a]. Statistik [Ein27]. statistische [Sch25a, Sch52f]. statistischen [Sch18d, Sch84c]. Stature [McL99]. Steigversuche [Sch15b]. Stellung [Sch57d]. Sterne [Sch24c]. Sternfarben [Sch25h]. stetige [Sch26a]. still [Kae48, Wra09]. Stochastic [MS88]. Stockholm [HSD34a, Wil36]. Storage [Ben02]. Störfall [RFS^{+}91]. storia [Pol96]. Störungstheorie [Sch26f]. Story [Acz01, Ano33, Jon08]. Strahlen [KS14]. Strahlung [Sch12e]. Strahlungshypothese [Sch24b]. Strahlungstheorie [Sch24a]. Streben [Sch04]. Structure [Ano35b, Hug89, Rüg88, Wil16, Aud87, SW32, Sch41d, Sch50a, Sch54g, Sch85b, Sch93, Bel51a, Bel51b, Cor52, McC51]. Struggling [Ben92]. Struktur [Aud87, Sch93]. Student [Fle18, Ish20]. students [Sch59a]. Studien [Sch19b, Sch12c]. Studies [HS51, Sch41c, SH51, Sch19b, Sch70, Sch12c, Sch52e]. study [Sch13f]. stuff [Haw11]. stupor [Cam88]. Stuttgart [Sch13d]. sub [CD03]. sub-femtosecond [CD03]. subfactors [Jon03]. subjective [Sch25h]. subjektiven [Sch25h]. subnuclear [Zic02]. subordinate [Sch21b]. subspaces [FF03]. Suche [Gri13a]. sudden [RGS32]. sujet [Sch35n]. Sulla [Sch37b, Sch86g, DL82b, HSBA60, Sch26m]. sull’interpretazione [DL82b]. Summary [BS56]. summer [GRY06]. Sun [Sch44a]. Superconducting [Tes86]. Superposition [MMKW96, Win13a, Win13b, FF03, Zei99]. superpositions [Sim03]. Supplement [Sch17d]. Supply [Sch13e]. suppressing [FF03]. Supra [Sch35m, Sch36b]. Supra-conductivity [Sch35m, Sch36b]. surface [Sch10]. Surfing [Lev17a]. Surprises [Jon03]. Survive [He20]. Swapping [Bou99]. Symmetric [SH51, PS51]. Symmetries [Wig67d]. Symmetry [MS45, Bar03]. Symposium [Ano11b, Ano11c, GRZ99, TU77, Ano11a, GRY06]. Symposiums [Ano11b]. Synthesis [BSM06, HY87]. System [Sch44c, Ein27, Göt92b, MZ13, Sch13c, Sch18e]. Systematics [Sch44d]. Systems [Ano03c, Ein27, RFS^{+}91, Sch13c, AP03, Ben02, Kur03b, Kur03a, Sch14f, Sch35g, Sch36c, Ste03]. Szilard [Fra05].

T [Bal68, Hei68, Kle68a, Kle68b]. table [Kur03a]. Tagessehen [Sch20e, Sch20f, Sch20g]. Tait [ABdB^{+}53]. Tale [Koc78, Car93]. talentierter [KPMN20]. Tales [Hob17, Whe17]. tanulmányok [Sch70]. Tarner [Sch40b, Sch58b]. taught [CG14]. teacher [Bus13]. teaching [Sch13f]. Techniker [Kai85a, Kai85b]. Technology [RW94, Mil97]. Teil [ARK^{+}99, RGS32, Sch12b, Sch13e, Sch26j, Sch35j]. Teilchen [SS97, Sch15b]. Teleportation [Bou99, Col98, Hil99]. tells [FS06]. Temperament [SMJ35a, SMJ35b, Bro36, Ano35b, C.35]. Temperatur [Sch22c]. temperature [Sch22c]. temporal [dB65]. temporelle [dB65]. temps [Ano59]. Ten [Jon08, Ano10]. ten-qubit [Ano10]. tentative [dB24]. tenth [Wes57]. teoria [Sch26m, Sch37b, Sch86g]. term [Sch21b]. Terms [Sch21b]. Test [Gav28, Gis02, Cam88, Fry02]. testament [Sch06, Sch06]. testi [Pol96]. testo [D’A09]. Tetrads [Sch43b]. Teubner [Kai85a, Kai85b, Sch12b, Sch12d, Sch13f, Sch13i, Sch14b, Sch17c]. text [D’A09]. textbook [Ish20, Sch26e, Sch13f]. Textbooks [NM12]. Thales [Lak96]. their [Cer93, Sch13a, Sch29f]. them [Pic08]. themes [Dar92]. Theodosius [Adl03]. Theorem [Cla02, Eke02, Shi02, Zei02, Asp02, Sch24b, Wei02, Gri22]. theoretic [Sch33e]. Theoretical [ARK^{+}99, Bit96c, D’A00, Plo10, Sch44j, Sch52e, CD03, D’A01b, RGS32, Sch26j, Sch30a, Sch34a, Sch12d, Sch13i]. Theoretiker [Sch48g]. Theoretische [RFS^{+}91, Sch30a, Sch34a, Sch12d, Sch13i]. theoretischer [RGS32, Sch26j]. Theorie [Sch12e, Sch12f, Sch13a, Sch13e, Sch13g, Sch14g, Sch14b, Sch15b, Sch19b, Sch20e, Sch20f, Sch20g, Sch26e, Sch33b, Sch32e, Sch38d, Sch14b, Sch20b, Sch20h]. Theorien [Ano35a, Deh36, Haa34, Sch25d]. Theories [Kae31, Sch44f, Sel02, Ano35a, Che92, Haa34, Sch25d, Che92]. Theorists [Sch48g]. Theory [Acz01, Aud87, BS35, BS41, Bro36, D’A92a, D’A01a, Eps26, Gav28, Han77b, HSD34a, Hen88, Hob17, HR74, Hun74, MR87b, PS51, Pos62, RFS^{+}91, Sau16, Sch13e, Sch26e, Sch26l, Sch35c, Sch35m, Sch36b, Sch37a, Sch40c, Sch44a, Sch44b, Sch44g, Sch44k, Sch47c, Sch47d, SH51, Sch53c, Sch53a, Sch54f, Sch94a, Sch94c, Sci61, Wes91, Whe57, Ano11a, BV09, Che99, D’A92b, Fin86, Ghi99, GHKO99, Hal15, HS51, Hit87, Lev17b, MR82a, MR82b, MR01, Nar03, Sch12e, Sch12f, Sch13a, Sch13g, Sch14g, Sch14b, Sch15b, Sch19b, Sch20e, Sch20f, Sch20g, Sch24a, Sch25g, Sch26m, Sch26f, Sch31e, Sch32e, Sch33b, Sch38d, Sch46a, Sch48c, Sch57b, Sch57c, Sch61a, Sch84d, Sch84e, Sch86g, Sea87]. theory [Wei14, dB24, dR97, vM92, Sch14b, Sch20b, Sch20h, Sch25e, Sch25g, Sch37b, Sch57a, Sch87i, Whe17]. Theory/Brückner [RFS^{+}91]. There [Sch52b, Sch52c, Bel87a]. Thermal [Lei02, Sch24f]. thermische [Sch24f]. Thermodynamic [Par88, Sch55f]. Thermodynamics [Sch52f, Sch44j, Sch46b, Sch52e, Sch89g, Sch33d, HF46, Wal54]. Thermodynamik [Sch33d, Sch52f]. Thesis [Han77b]. thickness [dB65]. thing [Sch12d]. Think [Sch49a]. Thinking [Plo10, MZP97]. Third [Sch26f]. Thirring [Ano07, Gaw97, Jon97, Lie97, Zei97, Ano07, Pie97]. Thomson [Lak96]. those [Sta04]. thought [Bit93, Moo89, Sch87d, Sta04, Ter90, Kil90, MF91, Ano90]. thoughts [Mel88]. Three [Ano33, Bal86, HSD34b, Loe84, Sch44e, Hor99, Sch25g, SF11]. three-color [Sch25g]. Three-Man [SF11]. three-particle [Hor99]. Time [RFS^{+}91, Rog10, Rüg88, Ste03, Aud87, Bel51a, Bel51b, Cor52, Kle02, Lev17b, McC51, Sch50a, Sch51c, Sch54e, Sch54g, Sch61f, Sch85b, Sch93, Vis91, Bun53, Fin53, LH53a, LH53b, Lon52, Nag52]. time-dependent [Vis91]. Time-Resolved [RFS^{+}91]. timer [Sch12d]. Timerding [Sch12d, Sch13i]. times [Cro01a, Kle05]. Toast [Sch35q]. today [Bla03]. Todes [Wes57]. Todestages [KS86]. together [Sch82a, Sch03]. token [AH01]. Tompkins [Pra93]. Ton [Sch23a]. Tonaufnahmen [Sch02]. Tonträger [Sch02]. Too [Hür22]. topological [Kib87]. totalité [Bit92]. totality [Bit92]. Töten [Bar00]. tradition [Wes83]. traduction [Sch33c]. Tragedy [Jon08]. Tragweite [Hen12, SvM11]. Traité [Sch14c]. Trajectories [Smi23]. trajectory [Sch23b]. transactional [Cra86]. transform [GM94]. transformation [Sch30c]. transforms [Osk03]. transition [Sch26a, dB65]. Translated [Hie68, Sch13b, Sco67a, Han62, Mic68, Sch14c]. Translation [Tri80]. transport [Frö03]. trapped [MZ13]. Traubenberg [RGS32]. Traveling [Dro20]. Travelling [Rod19]. treatise [Sch14a, Sch14c]. Tremor [Sch89h]. Trespassing [Gef14b]. Triads [Sch43b]. Trieste [GRY06]. trilogy [Wil99]. Trinity [Sch58b]. Trinkspruch [Sch35q]. true [Sch35p]. Truth [Del45, Bun53]. Tuesday [Ano06, Ano07]. turbulence [Gaw97]. Twentieth [D’A00, D’A01b]. Twenty [CW85]. Twenty-five [CW85]. twilight [Sch25h]. Two [FR13a, McL99, Sau16, Sch44c, Wig67c, Hal15, Kle02, Sch24a]. two-photon [Kle02]. Two-Wave-System [Sch44c]. Twofold [MS45]. type [Gaw97].

U [Sch13i, ARK^{+}99, Haa34, RFS^{+}91, Sch17c, Wil36]. U.P [Whe17]. Übergang [Sch26a]. Übersetzt [Sch13b, Sch14c]. Uhr [Ano06, Ano11b, Sch54e]. ultrafast [Kar03]. ultrashort [Kle02]. umanesimo [Sch88a]. Umberto [Pol96]. umfassendem [Sch04]. umgearbeitete [Sch14b]. Umkehrung [Ber88, Sch31f]. Unanwendbarkeit [Sch34c]. unbounded [Kae48]. Uncertainties [RFS^{+}91]. Uncertainty [Kae31, Meh87c, Sch30i, Sei05, CG14, DD13]. Uncovered [Hür22]. underdetermination [Che92]. undergraduate [Ish20]. Underlie [Sei05]. understand [FS13, Sch87c]. Understanding [Hob17, Sch50b, Whe17, Sch54i]. Undulatory [Sch26l]. Unfinished [DT18]. ung [Ano35a]. Unified [Sau16, Sci61, Hal15, Hit87]. Unifying [Kra92]. Union [Sch44e]. unit [Sch86a]. Unitary [Sch44a, Sch44b, Sch44g, Sch44k]. Univ [Bel51a, LH53a]. Univers [Sch41d]. universal [Sch89i]. universalen [Sch89i]. Universe [HW86, Kae31, Sch54e, Haw87, Sch38c, Sch39b, Sch41a, Sch41d]. universes [Sch56c, Sch10, Bas56, Dav57, Lin57, McC56, Rob56, Yan56]. Universität [Ano11b, Sch30a, Sch34a]. University [Ano11b, ABdB^{+}53, Bun53, Ish20, Kil90, Kle68a, Krö98, Sch30a, Sch34a, SS53, Ter90, Ano38, Hal87]. Unlikely [Acz01]. Unorthodoxy [DL82a]. Unpublished [Cus96, Sch75b, Sch87d, Sch95a]. Unravelling [FR13c]. Unschärfeprinzip [Sch30i]. Unsere [Sch53d, Sch53e, Sch54i]. unsigned [Sch29a]. unspeakable [Bel87b, Per02]. Unterricht [Sch13f]. Unveil [Cho00]. Upanishad [Sch59a]. Upgraded [He20]. Uran [Mül40]. uranium [Mül40]. Uranprojekt [KPMN20]. Ursache [Sch29d]. Ursprung [Sch24g]. usability [Mül40]. use [Car93, Sch13a, Sch13f]. using [FF03, SK03]. usw [Sch22a].

V [Sch31a, Wes91, Dat87]. vacation [GRY06]. Valen [Sch20c]. Valera [McC87b]. Válogatott [Sch70]. Value [Sch29f]. Vanishing [Sch42c]. vapors [Sch21c]. vector [dR02, Sch20c]. Veda [Sch82d, SS53, Sch82d, SS53]. Vedantic [Sch86b]. vedantische [Sch86b]. vein [Joh07]. Vektoranalysis [Sch20c]. velocity [HS51]. Verabsäumte [Sch30f]. verb [Haa34]. verbesserte [Sch13f]. verdaderas [Sch35p]. Verfahren [SM35]. Verhalten [RGS32]. Verhältnis [Sch25g, Sch26k]. Verlag [Ben34, Hen88, Klo63, Wil36]. Verlagsgesellschaft [Deh36, Kai85a, Kai85b]. verm [Haa34]. vermehrte [Sch13f]. verschiedenen [GSL26]. version [Ber88, Sch12d]. verstehen [Sch87c]. verstorben [Sch25f]. Versuch [Sch25e, Sch32c, Sch21b]. Versuche [Sch13c]. Verteidigung [Sch89i]. vertical [Sch12e]. Vertrieb [ARK^{+}99]. Verwaschene [Sch29k]. Verwendbarkeit [Mül40]. Verwendung [Sch13a]. vibrations [Sch39b]. vida [GHHL10, Moo96, Sch88c]. vie [Miq07, Sch86d, Sch86e, Sch86f]. Vienna [Ano06, Ano11b, BRZ88, GM94, GHKO99, Hen88, Sch56d, TU77, BZ02, BIT01, Hal87, KDK^{+}98, KDK^{+}99, Kir95, MR87a, MR87b, Sch56a]. Vier [Sch28e]. Vierfarben [Sch25g]. Vierfarben- [Sch25g]. Vierte [Sch14b, Sch26h]. View [Bur65, Sch88c, Sch89e, Sco67a, Tho66, Asp02, Frö11, Göt92a, Joh07, Mel88, Sch53d, Sch59b, Sch64a, Sch79, Sch82d, Sch83, Sch85a, Sch97, Sch06, Sch08a, Sch08b, Sch48d, Sch86b]. Views [Kle68b, Ano11a, Hal87]. vigencia [GHHL10]. viii [Hen88]. Vincent [Sch13j]. Vinzenz [Sch13j]. viru [Bru09]. visibilities [Hor99]. vision [Sch20e, Sch20f, Sch20g, Sch87d]. visione [Sch87d]. viszonya [Sch66b]. vita [Sch87d, Sch88a]. vivante [Sch86e]. Vol [Sch12b, Sch12d]. vollständig [Ein27]. Volume [Hen88, Wes91, Sch13i, SvM11, MR87b, MR01]. vom [Pla29, Sch25c, Sch52g, Sch23c]. vor [Sch13j, vM92]. vorgelegt [Sch14e]. Vorlesungen [Sch13f, Sch28e, Sch30a, Sch34a]. Vorstellung [Sch53d, Sch53e, Sch54i]. Vortrag [Asp02, Ben02, Ber02, Cla02, Eke02, Frö03, Fry02, Gal03, Ghi02, Gis02, tH02, Hor02, Jon03, Koc02, Lei02, Nar03, Osk03, Pen02, Per02, Raj02, Rau02, Sch52g, Sch98, Sel02, Shi02, Sut02, Wei02, Whi02, Zei02, Zic02, d’E02, Eng02, Jac02, Kle02, Ste02b, dR02]. Vorträge [Ano04a, Wil36, HSD34a]. Vorwort [Sch29l]. vs [Kra11].

W [Ben34, C.35, Cha13, Sch13a, Sch14c, Wil36]. Waddington [Adl03]. Waerden [Sch33e]. Waga [Sch87h]. Wahlvorschläge [KK75]. Wahrscheinlichkeitstheoretische [Sch19b]. Walter [Ano07, Bit93, Gaw97, Jon97, Kil90, Lie97, Paw12, Zei97, Pie97, Ter90]. Wandlung [Sch30c]. Wandlungen [Sch13b]. want [Sta04]. war [Gol90]. Wärme [Sch22c, Sch22a, Sch26e, Sch26j]. Wärmen [Sch24a]. Wars [Stu18]. Warum [Sch33f]. Was [Gil09, RF69, GB08, Sch29m, Sch30g, Sch51d, Sch52g, Sch79, Sch86h, Sch86i, SM89, Sch97, Sch98, Sch99b, Sch02, Sch08b]. wasserstoffähnlichen [Sch25c]. Wasserstoffs [Sch24h]. Wave [Cro01b, D’A92a, D’A01a, Fow29, Hen88, Hie68, Kle84, Kra82, Mic68, MB11a, NM12, Prz63, Prz67, RFS^{+}91, Rüg88, Sch40c, Sch42c, Sch44c, Sch55c, Sch55d, Sch55h, Sch75a, Smi23, Wes91, BRZ88, DL82b, Ein27, Fri89, Han79b, Hou16, Kox13, MS88, Meh87a, Meh87b, MR87a, MR87b, Meh88, MR01, Mil03, Pla28, Prz86, Prz11, RW11, Ren13, Sch27a, Sch27c, Sch28a, Sch28b, Sch28e, Sch35e, Sch35l, Sch38a, Sch53c, Sch55g, Sch78, Sch82a, Sch99a, Sch01, Sch03, Wes79, vM92, Aud87, Sch28c, Sch63]. Wave-particle [NM12]. waves [Ano35a, BES86, Dar11, Dar13, Haa34, Sch27b, SS97]. ways [Bal99]. Web [Ano12]. Weber [Sch12b]. Weber-Wellstein [Sch12b]. Webster [Sch21c]. Wednesday [GRY06]. Weg [RW11]. Wege [Sch28d, Sch29j, Sch32d]. Weglänge [Sch24d]. weiche [KS14]. Weird [Str11]. Weisbuch [RFS^{+}91]. weitgeöffneten [Sch20i]. Weizsäcker [Bur65]. Welcome [Ste02a]. Weldbilds [Sch48d]. Wellen [SS97]. Wellenfunktion [Sch38a]. Wellenmechanik [Ein27, Hen12, Pla28, Sch28e, SvM11, Aud87, PS63, Prz63, RW11, Sch27a, Sch27c, Sch34b, Sch35e, Sch55c, Sch63, Sch78, vM92, Kil65, Kle65, Klo63]. Wellstein [Sch12b]. Weltall [Sch54e]. Weltalter [Sch54e]. Weltanschauung [Sch90]. Weltansicht [Sch06, Sch85a, Sch89e, Sco67a]. Weltbegriffs [Sch30c]. Weltbild [Sch32c, Sch79, Sch97, Sch08b]. Weltkraftkonferenz [Sch56d]. Welttheorie [Sch37b]. Weltverstehens [Sch50b]. Werke [Sch13e]. Werner [Adl03, Ful15, Han62, Jah56, Ful15, Meh87c]. Wert [Sch29f]. wes [Haa34]. West [Arm46]. Western [AFG90]. Weyl [Sch44k, Sig96]. Where [Ano12, Yox79, Bro36]. which [HY87]. Whitehead [Adl03]. Who [MZP97, RF69, Sta04]. wichtigsten [Sta04]. wide [Sch20i]. wide-open [Sch20i]. Wie [Sch13j]. Wiederkehr [KS86, SKD^{+}87, Wes57]. Wien [Ano06, Ano11b, Klo63, OGT92, Sch56d, Sch56a, Ano07, KDK^{+}98, KDK^{+}99, OGT92, Sch27d]. Wießner [Sch13j]. Wieviele [RFS^{+}91]. Wigner [Wig67d]. Wild [LMOS15]. wilde [LMOS15]. Wilhelm [BG61, Sch27d]. Wili [Jah56]. Will [Sch36a]. William [Hei68, Kle68a, Kle68b, Bal68]. Wilson [Sch21c, Cec01]. Winkler [Ano11b]. wirklich [Sch86i]. Wirklichkeit [Gri13a]. Wirkung [Sch29d]. wissen [Sta04, Sch04]. Wissenschaft [Sti60, FS06, Sch30h, Sch45c, Sch45d]. Wissenschaften [Sch13c]. wissenschaftliche [Sch84f]. Wissenschaftlicher [ARK^{+}99]. Wissenschaftsakademie [LMOS15]. Wissenschaftstheorie [Sch90]. Within [Smi23]. Without [Gri22, Ghi99]. Wolfgang [ARK^{+}99, Köh76]. wollen [Sta04]. words [Ano11b, Hei63, LL99]. Work [Rüg88, Meh87a, Rau99, Sch13e, Sch31a]. works [Bor63a, DL82b, Sch13e]. Workshop [Ano03c, BRZ88, GHKO99]. World [Arr06, Bus20, Hür22, Lev17a, Sal14, Sch29h, Sch56d, Sch89e, Stu18, Göt92a, Gri84b, Haw11, LL99, Oes99, Sch32c, Sch37b, Sch38d, Sch48d, Sch64a, Sch79, Sch82d, Sch83, Sch85a, Sch87d, Sch92b, Sch97, Sch06, Sch08a, Sch08b, W^{+}87, WdC02, Sch88c, Sco67a, AFG90, Sch30c, Bur65, Tho66]. Worlds [Bru04, SR92]. Worte [Ano11b]. would [Pie11]. wplyw [PS33]. Write [Ano33]. writings [Sch87d, Sch01, Sco67b, W^{+}87, WdC02, Bal68, Hei68, Kle68a, Kle68b]. written [Dic85]. wrong [Bre84a]. wspólczesnej [PS33]. Wu [Zuk79].

X [Sch14d]. X-rays [Sch14d]. xii [Kil90]. xix [Hen88]. xvi [Bun53, Kle68a]. xx [Hie68].

Year [FR13c, Mic98b, Sch98]. Yearbooks [Jah56]. Years [Ano03a, Ano11c, ARK^{+}99, CW85, RY13, TU77, Ano04a, Ano11a, Frö11, GM94, GGW94, Hal87, Hit87, Kae48, Mic98a, Mic03, MO95, Sch48c, Sch61a, Whi02]. York [Bel51a, Bun53, Hen88, Hie68, Kil90, LH53a, Ter90, Whe17].

Zagadnienia [PS33]. zehnten [Wes57]. Zeit [Aud87, Sch54e, Sch93]. Zelle [SM89, Sch99b]. Zentralbibliothek [Ano06, Ano11b, KDK^{+}98, KDK^{+}99]. Zero [BS55, BS56]. Ziegler [Moo90]. Zittern [Sch89h]. Zufälle [Sch29d]. Zufallsreihen [Sch18c]. Zukunft [Sch50b]. zum [GS87, Hen12, SS97, Sch13f, Sch29a, Sch33a, Sch79, Sch97, Sch08b, SvM11, Sch30i]. Zur [Sch17d, Wes57, Ano06, GB08, Hen12, KS86, KPMN20, Kil65, KK75, Kle65, Klo63, Pla28, PS63, Prz63, RW11, Sch12a, Sch13e, Sch21b, Sch25g, Sch26a, Sch27a, Sch78, Sch84c, Sch84d, Sch84e, SKD^{+}87, SvM11, SM35, Sch12f, Sch14f, Sch14e, Sch14g, Sch15b, Sch17e, Sch26m, Sch31g, Sch57d, Sch89i]. Zurich [Hen88, OGT92, Sch89f, MR87a, MR87b, OGT92]. zwei [Sch24a]. Zweite [Sch26i]. zweiten [Sch33d]. zwischen [Sch24f].

References

Appleton:1953:SPP

[ABdB^{+}53] Sir Edward Appleton, David Bohm, Louis de Broglie, Richard Courant, Albert Einstein, Pascual Jordan, Theodore V. Kármán, S. S. Penner, Alfred Landé, Erwin Schrödinger, and Hermann Weyl, editors. Scientific papers presented to Max Born on his retirement from the Tait Chair of Natural Philosophy in the University of Edinburgh. Oliver and Boyd, Edinburgh, UK; London, UK, 1953. vi + 94 pp. LCCN QC71 .S35.

Ackrill:1955:RBG

[Ack55] J. L. Ackrill. Review: Nature and the Greeks by Erwin Schrödinger. The Philosophical Review, 64(2):317–318, April 1955. CODEN ???? ISSN 0031-8108 (print), 1558-1470 (electronic). URL http:

           //www.jstor.org/stable/2182600.

Aczel:2001:EGM

[Acz01] Amir D. Aczel. Entanglement: The Greatest Mystery in Physics: The Unlikely Story of How Scientists, Mathematicians, and Philosphers Proved Einstein’s Spookiest Theory. Four Walls Eight Windows, New York, NY, USA, 2001. ISBN 1-56858-232-3. xviii + 284 pp. LCCN QC174.12 .A29 2001. URL http://www.loc.gov/

catdir/enhancements/fy0831/2002069338-d.html.

Adler:2003:HPM

[Adl03] Mortimer J. Adler, editor. Henri Poincaré, Max Planck, Alfred North Whitehead, Albert Einstein, Sir Arthur Eddington, Niels Bohr, G. H. Hardy, Werner Heisenberg, Erwin Schrödinger, Theodosius Dobzhansky, C. H. Waddington, volume 56 of Great books of the Western world. Encyclopaedia Britannica, Chicago, IL, USA, 2003. ISBN 0-85229-531-6. xvi + 749 pp. LCCN AC1 .G72 1990.

Aris:1983:SSC

[ADS83] Rutherford Aris, H. Ted (Howard Ted) Davis, and Roger H. Stuewer, editors. Springs of Scientific Creativity: Essays on Founders of Modern Science. University of Minnesota Press, Minneapolis, MN, USA, 1983. ISBN 0-8166-1087-8. viii + 342 pp. LCCN QC7 .S77 1983.

Adler:1990:GBW

[AFG90] Mortimer Jerome Adler, Clifton Fadiman, and Philip W. Goetz, editors. Great books of the Western world. Encyclopaedia Britannica, Chicago, IL, USA, second edition, 1990. ISBN 0-85229-531-6 (set). 749 pp. LCCN AC1 .G72 1990.

Aycock:2001:ST

[AH01] John Aycock and R. Nigel Horspool. Schrödinger’s token. Software—Practice and Experience, 31(8):803–814, July 10, 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-bin/
abstract/78502907/START;   http://www3.interscience.wiley.
com/cgi-bin/fulltext?ID=78502907&PLACEBO=IE.pdf.

Ahlgren:1996:SCH

[Ahl96] Andrew Ahlgren. Schrödinger’s cat at hand. Science, 273(5277): 860, August 16, 1996. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/273/5277/860.4.full.pdf.

Ajaib:2015:FFS

[Aja15] Muhammad Adeel Ajaib. A fundamental form of the Schrödinger equation. Foundations of Physics, 45(12):1586–1598, December 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10701-015-9944-z.

Alda:1999:CSC

[Ald99] Alan Alda. Curiosity and Schrödinger’s Cat. Scientific American, 281(4):10, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.nature.com/scientificamerican/journal/v281/n4/pdf/
scientificamerican1099-10.pdf.

Anonymous:1933:NPG

[Ano33] Anonymous. Nobel Prize goes to Bunin, Russian; short-story write and poet, exile in France, is first of his nation to get award; three physicists named; profs. Dirac and Schroedinger share for 1933 — Heisenberg gets the 1932 grant. New York Times, ?? (??):23, November 10, 1933. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/docview/100827281.

Anonymous:1934:DSP

[Ano34] Anonymous. Dr. Schroedinger at Princeton. New York Times, ??(??):??, March 26, 1934. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Anonymous:1935:MQE

[Ano35a] Anonymous. Materiewellen und Quantenmechanik: eine Einführung auf Grund der Theorien von de Broglie, Schrödinger, Heisenberg und Dirac. (German) [Matter waves and quantum mechanics: an introduction the foundations of the theories of de Broglie, Schrödinger, Heisenberg and Dirac]. Nature, 135(3403):88, January 19, 1935. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v135/n3403/pdf/135088d0.pdf.

Anonymous:1935:RSA

[Ano35b] Anonymous. Review: Structure of the atom: Science And The Human Temperament, by Erwin Schrödinger and James Murphy. British medical journal, 2(3900):626, October 5, 1935. CO-DEN BMJOAE. ISSN 0007-1447. URL http://www.jstor.org/

stable/25345578.

Anonymous:1938:PSU

[Ano38] Anonymous. Prof. E. Schrödinger and the University of Graz. Nature, 141(3577):929, May 21, 1938. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v141/n3577/pdf/141929c0.pdf.

Anonymous:1945:RBLa

[Ano45a] Anonymous. Review: What Is Life? The Physical Aspect Of The Living Cell by E. Schrödinger. British medical journal, 1(4388): 188, February 10, 1945. CODEN BMJOAE. ISSN 0007-1447.

URL http://www.jstor.org/stable/20347693.

Anonymous:1945:RBLb

[Ano45b] Anonymous. Review: What is Life? The Physical Aspect of the Living Cell. by Erwin Schrödinger. American Naturalist, 79(785): 554–555, December 1945. CODEN AMNTA4. ISSN 0003-0147 (print), 1537-5323 (electronic). URL http://www.jstor.org/

stable/2457911.

Anonymous:1949:PS

[Ano49] Anonymous. Prof. E. Schrödinger. Nature, 163(4151):794, May 21, 1949. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v163/
n4151/pdf/163794b0.pdf.

Anonymous:1951:NPH

[Ano51] Anonymous. Der Nobelpreis für Physik: Heisenberg–Schrödinger 1932 Dirac 1933/Chadwick 1935/1936 Hess–Anderson. Physikalische Blätter, 7(9):404–409, September 1951. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http:/

/onlinelibrary.wiley.com/doi/10.1002/phbl.19510070904/
abstract.

Anonymous:1959:RBH

[Ano59] Anonymous. Review: Science et Humanisme. La Physique de notre temps by E. Schrödinger. Revue de métaphysique et de morale, 64e année(1):108–110, January/March 1959. CODEN ???? ISSN 0035-1571 (print), 2102-5177 (electronic). URL http://www.jstor.

org/stable/40900294.

Anonymous:1961:OES

[Ano61] Anonymous. Obituary: Erwin Schrödinger. Physics Today, 14 (2):90, February 1961. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.

org/resource/1/phtoad/v14/i2/p90_s1.

Anonymous:1987:PESa

[Ano87a] Anonymous. Photo of Erwin Schrödinger. Foundations of Physics, 17(11):ii, November 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889802.

Anonymous:1987:PESb

[Ano87b] Anonymous. Photo of Erwin Schrödinger. Foundations of Physics, 17(12):ii, December 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889591.

Anonymous:1987:SEF

[Ano87c] Anonymous, editor. Schrödinger, Erwin: Festschrift. (German) [Erwin Schrödinger: Festival book]. ORF, Vienna, Austria, 1987. ISBN ???? ???? pp. LCCN ????

Anonymous:1988:PES

[Ano88] Anonymous. Photo of Erwin Schrödinger. Foundations of Physics, 18(3):iii–iv, March 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01891491.

Anonymous:1990:BRSc

[Ano90] Anonymous. Book reviews — Schrödinger: life and thought. Notes and Records of the Royal Society of London, 44(2):313–315, July 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:1993:EES

[Ano93] Anonymous. Eröffnung des Erwin Schrödinger International Institute of Mathematical Physics: 1993/04/20. (German) [Opening of the Erwin Schrödinger International Institute of Mathematical Physics: 4/20/1993]. Technical report, Internationales Erwin-Schrödinger-Institut für Mathematische Physik, Vienna, Austria, April 20, 1993.

Anonymous:1996:ARS

[Ano96] Anonymous. Atomic realization of Schrödinger’s cat. Science, 272 (5265):1077–0, May 24, 1996. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/272/5265/1077.1.full.

Anonymous:1997:LAS

[Ano97] Anonymous. Life after Schrödinger. Physics World, 10(4):3, April 1997. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-

7058 (electronic). URL http://physicsworldarchive.iop.org/
full/pwa-pdf/10/4/phwv10i4a1.pdf.

Anonymous:1999:CSS

[Ano99] Anonymous. 23. Closing session: 6. Sept. 1998. In Greenberger et al. [GRZ99], page ?? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN Q44.4 V662 v. 7.

Anonymous:2003:YES

[Ano03a] Anonymous, editor. 10 Years Erwin Schrödinger Institute: a series of lectures on the occasion of the 10th anniversary of the Institute’s foundation; 13–14 March, 2003. Internationales Erwin-Schrödinger-Institut für Mathematische Physik, Vienna, Austria, 2003. ISBN ???? LCCN ????

Anonymous:2003:CBS

[Ano03b] Anonymous. Controlling the behavior of a Schrödinger’s cat. Science, 299(5614):1809, March 21, 2003. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/299/5614/1809.5.full.pdf.

Anonymous:2003:SWE

[Ano03c] Anonymous, editor. Second Workshop on Entanglement and Decoherence of Complex Quantum Systems: 5–7 September 2003. Internationales Erwin-Schrödinger-Institut für Mathematische Physik, Vienna, Austria, 2003. ISBN ???? LCCN ????

Anonymous:2004:JSG

[Ano04a] Anonymous. 75 Jahre Schrödinger-Gleichung: 1926–2001; Vorträge am ESI. (German) [75 years of the Schrödinger equation: 1926–2001, lectures at the ESI]. Report, Internationales Erwin-Schrödinger-Institut für Mathematische Physik, Vienna, Austria, 2004.

Anonymous:2004:SDH

[Ano04b] Anonymous. Schrödinger and the double helix. Physics World, 17 (3):19, March 2004. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.iop.
org/full/pwa-pdf/17/3/phwv17i3a24.pdf.

Anonymous:2004:SRS

[Ano04c] Anonymous. Scientific report for // ESI, Erwin Schrödinger International Institute for Mathematical Physics. Report, Internationales Erwin-Schrödinger-Institut für Mathematische Physik, Vienna, Austria, 2004.

Anonymous:2005:CLB

[Ano05] Anonymous. Critique de livre: Il était sept Fois La révolution: Albert Einstein et les autres. La Recherche, ??(386):11–??, May 1, 2005. CODEN RCCHBV. ISSN 0029-5671 (print), 1625-9955 (elec-

tronic). URL http://www.larecherche.fr/idees/livres/il-
etait-sept-fois-revolution-albert-einstein-autres-01-05-
2005-73924.

Anonymous:2006:EEA

[Ano06] Anonymous. Einladung zur Eröffnung der Ausstellung Ludwig Boltzmann 1844–1906: am Dienstag, 2006-06-06 um 18 Uhr im Erwin Schrödinger Institut, Boltzmanngasse 9, 1090 Wien; eine Ausstellung der Österreichischen Zentralbibliothek für Physik. (German) [Invitation to the opening of the exhibition Ludwig Boltzmann 1844–1906: on Tuesday, 2006-06-06 at 18 clock at the Erwin Schrödinger Institute, Boltzmanngasse 9, 1090 Vienna, an exhibition of the Austrian Central Library for physics]. Report, Internationales Erwin-Schrödinger-Institut für Mathematische Physik, Vienna, Austria, 2006.

Anonymous:2007:TFO

[Ano07] Anonymous, editor. Thirring Fest: on the occasion of the 80th birthday of Walter Thirring; Tuesday, May 15, 2007; at the Lecture Hall of the Facuity of Physics, Wien. Internationales Erwin-Schrödinger-Institut für Mathematische Physik, Vienna, Austria, 2007. ISBN ???? LCCN ????

Anonymous:2010:EDH

[Ano10] Anonymous. Experimental demonstration of a hyper-entangled ten-qubit Schrödinger cat state. Nature Physics, 6(5):331–335, March 14, 2010. CODEN NPAHAX. ISSN 1745-2473 (print),

1745-2481 (electronic). URL http://www.nature.com/nphys/
journal/v6/n5/full/nphys1603.html.

Anonymous:2011:ESY

[Ano11a] Anonymous, editor. 1887–1961 Erwin Schrödinger: 50 years after; symposium at the ESI, 2011, January 13–15; this symposium is dedicated to Erwin Schrödinger’s legacy in quantum theory and to his views on the interpretation of quantum mechanics. Internationales Erwin-Schrödinger-Institut für Mathematische Physik, Vienna, Austria, 2011. ISBN ???? LCCN ????

Anonymous:2011:ESA

[Ano11b] Anonymous. Erwin Schrödinger 1887–1961: eine Ausstellung der Österreichischen Zentralbibliothek für Physik; Ausstellungseröffnung und Eröffnung des Schrödinger-Symposiums; 13.1.2011, 19.30 Uhr, Kleiner Festsaal der Universität Wien; einleitende Worte: Rektor Georg Winkler; Festvortrag: Jürgen Renn, MPI Berlin. (German) [Erwin Schrödinger 1887–1961: an exhibition of the Austrian Central Library for Physics, and the opening of the exhibition opening Schrödinger Symposium, 01.13.2011, 19.30 clock, Small Ballroom of the University of Vienna; introductory words: President George Winkler lecture: Jürgen Renn, Max Planck Institute in Berlin]. Report, Internationales Erwin-Schrödinger-Institut für Mathematische Physik, Vienna, Austria, January 13, 2011.

Anonymous:2011:OCS

[Ano11c] Anonymous. Opening ceremony / symposium ≪Erwin schrödinger — 50 years after≫. Report, Internationales Erwin-Schrödinger-Institut für Mathematische Physik, Vienna, Austria, 2011. Mitschnitt einer gemeinsamen Veranstaltung des Erwin Schrödinger-Instituts für Mathematische Physik (ESI) und der Fakultät für Physik der Universität Wien am Donnerstag, dem 13. Jänner 2011 im Kleinen Festsaal der Universität Wien. [Recording of a joint meeting of the Erwin Schrödinger Institute of Mathematical Physics (ESI) and the Department of Physics, University of Vienna on Thursday, the 13th January 2011 in the Small Hall of the University Vienna].

Anonymous:2012:WLF

[Ano12] Anonymous. Web life: Faraday’s Cage is where you put Schrödinger’s Cat. Physics World, 25(1):??, January 2012. CO-DEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (elec-

tronic). URL http://physicsworldarchive.iop.org/full/
pwa-pdf/25/01/phwv25i01a36.pdf.

Anonymous:2015:EDS

[Ano15] Anonymous. Einstein’s dice and Schrödinger’s cat. Scientific American, 313(3):94, September 2015. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v313/
n3/full/scientificamerican0915-94c.html;   http://www.
nature.com/scientificamerican/journal/v313/n3/pdf/scientificamerican0915-
94c.pdf.

Akulin:2003:CCP

[AP03] Vladimir M. Akulin and Pierre Pillet. 4. Control and coherence protection in quantum systems: an example of Rydberg atoms: 5. September 2003. In Anonymous [Ano03c], page ?? ISBN ????

LCCN ???? URL http://permalink.obvsg.at/AC04339690.

Atmanspacher:1999:BSM

[ARK^{+}99] Harald Atmanspacher, Helmut Rechenberg, Horst Kant, Rolf Hempelmann, Reinhold Blümel, Gerhard Börner, Houjun Mo, Philippe Blanchard, Marius Grundmann, Dietrich Stauffer, and Christian Uebing. Bücher und Software: Meÿenn: Wolfgang Pauli. Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u. a., Band IV/Teil II: 1953–1954/Mladjenovic: The Defining Years in Nuclear Physics. 1932–1960s/Frenkel: Professor Friedrich Houtermans Arbeiten, Leben, Schicksal/Kärger and Heitjans: Diffusion in Condensed Matter/Friedrich: Theoretical Atomic Physics/ Peacock: Cosmological Physics/Grosse u Martin: Particle Physics and the Schrödinger Equation/Gaponenko: Optical Properties of Semiconductor Nanocrystals/Gershenfeld: The Nature of Mathematical Modeling/Origin 6.0: Vertrieb. Physikalische Blätter, 55 (9):79–83, September 1999. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/phbl.19990550919/abstract.

Armattoe:1946:GAW

[Arm46] Raphael Ernest Grail Armattoe. The golden age of West African civilization. Londonderry Sentinel, Londonderry, Northern Ireland, 1946. 98 + 18 pp. LCCN ???? Introduction by E. Schrödinger.

Arndt:2003:SWE

[Arn03] Markus Arndt. 8. Quantum interference and decoherence of complex molecules: 6. September 2003. In Anonymous [Ano03c], page ?? ISBN ???? LCCN ????

ArroyoCamejo:2006:SQG

[Arr06] Silvia Arroyo Camejo. Skurrile Quantenwelt. (German) [Crazy Quantum World]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 3-540-29720-0. xiii + 246 pp. LCCN QC174.12 .A77 2006. EUR 29.95,

SFR 51.00. URL http://www.springer.com/physics/quantum+
physics/book/978-3-540-29720-8.

Ash:1996:FMS

[AS96] Mitchell G. Ash and Alfons Söllner, editors. Forced migration and scientific change: emigré German-speaking scientists and scholars after 1933. Publications of the German Historical Institute. German Historical Institute, Washington, DC, USA, 1996. ISBN 0-521-49741-8 (hardcover). xviii + 301 pp. LCCN E184.G3 F73 1996. URL http://www.loc.gov/

catdir/description/cam027/95024894.html;   http://www.
loc.gov/catdir/samples/cam031/95024894.html;   http://
www.loc.gov/catdir/toc/cam026/95024894.html.

Aspect:2002:BTN

[Asp02] Alain Aspect. 11. Bell’s theorem: the naive view of en experimentalist: Vortrag am 11. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Atkins:1987:RLE

[Atk87] P. W. Atkins. The rich legacy of Erwin Schrödinger. Nature, 326 (6116):913–914, April 30, 1987. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v326/n6116/pdf/326913a0.pdf.

Audretsch:1987:WRZ

[Aud87] J. Audretsch. Wellenmechanik und Raum-Zeit-Struktur: Erwin Schrödinger und die Allgemeine Relativitätstheorie. (German) [Wave mechanics and space–time structure: Erwin Schrödinger and the General Theory of Relativity]. Physikalische Blätter, 48 (8):333–337, August 1987. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/phbl.19870430808/abstract.

Axon:1989:ISC

[Axo89] T. J. Axon. Introducing Schrödinger’s cat in the laboratory. American Journal of Physics, 57(4):317–321, 1989. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://scitation.aip.org/content/aapt/journal/ajp/
57/4/10.1119/1.16128.

Buonaiuti:1954:SN

[B^{+}54] Ernesto Buonaiuti et al., editors. Spirit and nature, volume 30:1 of Papers from the Eranos yearbooks; Princeton/Bollingen paperbacks; Bollingen series. Princeton University Press, Princeton, NJ, USA, 1954. ISBN 0-691-01841-3 (paperback), 0-691-09736-4. xvi + 498 + 3 pp. LCCN BF1003 .S73 1982. URL http://www.loc.

gov/catdir/description/prin021/82245044.html;   http:/
/www.loc.gov/catdir/toc/prin031/82245044.html. Reprinted

in 1982.

Balazs:1968:BGB

[Bal68] Nándor L. Balázs. Birth and growth of Ψ: Book review: William T. Scott, Erwin Schrödinger: An Introduction to His Writings. Physics Today, 21(3):89–90, March 1968. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v21/i3/p89_s1.

Ballentine:1986:CTP

[Bal86] L. E. Ballentine. Comment on “Three perspectives on Schrödinger’s cat” [Am. J. Phys. 52, 1089 (1984)]. American Journal of Physics, 54(10):947–948, 1986. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://scitation.

aip.org/content/aapt/journal/ajp/54/10/10.1119/1.14799.

See [Loe84].

Ball:1990:SCE

[Bal90] Philip Ball. Schrödinger’s cat ensnared. Nature, 347(6291):330– 331, September 27, 1990. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v347/n6291/pdf/347330a0.pdf.

Balibar:1992:CEE

[Bal92] F. Balibar. La correspondance entre Einstein et Schrödinger. (French) [The correspondence between Einstein et Schrödinger]. In Bitbol and Darrigol [BD92], pages 307–312. ISBN 2-86332-116-1. LCCN QC16.S265 E78 1992. 263.75F.

Ball:1999:MWS

[Bal99] Philip Ball. More ways to skin Schrödinger’s cat. Nature

News, May 27, 1999. URL http://www.nature.com/news/1999/
990527/full/news990527-4.html.

Ball:2000:HDK

[Bal00] Philip Ball. How decoherence killed Schrödinger’s cat. Nature

News, January 20, 2000. URL http://www.nature.com/news/
2000/000120/full/news000120-10.html.

Ball:2003:SCC

[Bal03] Philip Ball. Schrödinger’s cat comes closer. Nature News, October 1, 2003.

Ball:2011:FSC

[Bal11] Philip Ball. Fattening up Schrödinger’s cat. Nature News, April 5,

2011. URL http://www.nature.com/news/2011/110405/full/
news.2011.210.html.

Barbour:1959:BRE

[Bar59] Ian G. Barbour. Book review: Erwin Schrödinger, Mind and Matter. American Journal of Physics, 27(7):529, October 1959. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://ajp.aapt.org/resource/1/ajpias/v27/i7/p529_
s1.

Baranzke:2000:LTE

[Bar00] Heike Baranzke, editor. Leben — Töten — Essen: anthropologische Dimensionen. (German) [Life — Death — Food: anthropological dimensions]. Verlag von S. Hirzel, Stuttgart, Germany, 2000. ISBN 3-7776-1033-X. 422 pp. LCCN BD450. URL http://www.gbv.de/

dms/faz-rez/FD120001130649220.pdf.

Barone:2003:MJE

[Bar03] Antonio Barone. 16. More on the Josephson Effect as a probe of the order parameter symmetry: 7. September 2003. In Anonymous [Ano03c], page ?? ISBN ???? LCCN ????

Basri:1956:BRS

[Bas56] Saul A. Basri. Book review: E. Schrödinger, Expanding Universes. American Journal of Physics, 24(6):472–473, September 1956. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://link.aip.org/link/?AJP/24/472/2.

Born:1969:AP

[BBSR69] Max Born, R. J. (Roger John) Blin-Stoyle, and J. M. Radcliffe. Atomic Physics. Dover books on physics and chemistry. Dover, New York, NY, USA, eighth edition, 1969. ISBN 0-486-65984-4. xiv + 495 + 11 pp. LCCN QC776 .B5713 1989. US$11.95. URL http:/

/www.loc.gov/catdir/description/dover031/89012033.html.

Bertotti:1994:ESS

[BC94] B. Bertotti and Umberto Curi, editors. Erwin Schrödinger: scienziato e filosofo, volume 3 of Percorsi della scienza, storia, testi, problemi. Il poligrafo, Padova, Italy, 1994. ISBN 88-7115-034-1. LCCN QC16.S265 E77 1994. L30000. Papers presented at the congress held in Venice, 1987, now revised. Includes a selection of Schrödinger’s correspondence.

Bitbol:1992:ESP

[BD92] Michel Bitbol and Olivier Darrigol, editors. Erwin Schrödinger: philosophy and the birth of quantum mechanics = philosophie et naissance de la mécanique quantique. Editions Frontières, Gif-sur-Yvette, France, 1992. ISBN 2-86332-116-1. xiii + 460 pp. LCCN QC16.S265 E78 1992. 263.75F.

Ben-Dov:1995:AOB

[BD95] Yoav Ben-Dov. Analyses d’ouvrages: Erwin Schrödinger : Philosophy and the birth of quantum mechanics. [Recueil de 24 communications présentées au Colloque ^{≪}Erwin Schrödinger, philosophie et naissance de la mécanique quantique ^{≫}] par Michel Bitbol; Olivier Darrigol. Revue d’Histoire des Sciences, 48(3):370– 371, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.

org/stable/23633210.

Becker:1954:RBG

[Bec54] Oskar Becker. Review: Nature and the Greeks by Erwin Schrödinger. Gnomon: Kritische Zeitschrift für die Gesamte Klassische Altertumswissenschaft, 26(8):553–554, ???? 1954. CODEN GNMNAK. ISSN 0017-1417 (print), 0945-9790 (electronic). URL

http://www.jstor.org/stable/27681031.

Belinfante:1951:BRB

[Bel51a] Frederik J. Belinfante. Book review: Space–time Structure. Erwin Schrödinger . New York: Cambridge Univ. Press, 1950. 119 pp. $2.75. Science, 113(2937):421–422, April 13, 1951. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:

//www.sciencemag.org/content/113/2937/421.full.pdf.

Belinfante:1951:RBT

[Bel51b] Frederik J. Belinfante. Review: Space–time Structure by Erwin Schrödinger. Science, 113(2937):421–422, April 13, 1951. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL

http://www.jstor.org/stable/1678694.

Bell:1964:EPR

[Bel64] John S. Bell. On the Einstein–Podolsky–Rosen paradox. Physics, 1(3):195–200, 1964. ISSN ???? URL http://en.wikipedia.org/

wiki/EPR_paradox;   http://plato.stanford.edu/entries/
bell-theorem/;   http://scienceworld.wolfram.com/physics/
Einstein-Podolsky-RosenParadox.html;   http://www.ncsu.
edu/felder-public/kenny/papers/bell.html;   http://www.
physics.princeton.edu/~mcdonald/examples/QM/bell_physics_

1_195_64.pdf. Reprinted in [Bel87b, pp. 14–21].

Bell:1987:TQJ

[Bel87a] J. S. Bell. Are there quantum jumps? In Kilmister [Kil87b], pages 41–52. ISBN 0-521-37929-6 (paperback), 0-521-34017-9. LCCN QC16.S265 S36 1987; Q143.S3. URL http://www.loc.gov/

catdir/enhancements/fy0901/86028338-d.html;   http://www.
loc.gov/catdir/enhancements/fy0901/86028338-t.html. Pa-

pers presented at a conference, March 31–April 3, 1987, at Imperial College, London.

Bell:1987:SUQ

[Bel87b] John S. Bell. Speakable and unspeakable in quantum mechanics: collected papers on quantum philosophy. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-33495-0. xii + 212 pp. LCCN QC173.97 .B45 1987.

Beller:1992:SDG

[Bel92] M. Beller. Schrödinger’s dialogue with Göttinger–Copenhagen physicists. In Bitbol and Darrigol [BD92], pages 277–306. ISBN 2-86332-116-1. LCCN QC16.S265 E78 1992. 263.75F.

Beller:1999:QDM

[Bel99] Mara Beller. Quantum dialogue: the making of a revolution. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA and London, UK, 1999. ISBN 0-226-04181-6 (hardcover), 0-226-04182-4. xv + 365 + 8 pp. LCCN QC174.13

.B45 1999. URL http://www.loc.gov/catdir/description/
uchi052/99035499.htm;   http://www.loc.gov/catdir/toc/
uchi052/99035499.htm.

Belton:2005:GSK

[Bel05] Neil Belton. A game with sharpened knives. Weidenfeld and Nicolson, London, UK, 2005. ISBN 0-297-64359-2 (cased), 0-297-84814-3 (trade paperback). 328 pp. LCCN PR6102.E48 G36 2005.

Belton:2006:GSK

[Bel06] Neil Belton. A game with sharpened knives. Phoenix fiction. Phoenix, London, UK, 2006. ISBN 0-7538-1801-9. 328 pp. LCCN PR6102.E45.

Belton:2007:SGK

[Bel07] Neil Belton. Ein Spiel mit geschliffenen Klingen: Roman. (German) [A game with sharpened blade: a novel]. S. Fischer, Frankfurt am Main, Germany, 2007. ISBN 3-10-005605-1. 456 pp. LCCN ???? EUR 21.90, SFR 38.50. Translated from the English by Bernhard Robben.

Bennewitz:1934:NBM

[Ben34] ?. Bennewitz. Neue Bücher: Die moderne Atomtheorie. W. Heisenberg; E. Schrödinger; P. A. M. Dirac. Verlag S. Hirzel, Leipzig 1934. Preis geh. RM. 2,50. Angewandte Chemie, 47(37):655, 1934. CO-DEN ANCEAD. ISSN 1521-3757.

Benjamin:1948:RBL

[Ben48] A. Cornelius Benjamin. Review: What is Life? by Erwin Schrödinger. Philosophy and Phenomenological Research, 8(3): 481–483, March 1948. CODEN PPHRAI. ISSN 0031-8205 (print),

1933-1592 (electronic). URL http://www.jstor.org/stable/
2103222.

BenMenahem:1992:SRS

[Ben92] Y. BenMenahem. Struggling with realism: Schrödinger’s case. In Bitbol and Darrigol [BD92], pages 25–40. ISBN 2-86332-116-1. LCCN QC16.S265 E78 1992. 263.75F.

Bennett:1995:QIC

[Ben95] Charles H. Bennett. Quantum information and computation. Physics Today, 48(10):24–??, October 1995. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v48/i10/p24_s1.

Bennett:1999:QIS

[Ben99] Charles H. Bennett. 22. Quantum information: 6. Sept. 1998. In Greenberger et al. [GRZ99], page ?? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN Q44.4 V662 v. 7.

Bennett:2002:SRC

[Ben02] Charles Bennett. 21. Storage and retrieval of classical information in multiparticle quantum systems: Vortrag am 13. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Bergmann:1988:QMV

[Ber88] Otto Bergmann. A quantum mechanical version of the paper by E. Schrödinger “Über die Umkehrung der Naturgesetze”. Foundations of Physics, 18(3):373–378, March 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/article/10.1007/BF01891497.

Bertlmann:2002:MMC

[Ber02] Reinhold Bertlmann. 12. Magic moments: a collaboration with John Bell: Vortrag am 11. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Bose:1986:SQO

[BES86] S. N. Bose, A. Einstein, and E. Schrödinger. La statistica quantistica e le onde di materia. (Italian) [Quantum statistics and matter waves], volume 3 of Saggi di Scienze e Filosofia Naturale [Essays on Natural Philosophy and Natural Science]. Bibliopolis, Naples, Italy, 1986. ISBN 88-7088-128-8. 99 pp. Translated from the German by Paolo Bernardini.

Born:1928:PMG

[BEvL^{+}28] M. Born, A. Einstein, M. v. Laue, E. Schrödinger, and A. Sommerfeld. Planck-Medaille. (German) [Planck Medal]. Naturwissenschaften, 16(20):368, May 1928. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://www.

springerlink.com/content/gg356437362203j2/fulltext.pdf.

Born:1961:ES

[BG61] Max Born and G. Glaser. Erwin Schrödinger/Wilhelm Keil. Physikalische Blätter, 17(2):85–87, February 1961. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

URL http://onlinelibrary.wiley.com/doi/10.1002/phbl.
19610170205/abstract.

Bitbol:1992:EFT

[Bit92] M. Bitbol. Esquisses, forme et totalité (Schrödinger et le concept d’object). (French) [sketches, form, and totality (Schrödinger and the object concept)]. In Bitbol and Darrigol [BD92], pages 41–80. ISBN 2-86332-116-1. LCCN QC16.S265 E78 1992. 263.75F.

Bitbol:1993:AOB

[Bit93] Michel Bitbol. Analyses d’ouvrages: Schrödinger, life and thought par Walter Moore. Revue d’Histoire des Sciences, 46(2–3):309– 310, avril–septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633836.

Bitbol:1996:CBS

[Bit96a] Michel Bitbol. The controversy between Schrödinger and the Göttingen–Copenhagen physicists in the 1950’s. In Schrödinger’s Philosophy of Quantum Mechanics [Bit96b], chapter 1, pages 1– 33. ISBN 0-7923-4266-6, 94-009-1772-4. ISSN 0068-0346. LCCN Q174 .B67 vol. 188; QC173.98. URL https://link.springer.

com/chapter/10.1007/978-94-009-1772-9_1.

Bitbol:1996:SPQ

[Bit96b] Michel Bitbol. Schrödinger’s Philosophy of Quantum Mechanics, volume 188 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 0-7923-4266-6, 94-009-1772-4. ISSN 0068-0346. xi + 285 pp. LCCN Q174

.B67 vol. 188; QC173.98. URL http://www.loc.gov/catdir/
enhancements/fy0813/96042123-d.html;   http://www.loc.
gov/catdir/enhancements/fy0813/96042123-t.html;   https:
//link.springer.com/book/10.1007/978-94-009-1772-9.

Bitbol:1996:STP

[Bit96c] Michel Bitbol. Schrödinger’s theoretical project. In Schrödinger’s Philosophy of Quantum Mechanics [Bit96b], chapter 2, pages 34– 78. ISBN 0-7923-4266-6, 94-009-1772-4. ISSN 0068-0346. LCCN Q174 .B67 vol. 188; QC173.98. URL https://link.springer.

com/chapter/10.1007/978-94-009-1772-9_2.

Blackmore:2001:EMV

[BIT01] J. Blackmore, R. Itagaki, and S. Tanaka, editors. Ernst Mach’s Vienna 1895–1930: Or Phenomenalism as Philosophy of Science, volume 218 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 90-481-5752-8, 94-015-9690-5 (ebook). ISSN 0068-0346. LCCN ???? URL https://link.

springer.com/book/10.1007/978-94-015-9690-9.

Baker-Jarvis:2003:MBA

[BJK03] James Baker-Jarvis and Pavel Kabos. Modified de Broglie approach applied to the Schrödinger and Klein–Gordon equations. Physical Review A (Atomic, Molecular, and Optical Physics), 68 (??):042110, ???? 2003. CODEN PLRAAN.

Blatt:1999:QOS

[Bla99] Rainer Blatt. 12. Quantum optics with single atoms: 5. Sept. 1998. In Greenberger et al. [GRZ99], page ?? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN Q44.4 V662 v. 7.

Blatter:2000:SCN

[Bla00] Gianni Blatter. Schrödinger’s cat is now fat. Nature, 406(6791): 25–26, July 6, 2000. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v406/n6791/full/406025a0.html.

Blaauboer:2003:QDA

[Bla03] Miriam Blaauboer. 15. QUACS at Delft: from August 2002 to today and beyond: 7. September 2003. In Anonymous [Ano03c], page ?? ISBN ???? LCCN ????

Blinder:2001:KGS

[Bli01] S. M. Blinder. A kinder, gentler Schrödinger’s cat experiment. American Journal of Physics, 69(6):633–634, 2001. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://scitation.aip.org/content/aapt/journal/ajp/
69/6/10.1119/1.1287720.

Bloch:1976:HED

[Blo76a] Felix Bloch. Heisenberg and the early days of quantum mechanics. Physics Today, 29(12):23–27, December 1, 1976. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://physicstoday.scitation.org/doi/abs/10.1063/
1.3024633.

Bloch:1976:RHE

[Blo76b] Felix Bloch. Reminiscences of Heisenberg and the early days of quantum mechanics. Physics Today, 29(12):23–27, December 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/phtoad/

v29/i12/p23/s1. Reprinted in [WP85, pages 319–323].

Bunge:2001:SRS

[BM01] Mario Augusto Bunge and Martin Mahner, editors. Scientific realism: selected essays of Mario Bunge. Prometheus Books, Amherst, NY, USA, 2001. ISBN 1-57392-892-5. 438 pp. LCCN Q175.32

.R42 B87 2001. URL http://www.loc.gov/catdir/toc/fy032/
2001019544.html.

Born:1963:AAGb

[Bor63a] Max Born. Ausgewählte Abhandlungen. (German) [Selected works], volume 2. Vandenhoeck & Ruprecht, Göttingen, West Germany, 1963. viii + 706 pp. LCCN ???? Mit einem Verzeichnis der wissenschaftlichen Schriften (German) [With a table of contents of scientific writings].

Born:1963:ES

[Bor63b] Max Born. Erwin Schrödinger. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63a], chapter 93, pages 691–694. LCCN ???? Reprinted from Physik. Bl. 17, 85–87 (1961).

Bouwmeester:1999:QTE

[Bou99] Dik Bouwmeester. 20. Quantum Teleportation and Entanglement Swapping: 5. Sept. 1998. In Greenberger et al. [GRZ99], page ?? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN Q44.4 V662 v. 7.

Boudinot:2012:BA

[Bou12] Ryan Boudinot. Blueprints of the Afterlife. Black Cat, New York, NY, USA, 2012. ISBN 0-8021-7091-9 (paperback). 430 pp. LCCN PS3602.O888 B58 2012.

Barbour:1995:MPN

[BP95] Julian B. Barbour and Herbert Pfister, editors. Mach’s principle: from Newton’s bucket to quantum gravity, volume 6 of Einstein studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1995. ISBN 0-8176-3823-7, 3-7643-3823-7 (paperback). vii + 536 pp. LCCN QC137 .M33 1995. URL http://

www.gbv.de/dms/bowker/toc/9780817638238.pdf;   http://
www.zentralblattmath.org/zmath/en/search/?an=0846.01008.

Ballentine:1971:QMD

[BPW^{+}71] Leslie E. Ballentine, Philip Pearle, Evan Harris Walker, Mendel Sachs, Toyoki Koga, Joseph Gerver, and Bryce DeWitt. Quantum mechanics debate. Physics Today, 24(4):36–44, April 1971. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://www.physicstoday.org/resource/1/phtoad/v24/

i4/p36_s1. See original article [DeW70].

Braunizer:1992:ESS

[Bra92] R. Braunizer. Erwin Schrödinger: some elements of biography. In Bitbol and Darrigol [BD92], pages 3–8. ISBN 2-86332-116-1. LCCN QC16.S265 E78 1992. 263.75F.

Brandt:2009:HCD

[Bra09] Siegmund Brandt. The harvest of a century: discoveries of modern physics in 100 episodes. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-954469-7 (hardcover). xiv + 500 pp. LCCN QC7 .B64 2009.

Brenner:1984:SWN

[Bre84a] Sydney Brenner. Schrödinger wrong, von Neumann right. Web

interview (2m5s), 1984. URL https://www.webofstories.com/
play/13259?o=MS;   https://www.webofstories.com/play/sydney.
brenner/46.

Brenner:1984:SBC

[Bre84b] Sydney Brenner. Schrödinger’s belief in calculating an organism from chromosomes. Web interview (5m0s), 1984. URL https:

//www.webofstories.com/play/13260?o=MS;   https://www.
webofstories.com/play/sydney.brenner/47.

Brown:1936:RBS

[Bro36] G. Burniston Brown. Review: Where Is Science Going? by Max Planck, James Murphy; Atomic Theory and the Description of Nature by Niels Bohr; Science and the Human Temperament by Erwin Schrödinger, James Murphy. Philosophy, 11(43):366–367, July 1936. ISSN 0031-8191 (print), 1469-817X (electronic). URL http:

//www.jstor.org/stable/3746205.

Brush:1980:CCP

[Bru80] Stephen G. Brush. The chimerical cat: Philosophy of quantum mechanics in historical perspective. Social Studies of Science, 10 (4):393–447, November 1, 1980. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL http://www.jstor.

org/stable/284918;   https://journals.sagepub.com/doi/
pdf/10.1177/030631278001000401.

Bruce:2004:SRM

[Bru04] Colin Bruce. Schrödinger’s Rabbits: The Many Worlds of Quantum. National Academies Press, Washington, DC, USA, 2004. ISBN 0-309-54658-3. ix + 272 pp. URL http://www.nap.edu/

openbook.php?record_id=11002.

Brumfiel:2009:PPS

[Bru09] Geoff Brumfiel. Physicists propose ‘Schrödinger’s virus’ experiment. Nature News, September 10, 2009. URL http://www.

nature.com/news/2009/090910/full/news.2009.903.html.

Badurek:1988:MWI

[BRZ88] Gerald Badurek, Helmut Rauch, and Anton Zeilinger, editors. Matter wave interferometry: proceedings of the International Workshop on Matter Wave Interferometry in the Light of Schrödinger’s Wave Mechanics held in Vienna, Austria, 14–16 September 1987; on the occasion of the 100th anniversary of E. Schrödinger’s birth, Physica B + C. North-Holland, Amsterdam, The Netherlands, 1988. CODEN PHBCDQ. ISBN ???? ISSN 0378-4363 (print), 1873-2127 (electronic). LCCN ???? Published in Physica, 1511/2, 1988.

Born:1935:AFC

[BS35] Max Born and Erwin Schrödinger. The absolute field constant in the new field theory. Nature, 135(3409):342, March 2, 1935. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v135/
n3409/pdf/135342a0.pdf.

Broderick:1940:BAP

[BS41] T. S. Broderick and Erwin Schrödinger. Boolean algebra and probability theory. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 46(??):103–112, ???? 1940– 1941. CODEN PRIAAK. ISSN 0035-8975. URL http://www.

jstor.org/stable/20490751.

Bass:1955:MPM

[BS55] L. Bass and E. Schrödinger. Must the photon mass be zero? Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 232(1188):1–6, October 11, 1955. CO-DEN PRLAAZ. ISSN 0080-4630. URL http://www.jstor.org/

stable/99679.

Bass:1956:MPM

[BS56] L. Bass and Erwin Schrödinger. Must the photon mass be zero? [summary and discussion]. Il Nuovo Cimento, Supplemento, 4(10): 825–826, ???? 1956.

Baird:2006:PCS

[BSM06] Davis Baird, Eric Scerri, and Lee McIntyre. Philosophy of Chemistry: Synthesis of a New Discipline, volume 242 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346. vii + 362 pp. LCCN Q174 .B67 vol. 242 QD6. URL https://link.

springer.com/book/10.1007/1-4020-3261-7.

Buckingham:1987:QC

[Buc87] A. D. Buckingham. Quantum chemistry. In Kilmister [Kil87b], pages 112–118. ISBN 0-521-37929-6 (paperback), 0-521-34017-9. LCCN QC16.S265 S36 1987; Q143.S3. URL http://www.loc.

gov/catdir/enhancements/fy0901/86028338-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0901/86028338-t.html.

Papers presented at a conference, March 31–April 3, 1987, at Imperial College, London.

Bunge:1953:RRR

[Bun53] Mario Bunge. Review: Relativity — A Richer Truth, by Philipp Frank. Boston: The Beacon Press, 1950. $2.00. Pp. xvi + 142; Science and Humanism. Physics in Our Time, by Erwin Schrödinger, London and New York: Cambridge University Press, 1951. $1.75. Pp. 68. Science & Society, 17(3):266–269, Summer 1953. CO-DEN ???? ISSN 0036-8237 (print), 1943-2801 (electronic). URL

http://www.jstor.org/stable/40400200.

Bunge:2001:SCD

[Bun01] Mario Bunge. Schrödinger’s cat is dead (1999). In Mahner [Mah01], chapter 18, pages 256–262. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Burns:1955:RBG

[Bur55] Michael Burns. Review: Nature and the Greeks by Erwin Schrödinger. Science & Society, 19(2):186–189, ???? 1955. CO-DEN ???? ISSN 0036-8237 (print), 1943-2801 (electronic). URL

http://www.jstor.org/stable/40400338.

Burhoe:1965:BBV

[Bur65] Ralph Wendell Burhoe. Books: My View of the World, by Erwin Schrödinger; The Relevance of Science, Creation and Cosmogony, by C. F. von Weizsäcker. Bulletin of the Atomic Scientists, 21 (7):33–35, September 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bussey:2013:QDW

[Bus13] Peter J. Bussey. The Quantum Divide: Why Schrödinger’s Cat is either Dead or Alive, by Christopher C. Gerry and Kimberley M. Bruno, Scope: general interest. Level: non-specialists, teacher. Contemporary Physics, 54(2):138–139, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2020:BRS

[Bus20] Peter J. Bussey. Book review: Schrödinger’s mechanics: interpretation, by D. B. Cook, Singapore, World Scientific, 2019, 196 pp., £69 (hardback), ISBN: 978-1-78634-490-8. Scope: monograph; Level: specialist. Contemporary Physics, 61(1):64, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bacciagaluppi:2009:QTC

[BV09] Guido Bacciagaluppi and Antony Valentini, editors. Quantum theory at the crossroads: reconsidering the 1927 Solvay conference. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-81421-9 (hardcover). LCCN QC173.96 .B33 2009. URL http://

www.loc.gov/catdir/toc/ecip0817/2008019585.html.

Berliner:1929:LBF

[BWE^{+}29] A. Berliner, W. H. Westphal, J. Errera, Fr. Löwe, P. W. Danckwortt, et al. Literatur: Bericht über die Fortschritte der analytischen Chemie. (German) [Literature: Report on progress in analytical chemistry]. Zeitschrift für Analytische Chemie, 76(7–8):281–284, ???? 1929. CODEN ZANCA8.

ISSN 0372-7920. URL http://www.springerlink.com/content/
a73444t0177j5504/fulltext.pdf.

Bertlmann:2002:QUB

[BZ02] Reinhold A. Bertlmann and Anton Zeilinger, editors. Quantum [Un]speakables: From Bell to Quantum Information: Conference in Commemoration of the Physicist John S. Bell, 10–14 November 2000, Vienna. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-42756-2. LCCN ????

C:1935:RBH

[C.35] H. T. C. Review: Science and the Human Temperament by Erwin Schrödinger, James Murphy, W. H. Johnston. The Journal of Philosophy, 32(16):442–443, August 1, 1935. CODEN ???? ISSN 0022-362X (print), 1939-8549 (electronic). URL http://www.jstor.

org/stable/2016569.

Campa:1988:EEE

[Cam88] Riccardo Campa. El estupor de Epicuro: ensayo sobre Erwin Schrödinger. (Spanish) [The stupor of Epicurus: test of Erwin Schrödinger], volume 7 of Alianza ensayo. Alianza, Buenos Aires, Argentina, 1988. ISBN 950-40-0036-3. 205 pp. LCCN QC16.S265 C35 1988.

Carlson:1945:RBL

[Car45] A. J. Carlson. Review: What Is Life? by Erwin Schrödinger. The Scientific Monthly, 61(6):495–496, December 1945. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL

http://www.jstor.org/stable/18388.

Cartier:1993:FDC

[Car93] Pierre Cartier. From Feynman diagrams to the classification of knots: a tale about the use of physics in pure mathematics. Report, Internationales Erwin-Schrödinger-Institut für Mathematische Physik, Vienna, Austria, 1993.

Cartlidge:2013:HFS

[Car13] Edwin Cartlidge. How fat is Schrödinger’s cat? Physics World, 26(4):??, April 25, 2013. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL https://physicsworld.

com/a/how-fat-is-schrodingers-cat/.

Chatzidimitriou-Dreismann:2003:EDP

[CD03] C. Aris Chatzidimitriou-Dreismann. 11. Entanglement and decoherence of protons in condensed matter — sub-femtosecond neutron scattering result and theoretical considerations: 6. September 2003. In Anonymous [Ano03c], page ?? ISBN ???? LCCN ????

Ceccarelli:2001:SSR

[Cec01] Leah Ceccarelli. Shaping science with rhetoric: the cases of Dobzhansky, Schrödinger, and Wilson. University of Chicago Press, Chicago, IL, USA and London, UK, 2001. ISBN 0-226-09906-7 (cloth), 0-226-09907-5 (paperback). xi + 204 pp. LCCN

QH303.6 .C433 2001. URL http://www.loc.gov/catdir/bios/
uchi051/00012179.html;   http://www.loc.gov/catdir/description/
uchi051/00012179.html;   http://www.loc.gov/catdir/toc/
uchi051/00012179.html.

Cerjan:1993:NGM

[Cer93] C. Cerjan, editor. Numerical grid methods and their application to Schrödinger’s equation, volume 412 of NATO ASI series. Series C, Mathematical and physical sciences. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1993. ISBN 0-7923-2423-4. LCCN QC174.26.W28 N38 1993. Published in cooperation with NATO Scientific Affairs Division. Proceedings of the NATO Advances Research Workshop on Grid Methods in Atomic and Molecular Quantum Calculations Corte, Corsica, France, September 27–October 3, 1992.

Crease:2014:QMH

[CG14] Robert P. Crease and Alfred S. Goldhaber. The quantum moment: how Planck, Bohr, Einstein, and Heisenberg taught us to love uncertainty. W. W. Norton & Co., New York, NY, USA, 2014. ISBN 0-393-06792-0 (hardcover). vii + 332 pp. LCCN QC174.123 .C74 2014.

Chakraborty:2013:IQM

[Cha13] Aniruddha Chakraborty. Introduction to Quantum Mechanics: Schrödinger Equation and Path Integral, 2nd edn., by Harald J. W. Müller-Kirsten, Scope: review. Level: postgraduate. Contemporary Physics, 54(1):79, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chevalley:1992:CEB

[Che92] C. Chevalley. Le conflit de 1926 entre Bohr et Schrödinger: un exemple de sous-détermination des théories. (French) [The conflict of 1926 between Bohr and Schrödinger: an example of underdetermination of theories]. In Bitbol and Darrigol [BD92], pages 81–94. ISBN 2-86332-116-1. LCCN QC16.S265 E78 1992. 263.75F.

Chevalley:1999:DBB

[Che99] Catherine Chevalley. 5. On the difference between Bohr’s epistemology and the so-called Copenhagen interpretation of quantum theory: 4. Sept. 1998. In Greenberger et al. [GRZ99], page ?? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN Q44.4 V662 v. 7.

Cho:2000:PUS

[Cho00] Adrian Cho. Physicists unveil Schrödinger’s SQUID. Science, 287 (5462):2395, March 31, 2000. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/287/5462/2395.1.

Cohen:1983:PPP

[CL83] R. S. Cohen and L. Laudan, editors. Physics, Philosophy and Psychoanalysis: Essays in Honour of Adolf Grünbaum, volume 76 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. ISBN 94-009-7055-2, 94-009-7057-9 (e-book). ISSN 0068-0346. xviii + 339 pp. LCCN ???? URL https://link.springer.

com/book/10.1007/978-94-009-7055-7.

Clauser:2002:EHB

[Cla02] John Clauser. 3. Early history of Bell’s Theorem: Vortrag am 10. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Clary:2023:FMR

[Cla23] David C. Clary. Foreign membership of the Royal Society: Schrödinger and Heisenberg? Notes and Records of the Royal Society of London, 77(3):513–536, September ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Cohen:1992:SNS

[Coh92] R. S. Cohen. Some notes on Schrödinger and mysticism. In Bitbol and Darrigol [BD92], pages 95–120. ISBN 2-86332-116-1. LCCN QC16.S265 E78 1992. 263.75F.

Cohen:2005:SM

[Coh05] Jack Cohen. Schrödinger’s mousetrap. Nature, 434(7030):143, March 9, 2005. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v434/n7030/full/434143a.html.

Collins:1998:QTC

[Col98] Graham P. Collins. Quantum teleportation channels opened in Rome and Innsbruck. Physics Today, 51(2):18–??, February 1998. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://www.physicstoday.org/resource/1/
phtoad/v51/i2/p18_s1.

Collins:2000:SG

[Col00a] Graham P. Collins. Schrödinger’s games. Scientific American, 282 (1):28–29, January 2000. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v282/n1/pdf/scientificamerican0100-
28b.pdf.

Collins:2000:SS

[Col00b] Graham P. Collins. Schrödinger’s squid. Scientific American, 283 (4):23–24, October 2000. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v283/n4/pdf/scientificamerican1000-
23.pdf.

Cook:2018:SMIi

[Coo18] David Branston Cook. Schrødinger’s mechanics: an interpretation. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2018. ISBN 1-78634-491-2, 1-78634-492-0. xiii + 181 pp. LCCN QC174.26.W28. URL http://www.

worldscientific.com/worldscibooks/10.1142/q0143\#t=toc.

Corson:1952:BRE

[Cor52] E. M. Corson. Book review: Erwin Schrödinger, Space–time Structure. Physics Today, 5(6):19–20, June 1952. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

link.aip.org/link/?PTO/5/19/2.

Coulson:1962:IFM

[Cou62] C. A. Coulson. Interatomic forces: Maxwell to Schrödinger. Nature, 195(4843):744–749, August 25, 1962. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v195/n4843/pdf/195744a0.pdf.

Cramer:1986:TIQ

[Cra86] John G. Cramer. The transactional interpretation of quantum mechanics. Reviews of Modern Physics, 58(3):647–687, July 1986. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.58.647;   http://rmp.aps.org/abstract/
RMP/v58/i3/p647_1;   https://en.wikipedia.org/wiki/Transactional_
interpretation.

Cropper:2001:GPL

[Cro01a] William H. Cropper. Great physicists: the life and times of leading physicists from Galileo to Hawking. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-513748-5. xii + 500 pp. LCCN QC15 .C76 2001. URL http:

//www.loc.gov/catdir/enhancements/fy0637/2001021611-d.
html;   http://www.loc.gov/catdir/enhancements/fy0724/
2001021611-b.html.

Cropper:2001:WME

[Cro01b] William H. Cropper. Wave mechanics: Erwin Schrödinger and Louis de Broglie. In Great physicists: the life and times of leading physicists from Galileo to Hawking [Cro01a], pages 275–

292. ISBN 0-19-513748-5. LCCN QC15 .C76 2001. URL http:

//www.loc.gov/catdir/enhancements/fy0637/2001021611-d.
html;   http://www.loc.gov/catdir/enhancements/fy0724/
2001021611-b.html.

Csizmadia:1996:SCH

[CS96] András Málnási Csizmadia and Achilles J. Sanchez. Schrödinger’s cat at hand. Science, 273(5277):859–860, August 16, 1996. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.sciencemag.org/content/273/5277/857.6;
 http://www.sciencemag.org/content/273/5277/859.2.full.
pdf.

Cushing:1996:BRE

[Cus96] James T. Cushing. Book review: Erwin Schrödinger and Michel Bitbol: The Interpretation of Quantum Mechanics: Dublin Seminars (1949–1955) and Other Unpublished Essays. Isis, 87(3):570– 571, September 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
236045.

Cohen:1985:PTF

[CW85] Robert S. Cohen and Marx W. Wartofsky, editors. A Portrait of Twenty-five Years: Boston Colloquium for the Philosophy of Science 1960–1985, volume 89 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1985. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346. LCCN ???? URL https:

//link.springer.com/book/10.1007/978-94-009-5345-1.

DAgostino:1992:CCP

[D’A92a] Salvo D’Agostino. Continuity and completeness in physical theory: Schrödinger’s return to the wave interpretation of quantum mechanics. In Bitbol and Darrigol [BD92], pages 339–362. ISBN 2-86332-116-1. LCCN QC16.S265 E78 1992. 263.75F.

DAgostino:1992:SCC

[D’A92b] Salvo D’Agostino. Schrödinger on continuity and completeness in physical theory. Physis: Rivista Internazionale di Storia della Scienza. Nuova Serie, 29(2):539–561, 1992. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

DAgostino:2000:HIT

[D’A00] Salvo D’Agostino. A History of the Ideas of Theoretical Physics: Essays on the Nineteenth and Twentieth Century Physics, volume 213 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 0-7923-6094-X, 94-010-9034-3. ISSN 0068-0346. xviii + 381 pp. LCCN Q174 .B67 no. 213. URL https:

//link.springer.com/book/10.1007/978-94-010-9034-6.

DAgostino:2001:WMS

[D’A01a] Salvo D’Agostino. From the 1926 wave mechanics to a secondquantisation theory: Schrödinger’s new interpretation of wave mechanics and microphysics in the 1950’s. In A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics [D’A01b], chapter 13, pages 273–302. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346. LCCN

QC5.53; QC6. URL https://link.springer.com/chapter/10.
1007/978-94-010-9034-6_13.

DAgostino:2001:HIT

[D’A01b] Salvo D’Agostino. A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics, volume 213 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346. xviii + 381 pp. LCCN QC5.53; QC6.

DAgostino:2009:GSL

[D’A09] Salvo D’Agostino. Il ‘gatto’ di Schroedinger in una lettura del testo originale del celebre saggio. (Italian) [Schrödinger’s cat in a reading of the original text of the famous essay]. Physis: Rivista Internazionale di Storia della Scienza. Nuova Serie, 46(1–2):235– 252, 2009. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Darrigol:1992:SSP

[Dar92] O. Darrigol. Schrödinger’s statistical physics and some related themes. In Bitbol and Darrigol [BD92], pages 237–276. ISBN 2-86332-116-1. LCCN QC16.S265 E78 1992. 263.75F.

Darrigol:2011:FRW

[Dar11] Olivier Darrigol. A few reasons why Louis de Broglie discovered [de] Broglie’s waves and yet did not discover Schrödinger’s equation. Mitschnitt einer gemeinsamen Veranstaltung des Erwin Schrödinger-Instituts für Mathematische Physik (ESI) und der Fakultät für Physik der Universität Wien am Donnerstag, dem 13. Jänner 2011 im Kleinen Festsaal der Universität Wien [Recording of a joint event at the Erwin Schrödinger Institute for Mathematical Physics (ESI) and the Department of Physics, University of Vienna on Thursday, the 13th January 2011 in the Small Hall of the University of Vienna]., 2011.

Darrigol:2013:FRW

[Dar13] Olivier Darrigol. A few reasons why Louis de Broglie discovered Broglie’s waves and yet did not discover Schrödinger’s equation. In Reiter and Yngvason [RY13], pages 165–174. ISBN 3-03719-121-X (paperback), 3-03719-621-1 (e-book). LCCN QC16.S265 I58 2011.

Datzeff:1938:PBP

[Dat38] Assène Datzeff. Sur le problème des barrières de potentiel et la résolution de l’équation de Schrödinger. (French) [On the problem of potential barriers and the solution of Schrödinger’s equation ]. Masson et cie, Paris, France, 1938. 2 + 91 + 1 pp. LCCN QC174.2 .D38.

Datta:1987:SVM

[Dat87] Santanu Datta. Schrödinger v. Marx. Nature, 329(6140):578, October 21, 1987. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v329/n6140/pdf/329578d0.pdf.

Davies:1957:RBU

[Dav57] J. T. Davies. Review: Expanding Universes by Erwin Schrödinger. British Journal for the Philosophy of Science, 8(29):78–79, May 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://www.jstor.org/stable/685387.

deBroglie:1924:TTL

[dB24] Louis de Broglie. A tentative theory of light quanta. Philosophical Magazine (6), 47(??):446–458, ???? 1924. CODEN PHMAA4. ISSN 1941-5982 (print), 1941-5990 (electronic).

deBeauregard:1965:PCE

[dB65] O. Costa de Beauregard. Le · paradoxe · des correlations d’Einstein: et de Schrödinger et l’épaisseur temporelle de la transition quantique. (French) [The · paradox · of the correlations of Einstein: and Schrödinger and the temporal thickness of the quantum transition]. Dialectica: International Review of Philosophy of Knowledge, 19(3–4):280–289, December 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Daniel-Cressey:2009:SGF

[DC09] Daniel-Cressey. Schrödinger’s God. Nature News, March

16, 2009. URL http://blogs.nature.com/news/2009/03/
schroedingers_god.html.

Dou:2013:GHS

[DD13] Yan-Ni Dou and Hong-Ke Du. Generalizations of the Heisenberg and Schrödinger uncertainty relations. Journal of Mathematical Physics, 54(10):103508, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

dEspagnat:2002:MIJ

[d’E02] Bernard d’Espagnat. 6. My interaction with John Bell: Vortrag am 10. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Anton:1989:IQP

[Deg89] Peter Anton Degen. Interpretations of quantum physics, the mystical and the paranormal: Einstein, Schrödinger, Bohr, Pauli and Jordan. Ph.D. dissertation, Drew University, Madison, NJ, USA, December 1989. 319 pp.

Dehlinger:1936:BMQ

[Deh36] U. Dehlinger. Buchbesprechung: Materiewellen und Quantenmechanik. Eine Einführung auf Grund der Theorien von De Broglie, Schrödinger, Heisenberg und Dirac. 4. und 5. Aufl. Von Arthur Haas. Leipzig 1934, Akademische Verlagsgesellschaft. 299 S., 7 Abb. Preis geb. RM 7.80. Zeitschrift für Elektrochemie und Angewandte Physikalische Chemie, 42(1):43, 1936. CODEN ZEA-PAA. ISSN 0005-9021.

Delbruck:1945:RBL

[Del45] Max Delbrück. Review: What is Life? The Physical Aspect of the Living Cell and what is truth? Quarterly review of biology, 20(4):370–372, December 1945. CODEN QRBIAK. ISSN 0033-5770 (print), 1539-7718 (electronic). URL http://www.jstor.

org/stable/2809111.

DeWitt:1970:QMR

[DeW70] Bryce S. DeWitt. Quantum mechanics and reality. Physics Today, 23(9):30–35, September 1970. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.

org/resource/1/phtoad/v23/i9/p30_s1.

Daintith:1999:DS

[DG99] John Daintith and Derek Gjertsen, editors. A Dictionary of Scientists. Oxford paperback reference. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-585-11047-6 (e-book), 0-19-280086-8 (paperback). 586 pp.

LCCN Q141 .D52 1994. URL http://www.loc.gov/catdir/
enhancements/fy0637/99488304-d.html;   http://www.loc.
gov/catdir/enhancements/fy0637/99488304-t.html;   http:
//www.netLibrary.com/urlapi.asp?action=summary&v=1&bookid=
12306.

deGosson:2014:BJQ

[dG14] Maurice A. de Gosson. Born–Jordan quantization and the equivalence of the Schrödinger and Heisenberg pictures. Foundations of Physics, 44(10):1096–1106, October 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10701-014-9831-
z.pdf.

Dick:1985:ADL

[Dic85] Auguste Dick. Auguste Dick liest aus Briefen von Erwin Schrödinger. (German) [Auguste Dick reads from letters written by Erwin Schrödinger]. Zentralbibliothek für Physik in Wien, Vienna, Austria, 1985.

Dirac:1961:OPE

[Dir61] P. A. M. Dirac. Obituary: Prof. Erwin Schrödinger, For.Mem.R.S. Nature, 189(4762):355–356, February 4, 1961. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v189/n4762/pdf/189355a0.pdf.

Dirac:1964:ESH

[Dir64] P. A. M. Dirac. Equivalence of the Schrödinger and Heisenberg pictures. Nature, 204(4960):772, November 21, 1964. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v204/
n4960/pdf/204772a0.pdf.

DaCosta:1992:SP

[DKF92] N. C. A. Da Costa, D. Krause, and S. French. The Schrödinger problem. In Bitbol and Darrigol [BD92], pages 445–?? ISBN 2-86332-116-1. LCCN QC16.S265 E78 1992. 263.75F.

DeMaria:1982:SDU

[DL82a] M. De Maria and F. La Teana. Schrödinger’s and Dirac’s unorthodoxy in quantum mechanics. Fundamenta Scientiae, 3(??): 129–148, ???? 1982.

DeMaria:1982:PLS

[DL82b] Michelangelo De Maria and Francesco La Teana. I primi lavori di E. Schrödinger sulla meccanica ondulatoria e la nascita delle polemiche con la scuola di Goettingen–Copenhagen sull’interpretazione della meccanica quantistica. (Italian) [The first works of E. Schrödinger on wave mechanics and the birth of the controversy with the Göttingen–Copenhagen school’s interpretation of quantum mechanics]. Physis: Rivista Internazionale di Storia della Scienza, 24(1):33–54, 1982. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Dorling:1987:SOI

[Dor87]  Jon Dorling.  Schrödinger’s original interpretation of the

Schrödinger equation: a rescue attempt. In Kilmister [Kil87b], pages 16–40. ISBN 0-521-37929-6 (paperback), 0-521-34017-9. LCCN QC16.S265 S36 1987; Q143.S3. URL http://www.loc.

gov/catdir/enhancements/fy0901/86028338-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0901/86028338-t.html.

Papers presented at a conference, March 31–April 3, 1987, at Imperial College, London.

deRegt:1997:ESA

[dR97] Henk W. de Regt. Erwin Schrödinger, Anschaulichkeit, and quantum theory. Studies in History and Philosophy of Modern Physics, 28(4):461–481, December 1997. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219897000178.

deRegt:2001:ES

[dR01] Henk W. de Regt. Erwin Schrödinger. In Blackmore et al. [BIT01], pages 85–104. ISBN 90-481-5752-8, 94-015-9690-5 (e-book). ISSN

0068-0346. LCCN ???? URL https://link.springer.com/
chapter/10.1007/978-94-015-9690-9_4.

deRafael:2002:VMD

[dR02] Eduardo de Rafael. 16. from vector meson dominance to large-Nc QCD: Vortrag am 11. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Dronsfield:2020:BRT

[Dro20] Alan Dronsfield. Book review: Traveling with the Atom: a Scientific Guide to Europe and Beyond. Web site, March 20, 2020.

URL https://www.chemistryworld.com/review/traveling-
with-the-atom-a-scientific-guide-to-europe-and-beyond/
4011154.article.

Dirac:1983:ANP

[DS83] Paul Dirac and Erwin Schrödinger. 50th anniversary of the Nobel Prize for Schrödinger and Dirac. Fiz.-Mat. Spis. B′′lgar. Akad. Nauk., 25(58)(4):277–297, 1983. CODEN FMBMAC. ISSN 0015-3265. Translated from the French and German by D. Vachov.

Dale-Trotter:2018:BRQ

[DT18] Iain Dale-Trotter. Book review: The quantum heretics: What is Real: the Unfinished Quest for the Meaning of Quantum Physics by Adam Becker. Physics World, 31(12):??, December 2018. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL

https://physicsworld.com/a/the-quantum-heretics/.

Eberly:2009:SCH

[Ebe09] Professor J. H. Eberly. Schrödinger’s cat and her laboratory cousins (1984) by A. J. Leggett. Contemporary Physics, 50(1):241, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Eckart:1926:OCS

[Eck26] Carl Eckart. Operator calculus and the solution of the equations of quantum dynamics. Physical Review, 28(4):711–726, October 1926. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://link.aps.org/doi/10.1103/PhysRev.28.

711. The paper has a final note added in proof and dated 2 September 1926, referring to just-published work [Sch26k]. Eckart ends with “The conclusion of this paper may be stated, with Schroedinger, to be ‘that the wave-mechanics and the matrixmechanics are mathematically identical.’ ”.

Edelstein:1955:RBG

[Ede55] Ludwig Edelstein. Review: Nature and the Greeks by Erwin Schrödinger. Quarterly review of biology, 30(1):47–48, March 1955. CODEN QRBIAK. ISSN 0033-5770 (print), 1539-7718 (electronic).

URL http://www.jstor.org/stable/2814396.

Egler:1956:RSS

[Egl56] Frank E. Egler. Review: Schrödinger’s sources of science [Nature and the Greeks]. Ecology, 37(3):620–621, July 1956. CODEN ECOLAR. ISSN 0012-9658 (print), 1939-9170 (electronic). URL

http://www.jstor.org/stable/1930197.

Einstein:1927:BSW

[Ein27] Albert Einstein. Bestimmt Schrödingers Wellenmechanik die Bewegung eines Systems vollständig oder nur im Sinne der Statistik? (German) [Does Schrödinger’s wave mechanics determine the motion of a system completely, or only in the sense of statistics?]. Read before the Prussian Academy of Sciences, but never

published., May 5, 1927. URL http://alberteinstein.info/
vufind1/Record/EAR000034338.

Ekert:2002:BTQ

[Eke02] Artur Ekert. 29. The Bell Theorem in quantum cryptography: Vortrag am 14. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Englert:2002:QCE

[Eng02] Berge Englert. 22. quantification and characterisation of entanglement: Vortrag am 13. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Einstein:1935:CQM

[EPR35] Albert Einstein, Boris Podolsky, and Nathan Rosen. Can quantum mechanical description of physical reality be considered complete? Physical Review (2), 47(10):777–780, May 15, 1935. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v47/i10/p777_1.

Epstein:1926:SQT

[Eps26] Paul S. Epstein. Schrödinger’s quantum theory and the Stark Effect. Nature, 118(2969):444–445, September 25, 1926. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v118/
n2969/pdf/118444a0.pdf.

Einstein:1936:FL

[EST36] Albert Einstein, E. Schrödinger, and V. Tchernavin. The freedom of learning. Science, 83(2155):372–373, April 17, 1936. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.jstor.org/stable/1661860;   http://www.
sciencemag.org/content/83/2155/372.2.full.pdf.

Evans:2007:QMC

[ET07] James Evans and A. S. (Alan S.) Thorndike, editors. Quantum mechanics at the crossroads: new perspectives from history, philosophy and physics. The frontiers collection. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-32663-4. ISSN 1612-3018. x + 249 pp. LCCN QC174.12 .Q346 2007. URL http:

//www.loc.gov/catdir/enhancements/fy0824/2006934045-b.
html;   http://www.loc.gov/catdir/enhancements/fy0824/
2006934045-d.html;   http://www.loc.gov/catdir/toc/fy0708/
2006934045.html.

Everett:1957:RSF

[Eve57] Hugh Everett III. “relative state” formulation of quantum mechanics. Reviews of Modern Physics, 29(3):454–462, July 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
29.454;   http://rmp.aps.org/abstract/RMP/v29/i3/p454_

1. See assessment [Whe57].

Fara:2001:GPI

[Far01] Patricia Fara. Group portraits IV — the Seventh Solvay Conference [Bruxelles, 22–29 October 1933]. Endeavour, 25(4):137–138, December 2001. CODEN ENDEAS. ISSN 0160-9327 (print), 1873-1929 (electronic).

Farmelo:2002:IMB

[Far02] Graham Farmelo, editor. It must be beautiful: great equations of modern science. Granta, London, UK, 2002. ISBN 1-86207-479-8. xvi + 283 pp. LCCN Q125 .I88 2002.

Farmelo:2003:IMB

[Far03] Graham Farmelo, editor. It Must Be Beautiful: Great Equations of Modern Science. Granta, London, UK, 2003. ISBN 1-86207-555-7. xviii + 284 pp. LCCN Q125 .I88 2003.

Felber:2006:SBN

[Fel06] Adam Felber. Schrödinger’s ball: a novel. Random House Trade Paperbacks, New York, NY, USA, 2006. ISBN 0-8129-7442-5 (paperback). 244 pp. LCCN PS3606.E384 S34 2006. URL http:

//www.loc.gov/catdir/enhancements/fy0643/2005051652-s.
html;   http://www.loc.gov/catdir/enhancements/fy0645/
2005051652-d.html.

Fernandez:2003:CMC

[FF03] Ruth Garcia Fernandez and Michael Fleischhauer. 5. Creation and measurement of a coherent superposition of quantum states. Quasidecoherence free subspaces in quantum memories for photons: suppressing decoherence using collective excitations: 5. September 2003. In Anonymous [Ano03c], page ?? ISBN ???? LCCN ????

Finch:1953:RBH

[Fin53] Henry A. Finch. Review: Science and Humanism: Physics in Our Time by Erwin Schrödinger. Annals of the American Academy of Political and Social Science, 285(??):202–203, January 1953. ISSN 0002-7162 (print), 1552-3349 (electronic). URL http://www.

jstor.org/stable/1029173.

Fine:1986:SGE

[Fin86] Arthur Fine. The shaky game: Einstein, realism, and the quantum theory. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA and London, UK, 1986. ISBN 0-226-24946-8. xi + 186 pp. LCCN QC6 .F54 1986.

Flamm:1987:BIS

[Fla87] Dieter Flamm. Boltzmann’s influence on Schrödinger. In Kilmister [Kil87b], pages 4–15. ISBN 0-521-37929-6 (paperback), 0-521-34017-9. LCCN QC16.S265 S36 1987; Q143.S3. URL http://www.

loc.gov/catdir/enhancements/fy0901/86028338-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0901/86028338-t.html.

Papers presented at a conference, March 31–April 3, 1987, at Imperial College, London.

Fleisch:2018:SGS

[Fle18] Daniel A. Fleisch. A Student’s Guide to the Schrödinger Equation. Cambridge University Press, Cambridge, UK, 2018. ISBN 1-108-42426-0 (hardcover), 1-108-43960-8 (paperback). ???? pp. LCCN

QC174.26.W28 F5545 2018. URL https://www.cambridge.org/
us/academic/subjects/physics/general-and-classical-physics/
students-guide-schrodinger-equation.

Fowler:1929:RBP

[Fow29] R. H. Fowler. Review: Collected Papers on Wave Mechanics by E. Schrödinger; Selected Papers on Wave Mechanics by L. de Broglie and L. Brillouin; Four Lectures on Wave Mechanics by E. Schrödinger. Mathematical Gazette, 14(199):369–370, March 1929. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL http://www.jstor.org/stable/3606729.

Fernandez:2013:HST

[FR13a]  Bernard Fernandez and Georges Ripka.  Heisenberg and

Schrödinger, two sides of the same coin. In Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956 [FR13c], page 139. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (e-book). LCCN QC773 .F47 2013.

Fernandez:2013:SE

[FR13b] Bernard Fernandez and Georges Ripka. The Schrödinger equation. In Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956 [FR13c], pages 136–138. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (e-book). LCCN QC773 .F47 2013.

Fernandez:2013:UMA

[FR13c] Bernard Fernandez and Georges Ripka. Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (ebook). xviii + 522 pp. LCCN QC773 .F47 2013.

Frank:2005:ERG

[Fra05] Tibor Frank. Ever ready to go: The multiple exiles of Leo Szilard. Physics in Perspective (PIP), 7(2):204–252, June 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://adsabs.harvard.edu/abs/2005PhP...
..7..204F;   http://link.springer.com/article/10.1007/
s00016-004-0236-9;   http://www.springerlink.com/content/
1422-6960/.

Freistadt:1961:RBM

[Fre61] Hans Freistadt. Review: Mind and Matter by Erwin Schrödinger. Science & Society, 25(1):86–90, ???? 1961. CODEN ???? ISSN 0036-8237 (print), 1943-2801 (electronic). URL http://www.

jstor.org/stable/40400727.

French:2012:BRB

[Fre12] Steven French. Book review: What Is Life? The Intellectual Pertinence of Erwin Schrödinger. Isis, 103(1):191, March 2012. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/666429.

Frieden:1989:FIB

[Fri89] B. Roy Frieden. Fisher information as the basis for the Schrödinger wave equation. American Journal of Physics, 57(11):1004–1008, November 1989. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Froehlich:1938:SSE

[Fro38] H. Froehlich. A solution of the Schrödinger equation by a perturbation of the boundary conditions. Physical Review, 54(11):945–947, December 1, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Frohlich:2003:YES

[Frö03] Jürg Fröhlich. 4. Dissipative transport (recent results): Vortrag am 14. März 2003. In Anonymous [Ano03a], page ?? ISBN ???? LCCN ????

Frohlich:2011:MVQ

[Frö11] Jürg Fröhlich. A minimalist’s view of quantum mechanics, eighty five years after schrödinger’s “Annus mirabilis”. Mitschnitt einer gemeinsamen Veranstaltung des Erwin Schrödinger-Instituts für Mathematische Physik (ESI) und der Fakultät für Physik der Universität Wien am Donnerstag, dem 13. Jänner 2011 im Kleinen Festsaal der Universität Wien [Recording of a joint event at the Erwin Schrödinger Institute for Mathematical Physics (ESI) and the Department of Physics, University of Vienna on Thursday, the 13th January 2011 in the Small Hall of the University of Vienna]., 2011.

Fry:2002:NDT

[Fry02] Ed Fry. 14. A novel definitive test of Bell inequalities: Vortrag am 11. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Fischer:2006:EPF

[FS06] Ernst Peter Fischer and Klaus Sander. Ernst Peter Fischer erzählt, Paarläufe der Wissenschaft. (German) [Ernst Peter Fischer tells the pair runs of science]. Suppoš, Köln, Germany, 2006. ISBN

3-932513-68-1. LCCN ???? URL http://deposit.ddb.de/cgi-
bin/dokserv?id=2758637\%26prov=M\%26dok\_var=1\%26dok\
_ext=htm. 4 CDs.

Frohlich:2013:DWU

[FS13] Jürg Fröhlich and Baptiste Schubnel. Do we understand quantum mechanics — finally? In Reiter and Yngvason [RY13], pages 37–86. ISBN 3-03719-121-X (paperback), 3-03719-621-1 (e-book). LCCN QC16.S265 I58 2011.

Fukui:1987:OPC

[Fuk87] Kenichi Fukui. Orbital presentation of chemical reactions. In Kilmister [Kil87b], pages 102–111. ISBN 0-521-37929-6 (paperback), 0-521-34017-9. LCCN QC16.S265 S36 1987;

Q143.S3. URL http://www.loc.gov/catdir/enhancements/
fy0901/86028338-d.html;   http://www.loc.gov/catdir/enhancements/

fy0901/86028338-t.html. Papers presented at a conference, March 31–April 3, 1987, at Imperial College, London.

Fulde:2015:SQW

[Ful15] Peter Fulde. Schöpfer der Quantenphysik: Werner Heisenberg und Erwin Schrödinger. (German) [Creator of quantum physics: Werner Heisenberg and Erwin Schrödinger]. In Berlins wilde Energien: Porträts aus der Geschichte der Leibnizschen Wissenschaftsakademie. (German) [Berlin’s Wild Energies: Portraits from the History of the Leibniz Science Academy] [LMOS15], pages 436–460. ISBN 3-11-037598-2 (hardcover). LCCN AS182.B38 B477 2015.

Forstner:2020:BHP

[FW20] Christian Forstner and Mark Walker, editors. Biographies in the History of Physics: Actors, Objects, Institutions. Springer International Publishing, Cham, Switzerland, 2020. ISBN 3-030-48508-0 (print), 3-030-48509-9 (e-book), 3-030-48510-2, 3-030-48511-0. vi + 324 + 38 pp. LCCN QC16.E5 .B564 2020. URL http://www.

springerlink.com/content/978-3-030-48509-2.

Gumbrecht:2011:GMW

[G^{+}11] Hans Ulrich Gumbrecht et al., editors. Geist und Materie. What is life?: the intellectual pertinence of Erwin Schrödinger. Stanford University Press, Stanford, CA, USA, 2011. ISBN 0-8047-6915-X (cloth), 0-8047-6916-8 (paperback). ???? pp. LCCN QC16.S265 W4313 2011.

Gallavotti:2003:SEF

[Gal03] Giovanni Gallavotti. 1. Statistical equivalence in fluid mechanics models: Vortrag am 13. März 2003. In Anonymous [Ano03a], page ?? ISBN ???? LCCN ???? URL http://permalink.obvsg.

at/AC03695075.

Gao:2018:ECC

[Gao18] Shan Gao. Is an electron a charge cloud? A reexamination of Schrödinger’s charge density hypothesis. Foundations of Science, 23(1):145–157, March 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-017-9521-3.

Garcia:1953:RGB

[Gar53] Godofredo García. Die relativistischen Gleichungen von de Broglie und Schrödinger. Die Stabilität der Bewegungen. (German) [The relativistic equations of de Broglie and Schrödinger. The stability of the movements]. Actas Acad. Nac. Ci. Exact., Fis. Natur. Lima, 16(3–4):36–55, 1953.

Garwin:2005:SM

[Gar05] Laura Garwin. Schrödinger’s mousetrap. Nature, 433(7026):579, February 9, 2005. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v433/n7026/full/433579a.html.

Gaviola:1928:ETS

[Gav28] E. Gaviola. An experimental test of Schrödinger’s theory. Nature, 122(3081):772, November 17, 1928. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v122/n3081/pdf/122772a0.pdf.

Gawedzki:1997:FLO

[Gaw97] Krzysztof Gawedzki. Five lectures on the occasion on Walter Thirring’s 70th birthday. 4. Manybody type problems in a model of turbulence: 1997/04/10. Report, Internationales Erwin-Schrödinger-Institut für Mathematische Physik, Vienna, Austria, 1997.

Gumbrecht:2008:GMW

[GB08] Hans Ulrich Gumbrecht and Sabine Baumann. Geist und Materie — was ist Leben?: zur Aktualität von Erwin Schrödinger, volume 13 of Edition Unseld. Suhrkamp, Frankfurt am Main, Germany, 2008. ISBN 3-518-26013-8. 149 pp. LCCN ????

EUR 10.00. URL http://d-nb.info/988764180/04;   http://
www.gbv.de/dms/faz-rez/FD1200901072106665.pdf;   http:
//www.suhrkamp.de/buecher/geist\_und\_materie-hans\_ulrich\

_gumbrecht\_26013.html. Translated from English by Sabine Baumann.

Gee:2005:SM

[Gee05] Henry Gee. Schrödinger’s mousetrap. Nature, 434(7032):440–441, March 23, 2005. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v434/n7032/full/434440a.html.

Gefter:2014:SR

[Gef14a] Amanda Gefter. Schrödinger’s rats. In Trespassing on Einstein’s lawn: a father, a daughter, the meaning of nothing, and the beginning of everything [Gef14b], page ?? ISBN 0-345-53143-4 (hardcover), 0-345-53963-X (e-book). LCCN QC6 .G34 2014.

Gefter:2014:TEL

[Gef14b] Amanda Gefter. Trespassing on Einstein’s lawn: a father, a daughter, the meaning of nothing, and the beginning of everything. Bantam Books, New York, NY, USA, 2014. ISBN 0-345-53143-4 (hardcover), 0-345-53963-X (e-book). 418 pp. LCCN QC6 .G34 2014.

Goldhaber-Gordon:2005:SM

[GGGG05] Ilana Goldhaber-Gordon and David Goldhaber-Gordon. Schrödinger’s mousetrap. Nature, 433(7028):805, February 23, 2005. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v433/
n7028/full/433805a.html.

Galison:2002:QMS

[GGK02] Peter Galison, Michael Gordin, and David Kaiser, editors. Quantum Mechanics: Science and Society. Routledge, London, UK, 2002. ISBN 1-136-70972-X. 433 (est.) pp. LCCN ????

Gnaiger:1994:WCL

[GGW94] E. (Erich) Gnaiger, Frank N. Gellerich, and Markus Wyss, editors. What is controlling life?: 50 years after Erwin Schrödinger’s “What is Life?”, volume 205 of Modern trends in biothermokinetics (Innsbruck, Austria) (vol. 3). Innsbruck University Press, Innsbruck, Austria, 1994. ISBN 3-901249-17-6. 335 pp. LCCN QP517.B54 W48 1994.

Gumbrecht:2010:GMM

[GHHL10] Hans Ulrich Gumbrecht, Robert Pogue Harrison, Michael R. Hendrickson, and Robert B. Laughlin. Geist und Materie. Mente y materia: ¿Que es la vida?: sobre la vigencia de Erwin Schrödinger. (Spanish) [Mind and Matter. What is Life?: On the life of Erwin Schrödinger], volume 3069 of Conocimiento. Katz Editores, Buenos Aires, Argentina, 2010. ISBN 987-1566-28-X, 84-92946-12-1. 191 pp. LCCN QC16.S265 G4518 2010. Translation by Silvia Villegas.

Ghirardi:1999:DRP

[Ghi99] GianCarlo Ghirardi. 7. The dynamical reduction program: an example of a quantum theory without observers: 4. Sept. 1998. In Greenberger et al. [GRZ99], page ?? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN Q44.4 V662 v. 7.

Ghirardi:2002:JBD

[Ghi02] GianCarlo Ghirardi. 27. John Bell and the dynamical reduction program: Vortrag am 14. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Grosser:1999:NTG

[GHKO99] Michael Grosser, Günther Hörmann, Michael Kunzinger, and Michael Oberguggenberger, editors. Nonlinear theory of generalized functions: proceedings of the Workshop Nonlinear Theory of Generalized Functions — Erwin-Schrödinger-Institute, Vienna, October–December 1997, volume 401 of Chapman & Hall/CRC research notes in mathematics series. Chapman and Hall/CRC, Boca Raton, FL, USA, 1999. ISBN 0-8493-0649-3. LCCN QA324 .N66 1999.

Gilder:2009:AEW

[Gil09] Louisa Gilder. The Age of Entanglement: When Quantum Physics Was Reborn. Vintage Books, New York, NY, USA, 2009. ISBN 1-4000-9526-3. xviii + 443 pp. LCCN QC174.12 .G528 2009.

Gisin:2002:TRQ

[Gis02] Nicolas Gisin. 26. Test of relativistic quantum state collapse with moving reference frame: Vortrag am 14. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Gisin:2006:NAS

[Gis06] Nicolas Gisin. New additions to the Schrödinger cat family. Science, 312(5770):63–64, April 7, 2006. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/312/5770/63.full.pdf.

Golub:2023:HPF

[GL23] Robert Golub and Steve Keith Lamoreaux. The Historical and Physical Foundations of Quantum Mechanics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2023. ISBN 0-19-186123-5, 0-19-255536-7, 0-19-882218-9 (hardcover), 0-19-882219-7 (paperback). xiii + 747 pp. LCCN QC174.12 .G65 2023.

Gindikin:1994:YRT

[GM94] Simon G. (Semen Grigorevic) Gindikin and Peter Michor, editors. 75 years of Radon transform: proceedings of the conference held at the Erwin Schrödinger International Institute for Mathematical Physics in Vienna, August 31 September 4, 1992, volume 4 of Conference proceedings and lecture notes in mathematical physics. International Press, Cambridge, MA, USA, 1994. ISBN 1-57146-008-X. LCCN QA649 .A15 1994.

Gerlach:1929:BGM

[GNK^{+}29] Walther Gerlach, Lothar Nordheim, H. Kopfermann, W. Bothe, E. Schrödinger, et al. Besprechungen. (German) [Meetings]. Naturwissenschaften, 17(35):691–696, August 1929. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://www.

springerlink.com/content/m847550644672314/fulltext.pdf.

Goldhaber:1990:SPW

[Gol90] Maurice Goldhaber. Schrödinger in pre-war Germany. Nature, 343(6260):686, February 22, 1990. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v343/n6260/pdf/343686c0.pdf.

Gotschl:1992:ESW

[Göt92a] Johann Götschl, editor. Erwin Schrödinger’s world view: the dynamics of knowledge and reality, volume 16 of Theory and decision library. Series A, Philosophy and methodology of the social sciences. Kluwer Academic, Dordrecht, The Netherlands, 1992. ISBN 0-7923-1694-0. viii + 237 pp. LCCN QC16.S265 E78 1992.

Gotschl:1992:RPE

[Göt92b] Johann Götschl. The role of physics in Erwin Schrödinger’s philosophical system. In Bitbol and Darrigol [BD92], pages 121–150. ISBN 2-86332-116-1. LCCN QC16.S265 E78 1992. 263.75F.

Gottfried:1999:MOD

[Got99] Kurt Gottfried. 6. My ongoing debate with John Bell: 4. Sept. 1998. In Greenberger et al. [GRZ99], page ?? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN Q44.4 V662 v. 7.

Gribbin:1984:SSCb

[Gri84a] John R. Gribbin. In search of Schrödinger’s cat: quantum physics and reality. Bantam Books, New York, NY, USA, 1984. ISBN 0-553-34103-0 (paperback). xvi + 302 pp. LCCN QC173.98 .G75 1984.

Gribbin:1984:SSCa

[Gri84b] John R. Gribbin. In search of Schrödinger’s cat: the startling world of quantum physics explained. Wildwood House, London, UK, 1984. ISBN 0-7045-3071-6. xvi + 302 pp. LCCN QC173.98 .G75x 1984b. US$12.50.

Gribbin:1995:SKSb

[Gri95a] John R. Gribbin. Schrödinger’s kittens and the search for reality. Weidenfeld and Nicolson, London, UK, 1995. ISBN 0-297-81519-9 (hardcover). ix + 261 pp. LCCN ????

Gribbin:1995:SKSa

[Gri95b] John R. Gribbin. Schrödinger’s kittens and the search for reality: solving the quantum mysteries. Little, Brown and Company, Boston, Toronto, London, 1995. ISBN 0-316-32838-3. ix + 261 pp. LCCN QC174.12 .G747 1995.

Gribbin:2012:ESQ

[Gri12] John R. Gribbin. Erwin Schrödinger and the quantum revolution. Bantam Press, London, UK, 2012. ISBN 0-593-06776-2 (hardcover), 0-593-06865-3 (paperback), 1-4464-6571-3 (e-book). xii + 322 + 16 pp. LCCN QC16.S265 G75 2012.

Gribbin:2013:SNS

[Gri13a]  John Gribbin.  Auf der Suche nach Schrödingers Katze:

Quantenphysik und Wirklichkeit. (German) [On the search for Schrödinger’s cats: quantum physics and reality, volume 4030 of Serie Piper. Piper, München, Germany, 2013. ISBN 3-492-24030-5. 324 pp. LCCN ????

Gribbin:2013:ESQ

[Gri13b] John Gribbin. Erwin Schrödinger and the quantum revolution. Wiley, New York, NY, USA, 2013. ISBN 1-118-29926-4 (hardback), 1-118-33411-6 (pdf), 1-118-33188-5 (epub), 1-118-33519-8 (mobi). ???? pp. LCCN QC16.S265 G75 2012.

URL http://catalogimages.wiley.com/images/db/jimages/
9781118299265.jpg.

Grinwald:2022:SEC

[Gri22] P. M. Grinwald. Schrödinger’s equation as a consequence of the Central Limit Theorem without assuming prior physical laws. Foundations of Physics, 52(2):??, April 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://

link.springer.com/article/10.1007/s10701-022-00561-1.

Gallavotti:2006:BLI

[GRY06] Giovanni Gallavotti, Wolfgang L. Reiter, and Jakob Yngvason, editors. Boltzmann’s legacy: an international symposium at the Erwin Schrödinger Institute for Mathematical Physics; in commemoration of Ludwig Boltzmann’s death on September 5, 1906, during his summer vacation in Duino near Trieste; Wednesday, June 7, to Friday, June 9, 2006, ESI lectures in mathematics and physics. European Mathematical Society, Zúrich, Switzerland, 2006. ISBN 3-03719-057-4. LCCN QC174.7 .I5662 200.

Greenberger:1999:EEP

[GRZ99] Daniel Greenberger, Wolfgang L. Reiter, and Anton Zeilinger, editors. Epistemological & experimental perspectives on quantum physics: International Symposium, September 3–6, 1998, volume 7 of Vienna Circle Institute yearbook. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN Q44.4 V662 v. 7.

Guggenberger:1987:IEG

[GS87] Elisabeth Guggenberger and Erwin Schrödinger. Interviews von Elisabeth Guggenberger über Erwin Schrödinger mit …: Erwin Schrödinger zum 100. Geburtstag. (German) [Interviews by Elisabeth Guggenberger about Erwin Schrödinger with … Erwin Schrödinger’s 100th Birthday]. ORF, Vienna, Austria, 1987. ???? pp. Die Videoaufzeichnungen der Interviews wurden angefertigt anläßlich des 100. Geburtstages von Erwin Schrödinger für den am 1987-08-10 erstmals ausgestrahlten Fernsehfilm “Die Raupe kann den Schmetterling nicht verstehen”. [The video recordings of the interviews were made on the occasion of the 100th Birthday of Erwin Schrödinger for the television movie “The caterpillar has the Butterfly does not understand”, broadcast for the first time on 1987-08-10.].

Grotrian:1926:MVG

[GSL26] W. Grotrian, E. Schrödinger, and Lü. Mitteilungen aus verschiedenen Gebieten. (German) [Reports from different areas]. Naturwissenschaften, 14(8):144–148, February 1926. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://www.

springerlink.com/content/q6210hkj47w7q856/fulltext.pdf.

Haas:1934:MQE

[Haa34]  Arthur Haas.  Materiewellen und Quantenmechanik. Eine

Einführung auf Grund der Theorien von de Broglie, Schrödinger, Heisenberg und Dirac. 4. u. 5. verb. u. abermals wes. verm. Aufl. (German) [Matter waves and quantum mechanics. An introduction based on the theories of de Broglie, Schrödinger, Heisenberg and Dirac]. Akad. Verlagsges., Leipzig, Germany, 1934. viii + 299 + 7 pp.

Halpern:1987:ESV

[Hal87] Leopold Halpern. Erwin Schrödinger’s views on gravitational physics during his last years at the University of Vienna and some research ensuing from it. Foundations of Physics, 17(11): 1113–1130, November 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889804.

Halpern:2015:EDS

[Hal15] Paul Halpern. Einstein’s dice and Schrödinger’s cat: how two great minds battled quantum randomness to create a unified theory of physics. Basic Books, New York, NY, USA, 2015. ISBN 0-465-07571-1 (hardcover). x + 271 pp. LCCN QC174.17.C45 H35 2015.

Hanson:1962:RBP

[Han62] Norwood Russell Hanson. Review: The Physicist’s Conception of Nature by Werner Heisenberg (translated by Arnold J. Pomerans), Physics and Philosophy: The Revolution in Modern Science by Werner Heisenberg; Mind and Matter by Erwin Schrödinger. The Journal of Philosophy, 59(12):320–322, June 7, 1962. CODEN ???? ISSN 0022-362X (print), 1939-8549 (electronic). URL http:

//www.jstor.org/stable/2022902.

Hanle:1977:CAE

[Han77a] Paul A. Hanle. The coming of age of Erwin Schrödinger: His quantum statistics of ideal gases. Archive for History of Exact Sciences, 17(2):165–192, July 1977. CODEN AHESAN.

ISSN 0003-9519 (print), 1432-0657 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=17&issue=2&spage=165.

Hanle:1977:ESR

[Han77b] Paul A. Hanle. Erwin Schrödinger’s reaction to Louis de Broglie’s thesis on the quantum theory. Isis, 68(4):606–609, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230015.

Hanle:1979:IBH

[Han79a] Paul A. Hanle. Indeterminacy before Heisenberg: The case of Franz Exner and Erwin Schrödinger. Historical Studies in the Physical Sciences, 10(??):225–269, ???? 1979. CODEN HSPSAS. ISSN

0073-2672. URL http://www.jstor.org/stable/27757391.

Hanle:1979:SCS

[Han79b] Paul A. Hanle. The Schrödinger–Einstein correspondence and the sources of wave mechanics. American Journal of Physics, 47(7): 644–648, 1979. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://scitation.aip.org/content/
aapt/journal/ajp/47/7/10.1119/1.11950.

Haroche:1999:QEA

[Har99] Serge Haroche. 13. Quantum engineering with atoms and photons in a cavity: 5. Sept. 1998. In Greenberger et al. [GRZ99], page ?? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN Q44.4 V662 v. 7.

Hawking:1987:SEU

[Haw87] S. W. Hawking. The Schrödinger equation of the Universe. In Kilmister [Kil87b], pages 176–179. ISBN 0-521-37929-6 (paperback), 0-521-34017-9. LCCN QC16.S265 S36 1987;

Q143.S3. URL http://www.loc.gov/catdir/enhancements/
fy0901/86028338-d.html;   http://www.loc.gov/catdir/enhancements/
fy0901/86028338-t.html. See correction [].

Hawking:2011:DSM

[Haw11] Stephen Hawking, editor. The dreams that stuff is made of: the most astounding papers on quantum physics — and how they shook the scientific world. Running Press, Philadelphia, PA, USA, 2011. ISBN 0-7624-3434-1. xi + 1071 pp. LCCN QC173.98 .D74 2011.

Heisenberg:1996:FCS

[HBS96] Werner Heisenberg, Niels Bohr, and Erwin Schrödinger. Física cuántica: (Spanish) [Quantum physics]. Círculo de Lectores, Barcelona, Spain, 1996. ISBN ???? ???? pp. LCCN ????

Heisenberg:1961:MPb

[HBSA61] Werner Heisenberg, Max Born, Erwin Schrödinger, and Pierre Auger, editors. On modern physics. Clarkson N. Potter. Inc., New York, NY, USA, 1961. 108 pp. LCCN QC71 .O5.

Harman:2013:OSR

[HD13] Oren Solomon Harman and Michael R. Dietrich, editors. Outsider scientists: routes to innovation in biology. University of Chicago Press, Chicago, IL, USA and London, UK, 2013. ISBN 0-226-07837-X (hardcover), 0-226-07840-X (paperback), 0-226-07854-X (e-book). 374 pp. LCCN QH26 .O98 2013.

He:2020:CIC

[He20] Guang Ping He. Copenhagen interpretation can survive the upgraded Schrödinger’s cat Gedankenexperiment. Foundations of Physics, 50(7):715–726, July 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Heisenberg:1961:ESG

[Hei61a] Werner Heisenberg. Erwin Schrödinger 12.8.1887–4.1.1961. (German). In ????, editor, Bayerische Akademie der Wissenschaften. Jahrbuch 1961. (German) [Bavarian Academy of Science, Yearbook 1961], pages 194–196. Bayerische Akademie der Wissenschaften, Munich, West Germany, 1961.

Heitler:1961:ES

[Hei61b] Walter Heitler. Erwin Schrödinger. 1887–1961. Biographical Memoirs of Fellows of the Royal Society, 7(??):221–228, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL http://www.jstor.org/stable/769408.

Heisenberg:1963:GFE

[Hei63] Werner Heisenberg. Gedenkworte für Erwin Schrödinger. (German) [Memorial words for erwin schrödinger]. In ????, editor, Orden Pour le Mérite für Wissenschaft und Künste. Reden und Gedenkworte. Vol. 5. (German) [Pour le Mérite for Science and Art. Speeches and commemorative words. Vol 5], pages 27–35. Lambert Schneider, Heidelberg, West Germany, 1963. LCCN ????

Heilbron:1968:BRBa

[Hei68] John L. Heilbron. Book review: Erwin Schrödinger. An Introduction to His Writings by William T. Scott. Isis, 59(2):242–243, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228307.

Hendry:1988:BRJb

[Hen88] John Hendry. Book review: Jagdish Mehra & Helmut Rechenberg. The Historical Development of Quantum Theory. Volume 5. Erwin Schrödinger and the Rise of Wave Mechanics. Part 1. Schrödinger in Vienna and Zurich, 1887–1925 and Part 2. The Creation of Wave Mechanics; Early Response and Applications, 1925–1926. Berlin, Heidelberg, New York: Springer-Verlag, 1987. Pp. xix + 366 and viii + 615. ISBN 3-540-96284-0 and 96377-4. DM 148.00 (hardcover) and 98.00 (paperback). British Journal for the History of Science, 21(3):371–372, September 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

http://www.jstor.org/stable/4026844.

Hentschel:2012:BRE

[Hen12] Klaus Hentschel. Book review: Eine Entdeckung von ganz außerordentlicher Tragweite. Schrödingers Briefwechsel zur Wellenmechanik und zum Katzenparadoxon, 2 Bde. by Karl von Meyenn. Sudhoffs Archiv: Zeitschrift für Wissenschaftsgeschichte, 96(1): 118–119, ???? 2012. CODEN SUARAH. ISSN 0039-4564. URL

http://www.jstor.org/stable/43694434.

Heisenberg:1946:BES

[HF46] Werner Heisenberg and F. Firbas. Besprechungen: Erwin Schrödinger, Statistical Thermodynamics. (German) [Book Reviews]. Naturwissenschaften, 33(10):319, ???? 1946. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL

http://www.springerlink.com/content/wvk1r381221v8g10.

Hiebert:1968:BRB

[Hie68] Erwin N. Hiebert. Book review: Letters on Wave Mechanics: Schrödinger, Planck, Einstein, Lorentz. K. Przibram, Ed. Translated from the German with an introduction by Martin J. Klein. Philosophical Library, New York, 1967. xx + 75 pp., illus. $6. Science, 160(3828):664–666, May 10, 1968. CODEN SCIEAS.

ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/160/3828/664.full.pdf.

Hiley:1999:BIA

[Hil99] Basil Hiley. 8. The Bohm interpretation, active information and teleportation: 4. Sept. 1998. In Greenberger et al. [GRZ99], page ???? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN Q44.4 V662 v. 7.

Hittmair:1987:SUF

[Hit87] O. Hittmair. Schrödinger’s unified field theory seen 40 years later. In Kilmister [Kil87b], pages 165–175. ISBN 0-521-37929-6 (paperback), 0-521-34017-9. LCCN QC16.S265 S36 1987;

Q143.S3. URL http://www.loc.gov/catdir/enhancements/
fy0901/86028338-d.html;   http://www.loc.gov/catdir/enhancements/

fy0901/86028338-t.html. Papers presented at a conference, March 31–April 3, 1987, at Imperial College, London.

Hoaken:1996:SCH

[Hoa96] Paul C. S. Hoaken. Schrödinger’s cat at hand. Science, 273 (5277):860, August 16, 1996. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/273/5277/860.3.full.pdf.

Hobson:2017:TQU

[Hob17] Art Hobson. Tales of the Quantum: Understanding Physics’ Most Fundamental Theory. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2017. ISBN 0-19-067963-8 (hardcover), 0-19-938422-3. xvi + 287 pp. LCCN QC174.12 .H53 2017.

Hoffmann:1984:ES

[Hof84] Dieter Hoffmann. Erwin Schrödinger. Teubner, Stuttgart, Germany; Leipzig, Germany, 1984. ISBN 3-322-00586-0. 94 pp. LCCN ????

Horne:1999:CFV

[Hor99] Michael Horne. 17. Complementarity of fringe visibilities in threeparticle quantum mechanics: 5. Sept. 1998. In Greenberger et al. [GRZ99], page ?? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN Q44.4 V662 v. 7.

Horne:2002:IJB

[Hor02] Michael Horne. 8. Interactions with John Bell on the nonlocality problem: Vortrag am 10. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Houston:2016:BWS

[Hou16] Louis M. Houston. A de Broglie wave solution to the Maxwell– Schrödinger–Einstein equations. Theor. Math. Appl., 6(3):1–13, 2016. ISSN 1792-9687 (print), 1792-9709 (electronic).

Howard:1999:BBN

[How99] Don Howard. 4. A brief on behalf of Niels Bohr: 4. Sept. 1998. In Greenberger et al. [GRZ99], page ?? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN Q44.4 V662 v. 7.

Hund:1974:GQHa

[HR74] Friedrich Hund and Gordon Reece. The History of Quantum Theory. Harrap, London, UK, 1974. ISBN 0-245-50902-X. 260 pp. LCCN ???? Translation by Gordon Reece of Geschichte der Quantentheorie [?].

Hittmair:1951:SGT

[HS51] O. Hittmair and Erwin Schrödinger. Studies in the generalized theory of gravitation II: the velocity of light, volume A8 of Communications of the Dublin Institute for Advanced Studies. Institiúid Árd-Léinn Bhaile Átha Cliath (Dublin Institute for Advanced Studies), Dublin, Ireland, 1951. 15 pp. LCCN QC407 .H56.

Heisenberg:1960:DSF

[HSBA60] Werner Heisenberg, Erwin Schrödinger, Max Born, and Pierre Auger. Discussione sulla fisica moderna. (Italian) [Discussions on modern physics], volume 1952 of Quattro conferenze organizzate dalle “Rencontres internationales de Genève”. Biblioteca di cultura scientifica, 59. Paolo Boringhieri, Torino, 1960. v + 131 + 1 pp.

Heisenberg:1934:MAE

[HSD34a] Werner Heisenberg, Erwin Schrödinger, and P. A. M. (Paul Adrien Maurice) Dirac. Die moderne Atomtheorie: die bei der entgegennahme des Nobelpreises 1933 in Stockholm gehaltenen Vorträge. (German) [Modern Atomic Theory: the 1933 Nobel Prize in Physics lecture in Stockholm]. Verlag von S. Hirzel, Leipzig, Germany, 1934. 2 + 45 + 1 pp. LCCN QC173 .H38.

Heisenberg:1934:MQM

[HSD34b] Werner Heisenberg, Erwin Schrödinger, and P. A. M. (Paul Adrien Maurice) Dirac. Modern Quantum Mechanics: Three Nobel Reports. Technico-Theoretical State Press, Leningrad and Moscow, USSR, 1934. ???? pp.

Hewlett:1931:BGM

[HSE31] C. W. Hewlett, E. Schrödinger, and P. P. Ewald. Besprechungen. (German) [Meetings]. Naturwissenschaften, 19(48):966–968, November 1931. CODEN NATWAY. ISSN 0028-1042 (print),

1432-1904 (electronic). URL http://www.springerlink.com/
content/pqn3379r8n7r244q/fulltext.pdf.

Hazen:2010:GIS

[HT10] Robert M. Hazen and James Trefil, editors. Great ideas of science: a reader in the classic literature of science. Cognella, San Diego, CA, USA, 2010. 256 pp. LCCN Q161.2 .G74 2010.

Hughes:1989:SIQ

[Hug89] R. I. G. Hughes. The Structure and Interpretation of Quantum Mechanics. Harvard University Press, Cambridge, MA, USA, 1989. ISBN 0-674-84391-6 (hardcover), 0-674-84392-4 (paperback). xiii + 369 pp. LCCN QC174.12 .H82 1989.

Hund:1974:HQTb

[Hun74] Friedrich Hund. The History of Quantum Theory. Barnes and Noble, New York, NY, USA, 1974. ISBN 0-06-493060-2. 260 pp. LCCN QC173.98 .H8613 1974b. Translation by Gordon Reece of Geschichte der Quantentheorie [?].

Hurter:2022:TBS

[Hür22] Tobias Hürter. Too Big for a Single Mind: How the Greatest Generation of Physicists Uncovered the Quantum World. The Experiment, New York, NY, USA, 2022. ISBN 1-61519-920-9 (hardcover), 1-61519-921-7 (e-book). 357 pp. LCCN QC7 .H88813 2022. Translation to English by David Shaw of Das Zeitalter der Unschärfe, published by Klett-Cotta (2021).

Hey:1986:QU

[HW86] Anthony J. G. Hey and Patrick Walters. The Quantum Universe. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-26744-7 (hardcover), 0-521-31845-9 (paperback). vii + 180 pp. LCCN QC174.12 .H48 1987.

Hoch:1987:SOH

[HY87] P. K. Hoch and E. J. Yoxen. Schrödinger at Oxford: a hypothetical national cultural synthesis which failed. Annals of Science, 44(6): 593–616, November 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ishak:2020:BRSb

[Ish20] B. Ishak. Book review: A student’s guide to the Schrödinger equation, by Daniel A. Fleisch, Cambridge University Press, 2020, $24.99 (Paperback), ISBN 978-1-108-81978-7. Scope: guide, textbook. Level: advanced undergraduate. Contemporary Physics, 61 (4):298, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jackiw:2002:DCA

[Jac02] Roman Jackiw. 4. descendants of the chiral anomaly: Vortrag am 10. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Jahrreiss:1956:RBN

[Jah56] Walter O. Jahrreiss. Review: Spirit and Nature. Papers from the Eranos Yearbooks by Ernesto Buonaiuti, Friedrich Dessauer, C. G. Jung, Werner Kaegi, C. Kerenyi, Paul Masson-Oursel, Fritz Meier, Adolf Portmann, Max Pulver, Hugo Rahner, Erwin Schrödinger, Walter Wili. Quarterly review of biology, 31(2):163–164, June 1956. CODEN QRBIAK. ISSN 0033-5770 (print), 1539-7718 (electronic).

URL http://www.jstor.org/stable/2815242.

Jammer:1974:ESI

[Jam74a] Max Jammer. Early semiclassical interpretations. In The Philosophy of Quantum Mechanics: the Interpretations of Quantum Mechanics in Historical Perspective [Jam74b], chapter 2, pages 20–54. ISBN 0-471-43958-4. LCCN QC173.98 .J35.

Jammer:1974:PQM

[Jam74b] Max Jammer. The Philosophy of Quantum Mechanics: the Interpretations of Quantum Mechanics in Historical Perspective. Wiley, New York, NY, USA, 1974. ISBN 0-471-43958-4. xi + 536 pp. LCCN QC173.98 .J35.

Johansson:2007:IQM

[Joh07] Lars-Göran Johansson. Interpreting quantum mechanics: a realistic view in Schrödinger’s vein. Ashgate, Aldershot, England, 2007. ISBN 0-7546-5738-8 (hardcover). xii + 195 pp. LCCN

QC174.12 .J639 2007. URL http://www.loc.gov/catdir/toc/
ecip076/2006100194.html.

Johnson:2013:CMG

[Joh13] Jeffrey Allan Johnson. The case of the missing German quantum chemists: On molecular models, mobilization, and the paradoxes of modernizing chemistry in Nazi Germany. Historical Studies in the Natural Sciences, 43(4):391–452, September 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http:

//www.jstor.org/stable/10.1525/hsns.2013.43.4.391.

Jones:1997:FLO

[Jon97] Vaughan Jones. Five lectures on the occasion on Walter Thirring’s 70th birthday. 3. Planar algebras: 1997/04/10. Report, Internationales Erwin-Schrödinger-Institut für Mathematische Physik, Vienna, Austria, 1997.

Jones:2003:SSV

[Jon03] Vaughan Jones. 3. Surprises from subfactors: Vortrag am 13.März 2003. In Anonymous [Ano03a], page ?? ISBN ???? LCCN ????

Jones:2008:QTS

[Jon08] Sheilla Jones. The Quantum Ten: a Story of Passion, Tragedy, Ambition and Science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-536909-2. xii + 323 + 8 pp. LCCN QC174.12 .J66 2008.

Janossy:1987:MLJ

[JS87] Lajos Jánossy and Erwin Schrödinger. In memoriam: Lajos Jánossy 75, Erwin Schrödinger 100. MTA Központi Fizikai Kutató Intézete, Budapest, Hungary, 1987. ISBN 963-372-404-X. xx + 148 pp. LCCN QC16.J36 I5 1987.

kaltenbaek:2013:OSC

[kA13] Rainer kaltenbaek and Markus Aspelmeyer. Optomechanical Schrödinger cats — a case for space. In Reiter and Yngvason [RY13], pages 123–134. ISBN 3-03719-121-X (paperback), 3-03719-621-1 (e-book). LCCN QC16.S265 I58 2011.

Kaempffert:1931:HEU

[Kae31] Waldemar Kaempffert. How to explain the universe? Science in a quandary: One after another the theories put out by the scientists have been exploded and now science in its uncertainty has been forced to become idealistic and to drop the idea of a mechanical universe. New York Times, ???(??):120, January 11, 1931. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/99177884/13F390B7A1F783DA798.

Kaempffert:1948:RRB

[Kae48] Waldemar Kaempffert. The revolution that radium began: Fifty years after the Curies’ great discovery, nuclear physics is still a realm unbounded. New York Times, ??(??):SM13, SM25, SM27, December 26, 1948. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/108348269/.

Kaiser:1985:DHE

[Kai85a] Walter Kaiser. Dieter Hoffmann: Erwin Schrödinger. (Biographien hervorragender Naturwissenschaftler, Techniker und Mediziner, Band 66) Leipzig: BSB B. G. Teubner Verlagsgesellschaft 1984. 94 SS., Broschur 4,80 M. (DDR)/6,80 M. (außerhalb der DDR). Berichte zur Wissenschaftsgeschichte, 8(2):86, ???? 1985. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
bewi.19850080205/abstract.

Kaiser:1985:RDH

[Kai85b] Walter Kaiser. Rezension: Dieter Hoffmann: Erwin Schrödinger. (Biographien hervorragender Naturwissenschaftler, Techniker und Mediziner, Band 66) Leipzig: BSB B. G. Teubner Verlagsgesellschaft 1984. 94 SS., Broschur 4,80 M. (DDR)/6,80 M. (außerhalb der DDR). Berichte zur Wissenschaftsgeschichte, 8 (2):86, ???? 1985. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Karplus:1987:MDH

[Kar87] Martin Karplus. Molecular dynamics: from H+H2 to biomolecules. In Kilmister [Kil87b], pages 85–101. ISBN 0-521-37929-6 (paperback), 0-521-34017-9. LCCN QC16.S265 S36 1987;

Q143.S3. URL http://www.loc.gov/catdir/enhancements/
fy0901/86028338-d.html;   http://www.loc.gov/catdir/enhancements/

fy0901/86028338-t.html. Papers presented at a conference, March 31–April 3, 1987, at Imperial College, London.

Karlsson:2003:QCD

[Kar03] Erik B. Karlsson. 10. Quantum coherence and decoherence seen in ultrafast neutron scattering experiments on protons in condensed matter: 6. September 2003. In Anonymous [Ano03c], page ?? ISBN ???? LCCN ????

Katscher:1990:S

[Kat90] Friedrich Katscher. Schrödinger. Nature, 344(6266):484, April 5, 1990. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v344/
n6266/pdf/344484d0.pdf.

Kerber:2011:ESD

[KBG11] Gabriele Kerber, Nándor Balázs, and Peter Graf. Erwin Schrödinger 1887–1961: Dokumente, Materialien und Bilder. Österreichischen Zentralbibliothek für Physik, Vienna, Austria, second edition, 2011. ISBN 3-900490-13-9. 166 pp. LCCN ???? An exhibition of the Österreichischen Zentralbibliothek für Physik.

Kerber:1987:ESD

[KDK87] Gabriele Kerber, Auguste Dick, and Wolfgang Kerber. Erwin Schrödinger 1887–1961: Dokumente, Materialien und Bilder. Österreichischen Zentralbibliothek für Physik, Vienna, Austria, 1987. ISBN 3-900538-09-3. 158 pp. LCCN QC16.S265 D65 1987. An exhibition of the Österreichischen Zentralbibliothek für Physik. With contributions by Nándor Balázs and others.

Kerber:1998:ESAa

[KDK^{+}98] Gabriele Kerber, Auguste Dick, Wolfgang Kerber, Brigitte Kromp, and Guido Blechl. Erwin Schrödinger (1887–1961): Eine Ausstellung der Zentralbibliothek für Physik in Wien. (German) [Erwin Schrödinger (1887–1961): an exhibit of the Central Library of Physics in Vienna]. Web document, Zentralbibliothek für Physik, Vienna, Austria, September 3, 1998. 7 pp. URL http://www.zbp.

univie.ac.at/schrodinger/biographie.pdf.

Kerber:1998:ESAb

[KDK^{+}99] Gabriele Kerber, Auguste Dick, Wolfgang Kerber, Brigitte Kromp, Guido Blechl, and Karl von Meyenn. Erwin Schrödinger (1887– 1961): Eine Ausstellung der Zentralbibliothek für Physik in Wien. (German) [Erwin Schrödinger (1887–1961): an exhibit of the Central Library of Physics in Vienna]. Web document, Zentralbibliothek für Physik, Vienna, Austria, April 13,

1999. 37 pp. URL http://www.zbp.univie.ac.at/schrodinger/
biographie.pdf;   http://www.zbp.univie.ac.at/schrodinger/
ebibliographie/publications.htm.

Kennard:1931:CS

[Ken31] E. H. Kennard. The conditions on Schrödinger’s Ψ. Nature, 127 (3215):892–893, June 13, 1931. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v127/n3215/pdf/127892c0.pdf.

Konen:1925:BGM

[KGS25] H. Konen, W. Grotrian, and E. Schrödinger. Besprechungen. (German) [Meetings]. Naturwissenschaften, 13(33):707–711, August 1925. CODEN NATWAY. ISSN 0028-1042 (print),

1432-1904 (electronic). URL http://www.springerlink.com/
content/j2766038115313m1/fulltext.pdf.

Kuhn:1967:SHQ

[KHFA67] Thomas S. Kuhn, John L. Heilbron, Paul Forman, and Lini Allen. Sources for history of quantum physics: an inventory and report, volume 68 of Memoirs of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1967. ix + 176 pp. LCCN QC174.1 .S66.

URL http://www.amphilsoc.org/guides/ahqp/;   http://
www.amphilsoc.org/guides/ahqp/s-t.htm#schrodinger.

Kibble:1987:GFT

[Kib87] T. W. B. Kibble. Gauge fields, topological defects and cosmology. In Kilmister [Kil87b], pages 201–212. ISBN 0-521-37929-6 (paperback), 0-521-34017-9. LCCN QC16.S265 S36 1987;

Q143.S3. URL http://www.loc.gov/catdir/enhancements/
fy0901/86028338-d.html;   http://www.loc.gov/catdir/enhancements/

fy0901/86028338-t.html. Papers presented at a conference, March 31–April 3, 1987, at Imperial College, London.

Kiefer:2003:DCE

[Kie03] Claus Kiefer. 6. Decoherence — concepts and examples: 6. September 2003. In Anonymous [Ano03c], page ?? ISBN ???? LCCN ????

Kilmister:1965:RBW

[Kil65] C. W. Kilmister. Review: Briefe zur Wellenmechanik, by von Schrödinger, Planck, Einstein, and Lorentz, edited by K. Przibram. Mathematical Gazette, 49(369):346–347, October 1965. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

http://www.jstor.org/stable/3612915.

Kilmister:1987:I

[Kil87a] C. W. Kilmister. Introduction. In Kilmister [Kil87b], pages 1–3. ISBN 0-521-37929-6 (paperback), 0-521-34017-9. LCCN QC16.S265 S36 1987; Q143.S3. URL http://www.loc.gov/

catdir/enhancements/fy0901/86028338-d.html;   http://www.
loc.gov/catdir/enhancements/fy0901/86028338-t.html. Pa-

pers presented at a conference, March 31–April 3, 1987, at Imperial College, London.

Kilmister:1987:SCC

[Kil87b] C. W. (Clive William) Kilmister, editor. Schrödinger: centenary celebration of a polymath. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-37929-6 (paperback), 0-521-34017-9. LCCN QC16.S265 S36 1987; Q143.S3. URL http://www.loc.gov/

catdir/enhancements/fy0901/86028338-d.html;   http://www.
loc.gov/catdir/enhancements/fy0901/86028338-t.html. Pa-

pers presented at a conference, March 31–April 3, 1987, at Imperial College, London.

Kilmister:1990:BRB

[Kil90] C. W. Kilmister. Book review: Schrödinger. Life and Thought. Walter Moore. Cambridge University Press, New York, 1989. xii, 513 pp., illus. $39.50. Science, 247(4945):985–986, February 23, 1990. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/247/4945/
985.2.full.pdf.

Kimble:1999:CQC

[Kim99] Harry J. Kimble. 14. The cavity QED circus — flying photons, juggling atoms and fantastic finesse: 5. Sept. 1998. In Greenberger et al. [GRZ99], page ?? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN Q44.4 V662 v. 7.

Kirk:1955:RBG

[Kir55] G. S. Kirk. Review: Nature and the Greeks by Erwin Schrödinger. The Classical Review (New Series), 5(3–4):260–262, December 1955. CODEN ???? ISSN 0009-840X (print), 1464-3561 (elec-

tronic). URL http://www.jstor.org/stable/704341.

Kirchmayer:1995:MES

[Kir95] Hans Kirchmayer. From magnetism to eigenvalues — scientific links Dublin–Vienna. First of the Schrödinger Lecture Series held at Trinity College, Dublin, Ireland., Decem-

ber 6, 1995. URL http://www.tcd.ie/Physics/news/seminars/
Schrodinger/Lecture1/Lecture1.html.

Kirsten:1975:PPW

[KK75] Christa Kirsten and H.-G. Körber, editors. Physiker über Physiker: Wahlvorschläge zur Aufnahme von Physikern in die Berliner Akademie 1870 bis 1929 von Hermann von Helmholtz bis Erwin Schrödinger. (German) [Physicists about Physicists: Nominations for receiving physicists in the Berlin Academy 1870–1929], volume 1 of Studien zur Geschichte der Akademie der Wissenschaften der DDR. Akademie-Verlag, Berlin, Germany, 1975. ISBN ???? 299 pp. LCCN ???? With an introduction by Hans-Jürgen Treder.

Kira:2011:QSS

[KKS^{+}11] M. Kira, S. W. Koch, R. P. Smith, A. E. Hunter, and S. T. Cundiff. Quantum spectroscopy with Schrödinger-cat states. Nature Physics, 7(10):799–804, September 18, 2011. CODEN NPAHAX. ISSN 1745-2473 (print), 1745-2481 (electronic). URL http://www.

nature.com/nphys/journal/v7/n10/full/nphys2091.html.

Klein:1965:BRQ

[Kle65] Martin J. Klein. Book review: Quantum comments: Schrödinger — Planck — Einstein — Lorentz. Briefe zur Wellenmechanik, edited by Karl Przibram. Isis, 56(1):83–84, Spring 1965. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228464.

Klein:1968:BRB

[Kle68a] Martin J. Klein. Book review: Erwin Schrödinger. An Introduction to His Writings. William T. Scott. University of Massachusetts Press, Amherst, 1967. xvi + 175 pp. $6.50. Science, 159(3818): 967–968, March 1, 1968. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/159/3818/967.2.full.pdf.

Klein:1968:RVP

[Kle68b] Martin J. Klein. Review: The views of a physicist: Erwin Schrödinger. An Introduction to His Writings, by William T. Scott. Science, 159(3818):967–968, March 1, 1968. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

jstor.org/stable/1723568.

Klein:1984:RBP

[Kle84] Martin J. Klein. Review: Collected Papers on Wave Mechanics, 3rd ed. by E. Schrödinger. American Scientist, 72(2):192, March 1984. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (elec-

tronic). URL http://www.jstor.org/stable/27852538.

Kleppner:1996:MSC

[Kle96] Daniel Kleppner. Memoirs of Schrödinger’s cat. Physics Today, 49(11):11–13, November 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:

//scitation.aip.org/content/aip/magazine/physicstoday/
article/49/11/10.1063/1.881601.

Kleinpoppen:2002:CEU

[Kle02] Hans Kleinpoppen. 23. coherence effects and ultrashort time correlations of two-photon radiation of the metastable state of atomic hydrogen: Vortrag am 13. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Klein:2005:ESF

[Kle05] Etienne Klein. Il était sept fois la révolution: Albert Einstein et les autres . . .. (French) [The seven-times revolution: Albert Einstein and the others …]. Flammarion, Paris, France, 2005. ISBN 2-08-210343-9. 237 pp. LCCN QC16.E5.

Klose:1963:QBR

[Klo63] W. Klose. Quantentheorie: [book review:] E. Schrödinger, M. Planck, A. Einstein, and H. A. Lorentz, Briefe zur Wellenmechanik, Herausgegeben von K. Przibram. Wien: Springer-Verlag 1963. 67 S. ????, ??(??):454, ???? 1963.

Knight:1992:PSC

[Kni92] Peter Knight. Practical Schrödinger’s cats. Nature, 357(6378):438– 439, June 11, 1992. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v357/n6378/pdf/357438a0.pdf.

Kocher:1978:TSC

[Koc78] M. Kocher. The tale of Schrödinger’s cat. Physics Today, 31(5):15, May 1978. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://www.physicstoday.org/resource/1/
phtoad/v31/i5/p15_s1.

Kochen:2002:GQM

[Koc02] Simon Kochen. 17. Geometry and quantum mechanics: Vortrag am 13. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Kohler:1976:WKP

[Köh76] Wolfgang Köhler. Wolfgang Köhler papers 1862–1976. American Philosophical Society Library archive, 1976. URL http:

//amphilsoc.org/mole/view?docId=ead/Mss.B.K815-ead.xml.

8.0 Linear feet, Mss.B.K815. From the Web site: “After the dismissal of Jewish professors Köhler took a public stand in April 1933 with “Gespräche in Deutschland.” [(German) [Speech in Germany]] Published in Die Deutsche Allgemeine Zeitung, this was the last anti-Nazi article published openly in Germany, and Köhler expected to be arrested soon afterwards. In 1934 Köhler made his second trip to the United States to teach as the William James Lecturer at Harvard University (1934-1935), and as Visiting Professor at the University of Chicago (1935). At the end of his time in Chicago, Köhler decided to formally emigrate to the United States”.

Kox:2013:DBH

[Kox13] Anne J. Kox. The debate between Hendrik A. Lorentz and Schrödinger on wave mechanics. In Reiter and Yngvason [RY13], pages 153–164. ISBN 3-03719-121-X (paperback), 3-03719-621-1 (e-book). LCCN QC16.S265 I58 2011.

Kernbauer:2020:POM

[KPMN20] Alois Kernbauer, Manfred Popp, Frank Ulrich Müller, and Hermann Nicolai, editors. Paul Otto Müller: Schrödingers talentierter Schüler: seine Dissertation zur Allgemeinen Relativitätstheorie, seine Mitarbeit am Uranprojekt, volume 50 of Publikationen aus dem Archiv der Universität Graz. ADEVA, Akademische Drucku. Verlagsanstalt, Graz, Austria, 2020. ISBN 3-201-02051-6. 242 pp. LCCN ????

Kragh:1982:ESW

[Kra82] Helge Kragh. Erwin Schrödinger and the wave equation: the crucial phase. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 26(2):154–197, July 1982. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kragh:1992:UQR

[Kra92] Helge Kragh. Unifying quanta and relativity? Schrödinger’s attitude to relativistic quantum mechanics. In Bitbol and Darrigol [BD92], pages 315–338. ISBN 2-86332-116-1. LCCN QC16.S265 E78 1992. 263.75F.

Kragh:2011:QSV

[Kra11] Helge Kragh. Quantenspringerei: Schrödinger vs. Bohr. Report RePOSS #14, Centre for Science Studies, University of Aarhus, Denmark. Research group: History and philosophy of science, Århus, Denmark, February 2011. URL http://css.au.dk/

fileadmin/reposs/reposs-014.pdf.

Kragh:2013:QDB

[Kra13] Helge Kragh. A quantum discontinuity: the Bohr–Schrödinger dialogue. In Reiter and Yngvason [RY13], pages 135–152. ISBN 3-03719-121-X (paperback), 3-03719-621-1 (e-book). LCCN QC16.S265 I58 2011.

Kroner:1998:RMG

[Krö98] Hans-Peter Kröner. Rezensionen: Mitchell G. Ash/Alfons Söllner (Hrsgg.): Forced Emigration and Scientific Change. Emigré German-Speaking Scientists and Scholars after 1933. Cambridge: Cambridge University Press 1995. Berichte zur Wissenschaftsgeschichte, 21(2–3):196–197, ???? 1998. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Kohlrausch:1914:WBS

[KS14] K. W. F. Kohlrausch and Erwin Schrödinger. Über die weiche (b) Sekundärstrahlung von g-Strahlen. (German) [On the soft (b) secondary radiation of g-rays]. Sitzungsberichte der kaiserlichen Akademie der Wissenschaften in Wien. Mathematischnaturwissenschaftliche Klasse, Abteilung 2a, 123(??):1319–1367, ???? 1914.

Kohlrausch:1926:EMK

[KS26] K. W. F. Kohlrausch and Erwin Schrödinger. Das Ehrenfestsche Modell der H-Kurve. (German) [The Ehrenfest model of the Hcurve]. Physikalische Zeitschrift, 27(??):306–313, ???? 1926. CO-DEN PHZTAO. ISSN 0369-982X.

Kerber:1986:ESK

[KS86] Gabriele Kerber and Ulrike Smola. Erwin Schrödinger: 1887– 1961: Katalog zur Ausstellung anläßlich der 25. Wiederkehr seines Todestages. (German) [Erwin Schrödinger: 1887–1961, exhibition catalog on the occasion of the 25th anniversary of his death]. Zentralbibliothek für Physik, 1986.

Kienle:1932:BGM

[KSP32] H. Kienle, E. Schrödinger, and Peter Pringsheim. Besprechungen. (German) [Meetings]. Naturwissenschaften, 20(10):172–174, March 1932. CODEN NATWAY. ISSN 0028-1042 (print),

1432-1904 (electronic). URL http://www.springerlink.com/
content/k41867v411218261/fulltext.pdf.

Kurizki:2003:RTD

[Kur03a] Gershon Kurizki. 12. Round table discussion: “Coherence and entanglement of complex systems: bold dreams and realistic goals”:

6. September 2003. In Anonymous [Ano03c], page ?? ISBN ???? LCCN ????

Kurizki:2003:CEC

[Kur03b] Gershon Kurizki. 3. Controlled entanglement of complex systems: 5. September 2003. In Anonymous [Ano03c], page ?? ISBN ???? LCCN ????

L:1988:SOB

[L.88] K. L. Some other books of interest. Science, 239(4840):659, February 5, 1988. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.sciencemag.org/content/
239/4840/659.1.full.pdf.

Lakhtakia:1996:MMH

[Lak96] A. (Akhlesh) Lakhtakia, editor. Models and Modelers of Hydrogen: Thales, Thomson, Rutherford, Bohr, Sommerfeld, Goudsmit, Heisenberg, Schrödinger, Dirac, Sallhofer. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1996. ISBN 981-02-2302-1. xv + 424 pp. LCCN QC171.2 .M63

1996. URL http://www.worldscientific.com/worldscibooks/
10.1142/2788.

Largeault:1990:RBM

[Lar90] Jean Largeault. Review: L’esprit et la matière by Erwin Schrödinger and Michel Bitbol. Revue Philosophique de la France et de l’Étranger, 180(4):714–716, October/December 1990. CO-DEN ???? ISSN 0035-3833 (print), 2104-385X (electronic). URL

http://www.jstor.org/stable/41096389.

Largeault:1993:RBN

[Lar93] Jean Largeault. Review: La nature et les Grecs, ≪ La clöture de la reprèsentation ≫, by Erwin Schrödinger, Michel Bitbol. Revue Philosophique de la France et de l’Étranger, 183(4):740–741, October/December 1993. CODEN ???? ISSN 0035-3833 (print),

2104-385X (electronic). URL http://www.jstor.org/stable/
41097143.

Leggett:1984:SCH

[Leg84] A. J. Leggett. Schrödinger’s cat and her laboratory cousins. Contemporary Physics, 25(6):583–598, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Leggett:1990:SCA

[Leg90] Tony Leggett. Schrödinger’s cat among the neurons. Physics World, 3(3):63, March 1990. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://physicsworldarchive.

iop.org/full/pwa-pdf/3/3/phwv3i3a29.pdf.

Leggett:2000:NLS

[Leg00] Tony Leggett. New life for Schrödinger’s cat. Physics World, 13(8): 23–24, August 2000. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.iop.
org/full/pwa-pdf/13/8/phwv13i8a25.pdf.

Leggett:2009:SCH

[Leg09] A. J. Leggett. Schrödinger’s cat and her laboratory cousins. Contemporary Physics, 50(1):243–258, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Leggett:2013:SCH

[Leg13] Anthony J. Leggett. Schrödinger’s cat and her laboratory cousins. In Reiter and Yngvason [RY13], pages 87–108. ISBN 3-03719-121-X (paperback), 3-03719-621-1 (e-book). LCCN QC16.S265 I58 2011.

Leinaas:2002:TEA

[Lei02] Jon Magne Leinaas. 18. Thermal excitations of accelerated electrons: Vortrag am 13. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Lenhard:2014:DMC

[Len14] Johannes Lenhard. Disciplines, models, and computers: the path to computational quantum chemistry. Studies in History and Philosophy of Science Part A, 48(??):89–96, December 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368114000533.

Levin:2017:SQW

[Lev17a] F. S. (Frank S.) Levin. Surfing the Quantum World. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2017. ISBN 0-19-880827-5. vi + 284 pp. LCCN QC174.12

.L46 2017. URL https://global.oup.com/academic/product/
surfing-the-quantum-world-9780198808275.

Levin:2017:TCS

[Lev17b] Frank S. Levin. On theory and consensus in science: a journey in time. Web blog., October 17, 2017. URL https://blog.oup.

com/2017/10/theory-consensus-in-science/.

Lewis:1987:DBC

[Lew87] J. T. Lewis. Do bosons condense? In Kilmister [Kil87b], pages 136–145. ISBN 0-521-37929-6 (paperback), 0-521-34017-9. LCCN QC16.S265 S36 1987; Q143.S3. URL http://www.loc.gov/

catdir/enhancements/fy0901/86028338-d.html;   http://www.
loc.gov/catdir/enhancements/fy0901/86028338-t.html. Pa-

pers presented at a conference, March 31–April 3, 1987, at Imperial College, London.

Lark-Horovitz:1953:BRB

[LH53a] Karl Lark-Horovitz. Book review: Science and Humanism: Physics in Our Time. Erwin Schrödinger . New York: Cambridge Univ. Press, 1952. 68 pp. $1.75. Science, 118(3054):59–60, July 10, 1953. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/118/3054/
59.full.pdf.

Lark-Horovitz:1953:RBH

[LH53b] Karl Lark-Horovitz. Review: Science and Humanism: Physics in Our Time by Erwin Schrödinger. Science, 118(3054):59–60, July 10, 1953. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/1680023.

Lieb:1997:FLO

[Lie97] Elliott H. Lieb. Five lectures on the occasion on Walter Thirring’s 70th birthday. 1. Stability of matter: 1997/04/10. Report, Internationales Erwin-Schrödinger-Institut für Mathematische Physik, Vienna, Austria, 1997.

Lindsay:1957:BRE

[Lin57] R. B. Lindsay. Book review: Erwin Schrödinger, Expanding Universes. Physics Today, 10(1):38, January 1957. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

link.aip.org/link/?PTO/10/38/1.

Leibfried:2005:CSA

[LKS^{+}05] D. Leibfried, E. Knill, S. Seidelin, J. Britton, R. B. Blakestad, J. Chiaverini, D. B. Hume, W. M. Itano, J. D. Jost, C. Langer, R. Ozeri, R. Reichle, and D. J. Wineland. Creation of a six-atom Schrödinger cat state. Nature, 438(7068):639–642, December 1, 2005. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v438/
n7068/full/nature04251.html.

Levy-Leblond:1999:QWQ

[LL99] Jean-Marc Lévy-Leblond. 3. Quantum words for a quantum world?: 4. Sept. 1998. In Greenberger et al. [GRZ99], page ?? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN Q44.4 V662 v. 7.

Leibfried:2015:BWE

[LMOS15] Stephan Leibfried, Christoph Markschies, Ernst Osterkamp, and Günter Stock. Berlins wilde Energien: Porträts aus der Geschichte der Leibnizschen Wissenschaftsakademie. (German) [Berlin’s Wild Energies: Portraits from the History of the Leibniz Science Academy]. De Gruyter Akademie Forschung, Berlin, Germany, 2015. ISBN 3-11-037598-2 (hardcover). 517 pp. LCCN AS182.B38 B477 2015.

Lockwood:1989:MBQ

[Loc89] Michael Lockwood. Mind, brain, and the quantum: the compound ‘I’. Blackwell, Oxford, UK, 1989. ISBN 0-631-16183-X. xii + 365 pp. LCCN BF161 .L57 1989. US$50.00 (est.).

Lockwood:1992:WSS

[Loc92] M. Lockwood. What Schrödinger should have learnt from his cat. In Bitbol and Darrigol [BD92], pages 363–382. ISBN 2-86332-116-1. LCCN QC16.S265 E78 1992. 263.75F.

Loeser:1984:TPS

[Loe84] J. G. Loeser. Three perspectives on Schrödinger’s cat. American Journal of Physics, 52(12):1089–1093, December 1984. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://scitation.aip.org/content/aapt/journal/ajp/
52/12/10.1119/1.13765. See comment [Bal86].

Lonsdale:1952:RBH

[Lon52] Kathleen Lonsdale. Review: Science and Humanism: Physics in Our Time by Erwin Schrödinger. British Journal for the Philosophy of Science, 3(11):278–279, November 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://www.

jstor.org/stable/685276.

M:1947:BRBr

[M.47] W. M. M. Book review: What Is Life? by Erwin Schroedinger. Philosophy of Science, 14(2):174, April 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185850.

Maccone:2024:SCQ

[Mac24] Lorenzo Maccone. Schrödinger cats and quantum complementarity. Foundations of Physics, 54(1):??, February 2024.

CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s10701-023-00750-6.

Mahner:2001:SRS

[Mah01] Martin Mahner, editor. Scientific Realism: Selected Essays of Mario Bunge. Prometheus Books, Amherst, NY, USA, 2001. ISBN 1-57392-892-5. 438 pp. LCCN Q175.32 .R42.

Malaterre:2012:WLI

[Mal12] Christophe Malaterre. What is life? The intellectual pertinence of Erwin Schrödinger. International Studies in the Philosophy of Science, 26(2):229–231, 2012. CODEN ???? ISSN 0269-8595 (print),

1469-9281 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/02698595.2012.703484.

Matthew:1975:ES

[Mat75] J. A. D. Matthew. Erwin Schrödinger 1887–1961. Physics Education, 10(5):357–360, July 1975. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-6552 (electronic). URL http://iopscience.

iop.org/0031-9120/10/5/305.

Mathur:2005:SM

[Mat05] Neil Mathur. Schrödinger’s mousetrap. Nature, 433(7024):363, January 26, 2005. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v433/n7024/full/433363a.html.

Montwill:2011:ESW

[MB11a] Alex Montwill and Ann Breslin. Erwin Schrödinger — wave mechanics. In The quantum adventure: does God play dice? [MB11b], chapter 10, pages 133–148. ISBN 1-84816-647-8 (hardcover), 1-84816-648-6 (paperback), 1-84816-649-4 (e-book). LCCN QC174.12 .M66 2012.

Montwill:2011:QAD

[MB11b] Alex Montwill and Ann Breslin. The quantum adventure: does God play dice? World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2011. ISBN 1-84816-647-8 (hardcover), 1-84816-648-6 (paperback), 1-84816-649-4 (e-book). ix + 248 pp. LCCN QC174.12 .M66 2012.

McCrea:1951:RBT

[McC51] W. H. McCrea. Review: Space–time Structure by Erwin Schrödinger. Mathematical Gazette, 35(313):205–207, September 1951. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL http://www.jstor.org/stable/3611263.

McCrea:1956:RBU

[McC56] W. H. McCrea. Review: Expanding Universes by Erwin Schrödinger. Publications of the Astronomical Society of the Pacific, 68(403):373–374, August 1956. CODEN PASPAU. ISSN 0004-6280 (print), 1538-3873 (electronic). URL http://www.

jstor.org/stable/40673095.

McConnell:1987:SNO

[McC87a] James McConnell. Schrödinger’s nonlinear optics. In Kilmister [Kil87b], pages 146–164. ISBN 0-521-37929-6 (paperback), 0-521-34017-9. LCCN QC16.S265 S36 1987; Q143.S3. URL http://www.

loc.gov/catdir/enhancements/fy0901/86028338-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0901/86028338-t.html.

Papers presented at a conference, March 31–April 3, 1987, at Imperial College, London.

McCrea:1987:EVE

[McC87b] Sir William McCrea. Eamon de Valera, Erwin Schrödinger and the Dublin Institute. In Kilmister [Kil87b], pages 119–

135. ISBN 0-521-37929-6 (paperback), 0-521-34017-9. LCCN QC16.S265 S36 1987; Q143.S3. URL http://www.loc.gov/

catdir/enhancements/fy0901/86028338-d.html;   http://www.
loc.gov/catdir/enhancements/fy0901/86028338-t.html. Pa-

pers presented at a conference, March 31–April 3, 1987, at Imperial College, London.

McGregor:1999:SB

[McG99] H. R. McGregor. Schrödinger’s baby. W. Morrow and Co., New York, NY, USA, 1999. ISBN 0-688-16893-0. 249 pp. LCCN PR6063.C488 S38 1999.

McLay:1999:LME

[McL99] David B. McLay. Lise Meitner and Erwin Schrödinger: Biographies of two Austrian physicists of Nobel stature. Minerva, 37 (1):75–94, ???? 1999. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004536223625.

Mehra:1975:SCP

[Meh75] Jagdish Mehra, editor. The Solvay conferences on physics: aspects of the development of physics since 1911. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1975. ISBN 90-277-0635-2. xxxii + 415 pp. LCCN QC1.S792 M43.

Mehra:1987:ESRa

[Meh87a] Jagdish Mehra. Erwin Schrödinger and the rise of wave mechanics. I. Schrödinger’s scientific work before the creation of wave mechanics. Foundations of Physics, 17(11):1051–1112, November 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/
10.1007/BF01889803.

Mehra:1987:ESRb

[Meh87b] Jagdish Mehra. Erwin Schrödinger and the rise of wave mechanics. II. The creation of wave mechanics. Foundations of Physics, 17(12): 1141–1188, December 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889592.

Mehra:1987:NBD

[Meh87c] Jagdish Mehra. Niels Bohr’s discussions with Albert Einstein, Werner Heisenberg, and Erwin Schrödinger: the origins of the Principles of Uncertainty and Complementarity. Foundations of Physics, 17(5):461–506, May 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Mehra:1988:ESR

[Meh88] Jagdish Mehra. Erwin Schrödinger and the rise of wave mechanics. III. Early response and applications. Foundations of Physics, 18 (2):107–184, February 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://www.springerlink.

com/content/p144515610k37j02/. Part IV. Invited Papers Com-

memorating the Centenary of the Birth of Erwin Schrödinger, 12 August 1887.

Melgar:1988:PES

[Mel88] Miguel Ferrero Melgar. The philosophy of Erwin Schrödinger: A diachronic view of Schrödinger’s thoughts. Foundations of Physics, 18(3):357–371, March 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01891496.

Merton:1938:SSO

[Mer38] Robert K. Merton. Science and the social order. Philosophy of Science, 5(3):321–337, July 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.

org/stable/184838.

Merali:2008:RCS

[Mer08] Zeeya Merali. Reincarnation can save Schrödinger’s cat. Nature, 454(7200):8–9, July 2, 2008. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

news/2008/080702/full/454008a.html.

Moore:1991:BRS

[MF91] Walter Moore and Peter G. O. Freund. Book review: Schrödinger: Life and Thought. Physics Today, 44(2):92–??, February 1991. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Michels:1968:BRB

[Mic68] Walter C. Michels. Book review: Letters on Wave Mechanics; Albert Einstein, Erwin Schroedinger, Max Planck, H. A. Lorentz, translated by Martin J. Klein from a collection edited by K. Przibram. American Journal of Physics, 36(8):766–767, August 1968. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://link.aip.org/link/?AJP/36/766/2.

Michor:1998:FJG

[Mic98a] Peter W. Michor. Fünf Jahre 1993–1998. (German) [Five years 1993–1998]. Report, Internationales Erwin-Schrödinger-Institut für Mathematische Physik, Vienna, Austria, 1998.

Michor:1998:SRY

[Mic98b] Peter W. Michor. Scientific report for the 5 year period 1993–1997 including the prehistory 1991–1992. Report, Internationales Erwin-Schrödinger-Institut für Mathematische Physik, Vienna, Austria, 1998.

Michor:2003:SRY

[Mic03] Peter W. Michor. Scientific report for the years 1993–2002. Report, ESI, Erwin Schrödinger International Institute for Mathematical Physics, Vienna, Austria, 2003.

Mikhailov:1994:NSE

[Mik94] A. V. Mikhailov, editor. The nonlinear Schrödinger equation: proceedings of the Conference on The Nonlinear Schrödinger Equation; Chernogolovka, 25 July–3 August 1994, volume 87(1–4) of Physica D. North-Holland, Amsterdam, The Netherlands, 1994. ISBN ???? LCCN ????

Milburn:1997:SMQ

[Mil97] G. J. (Gerard J.) Milburn. Schrödinger’s machines: the quantum technology reshaping everyday life. W. H. Freeman, New York, NY, USA, 1997. ISBN 0-7167-3106-1. xvi + 188 pp. LCCN QC174.13 .M55 1997. URL http://

www.loc.gov/catdir/bios/hol057/96051024.html;   http:/
/www.loc.gov/catdir/description/hol055/96051024.html.

Miller:2003:EAS

[Mil03] Arthur I. Miller. Erotica, aesthetics and Schrödinger’s wave equation. In Farmelo [Far03], pages 110–131. ISBN 1-86207-555-7. LCCN Q125 .I88 2003.

Miquel:2007:QCQ

[Miq07] Paul-Antoine Miquel. Qu’est-ce que la vie ?. (French) [What is Life?]. J. Vrin, Paris, France, 2007. ISBN 2-7116-1780-7. 128 pp. LCCN ????

Monroe:1996:SCS

[MMKW96] C. Monroe, D. M. Meekhof, B. E. King, and D. J. Wineland. A Schrödinger cat superposition state of an atom. Science, 272(5265): 1131–1136, May 24, 1996. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/272/5265/1131.full.pdf.

Murphy:1995:WLN

[MO95] Michael P. (Michael Patrick) Murphy and Luke A. J. O’Neill, editors. What is life?: the next fifty years: speculations on the future of biology. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-45509-X (hardback). LCCN QH331 .W465 1995. URL http:

//www.loc.gov/catdir/description/cam026/94049438.html;
 http://www.loc.gov/catdir/toc/cam026/94049438.html.

Monteiro:2009:SD

[Mon09] Luiz Henrique Alves Monteiro. Schrödinger’s dog. Physics World, 22(9):52, September 2009. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://physicsworldarchive.

iop.org/full/pwa-pdf/22/09/phwv22i09a47.pdf.

Moore:1987:SEL

[Moo87] Walter Moore. Schrödinger’s entropy and living organisms. Nature, 327(6123):561, June 18, 1987. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v327/n6123/pdf/327561a0.pdf.

Moore:1989:SLT

[Moo89] Walter John Moore. Schrödinger: life and thought. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-35434-X. xi + 513 pp. LCCN QC16.S265 M66 1989. URL http://www.loc.

gov/catdir/description/cam023/88025807.html;   http://
www.loc.gov/catdir/toc/cam021/88025807.html.

Moore:1990:BRBa

[Moo90] Walter Moore. Book review: Janossy Lajos (1912–1978) es Erwin Schrödinger (1887–1961) levelse by Peter Kirdly; Maria Narayne Ziegler. Isis, 81(1):132–133, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/234134.

Moore:1994:LES

[Moo94] Walter John Moore. A life of Erwin Schrödinger. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-46934-1 (paperback). xiii + 349 pp. LCCN QC16.S265 M65 1994. URL http:

//www.loc.gov/catdir/description/cam026/93050147.html;
 http://www.loc.gov/catdir/toc/cam028/93050147.html.

Moore:1996:ESV

[Moo96] Walter John Moore. Erwin Schrödinger: una vida. (Spanish) [Erwin Schrödinger: a Life]. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-55593-0 (paperback). ix + 444 pp. LCCN QC16.S265 M6518 1996. URL http://www.loc.gov/

catdir/description/cam027/95051783.html;   http://www.
loc.gov/catdir/toc/cam023/95051783.html. Spanish transla-

tion of [Moo94].

Moore:2012:ESB

[Moo12] Walter J. Moore. Erwin Schrödinger: eine Biographie. Primus Verlag, Darmstadt, Germany, 2012. ISBN 3-86312-301-8. 423 pp. LCCN QC16.S265.

Moore:2015:ESBb

[Moo15] Walter Moore. Erwin Schrödinger: Eine Biographie. Konrad Theiss Verlag, Darmstadt, Germany, 2015. ISBN 3-8062-3189-3, 3-8062-3191-5, 3-8062-0045-9 (e-book). 554 pp. LCCN QC173.98.

Morse:1996:SCH

[Mor96] Randy Morse. Schrödinger’s cat at hand. Science, 273(5277):860, August 16, 1996. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.sciencemag.org/content/
273/5277/860.1.full.pdf.

Mehra:1982:QTPa

[MR82a] Jagdish Mehra and Helmut Rechenberg. The quantum theory of Planck, Einstein, Bohr and Sommerfeld: its foundation and the rise of its difficulties 1900–1925, volume 1 (part 1) of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 3-540-90642-8 (Berlin), 0-387-90642-8 (New York). 372 pp. LCCN QC 174.12. See other volumes in this series [?, ?, ?, MR82b, ?, ?].

Mehra:1982:QTPb

[MR82b] Jagdish Mehra and Helmut Rechenberg. The quantum theory of Planck, Einstein, Bohr and Sommerfeld: its foundation and the rise of its difficulties 1900–1925, volume 1 (part 2) of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90667-3 (New York), 3-540-90667-3 (Berlin). 375– 878 pp. LCCN QC 174.12. See other volumes in this series [?, ?, ?, MR82a, ?, ?].

Mehra:1987:ESRc

[MR87a] Jagdish Mehra and Helmut Rechenberg. Erwin Schrödinger and the rise of wave mechanics. Part 1, Schrödinger in Vienna and Zurich 1887–1925, volume 5, part 1 of The Historical Development of Quantum Theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-96284-0 (vol. 1, New York), 3-540-96284-0 (vol. 1, Berlin), 0-387-96377-4 (vol. 2, New York), 3-387-96377-4 (vol. 2, Berlin). xx + 366 pp. LCCN QC173.98 .M44 vol. 5. US$48.00, US$49.50. See other volumes in this series [?, ?, ?, MR82a, MR82b, ?].

Mehra:1987:ESRd

[MR87b] Jagdish Mehra and Helmut Rechenberg. The Historical Development of Quantum Theory. Volume 5. Erwin Schrödinger and the rise of wave mechanics. Part 1. Schrödinger in Vienna and Zürich 1887–1925, volume 5, part 1. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-95179-2 (paperback). xviii + 366 pp. LCCN ????

Mehra:2001:HDQ

[MR01] Jagdish Mehra and Helmut Rechenberg. The historical development of quantum theory. Volume 5. Erwin Schrödinger and the rise of wave mechanics. Part 2: The creation of wave mechanics. Early response and applications, 1925–1926. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 0-387-95180-6. vii + 368–980 pp. LCCN ???? URL

http://www.gbv.de/dms/bowker/toc/9780387951805.pdf.

McConnell:1943:SEP

[MS44] James McConnell and E. Schrödinger. The shielding effect of planetary magnetic fields. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 49(??):259–273, ???? 1943–1944. CODEN PRIAAK. ISSN 0035-8975. URL http://

www.jstor.org/stable/20488462.

Mautner:1944:IAC

[MS45] Friedrich Mautner and Erwin Schrödinger. Infinitesimal affine connections with twofold Einstein–Bargmann symmetry. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 50(??):223–231, ???? 1944–1945. CODEN PRI-

AAK. ISSN 0035-8975. URL http://www.jstor.org/stable/
20520643.

Marshall:1988:SOR

[MS88] Trevor Marshall and Emilio Santos. Stochastic optics: a reaffirmation of the wave nature of light. Foundations of Physics, 18 (2):185–223, February 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://www.springerlink.

com/content/r342562317734r75/. Part IV. Invited Papers Com-

memorating the Centenary of the Birth of Erwin Schrödinger, 12 August 1887.

Muller:1940:BVU

[Mül40] Paul Otto Müller. Eine Bedingung für die Verwendbarkeit von Uran als Sprengstoff. (German) [A condition for the usability of uranium as an explosive]. Report FA-002-482, Deutsches Museum Archiv, Berlin, Germany, May 1940.

McCartney:2003:PIP

[MW03] Mark McCartney and Andrew Whitaker, editors. Physicists of Ireland: passion and precision. Institute of Physics Publishers, Bristol and Philadelphia, 2003. ISBN 0-7503-0866-4 (hardback). xiv + 298 pp. LCCN QC15 .P47 2003. URL http://www.loc.gov/

catdir/enhancements/fy0806/2007298520-d.html;   http:/
/www.loc.gov/catdir/toc/fy0802/2007298520.html.

Muller:2013:DOS

[MZ13] Markus Müller and Peter Zoller. Digital and open system quantum simulation with trapped ions. In Reiter and Yngvason [RY13], pages 109–122. ISBN 3-03719-121-X (paperback), 3-03719-621-1 (e-book). LCCN QC16.S265 I58 2011.

Marshall:1997:WAS

[MZP97] I. N. Marshall, Danah Zohar, and F. David Peat. Who’s afraid of Schrödinger’s cat?: all the new science ideas you need to keep up with the new thinking. Morrow, New York, NY, USA, 1997. ISBN 0-688-11865-8. xxx + 402 pp. LCCN Q158.5 .M356 1997.

N:1946:RBL

[N.46] E. N. Review: What is Life? by Erwin Schrödinger. The Journal of Philosophy, 43(7):194–195, March 28, 1946. CODEN ???? ISSN 0022-362X (print), 1939-8549 (electronic). URL http://

www.jstor.org/stable/2019342.

Nagel:1952:RBH

[Nag52] Ernest Nagel. Review: Science and Humanism: Physics in Our Time by Erwin Schrödinger. The Journal of Philosophy, 49(16): 539–540, July 31, 1952. CODEN ???? ISSN 0022-362X (print),

1939-8549 (electronic). URL http://www.jstor.org/stable/
2021118.

Narnhofer:2003:YES

[Nar03] Heide Narnhofer. 2. Non-commutative ergodic theory: Vortrag am 13. März 2003. In Anonymous [Ano03a], page ???? ISBN ???? LCCN ????

Nef:1992:PCE

[Nef92] F. Nef. A propos d’une controverse entre Carnap et Schrödinger. (French) [On a controversy between Carnap and Schrödinger]. In Bitbol and Darrigol [BD92], pages 151–160. ISBN 2-86332-116-1. LCCN QC16.S265 E78 1992. 263.75F.

Neuser:1996:QGQ

[Neu96] Wolfgang Neuser, editor. Quantenphilosophie. (German) [Quantum Philosophy]. Verständliche Forschung. Spektrum Akademischer Verlag, Heidelberg, Germany; Berlin, Germany; Oxford, UK, 1996. ISBN 3-86025-366-2. 206 pp. LCCN ???? Foreword by Carl Friedrich von Weizsäcker and introduction by Wolfgang Neuser.

Niaz:2012:RWP

[NM12] Mansoor Niaz and Cecilia Marcano. Reconstruction of Waveparticle Duality and Its Implications for General Chemistry Textbooks. Springer briefs in education. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 94-007-4395-5 (paperback). viii + 46 pp. LCCN QC476.W38

N53 2012. URL http://www.springerlink.com/content/978-
94-007-4395-3.

Oeser:1999:EPM

[Oes99] Erhard Oeser. 21. Epistemological problems of measurement of quantum mechanics and the appearance of the classical world of macroscopic objects: 6. Sept. 1998. In Greenberger et al. [GRZ99], page ?? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN Q44.4 V662 v. 7.

Oberkofler:1992:ESB

[OGT92] Gerhard Oberkofler, Peter Goller, and Hans Thirring. Erwin Schrödinger: Briefe und Dokumente aus Zürich, Wien und Innsbruck. (German) [Erwin Schrödinger: Letters and Documents from Zürich, Wien, and Innsbruck]. Die Bibliothek, Innsbruck, Austria, 1992. 55 pp. LCCN ????

Ourjoumtsev:2007:GOS

[OJTBG07] Alexei Ourjoumtsev, Hyunseok Jeong, Rosa Tualle-Brouri, and Philippe Grangier. Generation of optical Schrödinger cats from photon number states. Nature, 448(7155):784–786, August 16, 2007. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v448/
n7155/full/nature06054.html.

Olby:1971:SPW

[Olb71] Robert Olby. Schrödinger’s problem: What is life? Journal of the History of Biology, 4(1):119–148, Spring 1971. CODEN JH-BIA9. ISSN 0022-5010 (print), 1573-0387 (electronic). URL http:

//link.springer.com/article/10.1007/BF00356979.

Omnes:1999:DIS

[Omn99] Roland Omnes. 9. Decoherence and irreversibility: 4. Sept. 1998. In Greenberger et al. [GRZ99], page ?? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN Q44.4 V662 v. 7.

Oppenheimer:1959:RBM

[Opp59] Jane Oppenheimer. Review: Mind and Matter by Erwin Schrödinger. American Scientist, 47(1):58A, ???? 1959. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL

http://www.jstor.org/stable/27827260.

Orlov:1999:QCG

[Orl99] Yuri Orlov. 18. Quantum and classical Gödelian indeterminism, measurement, and informational collapse into the past: 5. Sept. 1998. In Greenberger et al. [GRZ99], page ?? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN Q44.4 V662 v. 7.

Oskolkov:2003:SEO

[Osk03] Konstantin Oskolkov. 5. Schrödinger equation and oscillatory Hilbert transforms: Vortrag am 14.März 2003. In Anonymous [Ano03a], page ?? ISBN ???? LCCN ????

Ourjoumtsev:2006:GOS

[OTBLG06] Alexei Ourjoumtsev, Rosa Tualle-Brouri, Julien Laurat, and Philippe Grangier. Generating optical Schrödinger kittens for quantum information processing. Science, 312(5770):83–86, April 7, 2006. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/312/
5770/83.full.pdf.

Pagels:1982:CCQ

[Pag82] Heinz R. Pagels. The Cosmic Code: Quantum Physics as the Language of Nature. Simon and Schuster, New York, NY, USA, 1982. ISBN 0-671-24802-2. 370 + 8 pp. LCCN QC174.13 .P33 1982.

Park:1988:TAS

[Par88] James L. Park. Thermodynamic aspects of Schrödinger’s probability relations. Foundations of Physics, 18(2):225–244, February 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://www.springerlink.com/content/

v2666233mgk4t133/. Part IV. Invited Papers Commemorating the Centenary of the Birth of Erwin Schrödinger, 12 August 1887.

Pauling:1987:SCC

[Pau87] Linus Pauling. Schrödinger’s contributions to chemistry and biology. In Kilmister [Kil87b], pages 225–233. ISBN 0-521-37929-6 (paperback), 0-521-34017-9. LCCN QC16.S265 S36 1987;

Q143.S3. URL http://www.loc.gov/catdir/enhancements/
fy0901/86028338-d.html;   http://www.loc.gov/catdir/enhancements/

fy0901/86028338-t.html. Papers presented at a conference, March 31–April 3, 1987, at Imperial College, London.

Pawlak:2012:BSB

[Paw12] Alexander Pawlak. Erwin Schrödinger — Eine Biographie von Walter J. Moore. Physik Journal, ??(??):??, July 27, 2012. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (elec-

tronic). URL http://www.pro-physik.de/details/rezension/
2116281/Erwin_Schroedinger__Eine_Biographie.html. See

[Moo12].

Penrose:2002:QSR

[Pen02] Roger Penrose. 30. Quantum state reduction, gravitation and quanglement. Closing: Vortrag am 14. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Perlman:1964:ESH

[Per64] H. S. Perlman. Equivalence of the Schrödinger and Heisenberg pictures. Nature, 204(4960):771–772, November 21, 1964. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v204/
n4960/pdf/204771b0.pdf.

Perutz:1987:ESB

[Per87] M. F. Perutz. Erwin Schrödinger’s What is Life? and molecular biology. In Kilmister [Kil87b], pages 234–251. ISBN 0-521-37929-6 (paperback), 0-521-34017-9. LCCN QC16.S265 S36 1987;

Q143.S3. URL http://www.loc.gov/catdir/enhancements/
fy0901/86028338-d.html;   http://www.loc.gov/catdir/enhancements/

fy0901/86028338-t.html. Papers presented at a conference, March 31–April 3, 1987, at Imperial College, London.

Percival:1991:SQC

[Per91] Ian Percival. Schrödinger’s quantum cat. Nature, 351(6325): 357, May 30, 1991. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v351/n6325/pdf/351357a0.pdf.

Percival:2002:SUA

[Per02] Ian Percival. 28. Speakable and unspeakable after John Bell: Vortrag am 14. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Perez:2020:ESS

[Pér20] Enric Pérez. Erwin Schrödinger in the Second Spanish Republic, 1934–1935. In Forstner and Walker [FW20], chapter 4, pages 59–73. ISBN 3-030-48508-0 (print), 3-030-48509-9 (e-book), 3-030-48510-2, 3-030-48511-0. LCCN QC16.E5 .B564 2020. URL http://www.

springerlink.com/content/978-3-030-48509-2.

Pickover:2008:AHL

[Pic08] Clifford A. Pickover. Archimedes to Hawking: laws of science and the great minds behind them. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-533611-9 (hardcover). ix + 514 pp. LCCN Q175.32.R45 P53 2008. URL http:

//catdir.loc.gov/catdir/toc/ecip087/2007051167.html;
http://www.gbv.de/dms/goettingen/555633241.pdf;   http:
//www.oup.com/us/catalog/general/subject/Physics/?view=
usa&ci=9780195336115.

Pietschmann:1997:FLO

[Pie97] Herbert Pietschmann. Five lectures on the occasion on Walter Thirring’s 70th birthday. 5. contributions of scientists to music: 1997/04/12. Report, Internationales Erwin-Schrödinger-Institut für Mathematische Physik, Vienna, Austria, 1997.

Piermarocchi:2011:MSW

[Pie11] Carlo Piermarocchi. Materials spectroscopy: What would Schrödinger’s cat see? Nature Physics, 7(10):746, September 18, 2011. CODEN NPAHAX. ISSN 1745-2473 (print), 1745-2481 (elec-

tronic). URL http://www.nature.com/nphys/journal/v7/n10/
full/nphys2107.html.

Planck:1928:AWE

[Pla28] Max Planck. Über die Abhandlungen zur Wellenmechanik von Erwin Schrödinger. (German) [On the essays on wave mechanics by Erwin Schrödinger]. Deutsche Literaturzeitung, ??(??):59–62, ???? 1928. ISSN 0012-043X.

Planck:1929:AHS

[Pla29] Max Planck. Antrittsrede des Herrn Schrödinger, Erwiderung der Sekretars Herrn Planck. Offentliche Sitzung vom. 4 Juli 1929. (German) [Inaugural speech of Mr. Schrödinger, reply from Secretary Mr. Planck. Public meeting on 4 July 1929]. Sitzungsberichte der Preussischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):100–109, ???? 1929. CODEN SPW-PAI. ISSN 0371-2435.

Plotnitsky:2010:EPB

[Plo10] Arkady Plotnitsky. Epistemology and Probability: Bohr, Heisenberg, Schrödinger, and the Nature of Quantum-Theoretical Thinking, volume 161 of Fundamental theories of physics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 0-387-85333-2, 0-387-85334-0. xxxiii + 402 pp. LCCN QC173.98 .P56 2010. EUR 123.00, SFR 191. URL http://www.

gbv.de/dms/bowker/toc/9780387853338.pdf.

Polkinghorne:1959:BRE

[Pol59] J. C. Polkinghorne. Book review: Erwin Schrödinger, Mind and Matter. Physics Today, 12(1):40, January 1959. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/?PTO/12/40/2.

Polizzi:1996:BRB

[Pol96] Gaspare Polizzi. Book review: Bruno Bertotti — Umberto Curi (a cura di), Erwin Schrödinger scienziato e filosofo, Padova, Il poligrafo 1994, 198 pp. (≪Percorsi della scienza storia testi problemi≫, 3). Nuncius, 11(2):750–755, ???? 1996. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic).

URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00330.

Popper:1989:PS

[Pop89] Karl Popper. Popper on Schrödinger. Nature, 342(6248):337, November 23, 1989. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v342/n6248/pdf/342337a0.pdf.

Post:1962:RBT

[Pos62] H. R. Post. Review: Science Theory and Man by Erwin Schrödinger. British Journal for the Philosophy of Science, 13(49): 80–82, May 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-

3537 (electronic). URL http://www.jstor.org/stable/685233.

Prasad:1993:RRR

[Pra93] Rajendra Prasad. Review: Re-reading classics: A Mathematician’s Apology by G. H. Hardy; What Little I Remember by Otto Frisch; What Is Life?, with Mind and Matter and Autobiographical Sketches by Erwin Schrödinger; Mr. Tompkins in Paperback by George Gamow. Social Scientist, 21(3–4):98–101, April 1993.

ISSN 0970-0293. URL http://www.jstor.org/stable/3517634.

Przibram:1963:SPE

[Prz63] K. Przibram, editor. Schrödinger–Planck–Einstein–Lorentz: Briefe zur Wellenmechanik. (German) [Schrödinger–Planck– Einstein–Lorentz: Letters on Wave Mechanics]. Herausgegeben im Auftrage der österreichischen Akademie der Wissenschaften von K. Przibram. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1963. vi + 68 pp.

Przibram:1967:LWM

[Prz67] Karl Przibram. Letters on Wave Mechanics: Schrödinger, Planck, Einstein, Lorentz. Philosophical Library, New York, NY, USA, 1967. ISBN 0-85478-271-0. xv + 75 pp. LCCN QC174.2 .P73 1968. Translation by Martin J. Klein of the German original [Prz63].

Przibram:1986:LWM

[Prz86] Karl Przibram, editor. Letters on wave mechanics: Albert Einstein, Erwin Schrödinger, Max Planck and Hendrik A. Lorentz. Philosophical library, New York, NY, USA, 1986. ISBN ???? ???? pp. LCCN ????

Przibram:2011:AEL

[Prz11] Karl Przibram, editor. Albert Einstein: Letters on wave mechanics: correspondence with H. A. Lorentz, Max Planck, and Erwin Schrödinger. Open Road Integrated Media, New York, NY, USA, 2011. ISBN 1-4532-0468-7. xv + 101 pp. LCCN QC174.2.

Planck:1933:ZWN

[PS33] Max Planck and Erwin Schrödinger. Zagadnienia wspólczesnej nauki: indeterminizm, wplyw srodowiska na nauki przyrodnicze. (Polish) [Issues of modern science: indeterminism, environment and science natural]. Nakladem “Mathesis Polskiej”, Warszawa, Poland, 1933. ???? pp. LCCN ????

Papapetrou:1951:PCN

[PS51] A. Papapetrou and Erwin Schrödinger. The point-charge in the non-symmetric field theory. Nature, 168(4262):40–41, July 7, 1951. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v168/
n4262/pdf/168040a0.pdf.

Przibram:1963:BWG

[PS63] Karl Przibram and Erwin Schrödinger. Briefe zur Wellenmechanik. (German) [Letters on Quantum Mechanics]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1963. 67 pp. LCCN QC174.2 .P7.

Purrington:2018:HAC

[Pur18] Robert D. Purrington. The Heroic Age: the Creation of Quantum Mechanics, 1925–1940. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2018. ISBN 0-19-065517-8 (hardcover). xv + 403 pp. LCCN QC173.98 .P87 2018.

Pauling:1935:IQM

[PW35] Linus Pauling and E. Bright Wilson, Jr. Introduction to Quantum Mechanics: With Applications to Chemistry. McGraw-Hill, New York, NY, USA, 1935. xi + 468 pp. LCCN QC174.1 .P34. Reprinted by Dover, 1985.

Yin:2017:BQM

[qYL17] Zhang qi Yin and Tongcang Li. Bringing quantum mechanics to life: from Schrödinger’s cat to Schrödinger’s microbe. Contemporary Physics, 58(2):119–139, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rauscher:2011:OMP

[RA11] Elizabeth A. Rauscher and Richard L. Amoroso. Orbiting the moons of Pluto: complex solutions to the Einstein, Maxwell, Schrödinger, and Dirac equations, volume 45 of Series on knots and everything. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2011. ISBN 981-4324-24-8 (hardcover), 981-4324-25-6 (e-book). xvii + 391 pp. LCCN QC173.55 .R34 2011.

Raimond:2003:SWE

[Rai03] Jean-Michel Raimond. 13. Atom/mesoscopic field entanglement: a new insight into Rabi oscillation: 7. September 2003. In Anonymous [Ano03c], page ?? ISBN ???? LCCN ????

Rajaraman:2002:FCV

[Raj02] Ramamurti Rajaraman. 19. Fractional charge: Vortrag am 13. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Rauch:1999:QPW

[Rau99] Helmut Rauch. 16. Quantum physics at work: 5. Sept. 1998. In Greenberger et al. [GRZ99], page ?? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN Q44.4 V662 v. 7.

Rauch:2002:TMQ

[Rau02] Helmut Rauch. 13. Towards more quantum complete neutron experiments: Vortrag am 11. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Redei:2005:JNS

[Réd05] Miklós Rédei. John von Neumann selected letters, volume 27 of History of mathematics. American Mathematical Society, Providence, RI, USA, 2005. ISBN 0-8218-3776-1 (hardcover). ISSN 0899-2428. xxv + 301 pp. LCCN QA29.V66 A4 2005.

Reichenbach:1944:DSE

[Rei44a] Hans Reichenbach. Derivation of Schrödinger’s equation from de Broglie’s principle. In Philosophic Foundations of Quantum Mechanics [Rei44b], chapter 13, pages 66–71. LCCN QC174.1 .R4.

Reichenbach:1944:PFQ

[Rei44b] Hans Reichenbach. Philosophic Foundations of Quantum Mechanics. University of California Press, Berkeley, CA, USA, 1944. x + 182 pp. LCCN QC174.1 .R4.

Reichenbach:1998:DSE

[Rei98a] Hans Reichenbach. Derivation of Schrödinger’s equation from de Broglie’s principle. In Philosophic foundations of quantum mechanics [Rei98b], chapter 13, pages 66–71. ISBN 0-486-40459-5 (paperback). LCCN QC174.12 .R45 1998. URL http://www.loc.

gov/catdir/description/dover032/98040052.html.

Reichenbach:1998:PFQ

[Rei98b] Hans Reichenbach. Philosophic foundations of quantum mechanics. Dover, New York, NY, USA, 1998. ISBN 0-486-40459-5 (paperback). x + 182 pp. LCCN QC174.12 .R45 1998. URL http:/

/www.loc.gov/catdir/description/dover032/98040052.html.

Renn:2013:SGW

[Ren13] Jürgen Renn. Schrödinger and the genesis of wave mechanics. In Reiter and Yngvason [RY13], pages 9–36. ISBN 3-03719-121-X (paperback), 3-03719-621-1 (e-book). LCCN QC16.S265 I58 2011.

Raman:1969:WWI

[RF69] V. V. Raman and Paul Forman. Why was it Schrödinger who developed de Broglie’s ideas? Historical Studies in the Physical Sciences, 1(??):291–314, ???? 1969. CODEN HSPSAS. ISSN 0073-

2672. URL http://www.jstor.org/stable/27757299.

Rechenberg:1991:BBL

[RFS^{+}91] H. Rechenberg, A. Förster, W. Stamm, D. Stauffer, H. Rauch, W. Weppner, H. W. Schrötter, and W. Zinth. Buchbesprechung: Broglie, Lochak: Heisenberg’s Uncertainties and the Probabilistic Interpretation of Wave Mechanics/Spatschek: Theoretische Plasmaphysik Eine Einführung/Gaisser: Cosmic Rays and Particle Physics/Hüfner: Nur ein Störfall /Weisbuch: Complex Systems Dynamics/Audretsch u. Mainzer: Wieviele Leben hat Schrödingers Katze?/Herder Lexikon Physik/Kudo u. Fueki: Solid State Ionics/Kovetz: The Principles of Electromagnetic Theory/Brückner, Feller u. Grummt: Applications of Time-Resolved Optical Spectroscopy. Physikalische Blätter, 47(9):857–859, September 1991.

CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
phbl.19910470915/abstract.

RauschvonTraubenberg:1932:VSP

[RGS32] H. Rausch von Traubenberg, R. Gebauer, and E. Schrödinger. Über das Verhalten des Starkeffekts bei plötzlichen Feldänderungen (Experimenteller Teil von H. Rausch von Traubenberg und R. Gebauer; theoretischer Teil von E. Schrödinger). (German) [On the behavior of the Stark effect in sudden field changes (experimental part: H. Rausch Traubenberg and R. Gebauer, theoretical part: E. Schrödinger)]. Zeitschrift für Physik, 78(5–6):309– 317, May 1932. CODEN ZEPYAA. ISSN 0939-7922 (print),

1431-5831 (electronic). URL http://www.springerlink.com/
content/l14875ux1546551x/fulltext.pdf.

Rigden:2002:HEE

[Rig02] John S. Rigden. Hydrogen: the essential element. Harvard University Press, Cambridge, MA, USA, 2002. ISBN 0-674-00738-7. 280 pp. LCCN QD181.H1 R54 2002. URL http://catdir.loc.gov/

catdir/toc/fy022/2001051708.html.

Robertson:1956:RBU

[Rob56] H. P. Robertson. Review: Expanding Universes by Erwin Schrödinger. American Scientist, 44(4):300A, ???? 1956. CO-DEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic).

URL http://www.jstor.org/stable/27826853.

Rodgers:2019:TAS

[Rod19] Glen E. Rodgers. Travelling with the Atom: a Scientific Guide to Europe and Beyond. Royal Society of Chemistry, Cambridge, UK, 2019. ISBN 1-78801-528-2 (paperback), 1-78801-702-1 (e-book). xxxii + 551 pp. LCCN QC171.2 .R63 2020.

Rogers:2010:MIS

[Rog10] Kara Rogers, editor. The 100 Most Influential Scientists of All Time. The Britannica guide to the world’s most influential people. Britannica Educational Publishers, in association with Rosen Educational Services, New York, NY, USA, 2010. ISBN 1-61530-002-3 (library binding). 360 pp. LCCN Q162 .A15 2010.

Rowlands:2009:LC

[Row09] Peter Rowlands. The Liverpool cyclotrons. IOP History of Physics Newsletter, 25:31–46, February 2009. ISSN 1756-

168X. URL http://www.iop.org/activity/groups/subject/
hp/newsletter/archive/file_64509.pdf.

Ruger:1988:ACE

[Rüg88] Alexander Rüger. Atomism from cosmology: Erwin Schrödinger’s work on wave mechanics and space–time structure. Historical Studies in the Physical and Biological Sciences, 18(2):377–401, ???? 1988. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.jstor.org/stable/27757607.

Rapport:1964:P

[RW64] Samuel Rapport and Helen Wright, editors. Physics. New York University Press, New York City, NY, USA, 1964. xiii + 333 pp.

Renneberg:1994:STN

[RW94] Monika Renneberg and Mark Walker, editors. Science, Technology, and National Socialism. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-40374-X (hardcover), 0-521-52860-7 (paperback). xix + 422 pp. LCCN Q127.G3 S36 1994. URL http://

www.loc.gov/catdir/description/cam025/92041633.html;
http://www.loc.gov/catdir/samples/cam031/92041633.html;
 http://www.loc.gov/catdir/toc/cam028/92041633.html.

Renn:2011:SWW

[RW11] Jürgen Renn and Georg Winckler. Schrödingers Weg zur Wellenmechanik. (German) [Schrödinger’s road to wave mechanics]. Mitschnitt einer gemeinsamen Veranstaltung des Erwin Schrödinger-Instituts für Mathematische Physik (ESI) und der Fakultät für Physik der Universität Wien am Donnerstag, dem 13. Jänner 2011 im Kleinen Festsaal der Universität Wien [Recording of a joint event at the Erwin Schrödinger Institute for Mathematical Physics (ESI) and the Department of Physics, University of Vienna on Thursday, the 13th January 2011 in the Small Hall of the University of Vienna]., 2011.

Reiter:2013:ESY

[RY13] Wolfgang Reiter and Jakob Yngvason, editors. Erwin Schrödinger — 50 Years After. European Mathematical Society Publishing House, Zürich, Switzerland, 2013. ISBN 3-03719-121-X (paperback), 3-03719-621-1 (e-book). viii + 185 pp. LCCN QC16.S265 I58 2011.

Salam:1987:OPP

[Sal87] A. Salam. Overview of particle physics. In Kilmister [Kil87b], pages 180–200. ISBN 0-521-37929-6 (paperback), 0-521-34017-9. LCCN QC16.S265 S36 1987; Q143.S3. URL http://www.loc.gov/

catdir/enhancements/fy0901/86028338-d.html;   http://www.
loc.gov/catdir/enhancements/fy0901/86028338-t.html. Pa-

pers presented at a conference, March 31–April 3, 1987, at Imperial College, London.

Sallhofer:2001:PFD

[Sal01] Hans H. Sallhofer. Der Physikerstreit: ein fiktiver Diskurs mit Schrödinger und Heisenberg. (German) [The physicist dispute: a fictional discourse between Schrödinger and Heisenberg]. Universitas, München, Germany, 2001. ISBN 3-8004-1425-2. 312 pp. LCCN QC174.12 .S2555 2001.

Salamin:2014:SKA

[Sal14] Y. I. Salamin. Schrödinger’s Killer App: Race to Build the World’s First Quantum Computer, by Jonathan P. Dowling, Scope: general interest. Level: general readership. Contemporary Physics, 55(3): 252–254, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sauer:2016:BRP

[Sau16] Tilman Sauer. Book review: Paul Halpern. Einstein’s Dice and Schrödinger’s Cat: How Two Great Minds Battled Quantum Randomness to Create a Unified Theory of Physics. Isis, 107(2):427– 428, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schrodinger:1990:LMF

[SB90a] Erwin Schrödinger and Michel Bitbol. L’esprit et la matière. (French) [Mind and Matter]. Éditions du Seuil, Paris, France, 1990. ISBN 2-02-011469-0 (paperback). 250 pp. LCCN ????

Schrodinger:1990:LMP

[SB90b] Erwin Schrödinger and Michel Bitbol. L’Esprit et la matière; Précédé de L’Élision: Essai sur la philosophie d’E. Schrödinger. (French) [Mind and Matter: Preceded by The Elision: Essay on the philosophy of E. Schrödinger]. Éditions du Seuil, Paris, France, 1990. ISBN 2-02-011469-0 (paperback). 250 pp. LCCN ???? Translation, notes, and essay by Michel Bitbol.

Schrodinger:2011:LMF

[SB11] Erwin Schrödinger and Michel Bitbol. L’esprit et la matiére. (French) [Mind and matter]. Éditions du Seuil, Paris, France, 2011. ISBN ???? 322 pp. LCCN ???? Translation, notes, preface, and introduction by Michel Bitbol. Preceded by The Elision: Essay on the philosophy of E. Schrödinger, by Michel Bitbol.

Schrodinger:1910:LEO

[Sch10] Erwin Schrödinger. Über die Leitung der Elektrizität auf der Oberfläche von Isolatoren an feuchter Luft. (German) [On the conduction of electricity on the surface of insulators in moist air]. Sitzungsberichte der kaiserlichen Akademie der Wissenschaften in Wien. Mathematisch-naturwissenschaftliche Klasse, Abteilung 2a, 119(??):1215–1222, ???? 1910.

Schrodinger:1912:BKA

[Sch12a] Erwin Schrödinger. Beiträge zur Kenntnis der atmosphärischen Elektrizität. (German) [Contributions to the knowledge of atmospheric electricity]. Sitzungsberichte der kaiserlichen Akademie der Wissenschaften in Wien. Math.-nat. Kl., 121(??):2391–2406, ???? 1912.

Schrodinger:1912:EEW

[Sch12b] Erwin Schrödinger. Enzyklopädie der Elementarmathematik: Von Weber-Wellstein. III. Bd.: Angewandte Elementarmathematik. 1. Teil: Mathematische Physik. 2. Auflage. B. G. Teubner, Leipzig 1910. (German) [Encyclopedia of Elementary Mathematics: From Weber-Wellstein. III. Vol: Applied Elementary Mathematics. First Part: Mathematical Physics. Second Edition. B. G. Teubner, Leipzig, 1910]. Monatshefte für Mathematik, 23(1): A60, ???? 1912. CODEN MNMTA2. ISSN 0026-9255 (print),

1436-5081 (electronic). URL http://www.springerlink.com/
content/wh17m3776x09831m/fulltext.pdf.

Schrodinger:1912:SKD

[Sch12c] Erwin Schrödinger. Studien über Kinetik der Dielektrika, den Schmelzpunkt, Pyro und Piezoelektrizität. (German) [Studies on the kinetics of dielectrics, the melting point, pyro- and piezoelectricity]. Sitzungsberichte der kaiserlichen Akademie der Wissenschaften in Wien. Mathematisch-naturwissenschaftliche Klasse, Abteilung 2a, 121(??):1937–1972, ???? 1912.

Schrodinger:1912:TMR

[Sch12d] Erwin Schrödinger. Theoretische Mechanik: Von Robert Marcolongo. Autorisierte deutsche Bearbeitung von H. E. Timerding. I. Bd. Kinematik und Statik. B. G. Teubner, Leipzig 1911. (German) [Theoretical mechanics: From Robert Marcolongo. Authorized German version of HE timer thing. Vol I. kinematics and statics. B. G. Teubner, Leipzig, 1911]. Monatshefte für Mathematik, 23(1):A59–A60, ???? 1912. CODEN MNMTA2. ISSN 0026-9255 (print), 1436-5081 (electronic). URL http://www.springerlink.

com/content/h6u616q4844m4252/fulltext.pdf.

Schrodinger:1912:HDA

[Sch12e] Erwin Schrödinger. Über die Höhenverteilung der durchdringenden atmosphärischen Strahlung (Theorie). (German) [On the vertical distribution of penetrating atmospheric radiation (theory)]. Sitzungsberichte der kaiserlichen Akademie der Wissenschaften in Wien. Mathematisch-naturwissenschaftliche Klasse, Abteilung 2a, 121(??):2391–2406, ???? 1912.

Schrodinger:1912:KTM

[Sch12f] Erwin Schrödinger. Zur kinetischen Theorie des Magnetismus (Einfluß der Leitungselektronen). (German) [On the kinetic theory of magnetism (influence of conduction electrons)]. Sitzungsberichte der kaiserlichen Akademie der Wissenschaften in Wien. Mathematisch-naturwissenschaftliche Klasse, Abteilung 2a, 121 (??):1305–1328, ???? 1912.

Schrodinger:1913:AIT

[Sch13a] Erwin Schrödinger. Die Akkumulatoren, ihre Theorie, Herstellung, Behandlung und Verwendung: Von Prof. Dr. W. Bermbach. Leipzig 1911. (German) [The accumulators, their theory, manufacture, handling and use: By Prof. Dr. W. Bermbach. Leipzig, 1911]. Monatshefte für Mathematik, 24(1):A27, ???? 1913. CODEN MNMTA2. ISSN 0026-9255 (print), 1436-5081

(electronic). URL http://www.springerlink.com/content/
n3354794m4j91255/fulltext.pdf.

Schrodinger:1913:WMM

[Sch13b] Erwin Schrödinger. Die Wandlungen der Mechanik und die mechanische Naturerklärung: Von P. Duhem. (L’évolution de la mécanique). Übersetzt von Philipp Frank. Leipzig 1912. (German) [The changes in the mechanics and the mechanical explanation of nature: from Pierre Duhem. (L’évolution de la mécanique). Translated by Philip Frank. Leipzig 1912]. Monatshefte für Mathematik, 24(1):A25–A27, ???? 1913. CODEN MNMTA2. ISSN 0026-9255 (print), 1436-5081 (electronic). URL http://www.springerlink.

com/content/q0075538m6356330/fulltext.pdf.

Schrodinger:1913:GDM

[Sch13c] Erwin Schrödinger. Grundlagen des dezimalen metrischen Systems: von Méchain und Delambre und Versuche über die Länge des Sekundenpendels in Paris von Borda und Cassini. Ostwalds Klassiker der exakten Wissenschaften Nr. 181. Engelmann, Leipzig 1911. (German) [Fundamentals of the decimal metric system: from the Méchain and Delambre attempts and the length of the seconds pendulum in Paris by Borda and Cassini. Ostwald’s classics of the exact sciences, No. 181 Engelmann, Leipzig, 1911]. Monatshefte für Mathematik, 24(1):A54, ???? 1913. CODEN MNMTA2. ISSN 0026-9255 (print), 1436-5081 (electronic). URL http://www.

springerlink.com/content/hn236g71465t3070/fulltext.pdf.

Schrodinger:1913:GEP

[Sch13d] Erwin Schrödinger. Grundzüge der Experimentalphysik: von Prof. Hermann Roth; F. Grub, Stuttgart 1912. Mark 1,60. (German) [Fundamentals of experimental physics: Prof. Hermann Roth, F. Grub, Stuttgart 1912, 1.60 Marks]. Monatshefte für Mathematik, 24(1):A52–A53, ???? 1913. CODEN MNMTA2. ISSN 0026-9255 (print), 1436-5081 (electronic). URL http://www.springerlink.

com/content/q6qt17l26q522452/fulltext.pdf.

Schrodinger:1913:HGG

[Sch13e] Erwin Schrödinger. Hermann Graßmanns gesammelte mathematische und physikalische Werke: III. Bd. 1. Teil: Theorie der Ebbe und Flut, Prüfungsarbeit 1840 und Abhandlungen zur mathematischen Physik aus dem Nachlasse. Herausgegeben von Justus Graßmann und Friedrich Engel. 1911. (German) [Hermann Grassmann’s mathematical and physical collected works: III. Bd first Supply: Theory of the ebb and flow, audit work papers for 1840 and mathematical physics from the estate. Edited by Justus Grassmann and Friedrich Engel. 1911]. Monatshefte für Mathematik, 24(1):A38, ???? 1913. CODEN MNMTA2. ISSN 0026-9255 (print), 1436-5081 (electronic). URL http://www.springerlink.

com/content/lj432285638235tp/fulltext.pdf.

Schrodinger:1913:LPG

[Sch13f] Erwin Schrödinger. Lehrbuch der Physik: zum Gebrauche beim Unterricht, bei akademischen Vorlesungen und zum Selbststudium von E. Grimsehl. II. vermehrte und verbesserte Auflage. Teubner, Leipzig 1912. (German) [Textbook of physics: for the use in teaching, academic lectures and self-study of E. Grimsehl. Second enlarged and improved edition. Teubner, Leipzig 1912]. Monatshefte für Mathematik, 24(1):A53–A54, ???? 1913. CODEN MNMTA2. ISSN 0026-9255 (print), 1436-5081 (electronic). URL http://www.

springerlink.com/content/k15u7g72v1082194/fulltext.pdf.

Schrodinger:1913:NTA

[Sch13g] Erwin Schrödinger. Notiz über die Theorie der anomalen elektrischen Dispersion. (German) [Note on the theory of anomalous electric dispersion]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 15(22):1167–1172, November 30, 1913. CODEN VD-PEAZ. ISSN 0372-5448.

Schrodinger:1913:RGA

[Sch13h] Erwin Schrödinger. Radium-A-Gehalt der Atmosphäre in Seeham 1913. (German) [Radium-A content of the atmosphere in Seeham 1913]. Sitzungsberichte der kaiserlichen Akademie der Wissenschaften in Wien. Mathematisch-naturwissenschaftliche Klasse, Abteilung 2a, 122(100):2023–2067, December 1913.

Schrodinger:1913:TMR

[Sch13i] Erwin Schrödinger. Theoretische Mechanik: Von Robert Marcolongo. Autorisierte deutsche Bearbeitnng von H. E. Timerding. II. Band Dynamik und Mechanik der deformierbaren Körper. Teubner 1912. (German) [Theoretical mechanics: From Robert Marcolongo. Authorized by German U Bearbeitnng Timerding. Second volume dynamics and mechanics of deformable bodies. Teubner 1912]. Monatshefte für Mathematik, 24(1):A54–A55, ???? 1913. CODEN MNMTA2. ISSN 0026-9255 (print),

1436-5081 (electronic). URL http://www.springerlink.com/
content/qm661r7248245732/fulltext.pdf.

Schrodinger:1913:WSL

[Sch13j] Erwin Schrödinger. Wie schreitet das Licht im isotropen Medium vor?: Von Dr. Vinzenz Wießner. Dresden und Leipzig 1908. (German) [As the light falls from the isotropic medium: By Dr. Vincent Wießner Dresden and Leipzig 1908]. Monatshefte für Mathematik, 24(1):A25, ???? 1913. CODEN MNMTA2. ISSN 0026-9255 (print), 1436-5081 (electronic). URL http://www.springerlink.

com/content/gt5w11p363684nn3/fulltext.pdf.

Schrodinger:1914:THP

[Sch14a] E. Schrödinger. A treatise on hydromechanics: P. II. Hydrodynamics by A. S. Ramsey. Bell and Son, London 1913. Monatshefte für Mathematik, 25(1):A13, ???? 1914. CODEN MNMTA2. ISSN 0026-9255 (print), 1436-5081 (electronic). URL http://www.

springerlink.com/content/k17p1133g103mr07/fulltext.pdf.

Schrodinger:1914:TEE

[Sch14b] Erwin Schrödinger. Theorie der Elektrizität: Einführung in die Maxwellsche Theorie der Elektrizität, von Dr. A. Föppl. Vierte, umgearbeitete Auflage von Dr. M. Abraham. B. G. Teubner 1912. (German) [Theory of Electricity: Introduction to Maxwell’s theory of electricity, by Dr. A. Foppl. Fourth, revised edition of Dr. M. Abraham. B. G. Teubner 1912]. Monatshefte für Mathematik, 25(1):A12, ???? 1914. CODEN MNMTA2. ISSN 0026-9255 (print), 1436-5081 (electronic). URL http://www.springerlink.

com/content/xu1p23p63u06g62t/fulltext.pdf.

Schrodinger:1914:TCG

[Sch14c] Erwin Schrödinger. Traité de chimie générale: Von W. Nernst, übersetzt von A. Carvisy. A. Hermann, Paris 1911. (German) [Treatise on General Chemistry: From W. Nernst, translated by A. Carvisy. A. Hermann, Paris 1911]. Monatshefte für Mathematik, 25(1):A11–A12, ???? 1914. CODEN MNMTA2. ISSN 0026-9255 (print), 1436-5081 (electronic). URL http://www.springerlink.

com/content/w01254746311368m/fulltext.pdf.

Schrodinger:1914:SRE

[Sch14d] Erwin Schrödinger. Über die Schärfe der mit Röntgenstrahlen erzeugten Interferenzbilder. (German) [On the sharpness of interference images generated by X-rays]. Physikalische Zeitschrift, 15 (2):79–86, January 15, 1914. CODEN PHZTAO. ISSN 0369-982X.

URL http://hdl.handle.net/2027/uc1.c2831834?urlappend=
%3Bseq=111.

Schrodinger:1914:DEP

[Sch14e] Erwin Schrödinger. Zur Dynamik der elastischen Punktreihe: vorgelegt in der Sitzung am 25. Juni 1914. (German) [On the dynamics of the elastic point series: presented at the meeting on 25 June 1914]. Sitzungsberichte der kaiserlichen Akademie der Wissenschaften in Wien. Mathematisch-naturwissenschaftliche Klasse, Abteilung 2a, 123(2a):1679–1696, June 25, 1914.

Schrodinger:1914:DEG

[Sch14f] Erwin Schrödinger. Zur Dynamik elastisch gekoppelter Punktsysteme. (German) [On the dynamics of elastic coupled point systems]. Annalen der Physik (1900), 349(14):916–934, ???? 1914. ISSN 1521-3889.

Schrodinger:1914:TDG

[Sch14g] Erwin Schrödinger. Zur Theorie des Debyeeffekts. (German) [On the theory of the Debye effect]. Physikalische Zeitschrift, 15(10): 497–503, May 15, 1914. CODEN PHZTAO. ISSN 0369-982X.

URL http://hdl.handle.net/2027/uc1.c2831834?urlappend=
%3Bseq=535.

Schrodinger:1915:NKG

[Sch15a] Erwin Schrödinger. Notiz über den Kapillardruck in Gasblasen. (German) [Note on the capillary pressure in gas bubbles]. Annalen der Physik (1900), 351(3):413–418, ???? 1915. ISSN 1521-3889.

Schrodinger:1915:TFS

[Sch15b] Erwin Schrödinger. Zur Theorie der Fall- und Steigversuche an Teilchen mit Brown’scher Bewegung. (German) [On the theory of the rise and fall of particles in Brownian motion]. Physikalische Zeitschrift, 16(16):289–295, August 15, 1915. CODEN PHZ-

TAO. ISSN 0369-982X. URL http://hdl.handle.net/2027/
njp.32101054770928?urlappend=%3Bseq=337.

Schrodinger:1917:ENFa

[Sch17a] Erwin Schrödinger. Die Ergebnisse der neueren Forschung über Atom- und Molekularwärmen. (German) [The results of recent research on atomic and molecular heats]. Naturwissenschaften, 5(34): 537–543, August 24, 1917. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://www.springerlink.

com/content/r14m7uk7452u54m7/fulltext.pdf.

Schrodinger:1917:ENFb

[Sch17b] Erwin Schrödinger. Die Ergebnisse der neueren Forschung über Atom- und Molekularwärmen (Schluß). (German) [The results of recent research on atomic and molecular heats (final)]. Naturwissenschaften, 5(35):561–567, August 1917. CODEN NATWAY.

ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://www.

springerlink.com/content/32h4718241764341/fulltext.pdf.

Schrodinger:1917:DKM

[Sch17c] Erwin Schrödinger. Dynamik der Kristallgitter: Von Max Born. B. G. Teubner, Leipzig u. Berlin, 1915. (German) [Dynamics of crystal lattices: From Max Born. B. G. Teubner, Leipzig and Berlin, 1915]. Monatshefte für Mathematik, 28(1):A8–A10, ???? 1917. CODEN MNMTA2. ISSN 0026-9255 (print),

1436-5081 (electronic). URL http://www.springerlink.com/
content/n1g0l25nl1407k0x/fulltext.pdf.

Schrodinger:1917:NMA

[Sch17d] Erwin Schrödinger. Nachtrag zu meinen Aufsatz: Zur Akustik der Atmosphäre. (German) [Supplement to my article: On the acoustics of the atmosphere]. Physikalische Zeitschrift, 18(23): 567, December 1, 1917. CODEN PHZTAO. ISSN 0369-

982X. URL http://babel.hathitrust.org/cgi/pt?id=mdp.
39015009220800%3Bseq=615%3Bview=1up. See [Sch17e].

Schrodinger:1917:AAG

[Sch17e]  Erwin Schrödinger.  Zur Akustik der Atmosphäre. (Ger-

man) [On the acoustics of the atmosphere]. Physikalische Zeitschrift, 18(19):445–453, October 1, 1917. CODEN PHZ-

TAO. ISSN 0369-982X. URL http://hdl.handle.net/2027/
mdp.39015009220800?urlappend=%3Bseq=491. See supplement

[Sch17d].

Schrodinger:1918:EGG

[Sch18a] Erwin Schrödinger. Die Energiekomponenten des Gravitationsfeldes. (German) [The energy component of the gravitational field]. Physikalische Zeitschrift, 19(1):4–7, January 1, 1918. CODEN

PHZTAO. ISSN 0369-982X. URL http://babel.hathitrust.
org/cgi/pt?id=mdp.39015056706941%3Bseq=24%3Bview=1up.

Schrodinger:1918:D

[Sch18b] Erwin Schrödinger. Dielektrizität. (German) [Dielectric]. In Handbuch der Elektrizität und des Magnetismus, I, [1914]. (German) [Handbook of Electricity and Magnetism, I, [1914]], pages 157–231. Barth, Leipzig, Germany, 1918.

Schrodinger:1918:NOZ

[Sch18c] Erwin Schrödinger. Notiz über die Ordnung in Zufallsreihen. (German) [Note on the random order in decay series]. Physikalische Zeitschrift, 19(10):218–220, May 15, 1918. CODEN PHZ-

TAO. ISSN 0369-982X. URL http://hdl.handle.net/2027/
mdp.39015056706941?urlappend=%3Bseq=264.

Schrodinger:1918:ERA

[Sch18d] Erwin Schrödinger. Über ein in der experimentellen Radiumforschung auftretendes Problem der statistischen Dynamik. (German) [On an occurring problem of statistical dynamics in experimental research on radium]. Sitzungsberichte der kaiserlichen Akademie der Wissenschaften in Wien. Mathematischnaturwissenschaftliche Klasse, Abteilung 2a, 127(??):237–262, ???? 1918.

Schrodinger:1918:LAK

[Sch18e] Erwin Schrödinger. Über ein Lösungssystem der allgemein kovarianten Gravitationsgleichungen. (German) [On a solution of the system of generally-covariant gravitational equations]. Physikalische Zeitschrift, 19(2):20–22, January 15, 1918. CODEN PHZ-

TAO. ISSN 0369-982X. URL http://hdl.handle.net/2027/
mdp.39015056706941?urlappend=%3Bseq=42.

Schrodinger:1919:EFI

[Sch19a] Erwin Schrödinger. Der Energieinhalt der Festkörper im Lichte der neueren Forschung. (German) [The energy content of solids in the light of recent research]. Physikalische Zeitschrift, 20 (18):420–428; 450–455; 474–480; 497–503; 523–526, September 15, 1919. CODEN PHZTAO. ISSN 0369-982X. URL http://

hdl.handle.net/2027/hvd.hxgq91?urlappend=%3Bseq=659;
http://hdl.handle.net/2027/mdp.39015009220727?urlappend=
%3Bseq=508;   http://hdl.handle.net/2027/mdp.39015009220727?
urlappend=%3Bseq=542;   http://hdl.handle.net/2027/mdp.
39015009220727?urlappend=%3Bseq=568;   http://hdl.handle.
net/2027/mdp.39015009220727?urlappend=%3Bseq=593.

Schrodinger:1919:WSB

[Sch19b] Erwin Schrödinger. Wahrscheinlichkeitstheoretische Studien, betreffend Schweidler’sche Schwankungen, besonders die Theorie der Meßanordnung. (German) [Probabilistic studies concerning Schweidler fluctuations, particularly the theory of the measuring device]. Sitzungsberichte der kaiserlichen Akademie der Wissenschaften in Wien. Mathematisch-naturwissenschaftliche Klasse, Abteilung 2a, 128(??):177–237, ???? 1919.

Schrodinger:1920:EFA

[Sch20a] E. Schrödinger. Die Elektrizität: Von F. Adami. 2. Aufl. Günther, Bücher der Naturwissenschaft. 9. Bd. Leipzig, Philipp Reclam jun. (German) [Electricity: By F. Adami. Second Ed. Guenther, books of science. 9th Bd Leipzig, Philipp Reclam, Jr.]. Monatshefte für Mathematik, 30(1):A1, ???? 1920. CODEN MNMTA2. ISSN 0026-9255 (print), 1436-5081 (electronic). URL http://www.

springerlink.com/content/j43718717600tq44/fulltext.pdf.

Schrodinger:1920:TEC

[Sch20b] E. Schrödinger. Theorie der Erdgestalt: Von Clairaut. Herausgegeben von Jourdain und von Öttingen. Ostwalds Klassiker, Leipzig 1913. (German) [Theory of the Earth: From Clairaut. Edited by Jourdain and von Oettingen. Ostwald’s classics, Leipzig 1913]. Monatshefte für Mathematik, 30(1):A1, ???? 1920. CODEN MNMTA2. ISSN 0026-9255 (print),

1436-5081 (electronic). URL http://www.springerlink.com/
content/wq17p414719h8625/fulltext.pdf.

Schrodinger:1920:VVI

[Sch20c] E. Schrödinger. Vektoranalysis: Von S. Valen iner. Sammlung Göschen, 1912. 2. Aufl. (German) [Vector Analysis: by S. Valen iner. Goeschen Collection, 1912. Second Ed.]. Monatshefte für Mathematik, 30(1):A1, ???? 1920. CODEN MNMTA2. ISSN 0026-9255 (print), 1436-5081 (electronic). URL http://www.

springerlink.com/content/m1m136n72tt045m2/fulltext.pdf.

Schrodinger:1920:FGC

[Sch20d] Erwin Schrödinger. Farbenmetrik. (German) [Color metrics]. Zeitschrift für Physik, 1(5):459–466, October 1920. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://www.springerlink.com/content/j337555313upp81h/fulltext.
pdf.

Schrodinger:1920:GTFa

[Sch20e] Erwin Schrödinger. Grundlinien einer Theorie der Farbenmetrik im Tagessehen. (German) [Outline of a theory of color metrics in daylight vision]. Annalen der Physik (1900), 368(21):397–426, 1920. ISSN 1521-3889.

Schrodinger:1920:GTFb

[Sch20f] Erwin Schrödinger. Grundlinien einer Theorie der Farbenmetrik im Tagessehen. (German) [Outline of a theory of color metrics in daylight vision]. Annalen der Physik (1900), 368(21):427–456, 1920. ISSN 1521-3889.

Schrodinger:1920:GTFc

[Sch20g] Erwin Schrödinger. Grundlinien einer Theorie der Farbenmetrik im Tagessehen. (German) [Outline of a theory of color metrics in daylight vision]. Annalen der Physik (1900), 368(22):481–520, 1920. ISSN 1521-3889.

Schrodinger:1920:TPG

[Sch20h] Erwin Schrödinger. Theorie der Pigmente von größter Leuchtkraft. (German) [Theory of pigments of great luminosity]. Annalen der Physik (1900), 367(15):603–622, ???? 1920. ISSN 1521-3889.

Schrodinger:1920:KWB

[Sch20i] Erwin Schrödinger. Über die Kohärenz in weitgeöffneten Bündeln. (German) [On consistency in wide-open clusters]. Annalen der Physik (1900), 366(1):69–86, ???? 1920. ISSN 1521-3889.

Schrodinger:1921:IGP

[Sch21a] Erwin Schrödinger. Isotopie und Gibbssches Paradoxon. (German) [Isotopy and the Gibbs Paradox]. Zeitschrift für Physik, 5 (2–3):163–166, March 1921. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://www.springerlink.

com/content/r615226147x44628/fulltext.pdf.

Schrodinger:1921:VMD

[Sch21b]  Erwin Schrödinger.  Versuch zur modellmäßigen Deutung

des Terms der scharfen Nebenserien. (German) [Attempt at model-based interpretation of the term of the sharp subordinate series]. Zeitschrift für Physik, 4(3):347–354, September 1921. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831

(electronic). URL http://www.springerlink.com/content/
j266071843743571/fulltext.pdf.

Schrodinger:1921:WDQ

[Sch21c] Erwin Schrödinger. Webster, D. L.: Quantum emission phenomena in radiation, Found, C. G.: Ionization potentials of argon, nitrogen, carbon monoxide, helium, hydrogen and mercury and iodine vapors, Wilson, H. A.: On electromagnetic momentum [Articles Review]. Naturwissenschaften, 9(1):20–21, January 1921. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Schrodinger:1922:SWFb

[Sch22a] E. Schrödinger. Über die spezifische Wärme fester Körper usw. (German) [On the specific heat of solids, etc.]. Zeitschrift für Physik, 11(1):396, December 1922. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://www.

springerlink.com/content/m2v77142w6532484/fulltext.pdf.

Minor erratum for [Sch22c].

Schrodinger:1922:DBF

[Sch22b] Erwin Schrödinger. Dopplerprinzip und Bohrsche Frequenzbedingung. (German) [Doppler Principle and Bohr’s frequency condition]. Physikalische Zeitschrift, 23(15):301–303, August 1, 1922. CODEN PHZTAO. ISSN 0369-982X. URL http://hdl.handle.

net/2027/mdp.39015086723239?urlappend=%3Bseq=367.

Schrodinger:1922:SWFa

[Sch22c] Erwin Schrödinger. Über die spezifische Wärme fester Körper bei hoher Temperatur und über die Quantelung von Schwingungen endlicher Amplitude. (German) [On the specific heat of solids at high temperature and the quantization of finite-amplitude oscillations]. Zeitschrift für Physik, 11(1):170–176, December 1922. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831

(electronic). URL http://www.springerlink.com/content/
x0651j468414132l/fulltext.pdf.

Schrodinger:1923:TFG

[Sch23a] Erwin Schrödinger. Ton und Farbe. (German) [Sound and color]. Neue Zürcher Zeitung, ??(??):??, February 3, 1923.

Schrodinger:1923:BEQ

[Sch23b] Erwin Schrödinger. Über eine bemerkenswerte Eigenschaft der Quantenbahnen eines einzelnen Electrons. (German) [On a remarkable property of the quantum trajectory of a single electron]. Zeitschrift für Physik, 12(1):13–23, December 1923. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831

(electronic). URL http://www.springerlink.com/content/
j7117422717018jh/fulltext.pdf.

Schrodinger:1923:RGR

[Sch23c] Erwin Schrödinger. Vom Radium. (German) [On Radium]. Neue Zürcher Zeitung, ??(??):??, March 15, 1923.

Schrodinger:1924:BZA

[Sch24a] Erwin Schrödinger. Bemerkung zu zwei Arbeiten des Herrn Elemér Császár über Strahlungstheorie und spezifische Wärmen. (German) [Remark on two papers of Mr. Elemér Császár on radiation theory and specific heats]. Zeitschrift für Physik, 25(1): 173–174, December 1924. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://www.springerlink.

com/content/j467952024230755/fulltext.pdf.

Schrodinger:1924:BNS

[Sch24b] Erwin Schrödinger. Bohrs neue Strahlungshypothese und der Energiesatz (German) [Bohr’s new radiation hypothesis and the energy theorem]. Naturwissenschaften, 12(36):720–724, September 1924. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904

(electronic). URL http://www.springerlink.com/content/
k07m787384759035/fulltext.pdf.

Schrodinger:1924:ESG

[Sch24c] Erwin Schrödinger. Einiges über die Sterne. (German) [Some of the stars]. Neue Zürcher Zeitung, ??(??):??, October 28, 29, 30, 1924.

Schrodinger:1924:GFW

[Sch24d] Erwin Schrödinger. Gasentartung und freie Weglänge. (German) [Gas degeneracy and mean free path]. Physikalische Zeitschrift, 25 (??):41–45, ???? 1924. CODEN PHZTAO. ISSN 0369-982X.

Schrodinger:1924:KMA

[Sch24e] Erwin Schrödinger. Kann man Atome photographieren?. (German) [Can one photograph the atom?]. Neue Zürcher Zeitung, ??(??): ??, February 7, 1924.

Schrodinger:1924:TGZ

[Sch24f] Erwin Schrödinger. Über das thermische Gleichgewicht zwischen Licht- und Schallstrahlen. (German) [On the thermal equilibrium between light- and sound beams]. Physikalische Zeitschrift, 25(??): 89–94, ???? 1924. CODEN PHZTAO. ISSN 0369-982X.

Schrodinger:1924:UEA

[Sch24g] Erwin Schrödinger. Über den Ursprung der Empfindlichkeitskurven des Auges. (German) [On the origin of the sensitivity curves of the eye]. Naturwissenschaften, 12(45):925–929, November 1924. CODEN NATWAY. ISSN 0028-1042 (print),

1432-1904 (electronic). URL http://www.springerlink.com/
content/x710k61123640737/fulltext.pdf.

Schrodinger:1924:RWG

[Sch24h] Erwin Schrödinger. Über die Rotationswärme des Wasserstoffs. (German) [On the rotation heat of hydrogen]. Zeitschrift für Physik, 30(1):341–349, December 1924. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://www.

springerlink.com/content/r324527148578520/fulltext.pdf.

Schrodinger:1925:BSE

[Sch25a] Erwin Schrödinger. Bemerkungen über die statistische Entropiedefinition beim idealen Gas. (German) [Remarks on the statistical definition of entropy in an ideal gas]. Sitzungsberichte der Preussischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):434–441, ???? 1925. CODEN SPW-PAI. ISSN 0371-2435.

Schrodinger:1925:ERK

[Sch25b] Erwin Schrödinger. Die Erfüllbarkeit der Relativitätsforderung in der klassischen Mechanik. (German) [The satisfiability of the relativity requirement in the classical mechanics]. Annalen der Physik (1900), 382(11):325–336, ???? 1925. ISSN 1521-3889.

Schrodinger:1925:WSS

[Sch25c] Erwin Schrödinger. Die wasserstoffähnlichen Spektren vom Standpunkte der Polarisierbarkeit des Atomrumpfes. (German) [The hydrogen-like spectra from the standpoint of polarizability of the atomic core]. Annalen der Physik (1900), 382(9):43–70, ???? 1925. ISSN 1521-3889.

Schrodinger:1925:MHR

[Sch25d] Erwin Schrödinger. Marx, E.: Handbuch der Radiologie 6, Die Theorien der Radiologie [Book review]. (German) [Marx, E.: Handbook of Radiology 6, The theories of radiology]. Naturwissenschaften, 13(13):710–711, August 1925. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Schrodinger:1925:MVR

[Sch25e] Erwin Schrödinger. Michelsonscher Versuch und Relativitätstheorie. (German) [The Michelson experiment and Relativity theory]. Neue Zürcher Zeitung, ??(??):??, September 10, 1925.

Schrodinger:1925:OLV

[Sch25f] Erwin Schrödinger. Otto Lummer (verstorben). (German) [Otto Lummer (deceased)]. Neue Zürcher Zeitung, ??(??):??, July 23, 1925.

Schrodinger:1925:VVD

[Sch25g] Erwin Schrödinger. Über das Verhältnis der Vierfarben- zur Dreifarbentheorie. (German) [On the relation of the four-color theory to the three-color theory]. Sitzungsberichte der kaiserlichen Akademie der Wissenschaften in Wien. Mathematisch-naturwissenschaftliche Klasse, Abteilung 2a, 134(??):471–490, ???? 1925. Reprinted in [Sch94c].

Schrodinger:1925:SSQ

[Sch25h] Erwin Schrödinger. Über die subjektiven Sternfarben und die Qualität der Dämmerungsempfindung. (German) [On the subjective star colors and the quality of twilight sensation]. Naturwissenschaften, 13(18):373–376, May 1925. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://www.

springerlink.com/content/l17577p35122g524/fulltext.pdf.

Schrodinger:1925:FGC

[Sch25i] Erwin Schrödinger. Über Farbenmessung. (German) [On colorimetry]. Physikalische Zeitschrift, 26(??):349–352, ???? 1925. CODEN PHZTAO. ISSN 0369-982X.

Schrodinger:1926:SUM

[Sch26a] Erwin Schrödinger. Der stetige Übergang von der Mikro- zur Makromechanik. (German) [The continuous transition from microto macro-mechanics]. Naturwissenschaften, 14(28):664–666, July 1926. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904

(electronic). URL http://www.springerlink.com/content/
u8294161580651l7/;   http://www.springerlink.com/content/
u8294161580651l7/fulltext.pdf.

Schrodinger:1926:EIE

[Sch26b] Erwin Schrödinger. Die Energiestufen des idealen einatomigen Gasmodells. (German) [The energy levels of an ideal monatomic gas model]. Sitzungsberichte der Preussischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):23– 36, ???? 1926. CODEN SPWPAI. ISSN 0371-2435.

Schrodinger:1926:GGF

[Sch26c] Erwin Schrödinger. Die Gesichtsempfindungen. (German) [Facial sensations]. In Müller-Pouillets Lehrbuch der Physik, volume 2/1, pages 456–560. Friedrich Vieweg und Sohn, Braunschweig, Germany, 11 edition, 1926.

Schrodinger:1926:GJG

[Sch26d] Erwin Schrödinger. Guild, J.: Die geometrische Lösung von Farbenmischungsaufgaben [Article review]. (German) [Guild, J.: The geometrical solution of color mixture functions]. Naturwissenschaften, 14(8):146–147, February 19, 1926. CODEN NAT-WAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Schrodinger:1926:HKF

[Sch26e] Erwin Schrödinger. Herzfeld, K. F.: Kinetische Theorie der Wärme (Müller-Pouillets Lehrbuch der Physik 3/2, 11. Auflage) [Book review]. (German) [Herzfeld, KF: Kinetic Theory of Heat (Mueller-Pouillets textbook of Physics 3/2, 11th Edition)]. Physikalische Zeitschrift, 27(??):184–185, ???? 1926. CODEN PHZTAO. ISSN 0369-982X.

Schrodinger:1926:QED

[Sch26f] Erwin Schrödinger. Quantisierung als Eigenwertproblem (Dritte Mitteilung: Störungstheorie, mit Anwendung auf den Starkeffekt der Balmerlinien). (German) [Quantization as an eigenvalue problem (Third report: perturbation theory, with application to the Stark effect of the Balmer lines)]. Annalen der Physik (1900), 385 (13):437–490, ???? 1926. ISSN 1521-3889.

Schrodinger:1926:QEE

[Sch26g] Erwin Schrödinger. Quantisierung als Eigenwertproblem (Erste Mitteilung). (German) [Quantization as an eigenvalue problem (First report)]. Annalen der Physik (1900), 384(4):361–376, ???? 1926. ISSN 1521-3889.

Schrodinger:1926:QEV

[Sch26h] Erwin Schrödinger. Quantisierung als Eigenwertproblem (Vierte Mitteilung). (German) [Quantization as an eigenvalue problem (Fourth report)]. Annalen der Physik (1900), 386(18):109–139, ???? 1926. ISSN 1521-3889.

Schrodinger:1926:QEZ

[Sch26i] Erwin Schrödinger. Quantisierung als Eigenwertproblem (Zweite Mitteilung). (German) [Quantization as an eigenvalue problem (Second report)]. Annalen der Physik (1900), 384(6):489–527, ???? 1926. ISSN 1521-3889.

Schrodinger:1926:SWT

[Sch26j] Erwin Schrödinger. Spezifische Wärme (theoretischer Teil). (German) [Specific heat (theoretical part)]. In Handbuch der Physik, volume 10, pages 275–320. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1926.

Schrodinger:1926:VHB

[Sch26k] Erwin Schrödinger. Über das Verhältnis der Heisenberg–Born– Jordanschen Quantenmechanik zu der meinem. (German) [On the relation of Heisenberg–Born–Jordan quantum mechanics to mine]. Annalen der Physik (1900), 384(8):734–756, ???? 1926. ISSN 1521-3889.

Schrodinger:1926:UTM

[Sch26l] Erwin Schrödinger. An undulatory theory of the mechanics of atoms and molecules. Physical Review, 28(6):1049–1070, December 1, 1926. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/
PhysRev.28.1049.

Schrodinger:1926:EGG

[Sch26m] Erwin Schrödinger. Zur Einsteinschen Gastheorie. (German) [Italian: Sulla teoria del gas di Einstein] [On Einstein’s theory of gases]. Physikalische Zeitschrift, 27(??):95–101, ???? 1926. CODEN PHZ-TAO. ISSN 0369-982X.

Schrodinger:1927:AWG

[Sch27a] Erwin Schrödinger. Abhandlungen zur Wellenmechanik. (German) [Essays on wave mechanics]. Johann Ambrosius Barth Verlag, Leipzig, Germany, 1927. ix + 169 pp. LCCN QC174.2 .S3 1927.

Schrodinger:1927:EMG

[Sch27b] Erwin Schrödinger. Der Energieimpulssatz der Materiewellen. (German) [The energy momentum conservation of matter waves]. Annalen der Physik (1900), 387(2):265–272, ???? 1927. ISSN 1521-3889.

Schrodinger:1927:ENW

[Sch27c] Erwin Schrödinger. Energieaustausch nach der Wellenmechanik. (German) [Exchange of energy according to wave mechanics]. Annalen der Physik (1900), 388(15):956–968, ???? 1927. ISSN 1521-3889.

Schrodinger:1927:ESC

[Sch27d] Erwin Schrödinger. Erwin Schrödinger correspondence to Wilhelm Wien, 1925–1927. American Philosophical Society Library archive,

1927. URL http://www.amphilsoc.org/mole/view?docId=ead/

Mss.530.1.Ar2.7-ead.xml. 0.25 Linear feet, 17 items, photocopies, Mss.530.1.Ar2.7.

Schrodinger:1927:CGC

[Sch27e] Erwin Schrödinger. Über den Comptoneffekt. (German) [On the Compton effect]. Annalen der Physik (1900), 387(2):257–264, ???? 1927. ISSN 1521-3889.

Schrodinger:1928:CPW

[Sch28a] Erwin Schrödinger. Collected papers on wave mechanics. Blackie, Glasgow, Scotland, 1928. xiii + 146 pp. LCCN ????

Schrodinger:1928:FLW

[Sch28b] Erwin Schrödinger. Four lectures on wave mechanics: delivered at the Royal Institution, London, on 5th, 7th, 12th, and 14th March, 1928. Blackie, Glasgow, Scotland, 1928. viii + 53 pp. LCCN QA927 .S33 1928a.

Schrodinger:1928:MOF

[Sch28c] Erwin Schrödinger. La mécanique des ondes. (French) [Wave mechanics]. In Électrons et Photons. Rapports et Discussions du Cinquième conseil de Physique. (French) [Electrons and photons. Reports and discussions of the Fifth Council of Physics], pages 185– 213. Gauthier-Villars, Paris, France, 1928.

Schrodinger:1928:NWP

[Sch28d] Erwin Schrödinger. Neue Wege in der Physik. (German) [New directions in physics]. Elektrische Nachrichtentechnik, 5(??):485– 488, ???? 1928.

Schrodinger:1928:VVW

[Sch28e] Erwin Schrödinger. Vier Vorlesungen über Wellenmechanik: gehalten an der Royal Institution in London im März 1928. (German) [Four lectures on wave mechanics: held at the Royal Institution in London in March 1928]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1928. ???? pp. LCCN ????

Schrodinger:1929:AHM

[Sch29a] Erwin Schrödinger. Adresse an Hrn. Max Planck zum fünfzigjährigen Doktorjubiläum am 28. Juni 1929 [unsigned]. (German) [Address to Mr. Max Planck on the fiftieth anniversary of his doctorate on 28th June 1929]. Sitzungsberichte der Preussischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??): 341–342, ???? 1929. CODEN SPWPAI. ISSN 0371-2435.

Schrodinger:1929:AHSa

[Sch29b] Erwin Schrödinger. Antrittsrede des Hrn. Schrödinger. (German) [Inaugural address of Mr. Schrödinger]. Sitzungsberichte der Preussischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, C–CII(??):??, ???? 1929. CODEN SP-WPAI. ISSN 0371-2435.

Schrodinger:1929:AHSb

[Sch29c] Erwin Schrödinger. Antrittsrede des Hrn. Schrödinger. (German) [Inaugural address of Mr. Schrödinger]. Naturwissenschaften, 17 (37):732–733, September 1929. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Schrodinger:1929:GZK

[Sch29d] Erwin Schrödinger. Das Gesetz der Zufälle. Der Kampf um Ursache und Wirkung in den modernen Naturwissenschaften. (German) [The Law of Chance: The Problem of Causation in Modern Science]. Die Koralle, 5(??):417–418, ???? 1929.

Schrodinger:1929:RLG

[Sch29e] Erwin Schrödinger. Das Rätsel des Lichts. (German) [The mystery of light]. Die Koralle, 5(??):294–298, ???? 1929.

Schrodinger:1929:EWP

[Sch29f] Erwin Schrödinger. Der erkenntnistheoretische Wert physikalischer Modellvorstellungen. (German) [Conceptual Models in Physics and their Philosophical Value]. Jahresbericht des Physikalischen Vereins zu Frankfurt am Main 1928/29, ??(??):44–51, ???? 1929.

Schrodinger:1929:EQD

[Sch29g] Erwin Schrödinger. Die Erfassung der Quantengesetze durch kontinuierliche Funktionen. (German) [The detection of quantum laws by continuous functions]. Naturwissenschaften, 17(26):486– 489, June 1929. CODEN NATWAY. ISSN 0028-1042 (print),

1432-1904 (electronic). URL http://www.springerlink.com/
content/x6v70721745423jn/fulltext.pdf.

Schrodinger:1929:ENP

[Sch29h] Erwin Schrödinger. Eddington, A. S.: The Nature of the Physical World [Book review]. Naturwissenschaften, 17(35):694, August 1929. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Schrodinger:1929:EE

[Sch29i] Erwin Schrödinger. Einstein explained. World’s Work, ??(??): 52–55; 146, ???? 1929.

Schrodinger:1929:NWP

[Sch29j] Erwin Schrödinger. Neue Wege in der Physik. (German) [New directions in physics]. Elektrotechnische Zeitschrift, 50(??):15–16, ???? 1929.

Schrodinger:1929:VEG

[Sch29k] Erwin Schrödinger. Verwaschene Eigenwertspektra. (German) [Slurred eigenvalue spectra]. Sitzungsberichte der Preussischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):668–682, ???? 1929. CODEN SPWPAI. ISSN 0371-2435.

Schrodinger:1929:VGF

[Sch29l] Erwin Schrödinger. Vorwort. (German) [Foreword], volume III–IV, page ?? Verlag von S. Hirzel, Leipzig, Germany, 1929.

Schrodinger:1929:WNG

[Sch29m] Erwin Schrödinger. Was ist ein Naturgesetz?. (German) [What is a law of nature?]. Naturwissenschaften, 17(1):9–11, January 1929. CODEN NATWAY. ISSN 0028-1042 (print),

1432-1904 (electronic). URL http://www.springerlink.com/
content/q7576220557l5812/fulltext.pdf.

Schrodinger:1930:BRL

[Sch30a] Erwin Schrödinger. Book review: Lorentz, H. A.: Vorlesungen über theoretische Physik an der Universität Leiden 4. (German) [Lorentz, H. A.: Lectures on theoretical physics at the University of Leiden 4]. Deutsche Literaturzeitung, 4(??):43–44, January 1930.

Schrodinger:1930:GLG

[Sch30b] Erwin Schrödinger. Das gehetzte Licht. (German) [The harried light]. Uhu (Das neue Ullstein-Magazin), 6(??):106–112, ???? 1930.

Schrodinger:1930:WPW

[Sch30c] Erwin Schrödinger. Die Wandlung des physikalischen Weltbegriffs. (German) [The transformation of the concept of the physical world]. Lecture delivered at Deutsches Museum, Munich, 6th May 1930, 1930.

Schrodinger:1930:NEG

[Sch30d] Erwin Schrödinger. Naturwissenschaft und Ethik. (German) [Science and ethics]. Die Vossische Zeitung, ??(??):??, dez 25, 1930.

Schrodinger:1930:KBR

[Sch30e] Erwin Schrödinger. Über die kräftefreie Bewegung in der relativistischen Quantenmechanik. (German) [On force-free motion in relativistic quantum mechanics]. Sitzungsberichte der Preussischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, 24(??):418–428, ???? 1930. CODEN SPWPAI. ISSN 0371-2435.

Schrodinger:1930:VPG

[Sch30f] Erwin Schrödinger. Verabsäumte Pflichten. (German) [Neglected duties]. Reclams Universum, 47(??):263, ???? 1930.

Schrodinger:1930:WEE

[Sch30g] Erwin Schrödinger. Was ist eigentlich Elektrizität?. (German) [What exactly is electricity?]. Die Koralle, 6(??):110–112, ????

Schrodinger:1930:WKS

[Sch30h] Erwin Schrödinger. Wissenschaft–Kunst–Spiel. (German) [Science– art–game]. Die Koralle, 6(??):404; 410; 425–426, ???? 1930.

Schrodinger:1930:HUG

[Sch30i] Erwin Schrödinger. Zum Heisenbergschen Unschärfeprinzip. (German) [On Heisenberg’s Uncertainty Principle]. Sitzungsberichte der Preussischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):296–303, ???? 1930. CODEN SPW-PAI. ISSN 0371-2435.

Schrodinger:1931:BAH

[Sch31a] Erwin Schrödinger. Bemerkung zu der Arbeit des Herrn V. Fock: Die inneren Freiheitsgrade des Elektrons. (German) [Comment on the work of Mr. V. Fock: The internal degrees of freedom of the electron]. Zeitschrift für Physik, 70(11–12):808–810, November 1931. CODEN ZEPYAA. ISSN 0939-7922 (print),

1431-5831 (electronic). URL http://www.springerlink.com/
content/g02j511u11488684/fulltext.pdf.

Schrodinger:1931:IGS

[Sch31b] Erwin Schrödinger. Interviews with great scientists. Prof. Schrödinger. The Observer, ??(??):15–16, January 11, 1931.

Schrodinger:1931:JHS

[Sch31c] Erwin Schrödinger. Jeffreys, H.: Scientific Inference [Book review]. Naturwissenschaften, 19(48):967–968, November 1931. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Schrodinger:1931:SC

[Sch31d] Erwin Schrödinger. Science as culture. The International Forum, ??(??):10–11, January 1, 1931.

Schrodinger:1931:SRQ

[Sch31e] Erwin Schrödinger. Spezielle Relativitätstheorie und Quantenmechanik. (German) [Special Relativity theory and quantum mechanics]. Sitzungsberichte der Preussischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):238–247, ???? 1931. CODEN SPWPAI. ISSN 0371-2435.

Schrodinger:1931:UNG

[Sch31f] Erwin Schrödinger. Über die Umkehrung der Naturgesetze. (German) [On the reverse of the laws of nature]. Sitzungsberichte der Preussischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):144–153, ???? 1931. CODEN SPW-PAI. ISSN 0371-2435.

Schrodinger:1931:QEG

[Sch31g] Erwin Schrödinger. Zur Quantendynamik des Elektrons. (German) [On the quantum dynamics of the electron]. Sitzungsberichte der Preussischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):63–72, ???? 1931. CODEN SPW-PAI. ISSN 0371-2435.

Schrodinger:1932:AHS

[Sch32a] Erwin Schrödinger. [Antrittsrede des Hrn. Schrödinger]. (German) [Inaugural address of Mr. Schrödinger]. In E. Wildhagen, editor, Materie und Energie vom naturwissenschaftlichen Weltbild der Gegenwart. (German) [Matter and energy in the natural sciences: Current world view], pages 355–362. Deutsche Buchgemeinschaft, Berlin, Germany, 1932.

Schrodinger:1932:DEI

[Sch32b] Erwin Schrödinger. Dirac’sches Elektron im Schwerefeld. I. (German) [Dirac’s electron in the gravitational field. I]. Sitzungsberichte der Preussischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):105–128, ???? 1932. CODEN SPW-PAI. ISSN 0371-2435.

Schrodinger:1932:EWP

[Sch32c] Erwin Schrödinger. Eddington, A. S.: Das Weltbild der Physik und ein Versuch seiner philosophischen Deutung [Book review]. (German) [Eddington, A. S.: The world of physics and an attempt at philosophical interpretation]. Naturwissenschaften, 20(10):172– 173, March 1932. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Schrodinger:1932:NWP

[Sch32d] Erwin Schrödinger. Neue Wege in der Physik. (German) [New directions in physics]. In E. Wildhagen, editor, Materie und Energie vom naturwissenschaftlichen Weltbild der Gegenwart. (German) [Matter and energy from the natural sciences World view of the present], pages 348–354. Deutsche Buchgemeinschaft, Berlin, Germany, 1932.

Schrodinger:1932:TRE

[Sch32e] Erwin Schrödinger. Sur la théorie relativiste de l’électron et l’interprétation de la mécanique quantique. (French) [On the relativistic theory of the electron and the interpretation of quantum mechanics]. Annales de l’Institut Henri Poincaré, 2(??):269–310, ???? 1932. CODEN AIHPA2. ISSN 0365-320x (print), 2400-4855 (electronic).

Schrodinger:1932:IPG

[Sch32f] Erwin Schrödinger. Über Indeterminismus in der Physik. (German) [On indeterminism in physics]. Barth, Leipzig, Germany, 1932. ???? pp. LCCN ????

Schrodinger:1932:AKG

[Sch33a] Erwin Schrödinger. Anmerkungen zum Kausalproblem. (German) [Comments on the problem of causality]. Erkenntnis, 3(1):65–70, ???? 1932/1933. CODEN ???? ISSN

1876-2514. URL http://www.jstor.org/stable/20011666;
 http://www.springerlink.com/content/h47484u6264m1613/
fulltext.pdf.

Schrodinger:1933:LDT

[Sch33b] Erwin Schrödinger. L’électron de Dirac dans la theorie de la relativité générale. (French) [The Dirac electron in the theory of General Relativity]. In Comptes Rendus du Congrès International d’Électricité Paris 1932 (Première Section), pages 581–591. Gauthier-Villars, Paris, France, 1933.

Schrodinger:1933:MMO

[Sch33c] Erwin Schrödinger. Mémoires sur la mécanique ondulatoire: avec un avant-propos de l’auteur et des notes inédites spécialement écrites pour cette traduction. Alcan, Paris, France, 1933. xxvi + 234 pp. LCCN ????

Schrodinger:1933:ZHT

[Sch33d] Erwin Schrödinger. Über den zweiten Hauptsatz der Thermodynamik. (German) [On the Second Law of Thermodynamics]. Sitzungsberichte der Preussischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):165, ???? 1933. CO-DEN SPWPAI. ISSN 0371-2435.

Schrodinger:1933:VWB

[Sch33e] Erwin Schrödinger. van der Waerden, B. L.: Die gruppentheoretische Methode in der Quantenmechanik [Book review]. (German) [van der Waerden, B. L.: The group-theoretic method in quantum mechanics]. Physikalische Zeitschrift, 34(??):184, ???? 1933. CODEN PHZTAO. ISSN 0369-982X.

Schrodinger:1933:WAK

[Sch33f] Erwin Schrödinger. Warum sind die Atome so klein?. (German) [Why are atoms so small?]. Forschungen und Fortschritte, 9(??): 125–126, ???? 1933.

Schrodinger:1934:BRL

[Sch34a] Erwin Schrödinger. Book review: Lorentz, H. A.: Vorlesungen über theoretische Physik an der Universität Leiden 5. (German) [Lorentz, H. A.: Lectures on theoretical physics at the University of Leiden 5]. Deutsche Literaturzeitung, ??(??):1146, June 17, 1934.

Schrodinger:1934:GWG

[Sch34b] Erwin Schrödinger. Der Grundgedanke der Wellenmechanik. (German) [Der Grundgedanke der Wellenmechanik], pages 19–36. Verlag von S. Hirzel, Leipzig, Germany, 1934.

Schrodinger:1934:UGI

[Sch34c] Erwin Schrödinger. Über die Unanwendbarkeit der Geometrie im Kleinen. (German) [On the inapplicability of geometry in the small]. Naturwissenschaften, 22(31):518–520, August 1934. CODEN NATWAY. ISSN 0028-1042 (print),

1432-1904 (electronic). URL http://www.springerlink.com/
content/k31x5lq56381245q/fulltext.pdf.

Schrodinger:1935:GSQb

[Sch35a] E. Schrödinger. Die gegenwärtige Situation in der Quantenmechanik. (German) [The present situation in quantum mechanics]. Naturwissenschaften, 23(49):823–828, December 1935. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904

(electronic). URL http://www.springerlink.com/content/
hl6541202ukn4741/fulltext.pdf. English translation in [Tri80].

Schrodinger:1935:GSQc

[Sch35b] E. Schrödinger. Die gegenwärtige Situation in der Quantenmechanik. (German) [The present situation in quantum mechanics]. Naturwissenschaften, 23(50):844–849, December 1935. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904

(electronic). URL http://www.springerlink.com/content/
m489741169604131/fulltext.pdf. English translation in [Tri80].

Schrodinger:1935:CBN

[Sch35c] Erwin Schrödinger. Contributions to Born’s new theory of the electromagnetic field. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 150(870):465–477, June 1, 1935. CODEN PRLAAZ. ISSN 0080-4630. URL http://

www.jstor.org/stable/96399.

Schrodinger:1935:D

[Sch35d] Erwin Schrödinger. The decline. Time and Tide (University Supplement), ??(??):1757–1758, November 30, 1935.

Schrodinger:1935:GWG

[Sch35e] Erwin Schrödinger. Der Grundgedanke der Wellenmechanik. (German) [The basic idea of wave mechanics]. In Les Prix Nobel en 1933. (French) [The 1933 Nobel Prizes], pages 1–13. Norstedt & Söner, Stockholm, Sweden, 1935.

Schrodinger:1935:GSQa

[Sch35f] Erwin Schrödinger. Die gegenwärtige Situation in der Quantenmechanik. (German) [The present situation in quantum mechanics]. Naturwissenschaften, 23(48):807–812, November 1935. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904

(electronic). URL http://www.springerlink.com/content/
l50v426520375016/fulltext.pdf. English translation in [Tri80].

Schrodinger:1935:DPR

[Sch35g] Erwin Schrödinger. Discussion of probability relations between separated systems. Mathematical proceedings of the Cambridge Philosophical Society, 31(4):555–563, October 1, 1935. CO-DEN MPCPCO. ISSN 0305-0041 (print), 1469-8064 (electronic).

URL http://www.jiscjournalarchives.ac.uk/openurl.html?
ref=cup/PSP31_04/S0305004100013554a.pdf.

Schrodinger:1935:ERF

[Sch35h] Erwin Schrödinger. Equality and relativity of freedom. The Listener, ??(??):952–953, June 5, 1935. German: Gleichheit und Relativität der Freiheit.

Schrodinger:1935:ESA

[Sch35i] Erwin Schrödinger. Erwin Schrödinger [autobiographical sketch]. In Les Prix Nobel en 1933. (French) [The 1933 Nobel Prizes], pages 86–88. Norstedt & Söner, Stockholm, Sweden, 1935.

Schrodinger:1935:FPM

[Sch35j] Erwin Schrödinger. Frank, Ph. and von Mises, R.: Die Differentialund Integralgleichungen der Mechanik und Physik 2, Physikalischer Teil [Book review]. (German) [Frank, Ph. and von Mises, R.: The differential and Integral equations of mechanics and physics 2, physical aspects]. Naturwissenschaften, 23(29):516–517, July 1935. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Schrodinger:1935:GRF

[Sch35k] Erwin Schrödinger. Gleichheit und Relativität der Freiheit. (German) [Equality and relativity of freedom]. Die Auslese aus Zeitschriften des In- und Auslandes, 9(??):650–652, ???? 1935. Abbreviated translation of [Sch35h].

Schrodinger:1935:NMO

[Sch35l] Erwin Schrödinger. La nueva mecánica ondulatoria. (Spanish) [The new wave mechanics]. In Cursos de la Universidad Internacional de Verano en Santander. (Spanish) [Courses of the International Summer University in Santander], volume 1, pages 1–73. Signo, Madrid, 1935. Translated to Spanish by X. Zubiri.

Schrodinger:1935:PTS

[Sch35m] Erwin Schrödinger. Phenomenological theory of supra-conductivity. Nature, 137(3472):824, March 2, 1935. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v137/n3472/pdf/137824a0.pdf.

Schrodinger:1935:QRA

[Sch35n] Erwin Schrödinger. Quelques remarques au sujet des bases de la connaissance scientifique. (French) [Some remarks on the foundations of scientific knowledge]. Scientia (Rivista di Scienza), 57(??):

181–191, ???? 1935. URL http://amshistorica.unibo.it/7.

Schrodinger:1935:SAP

[Sch35o] Erwin Schrödinger. Science, art and play. The Philosopher (London), 13(??):11–18, ???? 1935.

Schrodinger:1935:SLV

[Sch35p] Erwin Schrödinger. ¿Son lineales las verdaderas ecuaciones del campo electromagnético?. (Spanish) [Are the true linear field equations electromagnetic?]. Anales de la Sociedad Española de Fisica y Quimica, 33(??):511–517, ???? 1935.

Schrodinger:1935:TGT

[Sch35q] Erwin Schrödinger. Trinkspruch. (German) [Toast]. In Les Prix Nobel en 1933. (French) [The 1933 Nobel Prizes], pages 79–81. Norstedt & Söner, Stockholm, Sweden, 1935.

Schrodinger:1936:IFW

[Sch36a] Erwin Schrödinger. Indeterminism and free will. Nature, 138 (3479):13–14, July 4, 1936. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v138/n3479/pdf/138013a0.pdf.

Schrodinger:1936:PTS

[Sch36b] Erwin Schrödinger. Phenomenological theory of supra-conductivity. Nature, 137(3472):824, May 16, 1936. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v137/n3472/pdf/137824a0.pdf.

Schrodinger:1936:PRB

[Sch36c] Erwin Schrödinger. Probability relations between separated systems. Mathematical proceedings of the Cambridge Philosophical Society, 32(3):446–452, October 1, 1936. CODEN MPCPCO. ISSN 0305-0041 (print), 1469-8064 (electronic).

Schrodinger:1937:RTP

[Sch37a] Erwin Schrödinger. Relativity theory of protons and electrons. Nature, 140(3548):742–744, October 30, 1937. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v140/n3548/pdf/140742a0.pdf.

Schrodinger:1937:EWS

[Sch37b] Erwin Schrödinger. Über die Eddington’sche Welttheorie — Sulla teoria dell’universo di Eddington. [Extract]. (German / Italian) [On Eddington’s world theory]. In Celebrazione del secondo centenario della nascita di Luigi Galvani. Congresso della Societa Italiana di Fisica, 17–20 Octobre 1937, A. XV, N. 12. (Italian) [Celebrating the bicentenary of the birth of Luigi Galvani. Congress of the Italian Society of Physics, 17–20 October 1937, A. XV, No. 12], page ?? Societa Italiana di Fisica, Bologna, Italy, 1937.

Schrodinger:1938:MWG

[Sch38a] Erwin Schrödinger. Die Mehrdeutigkeit der Wellenfunktion. (German) [The ambiguity of the wave function]. Annalen der Physik (Series 5), 424(1-2):49–55, ???? 1938. CODEN ANPYA2. ISSN 1521-3889.

Schrodinger:1938:ESR

[Sch38b] Erwin Schrödinger. Eigenschwingungen des sphärischen raumes. (German) [Natural oscillations of spherical space]. Commentationes Pontificiae Academiae Scientiarum, 2(9):321–364, ???? 1938. CODEN ????

Schrodinger:1938:MFP

[Sch38c] Erwin Schrödinger. Mean free path of protons in the Universe. Nature, 141(3566):410, March 5, 1938. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v141/n3566/pdf/141410a0.pdf.

Schrodinger:1938:TDM

[Sch38d] Erwin Schrödinger. Sur la théorie du monde d’Eddington. (French) [On the world theory of Eddington]. Il Nuovo Cimento (8), 15 (4):246–254, ???? 1938. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL http://www.springerlink.

com/content/9876040w8p143078/fulltext.pdf.

Schrodinger:1939:NNR

[Sch39a] Erwin Schrödinger. Nature of the nebular red-shift. Nature, 144 (3648):593, September 30, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v144/n3648/pdf/144593a0.pdf.

Schrodinger:1939:PVE

[Sch39b] Erwin Schrödinger. The proper vibrations of the expanding Universe. Physica, 6(7–12):899–912, July/December 1939. CO-DEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0031891439900911.

Schrodinger:1940:PPS

[Sch40a] E. Schrödinger. The philosophy of physical science. Nature, 145 (3672):402–403, March 16, 1940. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v145/n3672/pdf/145402a0.pdf.

Schrodinger:1940:EPP

[Sch40b] Erwin Schrödinger. Eddington, A. S.: The Philosophy of Physical Science (Tarner Lectures, 1938) [Book review]. Nature, 145(3672): 402–403, March 16, 1940. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v145/n3672/pdf/145402a0.pdf.

Schrodinger:1940:GTR

[Sch40c] Erwin Schrödinger. The General Theory of Relativity and wave mechanics. Wis- en natuurkundig Tijdschrift, 10(??):2–9, ???? 1940. CODEN WNTJAF. ISSN 0372-7416.

Schrodinger:1940:MDE

[Sch41a] Erwin Schrödinger. Maxwell’s and Dirac’s equations in the expanding Universe. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 46(??):25–47, ???? 1940–1941. CODEN PRIAAK. ISSN 0035-8975. URL http://

www.jstor.org/stable/20490746.

Schrodinger:1940:MDQ

[Sch41b] Erwin Schrödinger. A method of determining quantum-mechanical eigenvalues and eigenfunctions. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 46(??): 9–16, ???? 1940–1941. CODEN PRIAAK. ISSN 0035-8975. URL

http://www.jstor.org/stable/20490744.

Schrodinger:1941:FSS

[Sch41c] Erwin Schrödinger. Further studies on solving eigenvalue problems by factorization. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 46(??):183–206, ???? 1941. CODEN PRIAAK. ISSN 0035-8975. URL http://

www.jstor.org/stable/20490756.

Schrodinger:1941:SUR

[Sch41d] Erwin Schrödinger. La structure de l’univers en relation avec la structure corpusculaire. (French) [The structure of the Universe in relation to corpuscular structure]. Bulletin de la Société Philomathique de Paris, 123(??):26–30, ???? 1941.

Schrodinger:1941:PRB

[Sch41e] Erwin Schrödinger. Prof. Richard Bär [Obituary]. Nature, 147 (3731):536, May 3, 1941. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v147/n3731/pdf/147536b0.pdf.

Schrodinger:1941:FHE

[Sch42a] Erwin Schrödinger. The factorization of the hypergeometric equation. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 47(??):53–54, ???? 1941–1942. CO-DEN PRIAAK. ISSN 0035-8975. URL http://www.jstor.org/

stable/20488434.

Schrodinger:1941:NLO

[Sch42b] Erwin Schrödinger. Non-linear optics. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 47 (??):77–117, ???? 1941–1942. CODEN PRIAAK. ISSN 0035-8975.

URL http://www.jstor.org/stable/20488437.

Schrodinger:1941:SWE

[Sch42c] Erwin Schrödinger. On the solutions of wave equations for nonvanishing rest-mass including a source-function. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 47(??):1–23, ???? 1941–1942. CODEN PRIAAK. ISSN

0035-8975. URL http://www.jstor.org/stable/20488431.

Schrodinger:1942:DSP

[Sch43a] Erwin Schrödinger. Dynamics and scattering-power of Born’s electron. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 48(??):91–122, ???? 1942–1943. CODEN PRIAAK. ISSN 0035-8975. URL http://www.jstor.

org/stable/20488443.

Schrodinger:1943:PTT

[Sch43b] Erwin Schrödinger. Pentads, tetrads, and triads of meson-matrices. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 48(??):135–146, ???? 1943. CODEN PRI-

AAK. ISSN 0035-8975. URL http://www.jstor.org/stable/
20488445.

Schrodinger:1943:ESP

[Sch44a] Erwin Schrödinger. The Earth’s and the Sun’s permanent magnetic fields in the unitary field theory. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 49(??): 135–148, ???? 1943–1944. CODEN PRIAAK. ISSN 0035-8975.

URL http://www.jstor.org/stable/20488455.

Schrodinger:1943:GUT

[Sch44b] Erwin Schrödinger. The general unitary theory of the physical fields. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 49(??):43–58, ???? 1943–1944. CO-DEN PRIAAK. ISSN 0035-8975. URL http://www.jstor.org/

stable/20488450.

Schrodinger:1943:NES

[Sch44c] Erwin Schrödinger. A new exact solution in non-linear optics (twowave-system). Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 49(??):59–66, ???? 1943– 1944. CODEN PRIAAK. ISSN 0035-8975. URL http://www.

jstor.org/stable/20488451.

Schrodinger:1943:SMM

[Sch44d] Erwin Schrödinger. Systematics of meson-matrices. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 49(??):29–42, ???? 1943–1944. CODEN PRIAAK. ISSN

0035-8975. URL http://www.jstor.org/stable/20488449.

Schrodinger:1943:UTF

[Sch44e] Erwin Schrödinger. The union of the three fundamental fields (gravitation, meson, electromagnetism). Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 49 (??):275–287, ???? 1943–1944. CODEN PRIAAK. ISSN 0035-

8975. URL http://www.jstor.org/stable/20488463.

Schrodinger:1944:ACP

[Sch44f] Erwin Schrödinger. The affine connexion in physical field theories. Nature, 153(3889):572–575, May 13, 1944. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v153/n3889/pdf/153572a0.pdf.

Schrodinger:1944:PCU

[Sch44g] Erwin Schrödinger. The point charge in the unitary field theory. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 49(??):225–235, ???? 1944. CODEN PRI-

AAK. ISSN 0035-8975. URL http://www.jstor.org/stable/
20488459.

Schrodinger:1944:RFA

[Sch44h] Erwin Schrödinger. Rate of n-fold accidental coincidences. Nature, 153(3889):592–593, May 13, 1944. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v153/n3889/pdf/153592b0.pdf.

Schrodinger:1944:SLN

[Sch44i] Erwin Schrödinger. The statistical law in nature. Nature, 153 (3893):704–705, July 10, 1944. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v153/n3893/pdf/153704a0.pdf.

Schrodinger:1944:STC

[Sch44j] Erwin Schrödinger, editor. Statistical thermodynamics: course of seminar lectures. Delivered in January–March 1944, at the School of Theoretical Physics. Institiúid Ard-Léinn Bhaile Átha Cliath (Dublin Institute for Advanced Studies), Dublin, Ireland, 1944. 2 + 135 pp. LCCN QC311 .S38.

Schrodinger:1944:UFT

[Sch44k] Erwin Schrödinger. Unitary field theory: Conservation identities and relation to Weyl and Eddington. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 49(??): 237–244, ???? 1944. CODEN PRIAAK. ISSN 0035-8975. URL

http://www.jstor.org/stable/20488460.

Schrodinger:1944:WLP

[Sch44l] Erwin Schrödinger. What is life? The physical aspect of the living cell. Cambridge University Press, Cambridge, UK, 1944. viii + 91 pp. LCCN QH331 .S355 1945. Based on lectures delivered under the auspices of the Institute at Trinity College, Dublin, in February 1943.

Schrodinger:1945:AS

[Sch45a] Erwin Schrödinger. Austrian science. In Science in Austria. Leaflet, presented on the occasion of the meeting of British and Austrian scientists in support of the restoration of science in Austria, pages 12–13. Association of Austrian Engineers, Chemists and Scientific Workers in Great Britain, London, UK, 1945.

Schrodinger:1945:DAC

[Sch45b] Erwin Schrödinger. On distant affine connection. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 50(??):143–154, ???? 1945. CODEN PRIAAK. ISSN

0035-8975. URL http://www.jstor.org/stable/20520639.

Schrodinger:1945:OWGa

[Sch45c] Erwin Schrödinger. Österreichische Wissenschaft. (German) [Austrian science]. Kulturelle Schriftenreihe des FAM (Free Austrian Movement), London, ??(??):1–3, ???? 1945.

Schrodinger:1945:OWGb

[Sch45d] Erwin Schrödinger. Österreichische Wissenschaft. (German) [Austrian science]. Austro American Tribune. Anti-Nazi Monthly, 4 (??):7, ???? 1945. Slightly modified version of [Sch45c].

Schrodinger:1945:WLP

[Sch45e] Erwin Schrödinger. What is life? The physical aspect of the living cell. Cambridge University Press, Cambridge, UK, 1945. viii + 91 pp. LCCN QH331 .S355 1945. Based on lectures delivered under the auspices of the Institute at Trinity College, Dublin, in February 1943.

Schrodinger:1946:AFM

[Sch46a] Erwin Schrödinger. Affine Feldtheorie und Meson. (German) [Affine field theory and meson]. Verhandlungen der Schweizerischen Naturforschenden Gesellschaft, 126(??):53–61, ???? 1946. CODEN VSNGAY. ISSN 0080-7362. URL http://retro.seals.

ch/digbib/view2?pid=sng-005:1946:126::56.

Schrodinger:1946:STC

[Sch46b] Erwin Schrödinger. Statistical thermodynamics: a course of seminar lectures. Cambridge University Press, Cambridge, UK, 1946. 88 pp. LCCN ????

Schrodinger:1947:GNG

[Sch47a] Erwin Schrödinger. Der Geist der Naturwissenschaft. (German) [The spirit of science]. In Eranos-Jahrbuch 1946, volume 14, pages 491–520. Rhein, Zürich, Switzerland, 1947.

Schrodinger:1947:FAF

[Sch47b] Erwin Schrödinger. The final affine field laws. I. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 51(??):163–171, ???? 1947. CODEN PRIAAK. ISSN

0035-8975. URL http://www.jstor.org/stable/20488479.

Schrodinger:1947:FTPa

[Sch47c] Erwin Schrödinger. The foundation of the theory of probability: I. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 51(??):51–66, ???? 1947. CODEN PRI-

AAK. ISSN 0035-8975. URL http://www.jstor.org/stable/
20488471.

Schrodinger:1947:FTPb

[Sch47d] Erwin Schrödinger. The foundation of the theory of probability: II. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 51(??):141–146, ???? 1947. CODEN PRI-

AAK. ISSN 0035-8975. URL http://www.jstor.org/stable/
20488476.

Schrodinger:1947:RBM

[Sch47e] Erwin Schrödinger. The relation between metric and affinity. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 51(??):147–150, ???? 1947. CODEN PRI-

AAK. ISSN 0035-8975. URL http://www.jstor.org/stable/
20488477.

Schrodinger:1945:GAF

[Sch48a] Erwin Schrödinger. The general affine field laws. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 51(??):41–50, ???? 1945–1948. CODEN PRIAAK. ISSN

0035-8975. URL http://www.jstor.org/stable/20488470.

Schrodinger:1945:PPN

[Sch48b] Erwin Schrödinger. Probability problems in nuclear chemistry. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 51(??):1–8, ???? 1945–1948. CODEN PRI-

AAK. ISSN 0035-8975. URL http://www.jstor.org/stable/
20488467.

Schrodinger:1948:JQG

[Sch48c] Erwin Schrödinger. 2400 Jahre Quantentheorie. (German) [2400 years of quantum theory]. Annalen der Physik (Series 6), 438(1): 43–48, ???? 1948. CODEN ANPYA2. ISSN 1521-3889.

Schrodinger:1948:BWN

[Sch48d] Erwin Schrödinger. Die Besonderheit des Weldbilds der Naturwissenschaft. (German) [The specificity of the scientific world view].

Acta Physica Austriaca, 1(??):201–245, ???? 1948. CODEN APASAP. ISSN 0001-6713.

Schrodinger:1948:FAFa

[Sch48e] Erwin Schrödinger. The final affine field laws. II. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 51(??):205–216, ???? 1948. CODEN PRIAAK. ISSN

0035-8975. URL http://www.jstor.org/stable/20488482.

Schrodinger:1948:FAFb

[Sch48f] Erwin Schrödinger. The final affine field laws. III. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 52(??):1–9, ???? 1948. CODEN PRIAAK. ISSN 0035-

8975. URL http://www.jstor.org/stable/20488487.

Schrodinger:1948:TPG

[Sch48g] Erwin Schrödinger. Theoretiker und Praktiker. (German) [Theorists and practitioners]. Die Furche [The furrow], ??(??):??, March 27, 1948.

Schrodinger:1949:DET

[Sch49a]  Erwin Schrödinger. Do electrons think?  BBC audio

broadcast and typewritten transcript., 1949. URL https:/

/fedora.phaidra.univie.ac.at/fedora/get/o:168238/bdef:
Asset/view;   https://quantumlifescience.wordpress.com/
2014/04/10/do-electrons-think-erwin-schrodinger-bbc-1949/
;   https://www.youtube.com/watch?v=hCwR1ztUXtU.

Schrodinger:1949:GGP

[Sch49b] Erwin Schrödinger. Gedichte. (German) [Poetry]. Küpper vorm. Bondi, Godesberg, Germany, 1949. ???? pp. LCCN ????

Schrodinger:1950:STS

[Sch50a] Erwin Schrödinger. Space–time structure. Cambridge University Press, Cambridge, UK, 1950. ISBN 0-521-31520-4. 119 pp. LCCN ????

Schrodinger:1950:FUZ

[Sch50b] Erwin Schrödinger. The Future of Understanding — Die Zukunft des Weltverstehens, 1950. Three BBC Talks on September 16, 23, 30, (1950).

Schrodinger:1950:WEP

[Sch50c] Erwin Schrödinger. What is an elementary particle? Endeavour, 9(??):109–116, ???? 1950. CODEN ENDEAS. ISSN 0141-3058, 0013-7162, 0160-9327.

Schrodinger:1950:I

[Sch51a] Erwin Schrödinger. Irreversibility. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 53(??): 189–195, ???? 1950–1951. CODEN PRIAAK. ISSN 0035-8975.

URL http://www.jstor.org/stable/20488517.

Schrodinger:1951:CPC

[Sch51b] Erwin Schrödinger. A combinatorial problem in counting cosmic rays. Proceedings of the Physical Society, London, Section A, 64 (11):1040–1041, November 1951. CODEN PPSAAM. ISSN 0370-1298 (print), 1747-3829 (electronic).

Schrodinger:1951:SHP

[Sch51c] Erwin Schrödinger. Science and humanism: physics in our time. Cambridge University Press, Cambridge, UK, 1951. ix + 68 pp. LCCN ????

Schrodinger:1951:WEG

[Sch51d] Erwin Schrödinger. Was ist ein Elementarteilchen?. (German) [What is an elementary particle?]. Die Pyramide, ??(??):2–4; 24– 25; 44–46, ???? 1951.

Schrodinger:1951:WEP

[Sch51e] Erwin Schrödinger. What is an elementary particle? In The Smithsonian Institution’s Annual Report, pages 183–196. United States Government Printing Office, Washington, DC, USA, 1951.

Schrodinger:1951:DIA

[Sch52a] Erwin Schrödinger. On the differential identities of an affinity. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 54(??):79–85, ???? 1951–1952. CODEN PRI-

AAK. ISSN 0035-8975. URL http://www.jstor.org/stable/
20488525.

Schrodinger:1952:TQJa

[Sch52b] Erwin Schrödinger. Are there quantum jumps? Part I. British Journal for the Philosophy of Science, 3(10):109–123, August ??, 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.oxfordjournals.org/content/III/
10/109.full.pdf+html;   http://www.jstor.org/stable/685552.

Schrodinger:1952:TQJb

[Sch52c] Erwin Schrödinger. Are there quantum jumps? Part II. British Journal for the Philosophy of Science, 3(11):233–242, November ??, 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537

(electronic). URL http://bjps.oxfordjournals.org/content/
III/11/233.full.pdf+html;   http://www.jstor.org/stable/
685266.

Schrodinger:1952:DNE

[Sch52d] Erwin Schrödinger. Dirac’s new electrodynamics. Nature, 169 (4300):538, March 29, 1952. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v169/n4300/pdf/169538a0.pdf.

Schrodinger:1952:STC

[Sch52e] Erwin Schrödinger. Statistical thermodynamics: a course of seminar lectures delivered in January–March 1944, at the School of Theoretical Physics, Dublin Institute for Advanced Studies. Cambridge University Press, Cambridge, UK, second edition, 1952. 95 pp. LCCN ????

Schrodinger:1952:STG

[Sch52f] Erwin Schrödinger. Statistische Thermodynamik. (German) [Statistical Thermodynamics]. Barth, Leipzig, DDR, 1952. ???? pp. LCCN ????

Schrodinger:1952:WMV

[Sch52g] Erwin Schrödinger. Was ist Materie?: Vortrag vom 9. Dezember 1952. (German) [What is matter? Lecture of 9 December 1952]. TO BE FOUND., 1952.

Schrodinger:1953:RQT

[Sch53a] E. Schrödinger. Relativistic quantum theory. British Journal for the Philosophy of Science, IV(16):328–329, February 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

URL http://bjps.oxfordjournals.org/content/IV/16/328.
full.pdf+html;   http://www.jstor.org/stable/686081.

Schrodinger:1952:RFR

[Sch53b] Erwin Schrödinger. Relativistic Fourier reciprocity and the elementary masses. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 55(??):29–50, ???? 1952– 1953. CODEN PRIAAK. ISSN 0035-8975. URL http://www.

jstor.org/stable/20488546.

Schrodinger:1953:GTR

[Sch53c] Erwin Schrödinger. The General Theory of Relativity and wave mechanics. In Appleton et al. [ABdB^{+}53], pages 65–74. LCCN QC71 .S35.

Schrodinger:1953:LAM

[Sch53d] Erwin Schrödinger. L’image actuelle de la matière (Sommaire) — Unsere Vorstellung von der Materie. (German) [The current view of matter]. In L’homme devant la science, Texte des conférences et des entretiens organisés par les rencontres internationales de Genève 1952. (French) [The human face of science, lectures and text interviews organized by International meetings, Geneva, 1952], pages 31–54. Baconnière, Neuchâtel, Switzerland, 1953.

Schrodinger:1953:UVM

[Sch53e] Erwin Schrödinger. Unsere Vorstellung von der Materie. (German) [Our conception of matter]. Merkur, 7(??):131–145, ???? 1953.

Schrodinger:1953:WM

[Sch53f]  Erwin Schrödinger. What is matter?  Scientific Amer-

ican, 189(3):52–57, September 1953. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.nature.com/scientificamerican/journal/v189/n3/pdf/
scientificamerican0953-52.pdf. Abbreviated version of “Our

Conception of Matter”.

Schrodinger:1954:BPG

[Sch54a] Erwin Schrödinger. Begegnung; parabel. (German) [Meeting; parable]. Kontinente. Wege und Probleme der Gegenwart, 7(??):28, ???? 1954.

Schrodinger:1954:ECC

[Sch54b] Erwin Schrödinger. Electric charge and current engendered by combined Maxwell–Einstein-fields. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 56(??):

13–21, ???? 1954. CODEN PRIAAK. ISSN 0035-8975. URL

http://www.jstor.org/stable/20488559.

Schrodinger:1954:MLA

[Sch54c] Erwin Schrödinger. Measurement of length and angle in quantum mechanics. Nature, 173(4410):442, March 6, 1954. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v173/
n4401/pdf/173442a0.pdf.

Schrodinger:1954:NG

[Sch54d] Erwin Schrödinger. Nature and the Greeks, volume 1948 of Shearman lectures, delivered at University College, London. Cambridge University Press, Cambridge, UK, 1954. 97 pp. LCCN B118 .S36.

Schrodinger:1954:OIW

[Sch54e] Erwin Schrödinger. Orientierung im Weltall; Erdalter und Weltalter; Die Kohlenstoff-Uhr; Raum und Zeit. (German) [Orientation in space, age of the Earth and the Universe, the carbon clock, time and space]. In Orientierung im Weltall (Das Internationale Forum. Berichte und Stellungnahmen. 3). (German) [Orientation in space (The International Forum. Reports and statements. 3)], pages 7–31. Fontana-Verlag, Zürich, Switzerland, 1954.

Schrodinger:1954:RQT

[Sch54f] Erwin Schrödinger. Relativistic quantum theory. British Journal for the Philosophy of Science, 4(16):328–329, February ??, 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

URL http://www.jstor.org/stable/686081. Extract from a pri-

vate letter.

Schrodinger:1954:STS

[Sch54g] Erwin Schrödinger. Space–time structure. Cambridge University Press, Cambridge, UK, 1954. viii + 119 pp. LCCN ???? Reprint, with corrections, of [Sch50a].

Schrodinger:1954:SS

[Sch54h] Erwin Schrödinger. The spirit of science. In Buonaiuti et al. [B^{+}54], pages 322–341. ISBN 0-691-01841-3 (paperback), 0-691-09736-4. LCCN BF1003 .S73 1982. URL http://www.loc.

gov/catdir/description/prin021/82245044.html;   http:/
/www.loc.gov/catdir/toc/prin031/82245044.html. Reprinted

in 1982.

Schrodinger:1954:UVM

[Sch54i] Erwin Schrödinger. Unsere Vorstellung von der Materie. (German) [Our understanding of matter]. Naturwissenschaftliche Rundschau, 7(??):277–282, ???? 1954. CODEN NARSAC. ISSN 0028-1050. Abbreviated version of [Sch53d].

Schrodinger:1955:AGAb

[Sch55a] Erwin Schrödinger. Atomenergie. (German) [Atomic energy]. Sie und er, ??(??):20–22, January 27, 1955.

Schrodinger:1955:AGAa

[Sch55b] Erwin Schrödinger. Die Atomisten. (German) [The atomists]. Merkur, 9(??):815–824, ???? 1955.

Schrodinger:1955:BWG

[Sch55c] Erwin Schrödinger. Die Bedeutung der Wellenmechanik. (German) [The meaning of wave mechanics]. In Louis de Broglie und die Physiker. (German) [Louis de Broglie and the Physicist], pages 18–25. Claassen, Hamburg, Germany, 1955.

Schrodinger:1955:SMO

[Sch55d] Erwin Schrödinger. La signification de la mécanique ondulatoire. (French) [The meaning of wave mechanics]. In Louis de Broglie, Physicien et Penseur. (French) [Louis de Broglie, Physicist and Thinker], pages 16–32. Albin Michel, Paris, France, 1955.

Schrodinger:1955:PE

[Sch55e] Erwin Schrödinger. The philosophy of experiment. Il Nuovo Cimento (10), 1(10):5–15, ???? 1955. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL http://www.

springerlink.com/content/eu217373q03p027n/fulltext.pdf.

Schrodinger:1955:TRB

[Sch55f] Erwin Schrödinger. A thermodynamic relation between frequencyshift and broadening. Il Nuovo Cimento (10), 1(10):63–69, ???? 1955. CODEN NUCIAD. ISSN 0029-6341 (print),

1827-6121 (electronic). URL http://www.springerlink.com/
content/fh60g6uh31213262/fulltext.pdf.

Schrodinger:1955:WESa

[Sch55g] Erwin Schrödinger. The wave equation for spin 1 in Hamiltonian form [I]. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 229(1176):39–43, April 5, 1955. CODEN PRLAAZ. ISSN 0080-4630. URL http://www.

jstor.org/stable/99722.

Schrodinger:1955:WESb

[Sch55h] Erwin Schrödinger. The wave equation for spin 1 in Hamiltonian form. II. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 232(1191):435–447, November 22, 1955. CODEN PRLAAZ. ISSN 0080-4630. URL http://www.

jstor.org/stable/99821.

Schrodinger:1956:KAA

[Sch56a] Erwin Schrödinger. Die Krise des Atombegriffs (Antrittsrede Wien 1956). (German) [The crisis of the atomic concept (inaugural address, Vienna, 1956)]. Manuscript in National Library of Ireland., March 9, 1956.

Schrodinger:1956:NGK

[Sch56b] Erwin Schrödinger. Die Natur und die Griechen: Kosmos und Physik. (German) [Nature and the Greeks: Cosmos and Physics]. Rowohlt, Reinbek bei Hamburg, Germany, 1956. 145 pp. LCCN ????

Schrodinger:1956:EU

[Sch56c] Erwin Schrödinger. Expanding universes. Cambridge University Press, Cambridge, UK, 1956. viii + 93 pp. LCCN QB500 .S36.

Schrodinger:1956:FGE

[Sch56d] Erwin Schrödinger. Festrede, gehalten bei der Eröffnung der fünften Weltkraftkonferenz, Wien 1956. (German) [Speech, held at the opening of the fifth World Power Conference, Vienna 1956]. In Fünfte Weltkraftkonferenz, Wien 1956, Gesamtbericht. (German) [Fifth World Power Conference, Vienna 1956, Global Report], volume I, pages 277–283. ????, ????, 1956.

Schrodinger:1956:WLOa

[Sch56e] Erwin Schrödinger. What is life? and other scientific essays. Cambridge University Press, Cambridge, UK, 1956. viii + 263 pp. LCCN QH331 .S355 1956.

Schrodinger:1956:WLOb

[Sch56f] Erwin Schrödinger. What is life? and other scientific essays. Doubleday, Garden City, NY, 1956. ???? pp. LCCN QH331 .S355

Schrodinger:1957:AGA

[Sch57a] Erwin Schrödinger. Die Atomtheorie. (German) [Atomic theory]. In Lebendige Stadt. Almanach 1957. (German) [Vibrant city. 1957 almanac], pages 157–161. Amt für Kultur und Volksbildung der Stadt Wien, Vienna, Austria, ???? 1957.

Schrodinger:1957:STMa

[Sch57b] Erwin Schrödinger. Science theory and man. George Allen and Unwin ltd., London, UK, 1957. xxiv + 2 + 27–223 pp. LCCN Q171

.S3 1957a. URL http://www.zbp.univie.ac.at/schrodinger/
ebibliographie/publications.htm. Formerly published under

the title Science and the human temperament. Translated from German by James Vincent Murphy W. H. Jonston. Foreword by Ernest Rutherford.

Schrodinger:1957:STMb

[Sch57c] Erwin Schrödinger. Science theory and man. Dover, New York, NY, USA, 1957. ???? pp. LCCN ????

Schrodinger:1957:GSM

[Sch57d] Erwin Schrödinger. Zur Geistesgeschichte der Stellung der Menschen. (German) [On the intellectual history of the position of people]. Der Mittelschullehrer und die Mittelschule, 6(??):280–282, ???? 1957.

Schrodinger:1958:MPE

[Sch58a] Erwin Schrödinger. Might perhaps energy be a merely statistical concept? Il Nuovo Cimento (10), 9(10):162–170, ???? 1958. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121

(electronic). URL http://www.springerlink.com/content/
7h8348503v0295nx/fulltext.pdf.

Schrodinger:1958:MMT

[Sch58b] Erwin Schrödinger. Mind and matter: the Tarner Lectures, delivered at Trinity College, Cambridge in October 1956. Cambridge University Press, Cambridge, UK, 1958. vii + 104 pp. LCCN ????

Schrodinger:1959:CUG

[Sch59a] Erwin Schrödinger. Commilitonen! upanishad. (German) [Fellow students! upanishad]. Quo via fert? Wohin führt der Weg, 10, Wien: Maturanten des Döblinger Gymnasiums, Wien XIX. [Short letter and unpublished poem]., 1959.

Schrodinger:1959:LAD

[Sch59b] Erwin Schrödinger. L’immagine attuale della materia. (Italian) [The current view of matter]. In Discussione sulla Fisica Moderna. (Italian) [Discussions on Modern Physics], pages 35–57. Boringhieri, Torino, Italy, 1959. Reprinted in 1960, 1964 and 1980.

Schrodinger:1960:CJC

[Sch60a] Erwin Schrödinger. Co to jest czastka elementarna?. (Polish) [What is an elementary particle?]. Postepy Fizyki [Progress in Physics], 11(??):135–150, ???? 1960.

Schrodinger:1960:WM

[Sch60b] Erwin Schrödinger. What is matter? In Robert K. Fitzgerel and William F. Kieffer, editors, Supplementary Readings for Chemical Bond Approach, pages 2–8. Division of Chemical Education of the American Chemical Society, Easton, PA, USA, 1960. Reprinted from Journal of Chemical Education, 1935–1960.

Schrodinger:1961:EKH

[Sch61a] Erwin Schrödinger. A 2400 éves kvantumelmélet. (Hungarian) [2400 years of quantum theory]. Fizikai Szemle (Budapest), 11(4):101–104, ???? 1961. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic). URL http://www.kfki.hu/

fszemle/archivum/fsz6104/.

Schrodinger:1961:AES

[Sch61b] Erwin Schrödinger. Antrittsrede von Erwin Schrödinger (1887-1961). Erwiderung von Max Planck. (German) [Inaugural speech by Erwin Schrödinger (1887-1961). Reply from Max Planck]. In Ch. Kirsten and H.-G. Körber, editors, Physiker über Physiker 2. Antrittsreden, Erwiderungen bei der Aufnahme von Physikern in die Berliner Akademie. Gedächtnisreden. 1870–1929. (German) [Physicist on physicist: second inaugural speeches, responses during the meeting of physicists in the Berlin Academy. Memorial speeches. 1870–1929], volume 8 of Studien zur Geschichte der Akademie der Wissenschaften der DDR. Akademie der Wissenschaften der DDR, Berlin, DDR, 1961.

Schrodinger:1961:OIMa

[Sch61c] Erwin Schrödinger. Our image of matter. In On Modern Physics, pages 45–66. Clarkson N. Potter, New York, NY, USA, 1961.

Schrodinger:1961:OIMb

[Sch61d] Erwin Schrödinger. Our image of matter. In On Modern Physics, pages 45–66. Orion Press, London, UK, 1961.

Schrodinger:1961:RNG

[Sch61e] Erwin Schrödinger. Religion und Naturwissenschaft. (German) [Religion and science]. Physikalische Blätter, 17(3):105–110, March 1961. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
phbl.19610170301/abstract. Short part of [Sch55b].

Schrodinger:1961:SHP

[Sch61f] Erwin Schrödinger. Science and humanism: physics in our time. Cambridge University Press, Cambridge, UK, 1961. 68p. pp. LCCN ????

Schrodinger:1962:OIM

[Sch62a] Erwin Schrödinger. Our image of matter. In On Modern Physics, pages 45–66. Crowell-Collier Publishing Company, New York, NY, USA, 1962.

Schrodinger:1962:WLP

[Sch62b] Erwin Schrödinger. What is life? The physical aspect of the living cell. Cambridge University Press, Cambridge, UK, 1962. viii + 91 pp. LCCN QH331 .S355 1962. Based on lectures delivered under the auspices of the Institute at Trinity College, Dublin, in February 1943.

Schrodinger:1963:WGW

[Sch63] Erwin Schrödinger. Die Wellenmechanik. (German) [Wave Mechanics]. Battenberg, Stuttgart, West Germany, 1963. ???? pp. LCCN ????

Schrodinger:1964:MVW

[Sch64a] Erwin Schrödinger. My view of the world. Cambridge University Press, Cambridge, UK, 1964. vii + 109 pp. LCCN ????

Schrodinger:1964:OIM

[Sch64b] Erwin Schrödinger. Our image of matter. In Rapport and Wright [RW64], pages 230–246.

Schrodinger:1966:GNG

[Sch66a] Erwin Schrödinger. Der Geist der Naturwissenschaft. (German) [The spirit of science]. In Gibt es Grenzen der Naturforschung? (German) [Are there limits to scientific research], volume 253, pages 15–36. Herder-Bücherei, Freiburg, Germany; Basel, Switzerland; Vienna, Austria, 1966.

Schrodinger:1966:HBJ

[Sch66b] Erwin Schrödinger. A Heisenberg–Born–Jordan-féle kvantummechanika viszonya az enyémhez. (Hungarian) [The relation of Heisenberg–Born–Jordan quantum mechanics to mine]. Magyar Fizikai Folyóirat, 14(??):359–374, ???? 1966. CODEN MGFFAC. ISSN 0025-0104.

Schrodinger:1967:WLP

[Sch67] Erwin Schrödinger. What is life? The physical aspect of the living cell & Mind and matter. Cambridge University Press, Cambridge, UK, 1967. 178 pp. LCCN QH331 .S355 1967. Based on lectures delivered under the auspices of the Institute at Trinity College, Dublin, in February 1943.

Schrodinger:1970:VTH

[Sch70] Erwin Schrödinger. Válogatott tanulmányok. (Hungarian) [Selected studies]. Gondolat, Budapest, Hungary, 1970. ???? pp. LCCN ????

Schrodinger:1975:WM

[Sch75a] Erwin Schrödinger. 6. wave mechanics. In Mehra [Meh75], chapter 6, pages 149–151. ISBN 90-277-0635-2. LCCN QC1.S792 M43.

Schrodinger:1975:FUD

[Sch75b] Erwin Schrödinger. Fragment from an unpublished dialogue of Galileo. Hermathena. A Dublin University Review, 119(??):74–77, ???? 1975.

Schrodinger:1977:QEP

[Sch77] E. Schrödinger. Quantization as an eigenvalue problem. 1. Uspekhi Fizicheskikh Nauk, 122(4):621–632, ???? 1977. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic).

Schrodinger:1978:AWC

[Sch78] Erwin Schrödinger. Abhandlungen zur Wellenmechanik. Collected papers on wave mechanics. Chelsea Pub. Co, New York, NY, USA, second edition, 1978. ISBN 0-8284-0302-3. xiii + 146 pp. LCCN ????

Schrodinger:1979:WNB

[Sch79] Erwin Schrödinger. Was ist ein Naturgesetz?: Beiträge zum naturwissenschaftlichen Weltbild. (German) [What is a law of nature?: Contributions to a scientific world view]. R. Oldenbourg, München, Germany, third edition, 1979. ISBN 3-486-46273-3 (paperback). 146 pp. LCCN ????

Schrodinger:1982:CPW

[Sch82a] Erwin Schrödinger. Collected papers on wave mechanics: together with his four lectures on wave mechanics. Chelsea, New York, NY, USA, third edition, 1982. ISBN 0-8284-1302-9. xiii + 207 pp. LCCN QC174.2.

Schrodinger:1982:HRK

[Sch82b] Erwin Schrödinger. Hadô rikigaku keiseishi: Shurêdingâ no shokan to shôden. (Japanese) [????]. Misuzu Shobo, Tokyo, Japan, 1982. ISBN ???? ???? pp. LCCN ????

Schrodinger:1982:I

[Sch82c] Erwin Schrödinger. Irreversibility. In P. T. Landsberg, editor, The Enigma of Time, pages 46–52. Adam Hilger Ltd., Bristol, UK; Boston, MA, USA, 1982. Reprinted in 1984 and 1985.

Schrodinger:1982:MCD

[Sch82d] Erwin Schrödinger. Ma conception du monde: le Veda d’un physicien. (French) [My view of the world: the Veda of a physicist]. Mercure de France, Paris, France, 1982. ISBN 2-7152-0088-9. 164 pp. LCCN ????

Schrodinger:1983:MVW

[Sch83] Erwin Schrödinger. My view of the world. Ox Box Press, Woodbridge, CT, USA, 1983. ISBN 0-918024-30-7, 0-918024-29-3. viii + 109 pp. LCCN ????

Schrodinger:1984:GSQ

[Sch84a] Erwin Schrödinger. Die gegenwärtige Situation in der Quantenmechanik (1935). (German) [The present situation in quantum mechanics (1935)]. In K. Baumann and R. U. Sexl, editors, Die Deutungen der Quantentheorie. (German) [The interpretations of quantum theory], volume 11 of Facetten der Physik, pages 98–129. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1984.

Schrodinger:1984:GACe

[Sch84b] Erwin Schrödinger. Gesammelte Abhandlungen = Collected papers. Verlag der Österreichischen Akademie der Wissenschaften, Vienna, Austria, 1984. ISBN 3-7001-0573-8 (set). ???? pp. LCCN QC3 .S37 1984.

Schrodinger:1984:GACa

[Sch84c] Erwin Schrödinger. Gesammelte Abhandlungen: Collected papers. Band 1, Beiträge zur statistischen Mechanik: Contributions to statistical mechanics. Österreichische Akademie der Wissenschaften, Vienna, Austria, 1984. ISBN 3-7001-0574-6, 3-528-07131-1. xii + 547 pp. LCCN ???? Foreword by Walter Thirring.

Schrodinger:1984:GACb

[Sch84d] Erwin Schrödinger. Gesammelte Abhandlungen: Collected papers. Band 2, Beiträge zur Feldtheorie: contributions to field theory. Österreichische Akademie der Wissenschaften, Vienna, Austria, 1984. ISBN 3-528-07132-X. xiii + 699 pp. LCCN ???? Foreword by Walter Thirring.

Schrodinger:1984:GACc

[Sch84e] Erwin Schrödinger. Gesammelte Abhandlungen: Collected papers. Band 3, Beiträge zur Quantentheorie: Contributions to quantum theory. Österreichische Akademie der Wissenschaften, Vienna, Austria, 1984. ISBN 3-7001-0576-2, 3-528-07133-8. xii + 761 pp. LCCN ???? Foreword by Walter Thirring.

Schrodinger:1984:GACd

[Sch84f] Erwin Schrödinger. Gesammelte Abhandlungen: Collected papers. Band 4, Allgemein wissenschaftliche und populäre Aufsätze: General scientific and popular papers. Österreichische Akademie der Wissenschaften, Vienna, Austria, 1984. ISBN 3-7001-0577-0, 3-528-07134-6. xii + 649 pp. LCCN ???? Foreword by Walter Thirring.

Schrodinger:1985:MLM

[Sch85a] Erwin Schrödinger. Mein Leben, meine Weltansicht. (German) [My life, my world view]. P. Zsolnay, Vienna, Austria, 1985. ISBN 3-552-03712-8. 181 pp. LCCN B82 .S378 1985. Foreword by Auguste Dick.

Schrodinger:1985:STS

[Sch85b] Erwin Schrödinger. Space–time structure. Cambridge science classics. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-31520-4 (paperback), 0-511-58644-2 (e-book). viii + 119 pp. LCCN QC173.59.S65 S37 1985. URL http://

ebooks.cambridge.org/chapter.jsf?bid=CBO9780511586446;
 http://www.loc.gov/catdir/description/cam023/85005695.
html;   http://www.loc.gov/catdir/toc/cam022/85005695.
html.

Schrodinger:1986:APE

[Sch86a] Erwin Schrödinger. Das arithmetische Paradoxon — Die Einheit des Bewußtseins. (German) [The arithmetic unit — The paradox of consciousness]. In H.-P. Dürr, editor, Physik und Transzendenz. Die großen Physiker unseres Jahrhunderts über ihre Begegnung mit dem Wunderbaren. (German) [Physics and transcendence. The great physicists of our Century, their encounters with the supernatural], pages 159–170. Scherz, Bern, Switzerland; München, Germany; Vienna, Austria, 1986.

Schrodinger:1986:VGG

[Sch86b] Erwin Schrödinger. Die vedantische Grundansicht. (German) [The Vedantic basic view]. In Physik und Transzendenz. Die großen Physiker unseres Jahrhunderts über ihre Begegnung mit dem Wunderbaren. (German) [Physics and transcendence. The great physicists of our Century, their encounters with the supernatural], pages 189–192. Scherz, Bern, Switzerland; München, Germany; Vienna, Austria, 1986.

Schrodinger:1986:NRG

[Sch86c] Erwin Schrödinger. Naturwissenschaft und Religion. (German) [Science and Religion]. In Physik und Transzendenz. Die großen Physiker unseres Jahrhunderts über ihre Begegnung mit dem Wunderbaren. (German) [Physics and transcendence. The great physicists of our Century, their encounters with the supernatural], pages 171–183. Scherz, Bern, Switzerland; München, Germany; Vienna, Austria, 1986.

Schrodinger:1986:QCQa

[Sch86d] Erwin Schrödinger. Qu’est-ce que la vie ?: de la physique à la biologie. (French) [What is Life?: From physics to biology]. Éditions du Seuil, Paris, France, 1986. ISBN 2-02-020223-9 (paperback). 176 pp. LCCN ???? Translation from English by Léon Keffler. Foreword by Roger Penrose. Preface by Antoine Danchin. Postscript by Claude Debru.

Schrodinger:1986:QCQb

[Sch86e] Erwin Schrödinger. Qu’est-ce que la vie ?: l’aspect physique de la cellule vivante. (French) [What is Life?: The physical aspect of the living cell]. C. Bourgois, Paris, France, 1986. ISBN 2-267-00453-4 (paperback). 240 pp. LCCN ???? Translation from English by Léon Keffler. Preface by Antoine Danchin. Postscript by Claude Debru.

Schrodinger:1986:QCQc

[Sch86f] Erwin Schrödinger. Qu’est-ce que la vie?: de la physique à la biologie. Points Sciences, Paris, France, 1986. ISBN 2-02-020223-9. 235 pp. LCCN ???? Translation from English by Léon Keffler. Foreword by Roger Penrose. Preface by Antoine Danchin. Postscript by Claude Debru.

Schrodinger:1986:STD

[Sch86g] Erwin Schrödinger. Sulla teoria del gas di Einstein. (Italian) [On Einstein’s gas theory]. In P. Bernardini, editor, La statistica quantistica e le onde di materia. (Italian) [Statistical and quantum matter waves], page ?? Bibliopolis, Napoli, Italy, 1986.

Schrodinger:1986:WMG

[Sch86h] Erwin Schrödinger. Was ist Materie?. (German) [What is matter?]. In W. Greiner and G. Wolschin, editors, Elementare Materie, Vakuum und Felder. (German) [Elementary matter, vacuum, and fields], pages 28–35. Spektrum der Wissenschaft: Verständliche Forschung, Heidelberg, West Germany, 1986. Slightly abridged version of [Sch53e].

Schrodinger:1986:WWG

[Sch86i] Erwin Schrödinger. Was ist wirklich? — Die Gründe für das Aufgeben des Dualismus von Denken und Sein oder von Geist und Materie. (German) [What is real? — The reasons for the abandonment of the Dualism of mind and spirit or of spirit and matter]. In Physik und Transzendenz. Die großen Physiker unseres Jahrhunderts über ihre Begegnung mit dem Wunderbaren. (German) [Physics and transcendence. The great physicists of our Century, their encounters with the supernatural], pages 184–188. Scherz, Bern, Switzerland; München, Germany; Vienna, Austria,

Schrodinger:1986:WLP

[Sch86j] Erwin Schrödinger. What is life?: the physical aspect of the living cell and mind and matter. Cambridge University Press, Cambridge, UK, 1967 edition, 1986. ISBN 0-521-06223-3. 178 pp. LCCN ???? Reprint of [Sch67].

Schrodinger:1987:ENG

[Sch87a] Annemarie Schrödinger. Erinnerungen an den Nobelpreis. (German) [Memories of the Nobel Prize]. In Erwin Schrödinger 1887– 1961: Dokumente, Materialien und Bilder [KDK87], pages 84–?? ISBN 3-900538-09-3. LCCN QC16.S265 D65 1987. An exhibition of the Österreichischen Zentralbibliothek für Physik. With contributions by Nándor Balázs and others.

Schrodinger:1987:BLP

[Sch87b] Erwin Schrödinger. Beruht Leben auf physikalischen Gesetzen?. (German) [Is life based on physical laws?]. In B.-O. Küppers, editor, Leben = Physik + Chemie? Das Lebendige aus der Sicht bedeutender Physiker. (German) [Life = physics + chemistry? The living from the perspective eminent physicist], volume 599 of Serie Piper, pages 73–83. Piper, München, Germany; Zürich, Switzerland, 1987.

Schrodinger:1987:RKS

[Sch87c] Erwin Schrödinger. Die Raupe kann den Schmetterling nicht verstehen: [100. Geburtstag Erwin Schrödinger]. (German) [The caterpillar can not understand the butterfly [100th birthday of Erwin Schrödinger]]. In Anonymous [Ano87c], page ?? ISBN ???? LCCN ????

Schrodinger:1987:VDM

[Sch87d] Erwin Schrödinger. La mia visione del mondo. La mia vita. Poesie: [il pensiero filosofico e gli scritti inediti di un protagonista della fisica contemporanea]. (Italian) [My vision of the world. My life. Poems: [philosophical thought and unpublished writings of a leading figure in contemporary physics]]. Garzanti, Milano, Italy, 1987. ISBN ???? ???? pp. LCCN ???? Edited by Bruno Bertotti.

Schrodinger:1987:P

[Sch87e] Erwin Schrödinger. Parabel. Physikalische Blätter, 43(8):335, August 1987. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-

3722 (electronic). URL http://onlinelibrary.wiley.com/doi/
10.1002/phbl.19870430808/abstract.

Schrodinger:1987:PGPa

[Sch87f] Erwin Schrödinger. Parabel. (German) [Parable]. In P. Király and M. Ziegler-Nárayné, editors, In memoriam. Lajos Jánossy — 75. Erwin Schrödinger — 100, volume XIII. MTA Központi Fizikai Kutató Intézete, Budapest, Hungary, 1987. Poems of E. Schrödinger.

Schrodinger:1987:PGPb

[Sch87g] Erwin Schrödinger. Parabel. (German) [Parable]. In P. Király and M. Ziegler-Nárayné, editors, Jánossy Lajos (1912–1978) és Erwin Schrödinger (1887–1961) levelezése. (Hungarian) [Janossy Louis (1912–1978) and Erwin Schrödinger (1887–1961) correspondence]. MTA Központi Fizikai Kutató Intézete, Budapest, Hungary, 1987.

Schrodinger:1987:WSJ

[Sch87h] Erwin Schrödinger. Waga sekaikan (jiden) (Japanese). Kyôritsu Shuppan, Tokyo, Japan, 1987. ISBN 4-320-03240-3. ix + 258 pp. LCCN ????

Schweber:1987:AHD

[Sch87i] S. S. Schweber. An abundance of historical detail on quantum mechanics: The Historical Development of Quantum Theory, Jagdish Mehra and Helmut Rechenberg. Physics Today, 40 (9):83–84, September 1987. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.

org/resource/1/phtoad/v40/i9/p83_s1.

Schrodinger:1988:CCV

[Sch88a] Erwin Schrödinger. Che cosè la vita? Scienza e umanesimo. (Italian) [What is life? Science and humanism]. Sansoni, Firenze, Italy, 1988. ISBN ???? ???? pp. LCCN ????

Schrodinger:1988:MMO

[Sch88b] Erwin Schrödinger. Mémoires sur la mécanique ondulatoire. (French) [Memoirs of quantum mechanics]. J. Gabay, Sceaux, France, 1988. ISBN 2-87647-048-9 (paperback). xxvi + 234 pp. LCCN ???? Translated from German by A. Proca.

Schrodinger:1988:MCD

[Sch88c] Erwin Schrödinger. Mi concepción del mundo, seguido de Mi vida. (Spanish) [My View of the World, followed by My Life]. Tusquets Ed., Barcelona, Spain, 1988. ISBN 84-7223-462-2. 157 pp. LCCN ????

Schrodinger:1989:NGG

[Sch89a] Erwin Schrödinger. Die Natur und die Griechen. (German) [Nature and the Greeks], volume 21781 of Diogenes-Taschenbuch. Diogenes, Zürich, Switzerland, 1989. ISBN 3-257-21781-1. 171 pp. LCCN ????

Schrodinger:1989:GMG

[Sch89b] Erwin Schrödinger. Geist und Materie. (German) [Mind and Matter], volume 21782 of Diogenes Taschenbuch. Diogenes, Zürich, Switzerland, 1989. ISBN 3-257-21782-X. 146 pp. LCCN MLCS 95/10976 (Q).

Schrodinger:1989:GKG

[Sch89c] Erwin Schrödinger. Glückliche Kindheit. (German) [Happy childhood]. In SOS-Kinderdorf-Jahrbuch, pages 12–13. SOS-Kinderdorf, Innsbruck, Austria, 1989.

Schrodinger:1989:HGA

[Sch89d] Erwin Schrödinger. Herbst. (German) [Autumn]. In W. Moore, editor, Schrödinger, life and thought, page 444. Cambridge University Press, Cambridge, UK, 1989.

Schrodinger:1989:MLM

[Sch89e] Erwin Schrödinger. Mein Leben, meine Weltansicht. (German) [My Life, My World View], volume 21783 of Diogenes-Taschenbuch. Diogenes, Zürich, Switzerland, 1989. ISBN 3-257-21783-8. 181 pp. LCCN ????

Schrodinger:1989:PZL

[Sch89f] Erwin Schrödinger. Parabel; Zürich; Liebeslied; Geborgen; Der Entäuschte; On the shore; Juni. In W. Moore, editor, Schrödinger, life and thought, pages 6, 151, 408, 410, 413, 419–420, 447. Cambridge University Press, Cambridge, UK, 1989. Poems with English translations.

Schrodinger:1989:ST

[Sch89g] Erwin Schrödinger. Statistical thermodynamics. Dover, New York, NY, USA, 1989. ISBN 0-486-66101-6. 95 pp. LCCN QC311.5 .S36 1989. US$3.95. URL http://www.loc.gov/catdir/

description/dover032/89016929.html.

Schrodinger:1989:ZGT

[Sch89h] Erwin Schrödinger. Zittern. (German) [Tremor]. In W. Moore, editor, Schrödinger, life and thought, page 422. Cambridge University Press, Cambridge, UK, 1989. Unpublished poem.

Schrodinger:1989:VUB

[Sch89i] Erwin Schrödinger. Zur Verteidigung der universalen Betrachtungsweise. (German) [In defense of the universal approach]. In K. Stadler, editor, Lust am Forschen. Ein Lesebuch zu den Naturwissenschaften. (German) [Joy of research. A book on natural science], volume 1050 of Serie Piper. Piper, München, Germany; Zürich, Switzerland, 1989.

Schinagl:1990:DIE

[Sch90] Wolfgang Schinagl. Dokumentation und Information: Erwin Schrödingers Beiträge zu Philosophie, Wissenschaftstheorie und Weltanschauung. Berichte zur Wissenschaftsgeschichte, 13(2):115– 119, ???? 1990. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Schrodinger:1992:NGF

[Sch92a] Erwin Schrödinger. La nature et les Grecs. (French) [Nature and the Greeks]. Éditions du Seuil, Paris, France, 1992. ISBN 2-02-012800-4 (paperback). 218 pp. LCCN ???? Notes and translation from English by Michel Bitbol and Annie Bitbol-Hespériés, preceded by La clôture de la représentation by Michel Bitbol.

Schrodinger:1992:PQR

[Sch92b] Erwin Schrödinger. Physique quantique et représentation du monde. (French) [Quantum physics and the model of the world]. Éditions du Seuil, Paris, France, 1992. ISBN 2-02-013319-9 (paperback). 184 pp. LCCN ???? Introduction and notes by Michel Bitbol.

Schrodinger:1992:WLP

[Sch92c] Erwin Schrödinger. What is life?: the physical aspect of the living cell; with, Mind and matter; and Autobiographical sketches. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-42708-8 (paperback). viii + 184 pp.

LCCN QH331 .S357 1992. URL http://www.loc.gov/catdir/
description/cam024/91028415.htm;   http://www.loc.gov/
catdir/toc/cam028/91028415.htm. Foreword by Roger Penrose.

Schrodinger:1993:SRZ

[Sch93] Erwin Schrödinger. Die Struktur der Raum-Zeit. (German) [The structure of space–time]. Wiss. Buchges., Darmstadt, Germany, 1993. ISBN 3-534-02282-3. viii + 136 pp. LCCN ????

Schrodinger:1994:MMA

[Sch94a] E. Schrödinger. ‘Moe mirovozzrenie’ + Austrian philosopher on philosophy and quantum theory. Voprosy Filosofii, ??(10):68–89, ???? 1994. ISSN 0042-8744.

Schrodinger:1994:MMQ

[Sch94b] E. Schrödinger. ‘Moe mirovozzrenie’ + quantum mechanics, metaphysics and philosophy. Voprosy Filosofii, ??(9):70–94, ???? 1994. ISSN 0042-8744.

Schrodinger:1994:RCT

[Sch94c] E. Schrödinger. On the relationship of 4-color theory to 3-color theory. Color Research and Application, 19(1):40–47, ???? 1994. ISSN 0361-2317. Reprinted from Sitzungsber Kaiserl Akad Wiss Wien, Vol 134, Pg 471, 1925 [Sch25g].

Schrodinger:1995:IQM

[Sch95a] Erwin Schrödinger, editor. The interpretation of quantum mechanics: Dublin seminars (1949–1955) and other unpublished essays. Ox Bow Press, Woodbridge, CT, USA, 1995. ISBN 1-881987-08-6, 1-881987-09-4 (paperback). vi + 151 pp. LCCN QC174.12 .S35 1995. With introduction by Michel Bitbol.

Schrodinger:1995:PFR

[Sch95b] Erwin Schrödinger. The possibility of fulfillment of the relativity requirement in classical mechanics. In Barbour and Pfister [BP95], pages 147–158. ISBN 0-8176-3823-7, 3-7643-3823-7 (paperback). LCCN QC137 .M33 1995. URL http://www.

gbv.de/dms/bowker/toc/9780817638238.pdf;   http://www.
zentralblattmath.org/zmath/en/search/?an=0846.01008.

Schirmacher:1996:GES

[Sch96a] Wolfgang Schirmacher, editor. German Essays on Science in the 20th Century, volume 82 of The German library. Continuum, New York, NY, USA, 1996. ISBN 0-8264-0746-3 (hardcover), 0-8264-0747-1 (paperback). xviii + 314 pp. LCCN QC9.G3 G47 1996. Contains works by Albert Einstein, Werner Heisenberg, Max Planck, and others.

Schrodinger:1996:GSQ

[Sch96b] Erwin Schrödinger. Die gegenwärtige Situation in der Quantenmechanik. (German) [The present situation in quantum mechanics]. In Neuser [Neu96], pages 21–. ISBN 3-86025-366-2. LCCN ???? Foreword by Carl Friedrich von Weizsäcker and introduction by Wolfgang Neuser.

Schrodinger:1996:NGS

[Sch96c] Erwin Schrödinger. Nature and the Greeks, and, Science and humanism. Canto. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-57550-8. x + 172 pp. LCCN ???? Foreword by Roger Penrose.

Schrodinger:1996:SAP

[Sch96d] Erwin Schrödinger. Science, art, and play. In Schirmacher [Sch96a], pages 24–32. ISBN 0-8264-0746-3 (hardcover), 0-8264-0747-1 (paperback). LCCN QC9.G3 G47 1996. Contains works by Albert Einstein, Werner Heisenberg, Max Planck, and others.

Schrodinger:1997:WNB

[Sch97] Erwin Schrödinger. Was ist ein Naturgesetz?: Beiträge zum naturwissenschaftlichen Weltbild. (German) [What is a law of nature?: Contributions to a scientific world view]. R. Oldenbourg, München, Germany, fifth edition, 1997. ISBN 3-486-46275-X, 3-486-56293-2. 146 pp. LCCN ????

Schrodinger:1998:WMV

[Sch98] Erwin Schrödinger. Was ist Materie?: Vortrag aus dem Jahre 1952. (German) [What is matter? Lecture from the year 1952]. ????, ????, 1998. ISBN ???? ???? pp. LCCN ???? Präsentiert von Reinhard Schlögl; am 20.1.1998 im großen Hörsaal des Instituts für Experimentalphysik; Einführung: Prof. Flamm [Presented by Reinhard Schlögl on 01.20.1998 in the main auditorium of the Institute of Experimental Physics. Introduction: Prof. Flamm].

Schrodinger:1999:FIW

[Sch99a] Erwin Schrödinger. The fundamental idea of wave mechanics. Resonance, 4(2):92–103, ???? 1999. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X (electronic). URL http://www.

springerlink.com/content/x25558467pq78238/fulltext.pdf.

Schrodinger:1999:WLL

[Sch99b] Erwin Schrödinger. Was ist Leben?: Die lebende Zelle mit den Augen des Physikers betrachtet. (German) [What is life?: The living cell seen by the eyes of a physicist], volume 1134 of Piper, 0179-5147; Serie Piper. Piper, München, Germany, fourth edition, 1999. ISBN 3-492-21134-8. 156 pp. LCCN ???? Translated from English by L. Mazurczak. Introduction by Ernst Peter Fischer.

Schrodinger:2000:LNE

[Sch00] Erwin Schrödinger. Leben von negativer Entropie. (German) [Life with negative entropy]. In Baranzke [Bar00], pages 115–?? ISBN 3-7776-1033-X. LCCN BD450. URL http://www.gbv.de/dms/

faz-rez/FD120001130649220.pdf.

Schrodinger:2001:NMO

[Sch01] Erwin Schrödinger. La nueva mecánica ondulatoria y otros escritos. (Spanish) [The new wave mechanics and other writings], volume 15 of Clásicos del pensamiento. Biblioteca Nueva, Madrid, Spain, 2001. ISBN 84-7030-956-0. 117 pp. LCCN ???? Introduction by Juan Arana. Translation to Spanish by Xavier Zubiri and Juan Arana.

Schrodinger:2002:WMT

[Sch02] Erwin Schrödinger. Was ist Materie? Tonträger: Original-Tonaufnahmen (1952). (German) [What is Matter? Original recordings (1952)]. Suppoš, Köln, Germany, 2002. ISBN 3-932513-30-4. LCCN ???? Two CDs.

Schrodinger:2003:CPW

[Sch03] Erwin Schrödinger. Collected papers on wave mechanics: together with his four lectures on wave mechanics. AMS Chelsea Publishing, Providence, RI, USA, third edition, 2003. ISBN 0-8218-3524-6. xiii + 207 pp. LCCN QC174.2 .S3213 2003.

Schrodinger:2004:SNU

[Sch04] Erwin Schrödinger. Das Streben nach umfassendem Wissen. (German) [The quest for comprehensive knowledge]. In Stadler [Sta04], pages 74–?? ISBN ???? LCCN ????

Schrodinger:2006:MLM

[Sch06] Erwin Schrödinger. Mein Leben, meine Weltansicht: die Autobiographie und das philosophische Testament. (German) [My life, my world view: the autobiography and the philosophical testament], volume 34273 of Dtv. Dt. Taschenbuch-Verl., München, Germany, 2006. ISBN 3-423-34273-0. 181 pp. LCCN ???? EUR 10.00.

URL http://www.zentralblatt-math.org/zmath/en/search/?

an=1107.01009. Foreword by Auguste Dick.

Schrodinger:2008:MVW

[Sch08a] Erwin Schrödinger. My view of the world. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-09048-2. viii + 109 pp. LCCN ????

Schrodinger:2008:WNB

[Sch08b] Erwin Schrödinger. Was ist ein Naturgesetz?: Beiträge zum naturwissenschaftlichen Weltbild. (German) [What is a law of nature?: Contributions to a scientific world view]. Scientia nova. R. Oldenbourg, München, Germany, sixth edition, 2008. ISBN 3-486-58671-8. 146 pp. LCCN ???? EUR 24.80.

Schrodinger:2010:EU

[Sch10] Erwin Schrödinger. Expanding universes. Cambridge University Press, Cambridge, UK, 2010. ???? pp. LCCN QB500 .S36.

Schrodinger:2012:WLP

[Sch12] Erwin Schrödinger. What is life?: the physical aspect of the living cell; with, Mind and matter; and Autobiographical sketches. Canto classics. Cambridge University Press, Cambridge, UK, 2012. ISBN 1-107-60466-4 (paperback). viii + 184 pp. LCCN QH331 .S357 2012.

Sciama:1961:IES

[Sci61] D. W. Sciama. On the interpretation of the Einstein–Schrödinger unified field theory. Journal of Mathematical Physics, 2(4):472– 477, April 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v2/i4/p472_s1.

Scott:1967:BRE

[Sco67a] William T. Scott. Book review: Erwin Schrödinger: My View of the World. Translated by Cecily Hastings from the German Meine Weltansicht (1961). Isis, 58(1):141–143, Spring 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228418.

Scott:1967:ESI

[Sco67b] William T. (William Taussig) Scott. Erwin Schrödinger: an introduction to his writings. University of Massachusetts Press, Amherst, MA, USA, 1967. xv + 175 pp. LCCN QC16.S265 S37.

Seaton:1987:QTA

[Sea87] M. J. Seaton. Quantum theory and astronomy. In Kilmister [Kil87b], pages 213–224. ISBN 0-521-37929-6 (paperback), 0-521-34017-9. LCCN QC16.S265 S36 1987; Q143.S3. URL http://www.

loc.gov/catdir/enhancements/fy0901/86028338-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0901/86028338-t.html.

Papers presented at a conference, March 31–April 3, 1987, at Imperial College, London.

Segre:1985:HPR

[Seg85] Emilio Segrè. Historical perspective: Refugee scientists and nuclear energy. Annals of the New York Academy of Sciences, 452(1):xv– xix, 1985. CODEN ANYAA9. ISBN 0-89766-298-9, 0-89766-299-7 (paperback). ISSN 0077-8923 (print), 1749-6632 (electronic). Sixth International Conference on Collective Phenomena: reports from the Moscow Refusnik Seminar / edited by Inga Fischer-Hjalmars and Joel L. Lebowitz. Contributions from the Moscow Refusnik Seminar and from two International Conferences on Collective Phenomena, one held in Stockholm, Sweden, 1–2 December 1983, and the other in Tel Aviv, Israel, 31 May–1 June 1984.

Seiden:1961:RBM

[Sei61] Rudolph Seiden. Review: Geist und Materie by Erwin Schrödinger. Books Abroad, 35(2):169, ???? 1961. CODEN ???? ISSN 0006-

7431. URL http://www.jstor.org/stable/40115667.

Seiders:1996:SCH

[Sei96] Barbara A. B. Seiders. Schrödinger’s cat at hand. Science, 273 (5277):860, August 16, 1996. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/273/5277/860.2.full.pdf.

Seife:2005:DDP

[Sei05] Charles Seife. Do deeper principles underlie quantum uncertainty and nonlocality? Science, 309(5731):98, July 1, 2005. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/309/5731/

98.summary. See comments [WZ51].

Selleri:2002:TES

[Sel02] Franco Selleri. 24. Theories equivalent to Special Relativity: Vortrag am 14. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Sesma:2013:ESS

[Ses13] J. Sesma. Exact solution of the Schrödinger equation with a Lennard-Jones potential. Journal of Mathematical Chemistry, 51 (7):1881–1896, August 2013. CODEN JMCHEG. ISSN 0259-9791 (print), 1572-8897 (electronic). URL http://link.springer.

com/article/10.1007/s10910-013-0189-9.

Sloan:2011:CPB

[SF11] Phillip R. Sloan and D. Brandon Fogel, editors. Creating a physical biology: the Three-Man Paper and early molecular biology. University of Chicago Press, Chicago, IL, USA and London, UK, 2011. ISBN 0-226-76782-5 (hardcover), 0-226-76783-3 (paperback). ix + 319 pp. LCCN QH506 .C73 2011.

Schrodinger:1941:ES

[SH42] Erwin Schrödinger and James Hamilton. Exchange and spin. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 47(??):39–52, ???? 1941–1942. CODEN PRI-

AAK. ISSN 0035-8975. URL http://www.jstor.org/stable/
20488433.

Schrodinger:1951:SNS

[SH51] Erwin Schrödinger and Otto Hittmair. Studies in the Non-Symmetric Generalization of the Theory of Gravitation I, volume 6 of Communications of the Dublin Institute for Advanced Studies. Series A, Theoretical physics. Institiúid Ard-Léinn Bhaile Átha Cliath (Dublin Institute for Advanced Studies), Dublin, Ireland, ???? 1951. 28 pp. LCCN ????

Shimony:1983:RPB

[Shi83] Abner Shimony. Reflections on the philosophy of Bohr, Heisenberg, and Schrödinger. In Cohen and Laudan [CL83], chapter 11, pages 209–221. ISBN 94-009-7055-2, 94-009-7057-9 (e-book). ISSN 0068-

0346. LCCN ???? URL https://link.springer.com/chapter/
10.1007/978-94-009-7055-7_11.

Shimony:1986:RPB

[Shi86] Abner Shimony. Reflections on the philosophy of Bohr, Heisenberg, and Schrödinger. In Cohen and Wartofsky [CW85], chapter 20, pages 305–317. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN

0068-0346. LCCN ???? URL https://link.springer.com/
chapter/10.1007/978-94-009-5345-1_20.

Shimony:1999:PEP

[Shi99] Abner Shimony. 1. Philosophical & experimental perspectives on quantum physics: 3. Sept. 1998. In Greenberger et al. [GRZ99], page ?? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN Q44.4 V662 v. 7.

Shimony:2002:RRB

[Shi02] Abner Shimony. 2. Recollections and reflections on Bell’s Theorem: Vortrag am 10. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Sigurdsson:1996:PLC

[Sig96] Skúli Sigurdsson. Physics, life, and contingency: Born, Schrödinger, and Weyl in exile. In Ash and Söllner [AS96], pages 48–70. ISBN 0-521-49741-8 (hardcover). LCCN E184.G3 F73

1996. URL http://www.loc.gov/catdir/description/cam027/
95024894.html;   http://www.loc.gov/catdir/samples/cam031/
95024894.html;   http://www.loc.gov/catdir/toc/cam026/
95024894.html.

Simon:2003:SWE

[Sim03] Christoph Simon. 14. Towards quantum superpositions of a mirror: 7. September 2003. In Anonymous [Ano03c], page ?? ISBN ???? LCCN ????

Schmidt-Kaler:2003:SWE

[SK03] Ferdinand Schmidt-Kaler. 7. Quantum logic using ion crystals: 6. September 2003. In Anonymous [Ano03c], page ?? ISBN ???? LCCN ????

Schrodinger:1987:DMB

[SKD^{+}87] Erwin Schrödinger, Gabriele Kerber, Auguste Dick, Wolfgang Kerber, et al., editors. Dokumente, Materialien und Bilder zur 100. Wiederkehr des Geburtstages von Erwin Schrödinger. (German) [Documents, materials and pictures on the 100th return of the birthday of Erwin Schrödinger]. Fassbaender, Vienna, Austria, 1987. ISBN 3-900538-09-3. 158 pp. LCCN QC16.S265 D65 1987.

Sommerfeld:1935:VNA

[SM35] Arnold Sommerfeld and A. W. Maue. Verfahren zur näherungsweisen Anpassung einer Lösung der Schrödinger- an die Dirac-Gleichung. (German) [Method for the approximate adaptation of a solution of the Schrödinger equation to the Dirac equation]. Annalen der Physik (1900), 414(7):629–642, ???? 1935. ISSN 0003-3804 (print), 1521-3889 (electronic).

Schrodinger:1989:WLL

[SM89] Erwin Schrödinger and Ludwig Mazurczak. Was ist Leben?: Die lebende Zelle mit den Augen des Physikers betrachtet. (German) [What is life?: The living cell seen by the eyes of a physicist], volume 1134; 1134 of Serie Piper. Piper, München, Germany, 1989. ISBN 3-492-11134-3. ISSN 0179-5147. 156 pp. LCCN ???? Translation from English by L. Mazurczak. Introduction by Ernst Peter Fischer.

Smith:2023:EWT

[Smi23] Leslie Smith. Electron wave trajectories within Schrödinger’s hydrogen atom, and relativistic consequences. Foundations of Physics, 53(5):??, October 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-023-00722-w.

Schrodinger:1935:SHTa

[SMJ35a] Erwin Schrödinger, James Vincent Murphy, and Walter Henry Johnston. Science and the Human Temperament. G. Allen and Unwin Ltd., London, 1935. 151 pp. LCCN Q171 .S3 1935a. Foreword by Lord Rutherford of Nelson.

Schrodinger:1935:SHTb

[SMJ35b] Erwin Schrödinger, James Vincent Murphy, and Walter Henry Johnston. Science and the Human Temperament. W. W. Norton & Co., New York, NY, USA, 1935. xxiv + 1 + 27–192 pp. LCCN ???? Foreword by Lord Rutherford of Nelson.

Sofronieva:2014:ESP

[Sof14] Tzveta Sofronieva. Erwin Schrödinger’s poetry. Science & Education (Springer), 23(3):655–672, March 2014. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Schrodinger:1929:ASE

[SP29] Erwin Schrödinger and Max Planck. Antrittsrede von Schrödinger. Erwiderung von Planck [Sonderdruck]. (German) [Inaugural speech by Schrödinger. Reply of Planck [reprint]]. Sitzungsberichte der Preussischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):??, ???? 1929. CODEN SPWPAI. ISSN 0371-2435.

Spaldin:2005:SM

[Spa05] Nicola Spaldin. Schrödinger’s mousetrap. Nature, 434(7029):25, March 2, 2005. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v434/n7029/full/434025a.html.

Sprague:1955:RBG

[Spr55] Rosamond Kent Sprague. Review: Nature and the Greeks by Erwin Schrödinger. Classical Philology, 50(1):84–85, January 1955. CODEN ???? ISSN 0009-837X (print), 1546-072X (electronic).

URL http://www.jstor.org/stable/266148.

Sanchez-Ron:1992:MMW

[SR92] J. M. Sanchez-Ron. A man of many worlds: Schrödinger and Spain. In Bitbol and Darrigol [BD92], pages 9–24. ISBN 2-86332-116-1. LCCN QC16.S265 E78 1992. 263.75F.

Sychrava:1953:PKP

[SS53] Lev Sychrava and Erwin Schrödinger. Poznámky ke knize profesora Erwina Schroedingera “Věda a humanism” (Cambridge University Press, 1952). (Czech) [Notes on the book by Professor Erwin Schrödinger The Veda of a physicist]. Cambridge University Press, Cambridge, UK, 1953. 106 pp. LCCN ????

Schlogl:1997:WTP

[SS97] Reinhard Schlögl and Erwin Schrödinger. Wellen und Teilchen — Pionierleistungen der modernen Physik: Portrait des Nobelpreisträgers Erwin Schrödinger zum 110. Geburtstag. (German) [Waves and particles — pioneering achievements of modern physics: Portrait of the Nobel laureate Erwin Schrödinger on his 110th birthday]. Radiosendung 19.8.1997 [Radio broadcast 19.8.1997]., August 19, 1997.

Simon:1965:ESH

[ST65] J. M. Simon and J. G. Taylor. Existence of the Schrödinger and Heisenberg pictures. Nature, 205(4978):1305–1306, March 27, 1965. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v205/
n4978/pdf/2051305a0.pdf.

Stark:1936:PLD

[Sta36] Johannes Stark. Philipp Lenard als deutscher Naturforscher. (German) [Philipp Lenard as German natural scientist]. Nationalsozialistische Monatshefte, 71(??):106–111, January 1936.

Stadler:2004:WDA

[Sta04] Klaus Stadler, editor. Die wichtigsten Denkanstöße für alle, die mehr wissen wollen. (German) [The most important food for thought for those who want to know more]. ????, München, Germany, 2004. ISBN ???? ???? pp. LCCN ????

Stewart:1991:SC

[Ste91] Ian Stewart. Schrödinger’s catflap. Nature, 353(6343):384–385, October 3, 1991. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v353/n6343/pdf/353384a0.pdf.

Steinberger:2002:WOJ

[Ste02a] Jack Steinberger. 1. Welcome and opening: Jack Steinberger: Personal recollections; Mary Bell: Reminiscences of John Bell. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Stenholm:2002:IMH

[Ste02b] Stig Stenholm. 10. information and meaning — how physical are they?: Vortrag am 11. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Stenholm:2003:SWE

[Ste03] Stig Stenholm. 9. Time reversal in dynamic quantum systems: 6. September 2003. In Anonymous [Ano03c], page ?? ISBN ???? LCCN ????

Stewart:2005:SM

[Ste05] Ian Stewart. Schrödinger’s mousetrap. Nature, 433(7023):200–201, January 19, 2005. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v433/n7023/full/433200a.html.

Sticker:1960:RBG

[Sti60] Bernhard Sticker. Review: Materie und Geist (Die Wissenschaft Bd. 113) by Erwin Schrödinger. Sudhoffs Archiv für Geschichte der Medizin und der Naturwissenschaften, 44(1):96, ???? 1960. CODEN ???? ISSN 0365-2610. URL http://www.jstor.org/

stable/20774642.

Strickland:2011:WSC

[Str11] Jeffrey Strickland. Weird scientists — the creators of quantum physics. Lulu.com, ????, 2011. ISBN 1-257-97624-9. LCCN ????

Stuewer:2018:AIN

[Stu18] Roger H. Stuewer. The Age of Innocence: Nuclear Physics Between the First and Second World Wars. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2018. ISBN 0-19-186658-X, 0-19-882787-3 (hardback), 0-19-256290-8 (e-book). xv + 484 pp. LCCN QC773 .S78 2018.

Sutherland:2002:PCA

[Sut02] David Sutherland. 20. Precursors of the chiral anomaly: Vortrag am 13. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Sutcliffe:2008:WMK

[Sut08] Brian T. Sutcliffe. What mathematicians know about the solutions of Schrödinger Coulomb Hamiltonian. Should chemists care? Journal of Mathematical Chemistry, 44(4):988–1008, November 2008. CODEN JMCHEG. ISSN 0259-9791 (print), 1572-8897 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10910-008-9358-7.

Schrodinger:2011:EGA

[SvM11] Erwin Schrödinger and Karl von Meyenn. Eine Entdeckung von ganz außerordentlicher Tragweite Schrödingers Briefwechsel zur Wellenmechanik und zum Katzenparadoxon, Band 1. (German) [A discovery of quite extraordinary importance: Schrödinger’s letters on quantum mechanics and on the Cat Paradox, volume 1]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 3-642-04335-6, 3-642-04334-8. xix + 508, 881–929 pp. LCCN ????

Schrodinger:1935:BGM

[SvMFT35] E. Schrödinger, R. v. Mises, W. Frankenburger, and E. Tiede. Besprechungen. (German) [Meetings]. Naturwissenschaften, 23 (29):516–518, July 1935. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://www.springerlink.

com/content/h2u78015q5735847/fulltext.pdf.

Schrodinger:1932:SAS

[SW32] Erwin Schrödinger and V. F. Weisskopf. Sommerfeld, A.: Atombau und Spektrallinien, 5. Auflage [Book review]. (German) [Sommerfeld, A., Atomic structure and spectral lines, Fifth edition]. Naturwissenschaften, 20(19):332, May 1932. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Schrodinger:1932:BGM

[SWK^{+}32] E. Schrödinger, V. Weisskopf, F. Kirchner, A. Prey, H. Hartinger, et al. Besprechungen. (German) [Meetings]. Naturwissenschaften, 20(19):332–335, May 1932. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://www.springerlink.

com/content/p02657j78m3267u5/fulltext.pdf.

Symonds:1987:SWL

[Sym87] Neville Symonds. Schrödinger and What is Life? Nature, 327 (6124):663–664, June 25, 1987. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v327/n6124/pdf/327663b0.pdf.

Taubes:1996:SAD

[Tau96] Gary Taubes. Schizophrenic atom doubles as Schrödinger’s cat — or kitten. Science, 272(5265):1101–0, May 24, 1996. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:

//www.sciencemag.org/content/272/5265/1101.full.pdf.

Temkin:1954:RBG

[Tem54] Owsei Temkin. Book review: Nature and the Greeks by Erwin Schrödinger. Isis, 45(4):389–390, December 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226788.

Teresi:1990:BRL

[Ter90] Dick Teresi. Book review: The Lone Ranger of quantum mechanics: Walter Moore, Schrödinger: Life and Thought, New York:

Cambridge University Press. New York Times, ??(??):??, January 7, 1990. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X,

1553-8095. URL http://www.nytimes.com/1990/01/07/books/
the-lone-ranger-of-quantum-mechanics.html.

Tesche:1986:SCR

[Tes86] C. D. Tesche. Schrödinger’s cat: A realization in superconducting devices. Annals of the New York Academy of Sciences, 480(??): 36–50, December 30, 1986. CODEN ANYAA9. ISSN 0077-8923 (print), 1749-6632 (electronic).

Tesche:2000:SC

[Tes00] Claudia Tesche. Schrödinger’s cat is out of the hat. Science, 290(5492):720–721, October 27, 2000. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/290/5492/720.

Hooft:2002:QMD

[tH02] Gerard ’t Hooft. 9. Quantum mechanics and determinism at the Planck scale: Vortrag am 11. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Thirring:1987:CSE

[Thi87] W. E. Thirring. Consequences of the Schrödinger equation for atomic and molecular physics. In Kilmister [Kil87b], pages 65– 84. ISBN 0-521-37929-6 (paperback), 0-521-34017-9. LCCN QC16.S265 S36 1987; Q143.S3. URL http://www.loc.gov/

catdir/enhancements/fy0901/86028338-d.html;   http://www.
loc.gov/catdir/enhancements/fy0901/86028338-t.html. Pa-

pers presented at a conference, March 31–April 3, 1987, at Imperial College, London.

Thirring:1999:MPS

[Thi99] Walter Thirring. 11. Macroscopic purification of states by interactions: 4. Sept. 1998. In Greenberger et al. [GRZ99], page ?? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN Q44.4 V662 v. 7.

Thirring:2011:ESP

[Thi11] Walter Thirring. Erwin schrödinger – personal reminiscences. Mitschnitt einer gemeinsamen Veranstaltung des Erwin Schrödinger-Instituts für Mathematische Physik (ESI) und der Fakultät für Physik der Universität Wien am Donnerstag, dem 13. Jänner 2011 im Kleinen Festsaal der Universität Wien [Recording of a joint event at the Erwin Schrödinger Institute for Mathematical Physics (ESI) and the Department of Physics, University of Vienna on Thursday, the 13th January 2011 in the Small Hall of the University of Vienna]., 2011.

Thirring:2013:ESP

[Thi13] Walter Thirring. Erwin Schrödinger — personal reminiscences. In Reiter and Yngvason [RY13], pages 1–8. ISBN 3-03719-121-X (paperback), 3-03719-621-1 (e-book). LCCN QC16.S265 I58 2011.

Thorndike:1955:BRE

[Tho55] A. M. Thorndike. Book review: Erwin Schrödinger, Nature and the Greeks. Physics Today, 8(2):19, February 1955. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/?PTO/8/19/1.

Thomson:1966:RBV

[Tho66] G. P. Thomson. Review: My View of the World by E. Schrödinger. British Journal for the Philosophy of Science, 16(64):348–349, February 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-

3537 (electronic). URL http://www.jstor.org/stable/686840.

Trimmer:1980:PSQ

[Tri80] John D. Trimmer. The present situation in quantum mechanics: a translation of Schrödinger’s “cat paradox” paper. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 124(5):323–338, October 10, 1980. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL http://www.jstor.org/stable/986572. English transla-

tion of [Sch35f, Sch35a, Sch35b].

Thirring:1977:SEP

[TU77] Walter E. Thirring and Paul Oskar Urban, editors. The Schrödinger equation: proceedings of the International Symposium “50 Years Schrödinger Equation” in Vienna, 10th–12th June, 1976, volume 17 of Acta physica Austriaca. Supplementum. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. ISBN 0-387-81437-X. LCCN QC174.26.W28 I57 1976.

Tuck:2011:MYH

[Tuc11] Lily Tuck. I married you for happiness. Atlantic Monthly Press, New York, NY, USA, 2011. ISBN 0-8021-1991-3. 193 pp. LCCN PS3570.U236 I2 2011. URL http://www.loc.gov/

catdir/enhancements/fy1205/2011276837-b.html;   http:/
/www.loc.gov/catdir/enhancements/fy1413/2011276837-d.html.

Tuthill:2005:SM

[Tut05] Peter Tuthill. Schrödinger’s mousetrap. Nature, 434(7031):277, March 16, 2005. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v434/n7031/full/434277a.html.

Terhal:2003:QEM

[TWD03] Barbara M. Terhal, Michael M. Wolf, and Andrew C. Doherty. Quantum entanglement: a modern perspective. Physics Today, 54(4):46–52, April 2003. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.

org/resource/1/PHTOAD/v56/i4. This paper discusses the ori-

gin of the concept of quantum entanglement, and the famous Schrödinger’s cat experiment, in Schrödinger’s famous three papers [Sch35f, Sch35a, Sch35b], and the later work of John Bell [Bel64]. Those papers were a response to the famous EPR paper [EPR35] that first raised the paradox described by entanglement, and that caused much debate with Bohr, Schrödinger, and others. See also [Eve57, Whe57, Vil86].

vanCalmthout:2017:RQSc

[vCK17] Martijn van Calmthout and Leo P. Kouwenhoven. Real Quanta: Simplifying Quantum Physics for Einstein and Bohr. Dundurn Press Limited, Toronto, ON, Canada, 2017. ISBN 1-4597-4049-1 (paperback), 1-4597-4050-5 (PDF), 1-4597-4051-3 (epub). 192 pp. LCCN QC174.123 .C3513 2017.

Venema:2005:SMF

[Ven05] Liesbeth Venema. Schrödinger’s mousetrap. Nature, 433(7025): 463, February 2, 2005. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v433/n7025/full/433463a.html.

Villars:1986:PSC

[Vil86] C. N. Villars. The paradox of Schrödinger’s cat. Physics Education, 21(4):232–237, July 1986. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-6552 (electronic).

Visscher:1991:FEA

[Vis91] P. B. Visscher. A fast explicit algorithm for the time-dependent Schrödinger equation. Computers in Physics, 5(6):596–??, November/December 1991. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/
10.1063/1.168415.

vonMeyenn:1992:SBA

[vM92] Karl von Meyenn. Schrödinger’s Beiträge zu Atomtheorie vor der Entstehung seiner Wellenmechanik. (German) [Schrödinger’s contributions to atomic theory before the creation of his wave mechanics]. In Bitbol and Darrigol [BD92], pages 193–236. ISBN 2-86332-116-1. LCCN QC16.S265 E78 1992. 263.75F.

Wilber:1987:CCE

[W^{+}87] Ken Wilber et al., editors. Cuestiones cuánticas escritos místicos de los físicos más famosos del mundo. (Spanish) [Quantum issues: mystical writings of the world’s most famous physicists]. Kairós, Barcelona, Spain, 1987. ISBN 84-7245-172-0. 298 pp. LCCN ????

Wall:1954:BRE

[Wal54] C. N. Wall. Book review: Erwin Schrödinger, Statistical Thermodynamics. American Journal of Physics, 22(5):347, May 1954. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://link.aip.org/link/?AJP/22/347/1.

Wolfson:2000:ERQ

[WB00] Richard Wolfson and Andreas Burgstaller. Einstein’s relativity and the quantum revolution: modern physics for non-scientists. Great courses, Science and mathematics. Teaching Computers, Chantilly, VA, USA, second edition, 2000. ISBN 1-56585-565-5. LCCN QC173.58 .W65 2000d. US$254.95. 4 videodiscs (720 min.).

Wilber:2002:CCE

[WdC02] Ken Wilber and Pedro de Casso. Cuestiones cuánticas escritos místicos de los físicos más famosos del mundo. (Spanish) [Quantum issues: mystical writings of the world’s most famous physicists].

Kairós, Barcelona, Spain, seventh edition, 2002. ISBN 84-7245-172-0. 296 pp. LCCN ????

Weihs:2002:BTS

[Wei02] Gregor Weihs. 25. Bell’s theorem for space-like separation and GHZ: Vortrag am 14. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Weinberger:2014:NBD

[Wei14] P. Weinberger. Niels Bohr and the dawn of quantum theory. Philosophical Magazine, 94(27):3072–3087, August 2014. CODEN PH-MAA4. ISSN 0031-8086.

Westphal:1957:CSJ

[Wes57] Wilhelm H. Westphal. Clara von Simson 60 Jahre/Zur zehnten Wiederkehr des Todes von Max Planck/Der Pour le Mérite auch für Erwin Schrödinger. (German) [Clara von Simson 60th birthday/the tenth anniversary of the death of Max Planck/the Pour le Mérite for Erwin Schrödinger]. Physikalische Blätter, 13(10):470– 471, 1957. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Wessels:1979:SRW

[Wes79] Linda Wessels. Schrödinger’s route to wave mechanics. Studies in History and Philosophy of Science Part A, 10(4):311–340, December 1979. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368179900189.

Wessels:1983:ESD

[Wes83] Linda Wessels. Erwin Schrödinger and the descriptive tradition. In Aris et al. [ADS83], pages 254–278. ISBN 0-8166-1087-8. LCCN QC7 .S77 1983.

Wessels:1991:BRB

[Wes91] Linda Wessels. Book review: The Historical Development of Quantum Theory. Volume V: Erwin Schrödinger and the Rise of Wave Mechanics by Jagdish Mehra; Helmut Rechenberg. Isis, 82(2):404– 405, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234908.

Walsby:2017:SCS

[WH17] A. E. Walsby and M. J. S. Hodge. Schrödinger’s code-script: not a genetic cipher but a code of development. Studies in History and Philosophy of Biological and Biomedical Sciences, 63 (??):45–54, June 2017. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848616301042.

Whaley:2011:QCE

[Wha11] K. Birgitta Whaley. Quantum coherence and entanglement in biology. Mitschnitt einer gemeinsamen Veranstaltung des Erwin Schrödinger-Instituts für Mathematische Physik (ESI) und der Fakultät für Physik der Universität Wien am Donnerstag, dem 13. Jänner 2011 im Kleinen Festsaal der Universität Wien [Recording of a joint event at the Erwin Schrödinger Institute for Mathematical Physics (ESI) and the Department of Physics, University of Vienna on Thursday, the 13th January 2011 in the Small Hall of the University of Vienna]., 2011.

Wheeler:1957:AER

[Whe57] John A. Wheeler. Assessment of Everett’s “relative state” formulation of quantum theory. Reviews of Modern Physics, 29(3):463–465, July 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.29.463;   http://rmp.aps.org/abstract/
RMP/v29/i3/p463_1. See [Eve57].

Wheeler:2017:BRT

[Whe17] Nicholas Wheeler. Book review: Tales of the Quantum: Understanding Physics’ Most Fundamental Theory. Hobson, Art. 303 pp. Oxford U.P., New York, 2017. Price: $29.95 (hardcover). ISBN 978-0-19-067963-7. American Journal of Physics, 85(9):722–723, September 2017. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Whitaker:2002:EEY

[Whi02] Andrew Whitaker. 5. Education and early years: Vortrag am 10. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Wightman:1955:RBG

[Wig55] William P. D. Wightman. Review: Nature and the Greeks by Erwin Schrödinger. British Journal for the Philosophy of Science, 5(20):342–343, February 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://www.jstor.

org/stable/685739.

Wigner:1967:PM

[Wig67a] Eugene Wigner. The problem of measurement. In Symmetries and reflections; scientific essays of Eugene P. Wigner [Wig67d], page ?? LCCN Q171 .W65.

Wigner:1967:RMB

[Wig67b] Eugene Wigner. Remarks on the mind–body problem. In Symmetries and reflections; scientific essays of Eugene P. Wigner [Wig67d], page ?? LCCN Q171 .W65.

Wigner:1967:TKR

[Wig67c] Eugene Wigner. Two kinds of reality. In Symmetries and reflections; scientific essays of Eugene P. Wigner [Wig67d], page ?? LCCN Q171 .W65.

Wigner:1967:SRS

[Wig67d] Eugene Paul Wigner. Symmetries and reflections; scientific essays of Eugene P. Wigner. Indiana University Press, Bloomington, IN, USA, 1967. viii + 280 pp. LCCN Q171 .W65.

Willers:1936:BWHb

[Wil36] R. A. Willers. Buchbesprechung: W. Heisenberg, E. Schrödinger, P. A. M. Dirac, Die moderne Atomtheorie. Die bei der Entgegennahme des Nobelpreises 1933 in Stockholm gehaltenen Vorträge. 45 S. u. 6 Fig. Leipzig 1934, Verlag von S. Hirzel. Preis kart. 2,50 M. Zeitschrift für Angewandte Mathematik und Mechanik, 16(2): 126, 1936. CODEN ZAMMAX. ISSN 1521-4001.

Wilson:1999:SCT

[Wil99] Robert Anton Wilson. Schrödinger’s cat trilogy. Orbit, London, UK, 1999. ISBN 1-85723-251-8. 545 pp. LCCN ????

Williams:2016:IME

[Wil16] Nicola Williams. Irene Manton, Erwin Schrödinger and the puzzle of chromosome structure. Journal of the History of Biology, 49(3):425–459, August 2016. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (electronic). URL http://

link.springer.com/article/10.1007/s10739-015-9424-5;

http://link.springer.com/content/pdf/10.1007/s10739-015-
9424-5.pdf.

Wineland:2013:NLS

[Win13a] David J. Wineland. Nobel Lecture: Superposition, entanglement, and raising Schrödinger’s cat. Reviews of Modern Physics, 85(3):1103–1114, July 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.85.1103;   http://
rmp.aps.org/abstract/RMP/v85/i3/p1103_1.

Wineland:2013:SER

[Win13b] David J. Wineland. Superposition, entanglement, and raising Schrödinger’s cat. Annalen der Physik (Berlin), 525(10–11):739– 752, 2013. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/andp.201300736/abstract.

Woodyard:2006:NPS

[Woo06] Jack R. Woodyard. A new paradigm for Schrödinger and Kohn. In Philosophy of Chemistry: Synthesis of a New Discipline [BSM06], chapter 13, pages 245–269. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346. LCCN Q174 .B67 vol. 242

QD6. URL https://link.springer.com/chapter/10.1007/1-
4020-3261-7_13.

Weart:1985:HP

[WP85] Spencer R. Weart and Melba Phillips, editors. History of physics, volume 2 of Readings from Physics Today. American Institute of Physics, Woodbury, NY, USA, 1985. ISBN 0-88318-468-0 (paperback). 375 pp. LCCN QC7 .H694 1985.

Wrachtrup:2009:QPS

[Wra09] Jörg Wrachtrup. Quantum physics: Schrödinger’s cat is still alive. Nature Physics, 5(4):248–249, April 2009. CODEN NPAHAX. ISSN 1745-2473 (print), 1745-2481 (electronic). URL http://www.

nature.com/nphys/journal/v5/n4/full/nphys1245.html.

Woertche:1951:SNC

[WZ51] Heinrich J. Woertche and Marek Zukowski. Schrödinger’s nuclear cat. Science, 309(5731):98–??, ???? 1951. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/309/5731/98/reply. See [Sei05].

Yam:1997:BSC

[Yam97] Philip Yam. Bringing Schrödinger’s Cat to life. Scientific American, 276(6):124–129, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.nature.com/scientificamerican/journal/v276/n6/pdf/
scientificamerican0697-124.pdf.

Yang:1956:RBU

[Yan56] C. N. Yang. Review: Expanding Universes by Erwin Schrödinger. Science, 124(3217):370, August 24, 1956. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

jstor.org/stable/1749773;   http://www.sciencemag.org/
content/124/3217/369.

Yang:1987:SRM

[Yan87] Chen Ning Yang. Square root of minus one, complex phases and Erwin Schrödinger. In Kilmister [Kil87b], pages 53–64. ISBN 0-521-37929-6 (paperback), 0-521-34017-9. LCCN QC16.S265 S36 1987;

Q143.S3. URL http://www.loc.gov/catdir/enhancements/
fy0901/86028338-d.html;   http://www.loc.gov/catdir/enhancements/

fy0901/86028338-t.html. Papers presented at a conference, March 31–April 3, 1987, at Imperial College, London.

Yang:2013:SPI

[Yan13a] Chen Ning Yang, editor. Selected papers II, with commentaries. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2013. ISBN 981-4449-00-8 (hardcover), 981-4449-01-6 (paperback), 981-4449-02-4 (e-book). x +

34 pp. LCCN QC21.3. URL http://www.worldscientific.com/
worldscibooks/10.1142/8640.

Yang:2013:SRM

[Yan13b] Chen Ning Yang. Square root of minus one, complex phases and Erwin Schrödinger. In Selected papers II, with commentaries [Yan13a], pages 94–105. ISBN 981-4449-00-8 (hardcover), 981-4449-01-6 (paperback), 981-4449-02-4 (e-book). LCCN

QC21.3. URL http://www.worldscientific.com/doi/abs/10.
1142/9789814449021_0014.

Yngvason:2003:ORJ

[YKZ03] Jakob Yngvason, Gershon Kurizki, and Anton Zeilinger. 1. Opening remarks / Jakob Yngvason; Gershon Kurizki; Anton Zeilinger.

Closing remarks / Gershon Kurizki: 5. September 2003. In Anonymous [Ano03c], page ?? ISBN ???? LCCN ????

Yngvason:1999:CCL

[Yng99] Jakob Yngvason. 19. Causality concepts in local quantum physics: 5. Sept. 1998. In Greenberger et al. [GRZ99], page ?? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN Q44.4 V662 v. 7.

Yoxen:1979:WDS

[Yox79] Edward J. Yoxen. Where does Schrödinger’s What Is Life? belong in the history of molecular biology? History of Science (UK), 17 (1):17–52, March 1979. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
17/1/17.full.pdf+html.

Zegers:2005:SM

[Zeg05] Remco Zegers. Schrödinger’s mousetrap. Nature, 433(7027):689, February 16, 2005. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v433/n7027/full/433689a.html.

Zeilinger:1997:FLO

[Zei97] Anton Zeilinger. Five lectures on the occasion on Walter Thirring’s 70th birthday. 2. Experimenting in Hilbert space with photons: 1997/04/10. Report, ESI, Erwin Schrödinger International Institute for Mathematical Physics, Vienna, Austria, 1997.

Zeilinger:1999:QSI

[Zei99] Anton Zeilinger. 2. Quantum superposition and information: 4. Sept. 1998. In Greenberger et al. [GRZ99], page ?? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN Q44.4 V662 v. 7.

Zeilinger:2002:BTQ

[Zei02] Anton Zeilinger. 15. Bell’s Theorem and quantum information: Vortrag am 11. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Zichichi:2002:JBC

[Zic02] Antonino Zichichi. 7. John Bell and the 10 challenges of subnuclear physics: Vortrag am 10. November 2000. In Bertlmann and Zeilinger [BZ02], page ?? ISBN 3-540-42756-2. LCCN ????

Zoller:1999:QRC

[Zol99] Peter Zoller. 15. Quantum repeaters for communication: 5. Sept. 1998. In Greenberger et al. [GRZ99], page ?? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN Q44.4 V662 v. 7.

Zukav:1979:DWL

[Zuk79] G. Zukav. The dancing Wu Li masters: an overview of the new physics. Rider and Hutchinson, London, UK, 1979. ISBN 0-09-139400-7, 0-09-139401-5 (paperback), 0-7126-3817-2 (paperback). 352 pp. LCCN ????

Zurek:1999:DCS

[Zur99] Wojciech H. Zurek. 10. Decoherence, chaos and the second law: 4. Sept. 1998. In Greenberger et al. [GRZ99], page ?? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN Q44.4 V662 v. 7.

Zurek:2000:QCS

[Zur00] Wojciech H. Zurek. Quantum cloning: Schrödinger’s sheep. Nature, 404(6774):130–131, March 9, 2000. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v404/n6774/full/404130b0.html.

Zurek:2003:EAI

[Zur03] Wojciech Zurek. 2. Environment — assisted invariance, ignorance, and information in quantum physics: 5. September 2003. In Anonymous [Ano03c], page ?? ISBN ???? LCCN ????